A Bibliography of The American Statistician

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 October 2024 Version 1.31

Title word cross-reference

(X_{1} < X_{2} < X_{k}) [HL96]. .338 [Fre07c]. .833 [Fre07c]. 1: m [Lev97]. 1 × 1 × r [Dun03]. 2^{k} [SM79]. 2 × 2 [Che11, MW92, MH93, SVR11, Kra88]. 2 × k [GK99]. 5 [Bar02, Len01c]. −X ± t_{α}(n − 1)(s/{n}) [Zuo10]. α [HB03]. χ^{2} [SP85]. ED100p [CCW^{+}86]. F [Chr03, Gue79, Lau65, LD98, ML92, Pei93, PS98c, Sho03, ZPM^{+}11, SM79, FK86]. f′/f [JH83]. gh [HR06]. H_{0} [Bet00]. I [Pav88]. k [BP02, Tar07]. L [Hos92]. L1 [All97]. n [BHO00, Kah87]. O [KD87]. ⊗ [Hu07]. P [BS11, Don99, MTA08, SSC99, Sch96a, Ber96a, BH02, Bro10, HB03, MW92, Rya09, SBB01]. P (X > Y) [Bro10]. pk [Bec96]. pn − k [Bec96]. QR [Ans85]. R [BK92]. R^{2} [And94, Mag90, SW91]. R^{2} > r_{yx_{1}}^{2} + r_{yx_{2}}^{2} [Ham87]. s [HJM08]. S^{2} [Zeh91, Duf86]. t [BM01, Ber81, BHO00, Box81, Bro10, CA97, DG97, Eis79, GJLW05, Gro90, HJM08, KG90, LaM94, LD98, MC91, Zuo10]. U [Sch91, Len83]. X [Spi87, Spi86]. x_{2} [Sch93a]. x_{2}|x_{1} [Sch93a]. Y [Spi87, Spi86]. Z [BHO00, Eis79, HJM08, Shi88b].

-Distance [All97]. -Means [Tar07].

-statistics [KD87]. -tables [Lau65]. -Test [Bro10]. -Value [BS11]. -Values [MTA08].

1.4 [Cow04c]. 11 [Hil05d]. 11th [Mil03]. 12 [Hil03c]. 12.0 [vH04]. 12.01 [Hil04b]. 12.2 [Hil08f]. 14th [RvE97]. 1913-1994 [CR95]. 1919 [Ben93]. 19th [LR06]. 1st [Mil03].

2} [Won98]. 2.0 [Maa99]. 2000

[ESL09, WJM^{+}03]. 21st [McD99]. 240 [LDL^{+}94]. 2nd [Cle03b, Coh02, Dob02, Fra00, Gen03, Gup02, Hai03, Hol99, Hyu01, Jer03, KK99, Kir00, Kre01, Led03, Mac99, Man03, Ste01, Utt02, Wil02].

3.0 [Hil03c]. 3.1 [Hil05i, Spi90]. 38 [SYC88, Zel85]. 39 [CH86, Kar86, Kar87]. 3rd [ACA03, Can01, Hol99, Jam02, Jam03, Joh99, Mil03, Pie01].

4.0 [Hil06b]. 4.03 [Alt02]. 40 [Spi87]. 43 [Ang90, GY90]. 47 [WS94]. 48 [DN94a]. 4th [Cow04a, Gon03, Kaf99, Wal00b].

5} [Gol99, Tan99, Hil07a]. 5.0 [Mac93]. 5.1 [Hil02e]. 5.2 [Hil06a]. 5.5 [HH06]. 52 [Ano98d]. 5th [Cle00b]. 5V [KDV94].

62004a [Zel85]. 6th [Pil98].

7 [Hil07f]. 7th [Obr98].

8.0 [Yal07]. 8.02 [Hil03c]. 83 [NO98]. 85b [SYC88]. 87b [Kar86, Kar87]. 87c [Zel85].

9 [Hil07b]. 9.0 [Hil05a, Hil05c, Hil06b]. 9th [Pil98].

A. [Wan91]. Abandon [Dal90]. Absolute [Bas97, BS03, SGC90, Bly80]. Abuse [HH01]. Abuses [Pau92]. Academia [Bic95]. Academic [CRV95, Moo01]. Academics [Sta01]. accelerated [MS87]. accelerated-life-testing [MS87]. Accent [Ano91a, Ano91b, Ano92, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b]. Accepted [GKR94, PG97]. accessible [DK95]. Accident [HS03]. Account [Bro11]. Accuracy [Alt02, CLS07, KP11, TM99, WSF10]. Accurate [LCD10]. Achieving [Cle06a]. Acknowledgement [DN94a]. Across [MW92]. Action [Dem75, Lyn03, Mil04c]. Active [Kva00, Maa99]. ActivEpi [Sch04f, Sch04f]. Activities [SMC^{+}07, S^{+}90]. Activity [Die93]. ActivStats [Maa99, MJ04]. Actuarial [Wüt08]. Acyclic [HKS06]. Addelman [DS81]. Adding [AC00a, Gre10, HR10]. Additive [Hil93]. Address [CJ92]. Adjacent [DM01]. Adjust [KP07]. Adjusted [Ann06a, BM01, GCGG95, LM03, LGH09, SM91]. Adjustment [And09, BSTM09, LK99, Pie80]. Ado [HK07]. Advanced [RSW99]. Advantage [HS94]. Adventures [Cle03d, Pic06]. Advertisements [DWBQ98]. Advertising [Dan92]. Advice [RSH01, Rub04b, VM03]. Advisory [Eld90]. Afraid [Tra07]. After [ZR10]. Against [Hel95, Jam99, Sil97]. Age [BB04, Cle05a, Cle05b]. Agencies [Gal03, Eld90, Gon94]. agents [CCW^{+}86]. Aggregation [Lic90]. agis [Fox10]. Agree [Che11]. Agreeing [LS01]. Agreement [LR06, dM07]. Agricultural [ACA03]. Agriculture [Rob07a]. Aid [Lee07]. Aided [GPA97]. Alai [CG95a]. Albert [Lea04]. Alfonso [FK87]. Algebra [Bag95, Dav06, Nas95]. Algebraically [Sha91b]. Algorithm [BS95, CG95b, DM01, Gan95, Hit03, Mil96, Nav97b, Pos93, Tar07, Ans85]. Algorithms [CGL83, HMR^{+}01, HL04, LR01, LDM10, MS98]. alias [KP79]. Alignment [Faw90]. All-Time [Ann06a]. Allegations [BB04]. allocation [WC87]. Altered [KKO^{+}06]. alternate [Far89]. Alternative [BH98, Bet00, BH02, Coh97, FMA^{+}09, Han79, Hay11, Ros00a, Har76, SSM80]. Alternatives [MLL10, MKV92, Suk92]. always [Ham87]. Amelioration [MB06]. Amer [Ang90, CH86, DN94a, GY90, SYC88, Spi87, Zel85]. Amer. [Kar86, Kar87, WS94]. America [NOMLT95, Tor07]. American [Alb03c, Alb04, Alb05a, Alb05b, LDL^{+}94, Ano98b, HBF90, MMR90, SMC^{+}07, Wal05, Wio90, ZH90]. Among [Bol03, SR97, XN03, Lee86]. Amstat [DWBQ98]. Analogues [VBB94]. Analyses [Gup02, HV95, KBH09, Lac00, Lun07b, Man03, MLL10, OH08a, OH08b, PS94, WW11, WASB10, SSH81]. Analysis [AC06b, Ank98, Ann07, Atk95, Bon02, BGF^{+}08, Car00, Cav07b, CGL83, CT91, Cle99a, Cle08a, Cle08b, Coh97, CW99, Cor00, Cow04b, DD06, Dav08b, Dav08d, DLM01, Dev07a, Dev07c, Dla05, FM92, FN92, Fre08, GL00, Gen03, GGR05, GW94, Hai03, HG97, HLW09, HHB91, HH01, Hol99, Hsi95, Hu07, IM01, Jam03, Joh99, JK99, II07, KT94, KDV94, Lar92, LDL^{+}94, Len01a, Lis08, Mac01, McG05, MB06, MV94, Men00, Mil03, NK96, NL95, Nel07, Nor95, OW91, OCE93, Ost02, Ost03, OSC00, Owe01, Pei93, PH02, PR98, Pou02, Pri05b, RHCR06, Rob07b, Rue04, Rum99, Rum01, Sap91, Sea94, She98, SC99, SW90, Smi98, Tho95, Van96b, Vic08, War02, Wie06, Wil02, Woo08, Wul03, Zim07, vH04, ABS81, Alb92, Ber87, Cas85, Hoe85, Man82, Mon79, Rav85]. analysis [Sch85]. Analytic [SVR11]. Analytics [Wie08]. Analyzing [Cot98, Faw90, Hal08, Han05, Kel04, Pan02, Was06]. Animal [ES93]. annotated [Wio90]. Anomalies [God91, HBC^{+}91]. AnotherWorld [Sne08]. ANOVA [AC04b, BSTM09, BK96, Hei81, Her86, Hin96, Rog84, Saw92, SD87, Sch93c, Ver99]. ANOVA-Type [BSTM09]. Answer [BHA02, Tuk49a]. Any [O’C90]. AP [FHP^{+}11]. AppDev [Pra08b]. Appear [Wri90]. APPL [GEL01]. Apple [SMP97]. Application [DM01, DO04, DuM99, GE11, Hag09, HBC^{+}91, RS93, Rie90]. Applications [Abb02, Bar09, Ber97b, BTV90, BHO98, Dev01, DB97, Epp93, GKT98, GV09, Hai07, Hu03, LHD00, LGXB11, Lun04a, Mor07, NS99, RH02, Rid09, Sta08, Tan02, Ver01, WTH95, Wie07, ZNY11, vDK02, BH86, Hog79, Knü94, RC88, DH05]. Applied [And09, Bon98, BGF^{+}08, Cow03, Dev07a, Fot02, GH91, Gon03, Hol99, Jam02, Joh99, KK99, Kir00, MGCR10, Mis06, RT01, SM97, Sch10a, Sim07, SW90, Wan10, Wil02, Woo08, Ano00b]. Applying [LG00]. Appraise [HRS10]. Approach [Alb97b, Bri96, Chu99, Cle08b, EGCD91, FS93, FN92, GV09, Gro91, Har08, Hil96b, Lee08, Len01b, LR06, Lor90, LGH09, MSS91, Mea03, MS02, Oks91, Owe01, Ree02, Ric10, Sim07, TTT09, Van96b, Zim07, dML09, GP89, MS87, Rup87, Sha82, Fra00, MKV92]. Approaches [Men04, SB09, Gol86]. Approximate [AC98a, AW91a, BH93, DH11, FLP07, LT96, LC99, ME95, Rya09]. Approximation [BS98, CLPC08, GC99, Kuo00, PA91, Pro08, Rea98, She84]. Approximations [Huz99, JB04, LJ09, OW10, Gue79]. Arab [Bro11]. Arguments [GM96]. ARIMA [HT02, McL99]. Arising [Jon99]. Arithmetic [Ste96]. Arithmetic-Geometric-Harmonic [Ste96]. ARMA [McL99]. Arnold [Gei84]. Arrival [BGG98]. Art [Arn07, Gla04]. Articles [Pet99]. Artifacts [JSM95]. Artificial [DG97]. ASA [Ano90a, Ano90b, Nor90]. Ask [Sim98]. Aspect [BWM98]. Aspects [CS94, IJW06, JK90]. Assess [Loo95]. Assessing [BUK^{+}08, ESL09, GKT91, Haz03, McC98, McC99, PC02, Tob04]. Assessment [FS01, GZK^{+}11, Grö09, KBH09, Kra88, Rus98, MS87]. Assessments [Tob04]. Assessors [TM99]. Assisted [Mil04a, MH76]. Associated [Wil08]. Association [Lee93, Smi11, Rav86, HBF90, MMR90, SMC^{+}07, Wal05, Wio90, ZH90]. Association-Not-Causation [Smi11]. associations [Kra88]. Assumptions [AL07, HM93, SF00, SD87]. Assurance [Hyd01]. Asymmetric [BM01, DR01]. Asymmetry [DR07, CW88]. Asymptotic [Sch91, CCW^{+}86, Har80]. Atlas [SW05]. Attention [Alt02]. Auctions [HJS06]. Audio [Pos90]. Audits [MSL^{+}08]. Augmenting [CH96]. Australia [Lud05]. author [Efr86, Gei84, Zel88]. authors [PS89]. Autocovariance [Per93].

Automatic [OL96]. Autoregressive [WGHE09, YW09, ME84]. Average [RS91b, SD01, YW09]. Averages [JK90, AT86, SP86]. averaging [Kru87]. Avoid [AC04a]. Awareness [CF91, Gon94]. AWAY [BH98].

B [BI09]. B-Splines [BI09]. Babies [CS05, CP99]. Bachelor [BGNP01]. Back [SL92, Sch03c, SW90]. Back-of-the-Envelope [SL92]. Bacon [DH05]. Bad [Ver99]. Badly [Hog91]. Bag [Cle05c]. Bagged [Ber07]. Bagplot [RRT99]. Balance [GE11, HRS10, Sch93c]. Balanced [BHA02]. Balkrishna [Ful80]. Ball [Gun93]. Ballots [WJM^{+}03]. Bancroft [FK87]. Bands [Ros00a]. Baseball [Ann06a, Bor99, Coc05b, HV95, Lac90, Lea04, RMY96, SG00, WC97]. Based [Ano98j, APT07, BR97, BN98, Bas97, Bis98, BA02, CF01, CW11, Czi98, DS98, Ede90a, Fre05, Grö07, HQ08, KBH09, Kot91, LG05, LS97, Mag90, MSL^{+}08, MV94, ME95, MLL10, PS98a, Pet99, RT01, SO94, Sho95, SS98, Spu98, SV03, Var98, Wal00a, Wan97, YB10, dML09, AO89, Aso80, Bel87, Coh97]. Basic [AC03, Coc05a, Dob02, Dob06a, Low98, Mil03, Pet99, Smi98, Wal00b, War02, Wel98]. Basics [Ann05c]. Basis [Bro01, Dem75]. Basketball [Car96, Gil00, RHCR06, SMH91, SSH96, SS99, War95, Wes90]. Basu [BHO98]. Batting [Ann06a]. Bayes [Alb97a, Alb97b, Ber97a, BM88a, Cas85, Chr05, Fre97, Gel08, LS99a, Lin97, Lit06, Moo97a, Moo97b, NVB09, Pau81, Sch97, Sho97, SS02, TGM08, Wit97, Zel88]. Bayesian [CH86, ABS81, Alb92, Alb93, AR08, BM02a, BCJ97, Ber96c, Ber97b, CH85, CU92, CB95b, DD06, DiP04, DuM99, Edw86, Efr86, Fre07b, GS08, Gel98, GGR05, GJLW05, Gre10, HMR^{+}01, IBP05, Kah87, KB07, Lee08, LS97, LP76, Lon10, MR09, Mat99, NS97, Ree02, Rum06, Sho98, Sin08, SG92, UJ08, ZR10]. Bayesians [Ros96]. Be [Fre07a, FG00, Kon95, Olk92, SJD10, BvE88, BHO00, Fre07b, JW04, MA84, SS99]. Beam [DT00]. Bearing [Gun93]. Been [Tra07]. Beginner [Nel01]. Beginners [Alb97a, Ber97a, Fre97, Lin97, Moo97a, Moo97b, Rub04b, Sch97, Sho97, Wit97]. Beginnings [Dav08a]. Behavior [CL02, Mag98]. Behavioral [Kre01, Lee08]. Behrens [DO04]. Being [FS06, LSO01, WSF02]. Believed [Fre07b]. Bend [WWD03]. Benefit [Hai03]. Benford [CG07, AP10, BS10b, Few09, LSE00, Var72]. Bernard [AS86]. Bernardo [Gei84]. Bernoulli [Ber96c, LT96, LP76, NW86, VBB94, War97]. Best [Ann06a, BR97, BH98, DB97, Eli00, Gur06, Sea97, SM91, BvE88, Kot86, MA84, PS89, SS84]. Best-of-Seven [BH98]. Bet [CW93]. Beta [Bai92, Cha95, DH05, Pos93]. Beta-Binomial [DH05]. Bets [Alb02]. Better [AC98a, Fre07c, Men10, SJD10, WE08, WGHE09]. Between [BM10, DS08, DHW11, EW02, FHP^{+}11, GS06, IBP05, JK99, LL97, LDM10, Nav97a, OW91, Sap91, SG01, Sch93a, Sch96b, Sha91b, Shi01, SF93, Aso80, FK86, VM89]. Between-Group [LDM10]. Beyond [Ann05c, Bel05, DS02, MH93, Nel07, PR10, Smi11]. Bias [Bar07, Duf86, ESL09, Gad01, GN02, HLP03, JS03, KH04, MR09, Sal99, Sto98]. Biased [SP94]. Biases [Ros01, vH04]. bibliography [Wio90]. Bimodal [SWW02]. Bin [Wan97]. Binary [CW92, EP91, Fay02, Jor94, Lee93, Mat99, PPS96, Stu90, Gei84, HTM89, Zel85]. Binomial [Afs08, AC98a, BRRB01, BCJ97, CLPC08, CL02, DH05, DN94b, FM08, HM01, Hwa09, KMS10, LJ09, LC99, LTN10, Lin92, Lon10, Pro08, Pul90, Ros07, SL92, Tie94, VBB94, AS84, Bly80, DN94a, Kah87, Knü94, She84]. Bioassay [DO04]. biographical [Bol84].

Bioinformatics [Fos08]. Biological [GPA97]. Biology [ACA03, Rob07b]. Biomedical [LD98, Rue04]. Biosciences [Cle01]. Biostatistical [AW91b, NOMLT95]. Biostatisticians [DAF^{+}94, Sch10a]. Biostatistics [BV11, Der93, Fer04, HFDJ08, KYA^{+}07, Kle95, Mil01, Zha07b]. Biotechnology [Kha04]. Birch [Cox84]. Birth [Nun92, SD04]. Birthday [Nun92, Sch10b]. Bit [Tur07]. Bivariate [Beh90, Cai94, DH05, GM91, Gre96, JH96, MS02, RRT99, GM86]. Black [Bro07, Tal07, WH07]. Blind [Sto98]. BLiP [LT97]. Block [FM92, GL00, Gat95, Pad08, Mur86]. Blocking [Faw90, Old95]. Blocks [MG97]. BLUE [BF91]. BLUEs [VAD90]. BLUP [Sea97]. BMDP [KDV94]. BMDP-5V [KDV94]. Board [Coc05a]. Boards [Gib90]. Bonaparte [Gun98]. Bonferroni [KG90, Sch84]. Bonus [Kes00]. Book [Abb02, AC02, ACA03, AC03, Ank98, Bar98, Bar99, Bat04, Bon98, Bon02, Can98, Cle99a, Cle03d, Cle03c, Cle03b, Cle03a, Coh02, Cot98, Cow03, Cow04b, Cow04a, Dev02, Dix07, Dob02, Dob03b, Dob03a, Fot02, Gen03, Gol98, Gol99, Gon03, Gup02, Hai03, Hay99, Hol99, Hu03, Jam99, Jam02, Jam03, Jer03, KK99, Kaf99, Led03, Len04, Low98, Lun02a, Lun02b, Maa99, Mac99, Man03, Mil03, Mil04a, Mil04b, Min07, Nag98, NO98, Obr98, Par98, Pil98, RH02, Rob03, Rob08, Row02, Rus98, SP98, Sch99, Sho98, Smi98, Tan02, Tan99, Ter03, Utt02, Wal02, Wil02, Won98, Wul03]. Books [Ano97d, Ano97e, Ano97f, Ano97g, Ano98f, Ano02e, Ano09g, Ano09h, Ano09i, Ano09j, Ano10f, Ano10g, Ano10h, Ano11i, Ano11f, Ano11g, Ano11h, Cle01, Col01, Dev01, Hai01, Hil01, Hyd01, Hyu01, Jam01b, Jam01a, Kre01, Len01b, Len01c, Len02, Mac01, Mil01, Pou02, Ree02, Rum01, Sch93c, Sha01, Str02, War02]. Boolean [Hel88]. Boosters [MZ92]. Bootstrap [All97, AT04, BS98, DK07, FF02, JS03, Pol99, RZR05, TKM04, EG83, Yan88]. Borel [SS01, Tom84]. Both [Atk94, Spi86, Spi87, Tuk80]. Bottles [TH00]. Bound [Gre96, Lev97, PN95]. Bounded [JH96]. Bounds [Ass96, CT91, DH11, FP81, Nav97a, Pol99, TM95, WB91, Wri90]. Bowditch [RA04]. Bowler [DPS04]. Bowling [CS94]. Box [DR07, HN98, FLP07, MTL78, RC88]. Box-and-Whiskers [DR07]. Boxer [Gel06]. Boxplot [RRT99]. Bragged [Ber07]. Breakdown [Bas91]. Breaking [CG07]. Breviary [SW05]. Bribery [CG95a]. Bridge [Dun94, Shu01, Per98]. Bridges [FHP^{+}11]. Brief [Ano03a, Ano04b, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, DL04, DL05, DAC05, DAC06, RACK07, RD05, BR78, MMR90]. Bringing [WC97]. Bronchial [HHB91]. BS [RSH01]. Buffett [FS06]. Buffon [Rie90]. Build [Bro08]. Building [FHP^{+}11, MG97]. Built [Sea97]. Built-In [Sea97]. Bulbs [MD96]. Business [AR08, Bon98, BO92, Chi08, Cle08b, CW10, Dev06c, Dev07a, Dev07b, Edw07, Hil96b, LH02, Mil01, Owe01, PS98a, Rum98, Sim07, Skl07, Smi98, SB09, Wal02].

C [Mon06]. Cables [Per98]. Calculated [Alb02]. Calculating [Bro01, Edw96, FT10, LDM10, Men04, Tie94, Tie98, TM99]. Calculation [AC00b, HL96, Lor90, Gue79]. Calculations [Bro01, HH01, PB03, SRC07]. Calculator [Mon06, Bar98, NO98]. Calculus [Hag09, Hwa09, SB02, Wan02]. Calibrated [Lit06]. Calibration [Fin06, FS93, SBB01, YHZ08, Zha00]. California [Tob04]. Called [Fre06]. Can [AP10, Ann06a, Coc05a, Fre07a, GN02, HR10, Olk92, Sch03c, SS99, Tra07, Utt03a, VM07, Wat91]. Candid [WSF02]. Canonical [Ren92]. Can’t [GN02, Ros00b]. Capital [Rum07]. Capstone [LRF11]. Carcinogenesis [HHB91]. Care [AW91b, JS11]. Career [BK93]. Carl [Bec82]. Carlo [BS98, FF02, KCGN98, KBH09, LH72, Rie90]. Carolina [Mon04, Ste08]. Case [BA02, Cal03, FT93, HEJ11, Kos05, PS98a, PHSN08, Shu01, SB09, Tob04, Wie99, WASB10, ZRK^{+}11, DK95]. Case-Based [PS98a]. Case-Control [PHSN08, ZRK^{+}11]. Casebook} [Smi98]. Casino [CW93]. CAST [Mil04a]. Catastrophic [MZ92]. Categorical [Fri95, Gan95, God91, LBTM03, MGLK99, Ost02, Ost03, Sim98, Was06, vS96]. Categories [MFS^{+}03]. Cauchy [AB92]. Causal [Old95]. Causality [Ros05]. Causation [Smi11]. Cautionary [AH02, HL98, HS92, Hsi95, Kah87]. Caveats [Gam09]. CD [Mil03, Bon02, Sch04f, Sne98]. CD-ROM [Mil03, Sch04f, Sne98]. Ceiling [AB03]. Celebration [Ano90a, Ano90b]. cells [Mur86]. Cement [PR98]. Censored [BS95, BH08, FM08, WT92]. Censoring [Man07, SD01]. Census [Cra08, Oko99, Tor07, CHS04]. Censusing [RS93]. Center [AW91b, BCJ97, Ede90a, WKB^{+}95, KM94]. Centering [Bel84]. Central [Bel01, DR99, Inl10, Jol95, Kon11, Sko08, PG82, Yan88]. Century [DL00, LR06, McD99, Cle03c]. certain [Kap79, KHK89, MS87]. Certification [Imr94b]. Chain [KCGN98]. Chains [AKLS10]. Challenge [AC06a, Bra07, Cle06a, Fin06, Hol07, Nor07, Rus98]. Challenges [Cha99, CRV95, Kir91]. Championship [Ano04h, Bil04, Col04, Har04, Mas04, Mea04, Ste04]. Chance [AC05b, Cle06b, Gue06, Hai01, Mac01, Sch06b]. Chances [Cle03d, Der93]. Change [BA06, Zho01, vS96, Har76]. changed [Boa94]. Changing [Sti10, Tor07]. Characteristic [BK07, DG07, Epp93, Hal08, Luc97, Wal95a, WTH95, Ger79]. Characteristics [MT11]. Characterization [BF91, BK99, Hu08, Nai87, GY89, GY90, Kel94]. Characterize [McC94]. Characterizing [EKG02, FRD07, SBL97]. Charles [MS01]. Chart [RHCR06, Wil06a]. Charts [Bor02, CG95a, LHD00, SF00]. Cheating [Kva96]. Chebyshev [SYC88, AC00b, SYM84, SS97]. Chemical [KM92, RKW98]. chemistry [RC88]. Chernoff [Nav97a, PN95]. Chi [Bai92, Che11, Dri99, Lev97, Mir01, WN07]. Chi-Square [Che11, Dri99]. Chi-Squared [Bai92, Lev97, Mir01]. Chicago [AC05a, AC06b]. Chip [Lee07]. Chisini [GV09]. Chocolate [Lee07]. Choice [BHA02, Wal00a, Wan97]. Choice-Based [Wal00a]. Cholesky [EW02]. Choosing [FT10, LS08, NP97]. Churchill [CR95]. Citation [Sch10a]. Citizenry [Sch90]. Citizens [Utt03b]. Claims [Utt02]. Class [Ahm96, DiP04, FK97, Gel98, Gro98, HLW09, NVB09, Mee86]. Class-Participation [Gel98]. Classes [AL07, Mag96, SB09]. Classical [HB03, WMAY09]. Classification [CW11, DS98, HV03, SS98, Han79]. Classification-Based [CW11]. Classroom [BG08, Kva96, Shu06, Ter05, SI82]. Clear [Ros00b]. Client [Sow03]. Clients [JW04]. Clinical [AW91b, CRM01, San06, WKB^{+}95, Ber87, RC88]. Clinicians [Lis08, Bar99]. Closed [Wan10]. Closeness [FK97]. closer [Has87]. CLT [AP10]. Cluster [HLW09, WF09, SR83]. Clustered [Pan02]. Clustering [Tar07]. Clusters [Moe05a, Moe05b]. Cocaine [Shu91]. Cochran [DM81]. Coefficient [AH95, CW92, DR01, Mir01, RvE97, Shi01, Tju09, Tur97, VM90, Van96a, Wri92, ZM94, vM93, GP89]. Coefficients [Bri94, KG90, LM03, Men00, Men04, Sha91b, SR97, Shi01, Tju09, Wri90]. Cognitive [LG00]. Coherent [SRMP08]. Cohort [WASB10]. Coin [CS05, GN02, Gel06, Kon95, KB07, Shu06].

Coincidence [Ste08]. Coincidences [Han92b, Sch10b]. Collaborators [Ano95d, Ano96c, Ano97a, Ano98a, Ano99a, Ano00a, Ano01a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a]. Collapse [MFS^{+}03]. Collapsing [Sha82]. Collection [FT93, War02]. Collections [BM03, JR92, SGK00]. Collector [Daw91, Kuo00, Rea98]. College [Alb03a, Ano04h, Bil04, Col04, Dun04, FHP^{+}11, Har04, Mas04, Mea03, Mea04, Par04, Ste04]. Colleges [LW98]. Collide [Hal10]. Collinearity [FK09]. color [Bea83]. Column [Alb03c]. combination [CCW^{+}86]. combinations [Ham84]. Combinatorial [Gla04]. Combining [CRM01, Peñ97, SC94, SMC^{+}07]. Come [SATV94]. Comment [CW92, GR09, HS09, Joh09, MG09, MD97, NL09, PG97, Sey91]. Comments [HSF^{+}94, PR94, Tuk85a, Gei84, PS89, Zel88]. Commission [RWM^{+}72]. Committees [Eld90]. Common [AC04a, Dav95, Dav98a, DHW11, MW92, OCE93, Tie98, Lee86]. Communication [Rie06]. Community [AS99, RT01]. Community-Based [RT01]. Compact [MY05]. Companion [Cow03, Sch04f, Sne98]. Companions [Mil01]. Comparative [Nor95, OCE93]. Compare [Bri96, LC99, AP10]. Comparing [ALTY00, Cal03, GCGG95, Han92a, Hos92, Kha92, ME11, Sch10b, She97, WS93, WS94]. Comparison [AW06, Bar02, Git94, HT02, HL10, HL01, KB07, LT96, Loc93, LDM10, Mat99, MV94, MW92, Mon97, Nav97a, Nay94, Sav90, TGM08, tWGB01, Aso80, BvE88, Pav88, SI82]. Comparisons [And94, BA06, Day98, PF95, ZPH99, DK87, HT83]. Competitive [ST95, SB98]. Complementary [Pie92]. complete [SR83]. completely [HB96]. Complex [Coh97, KG90, Rid09]. Complex-Step [Rid09]. Complications [DR99]. Component [CM92, LDM10, NK96, RS91a, AB92]. Components [BK99, GR04, HL98, JK99, Ren92]. Computation [Nie05]. Computational [AG04, Car00, Gen04, Har08, Lan04, SSH81]. Computations [Bar09, Kha91, LH72]. Compute [GV09]. Computer [Bag95, CF01, Car08, Cow06, GPA97, GC04, LDL^{+}94, Mil04a, Nas95, NL95, Owe01, Rob07b, Sch93c, Sul96, AAF72, MH76, PK82]. Computer-Aided [GPA97]. Computer-Assisted [Mil04a, MH76]. Computer-Based [CF01]. Computer-Intensive [Rob07b]. Computers [Alb02]. Computing [Alt02, Ano90c, Ano90d, Ano91c, Ano91d, Ano91e, Ano91f, Ano93b, Ano95h, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano98h, Ano98i, Bel01, CC01, Cha99, Cow04b, Dal90, DH11, FF02, FS01, Gen04, GEL01, GKT98, HM01, Hil02c, Hil02d, Hil04f, Hil07g, HL01, Hun05, Hyn96, LR01, MHT90, MS02, Mon04, Nag98, NL10, SA00, Sea94, Sko08, TM95, Upt01, Wan02, Zim00, CGL83]. Concentration [GH91]. Concept [Sul96]. Conceptions [Alb03a]. Concepts [Bar99, BM02b, BO92, Cha99, DS98, Dun99, Fra92, IM01, Kaf99, Kha91, Kre01, Len02, Lun02b, MMCsS96, Mor07, Pau92, SS98, dML09]. Conceptual [Bro08, Fri95, Kah06]. Concerning [Sen08, SS09]. Concerns [HEJ11]. Concise [Cra05]. Concordance [RS91b]. Concrete [Loo97]. Concurrency [HJS06]. Condition [Hu08, McL99]. Conditional [BL94, BW08, Gre91, GAR92, HHvdWZ06, MH93, OI07, Pan02, Pau08, PP98, Upt00, VSL01, Kap79]. Conditionality [Hel95, Pri11]. Conditionally [GM91]. Conditionals [MS94]. Conditioning [Bel84]. Conditions [CBH09, MM90, MA84]. Conferences [LH02]. Confessions [Kon95]. Confidence [Afs08, AC00a, And09, AS98, Bar02, Bar09, Ber96a, BR03, BS03, CC01, CRM01, CL07, Cle03b, Dou93, Ede90a, Fer90, FS93, Fre10, GE11, HM01, HL10, Juo93, LD94, LR08, LTN10, LK99, LDM10, ME95, MW92, MG04, New01, Nic94, Pes90, PE05, Pol99, Ros03, Ros00a, Rya09, SL92, SG01, Sch96b, TM95, TL93, Van96a, Ver99, VH05, WBT01, WS01, WL99, Wri90, ZG00, Zuo10, Ang89, Ang90, AS84, Ber81, Buo87, Car86, CCW^{+}86, Khu93]. Confidentiality [KKO^{+}06]. Confirmatory [Woo08, Tuk80]. Conflict [MR09, Spi07]. Confounder [JHFK04]. Confounding [Lyn03, Old95]. Confuse [AC05b]. Confusion [HB03]. Connection [Sap91, Sch96b, Wri92, VM89]. Connections [LL97]. Consequences [BHA02]. Conservative [Nic94, WSF02]. Considerations [NK96]. Consistency [HS91, YB10]. Consistent [LE00]. Consonance [NP97]. Constants [Gre10]. Constrained [MV07]. Constructed [LaM94]. Constructing [BK96, God91]. Construction [And09, JT96, LTN10, Nic94, Hid86]. Consultant [ES93, WKB^{+}95]. Consultants [GB06]. Consulting [Eng93, JW04, KM94, Kir91, Loh95, NOMLT95, Wil92, AC02]. contaminated [Riv81]. Contemporary [Wal02]. Context [Bol03, Gor11, MCM^{+}11]. Contingency [GK99, Gre91, Gre10, MW92, MH93, Mir01, Sha82]. Continuity [LC99]. Continuous [LBTM03, Rak06, Sch10b, Mul88]. Contour [DG07]. Contract [Shu01]. Contrast [AP10]. Contrasts [Lan06b, Sch85]. Contribution [Rie90]. Contributions [Dav11, Kha04]. Control [AC08b, BO92, CG95a, Git94, JHFK04, Kas98, LC09, PHSN08, PF95, SF00, Sto98, Str90, ZRK^{+}11]. Controversies [Kaf99, Utt02]. Controversy [Vos99]. Convergence [DHW11, Mil96, TKM04, dML09]. Conversation [Ano00g]. Conversion [Hil96a, Hil07a]. Cook [OH95]. Cookies [Lee07]. Cooperate [Hal10]. Cooperative [Die93]. Coordination [WKB^{+}95]. copulas [GM86]. COPY [Hal92]. Core [Hog94]. Cores [Der93]. Corner [Ano98j, Bis98, BHM01, CF01, CR01, Czi98, HJP^{+}01, HM01, HGA01, IM01, RT01, Spu98, Var98, Zho01]. Corrected [ML92]. Correction [Ang90, Ano98c, Ano03b, Ano03c, Ano04c, Ano05b, Ano05c, Ano06b, Ano10b, CH86, GY90, JS03, Kar86, Kar87, LC99, SYC88, Spi87, Upt00, WS94, Zel85]. Corrections [Ano02a, Ano03d, Ano04d, Ano08d, Dav98a]. Correlated [Ham87, HR10, Kel04, LSO01, PLC00, PPS96, SC09, VBB94]. Correlation [And09, BS79, Dav09, DGC97, DH11, DR01, Fri02, GK99, Gre96, HCW07, KAB^{+}97, Muk10, MC96, Nel98b, RM94, RvE97, SR97, Shi01, TO93, Tur97, VM90, VAD90, Wri92, ZM94, Zim00, FK86, Rav85, SP86, Sti86]. Correlations [Lan06b, MS02]. Correspondence [Tho95]. Corrgrams [Fri02]. Corroborate [CG95a]. Could [SJD10]. Count [Gla04]. Counterexamples [Hel95, JT96, Muk10]. Counterfactual [SGM07]. Counterparts [Epp93]. Counting [Edw96]. Counts [Wan10]. Coupon [Daw91, MS01]. Course [AC08b, BHM01, BO92, Cha04, Cra05, DNG98, Dav07, Dev02, Dev07a, Gel05, HM01, LRF11, Len01b, LP00, Mac01, Mac99, MGCR10, Mar07, Mru07, Pil98, Pou02, Pri05b, Rob03, SO94, Ste90, SV03, Wal00b, Wul03, Row02]. Courses [AC98b, Bir95, Bra96a, CF01, CRV95, Eng93, FR04, GZK^{+}11, Gen04, Hun05, Kle95, LJLC05, Led95, LW98, Par04, PS98a, SM97, SR96, WC97]. Court [GKR94, HS94, PG97]. Cousins [BSTM09]. Covariance [FM92, Hay11, MS98, Muk10, Van97]. covariances [Tam84]. Covariate [AH02, AFV96, HRS10, SJI05]. Covariates [Zho01]. Coverage [LS08, Pol99]. Cox [FLP07, Hal10, Hsi95, MM80, RC88, Zho01]. CQI [Hog99a]. Craig [DK95]. Cramer [Kag01, Nay02]. Cramer-Rao [Kag01]. Crash [Dav07, SR04]. Crisis [Dun04]. Criteria [Day98, GR04, HCW07, VH05, BvE88]. Criterion [NP97, Eat78]. Critical [HV95, Har80]. Cross [HRS10, MCM^{+}11, EG83]. Cross-Match [HRS10]. Cross-Validation [MCM^{+}11, EG83]. Crossing [FRD07, He97]. Crossover [PS94]. Crude [LGH09]. Cryptology [Bro11]. Cumulants [BL94, Smi95]. Cumulative [Tie94, Tie98]. Curious [Per93, Sil97]. Current [Hil03c, WG11, WY04, Gol86]. Curricula [NL10]. Curriculum [Ano98j, BN98, Bis98, BGNP01, Czi98, MHT90, SO94, Spu98, Var98]. Curves [BK07, FT93, GKT91, Hal08, He97, JR92, Mon95]. Cuthbert [Hun98]. Cutter [Sto98].

Daily [BG08]. Dallal [LDL^{+}94]. damage [MS87]. damage-assessment [MS87]. Damned [AC05b]. Daniel [Hun98]. Dare [Dal90]. Data [Alb97b, AW01, AC07c, Ann07, Arn07, Bar98, Bec96, BS10b, Bon02, BH08, BGF^{+}08, Can98, Car00, Cav07b, Cha99, CF07, Cle99a, Cle08a, Cle08b, Coh97, Cow04b, Dav08b, DGC97, Dev07c, Dla05, DT00, DuM99, Dun94, FM08, Fri95, GL00, Gen03, GR04, Gol92, Gon94, GR96, GC04, Han98, Han05, HBC^{+}91, HDE^{+}03, HQ08, HH01, Hsi95, Kah06, KKO^{+}06, Kel04, KK91, Kha08, KG90, KG98, Kot91, LW90, Len01c, MD96, Mac01, McN92, Mil03, Mil04c, MGLK99, Nie11, Ost02, Ost03, OH08a, OH08b, OSC00, Owe01, PLC00, Pan02, PP10, PB90, PR98, Pri05b, Rak06, Ree02, RHCR06, Ric08, Rob07b, Rue04, Rum98, Rum99, Rum01, Sch10a, Sch99, Sch93c, SS08, Sea94, SMVD09, SC99, SW90, Spa08, SJD10, Str90, Stu90, Su08, WWD03, WT92, Wal98, Wan97, War02, Was06]. Data [Wie06, Wie08, Wil02, tWGB01, YB10, Zim00, Zim07, Cas85, HTM89, KD87, Pav88, Rav85, SSH81, Cle99a, Mon97]. Data-Based [Wan97]. Data-Oriented [Alb97b]. Databases [OI07]. Datasets [Bra08a]. Dawid [Sen08]. Day [Cha04, Sch10b]. DBMS [Hal92]. DBMS/COPY [Hal92]. Deal [MCDD91a]. Dealing [Kir91]. Death [Gue06, Lud05, SD04]. Deaths [Cle03d]. Debate [CR01]. Debating [BB04]. Deception [Gun98]. Decision [Gel98, HS91, Lon10, Ric10, Rum98, Sch99, SRW95, Vic08, WY04]. Decision-Theoretic [Ric10]. Decisions [CLS07, Cle08b, SRW95]. Decline [VD05]. Decomposition [EW02, Grö07, EW85, Man82]. Deem [JW04]. Deeper [Men10, Mil04b]. Defectives [Wri90]. Deficiencies [GM96]. defined [BG89]. Defying [Han92b]. Degree [JW04, Lev97]. Degrees [AW91a, BGNP01, FMA^{+}09, Lee93, LW98, Sul96]. Degrees-of-Freedom [FMA^{+}09]. Delivery [Pra08b]. Delta [HL10, Oeh92, SRMP08, BG89, Cox84]. Demeaning [Bel84]. Deming [Boa94, Git94, Hah95, Hah02, Sne08, SF00]. Demonstrate [SF00]. Demonstrating [Old95]. Demonstration [Lee95, Lev93]. Demonstrations [Gel98, Mil04b]. denominator [FK86]. Density [BH93, CC06, FRD07, HN98, Hyn96, SA03, TK76]. Department [CRV95, Sti08]. Dependence [DHW11, FS01, GB03, MG04, Shu01, WSS93, FK86]. Dependency [Sab02]. Dependent [Beh90, Dav93, SK93, WASB10, Duf86, Pav88]. depending [GD89]. Depends [Sim98]. Depicting [Wai96]. Derivation [BL94, GS02, Sol00, VM90, WS87]. Derivations [Bet00, Sam92]. Derive [Kuk09, LC99, VE06, Wal95b, Nas95]. Derived [BSY92, MS95]. Deriving [BC92, dCS86]. Describing [Gro91]. Description [BCB00, Sch93a, Sch85]. Descriptive [Dav08d]. Design [AC98b, AC07a, Ank98, Arn06, BM02a, Ber06, BGF^{+}08, DLM01, DO04, Eas04, Gam09, GM96, Gun93, HS03, IM01, Law96, Len01a, Len01b, LE93, Lis08, Mat99, Mru07, MKV92, OF03, Pie01, PS94, Pra08a, RKW98, Rob07a, Rub04a, Tan02, Wul03, Box80, Hid86, Par98]. Designed [Czi99]. Designing [AC98b, FF02]. Designs [CH96, Faw90, FM92, Gat95, HMR^{+}01, LS97, Pad08, GS86, Hei81, Her86, Mur86, WC87]. Desire [CW10]. Desired [Fox10, Men09, Stu09, vC10]. Detect [AW01]. Detecting [Bar07, Bor97, DGC97, GB03]. detection [Kra88]. Determination [Afs08, Bro93, CW92, IBP05, Len01d, LM03, Men00, Sha91b, Shi01, Tju09]. Determine [LB00]. Determining [Che11]. Developed [Coh97]. Developing [AC06a, AT04, PS98a]. Development [BGF^{+}08, Zam08, RC88]. Developments [Ano90c, Ano91c, Ano91d, Pie80]. deviance [Has87]. Deviant [Olk92]. Deviates [Lee93]. Deviation [HJP^{+}01, MN97, O’C90]. Deviations [BS03, SGC90]. Device [Old95]. DHEA [EKG02]. Diagnostic [MC87, Vic08, Sti86]. Diagnostics [Bel84, FK09, Spi86, Spi87]. Diagram [Upt01]. Diagrams [LL97, Loh95]. Dialogue [BM10]. Diamond [LBTM03]. Diamond-Shaped [LBTM03]. Dice [KK91, PP10]. Dicing [Gue06]. Did [Ben93]. Difference [GS06, KMS10, SF93]. Differences [AC00a, FP81, Man07, SG01, Sho95, WB04]. Different [BLS00, Wie99]. Differential [BSY92, Hwa09, Wan02]. Differentiation [Rid09]. Differently [PF95]. Differing [KG95]. Difficult [Wat91]. Difficulties [CLS04]. Digesting [Bra96b]. Digit [Rod04]. Digital [BGF^{+}08]. Digits [Jad00, Ley96]. Dilemma [MCDD91a]. Dimension [GR04]. Dimension-Reducing [GR04]. Dimensional [CW04, EP91, LT97, ZCM95, Mon79]. Dimensions [TKM04]. Dirac [AT99]. Direct [HT02]. Directed [HKS06]. Direction [Hig99b]. directions [Tuk85a]. Discard [SJD10]. Discipline [BM02b, Hig99b, Kot10]. Disciplines [KC09]. Discovery [Bar98, Ber06, Cle99a, Kha04, Pic06, Ree02, Ric08]. Discrete [BSY92, FM84, JSM95, LR08, Lin92, MF07, Nav97a, Pes90, Sch10b, Ter05, TL93, VH05, Wal00a, WW97, BL87, EG87, GD89, KP79, Mee86]. Discriminant [LGH09, OW91, Ren92, Sap91]. Discriminate [OW91]. Discrimination [BB04, Tju09]. Discussion [Bel05, Ber03, Bil04, Bis98, Bra07, Bry99, Car03, Cob99, Col04, Czi98, Fre97, Gel04, Gro99, Har04, KCGN98, Lin97, LS99b, Mad99, Mas04, Mea04, New99, Nor07, OS99, Sch97, Sho97, Spu98, Tuk85b, Var98, Wit97, Efr86]. Discussions [SB09]. Disease [MB06, Wan10]. Dispersion [MY05, Sho95]. Display [BM03, LBTM03, MC96, Pri05b, SC99]. Displaying [Ber96b, Faw90, Jac08, JR92, KAB^{+}97]. Displays [ALTY00, DT00, Fri02, HT02]. Dissimilar [CF07]. Distance [All97, HFDJ08]. distinct [Aso80]. Distinguish [Wal95a]. Distinguishing [HB96]. Distributed [JB04, tWGB01]. Distribution [ALL10, AS98, BTV90, CLPC08, Cha95, CA97, DH05, Dav05, Dun99, Ede90a, FM84, GKT91, Gho96, GS02, Gro91, GS86, Hag09, HEJ11, HR06, Jol95, Kon87, LT97, LM08, LJ09, Ley96, LR01, LB00, Mad98, McC94, MF07, Mon95, OSC00, Pos93, RMY96, SA00, SS91, Sko08, SRMP08, Ste92, Wal95a, WTH95, Wil08, YB10, ZPM^{+}11, vDK02, Box81, Bur86, CH85, CH86, Dav80, Ger79, Gle89, Kap79, Kel94, Knü94, KP79, MN88, She84, WS87, dCS86]. Distribution-Fitting [Mad98]. Distribution-free [GS86, Kon87]. Distributional [Bai92, Hos92, Jon99, MT11, Kap79]. Distributions [BS99, BM03, BM94, BSY92, Dan92, Dri99, Fer90, GM91, Jam00, JH96, LS01, Lee93, LSE00, LC99, Lin92, MS02, OI07, PA91, Pes90, Rod04, SS91, SBL97, SS97, SC97, SS02, Tie98, VE06, Wel98, WW97, Wie07, XN03, ZG00, AB92, BL87, Gei84, GM86, GD89, Gle89, Lee86, LS88, Mee86, Mul88, Zel85]. Diurnal [EKG02]. Divisions [Gro90]. Do [Che11, GB06, Men09, Mon06, PF95, SG00, Stu09, TS03, Wat91, vC10, Los04a]. DOE [BA02]. Does [BHO00, FDC09, HS94, MD96, McC94, MFS^{+}03, Wal95a, Bar90]. Doing [Sta04]. Domains [Tal07]. Dose [Stu90, MS87]. Dose-Response [Stu90, MS87]. Dot [Wil99]. Double [Edw96, Rya09, AJ88]. Double-Elimination [Edw96]. Doughnut [Bor02]. Dr [Git94]. Draft [DM01]. Drawing [CLS04]. Driven [Eas10, Zim07]. Drug [Bor97, Kha04]. Drug-Discovery [Kha04]. Drugs [BHM01]. Dual [Nad09, Knü94]. Duelling [Cle03d]. Dummy [Coh91]. Dyadic [Ann07]. Dynamic [AC07c, Atk95, Col01, HBC^{+}91, IM01, Bol86]. Dynamics [GM96, VY85].

Early [APT07, Bet00, Bro11, Rie90, Sch81, Noe84]. Easier [Dav98b, Wat91]. Easy [LC09, Len01c, SR96]. Eckler [Wan91]. Ecological [For05]. Econometric [KT94]. Economic [Dev07a]. Economics [Bon98, Chu99, Dev06c, Edw07, Gel08, RH02, Sha01]. Ed [Utt02, ACA03, Can01, Cle03b, Coh02, Cow04a, Dob02, Fra00, Gen03, Gon03, Hai03, Jam03, Jer03, Kir00, Kre01, Led03, Man03, Mil03, Pie01, Pil98, Ste01, Obr98, Gup02]. Edition [Hyu01, Jam01a, Sha01, Wil02, Cle00b, Joh99, Wal00b, Gol98, Hol99, Kaf99, Bon02, Jam02, KK99, Mac99]. Editor [Ano97c, Ano98e, Ano99d, Ano00d, Ano00e, Ano00f, Ano01b, Goo65, Hig99a, Ost11d, Ost11a, Ost11b, Ost11c, Wes07, Wes08, Alb03b, Alb04, Alb05a, Alb05b, Ano95e, Ano95f, Ano95g, Ano96d, Ano96e, Ano96f, Ano96g, Ano98b, Ano98g, Ano99b, Ano02b, Ano02c, Ano02d, Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano04g, Ano05d, Ano05e, Ano06c, Ano06d, Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano09d, Ano09f, Ano09e, Ano10d, Ano10e, Coh10, Gol97a, Gol97b, HV09, Hil97a, Hil97b, Hil98, Hil99, Hil00, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil08b, Hil09, HN10, HB05b]. Editor [Lan97, Lan99, Lar94, Lar96, Ost09a, Ost09b, Sch93b, Sto01, Sto02, SA02, Stu09, Tuy09, Var72]. Editorial [Ano95d, Ano96c, Ano97a, Ano98a, Ano99a, Ano00a, Ano01a, Ano04a, Ano05a, Ano06a, Ano07a, Ano08a, Ano09a, Ano10a, Ano11a, Gib90, Wes06]. Editors [Ano11c, Ano11d, Ano11b, Ano11e]. EDsub100p [Stu90]. Educated [Utt03b]. Educating [Sow03]. Education [Cle06a, Dun04, FHP^{+}11, Gel08, Hog91, Kre01, LH02, MCGM95, RS91a, SS04, Sne93, Wil06b, Rus98]. Edwards [Boa94, Sne08]. Effect [AB03, Bro10, Dor96, Gam09, GR96, HR10, KP07, Kva00, Mat99, MG04, SRW95, Tim04, YT01]. Effective [AC00a, FMA^{+}09, Len01d, LH02, Pos93, WBT01, SI82]. Effectively [Bir95]. Effectiveness [MM90]. Effects [ALL10, BH98, BK96, CK01, CJ92, GL00, LBTM03, LDM10, MPB97, MB00, Ros01, Ros03, SG91, SGK00, SP86, Sea94, Sea97, SVR11, SF00, Alb92, GS86, SD87, VY85].

Efficacy [NS97]. Efficiency [GT95, YT01, Tuk49a]. Efficient [CW11, DO04, Leh80, NVB09, JH83]. Efroymson [Mil96]. Eggs [DH05]. Egon [Dav81]. Eight [Mor98]. Einstein [Igl07]. Eisenhart [CR95]. Elasticity [Ass96]. Election [CF08, WJM^{+}03]. Electric [Pal99]. Electronic [BM94, HS03, Mil01, Sne98, SATV94]. Elemental [MG97]. Elementary [Ber97b, Bro08, Cha08, Hu92, Lun04b, Mil03, Pil98, Rum01, Sam92, Sho98, VWB^{+}10, YA09, Cox84, GY89, GY90]. Elements [Hyu01, RH02, SP98, Sch04e, BH86]. Eleven [Mir01]. Elicitation [BI09]. Eliciting [SGM07]. Elimination [Edw96]. Else [AC05b, LB00]. Elteren [MLL10]. Elusive [Hel99]. Embracing [Wil94]. Emerging [HHJL93]. Empirical [Epp93, MCM^{+}11, Mon95, Cas85]. Employers [RSH01]. Encounters [Mac01]. Engineering [AC08a, Ano98j, BN98, Bis98, Czi98, Hog94, Hu03, Jam00, Kva00, SP98, Spu98, Tan02, Var98, War02, ZH90, vDK02, Kir00]. Engineers [Bis91, Car00, Cle99b, Cle08a, Hyd01, Kie06, Ste01, Kir00]. Enhance [RT01, Sow03]. Enough [BH08]. Ensembles [Zhu08]. Enterprise [Dix07, Spe08]. entropy [Shi88a]. Entry [DM01]. Enumeration [Hog92]. Envelope [SL92]. Environment [HHJL93, RKW98, SM09]. Environmental [SG91]. Environments [Nag98]. Epidemiology [Gil08, Kle95, Ost98, SGCS04]. ePrint [Hil07a]. Equal [Sho03]. Equality [RZR05, PS89]. Equating [Wal95b]. Equation [HL99]. Equations [Bar99, tWGB01, Fre83, Wom85]. Equiponderant [LBTM03]. Equitable [BS00a]. Equivalence [BRRB01, EW02, LW90, LC09, ML92, St.90]. equivalences [SSH81]. Equivariant [Bas91]. Era [BV11]. Errata [Ano09b]. Error [Ass96, AB03, Far90, Gat95, GAR92, Kaz95, KBH09, LK99, Nay02, VWB^{+}10, VD05, Wai96, Bar76, Bly80, Pav88, Spi86, Spi87]. Errors [AC04a, Bas97, CR96, Dem06, Fre06, Fre05, Gad01, Gor11, HR10, HB03, Koe09, LE00, SC09, Str90, Duf86]. Errors-in-Variables [CR96]. Essence [Dob06a]. Essential [BUK^{+}08, Dev06a, Can01]. Essentials [Edw07, Mil03]. Establishing [AW91b, SP85]. estimability [EW85]. Estimable [EGCD91, PE05]. Estimate [Bar09, Kva96, WE08]. estimated [SYM84, SYC88]. Estimates [BC92, Bly93, EM93, JP01, KG95, SK93, SA03, YB10, BM88a, Tuk49a]. Estimating [AW91a, AH95, Bor97, HL99, Hou97, KU96, LGH09, PP10, Per98, Rus91, Sch91, SD04, Sol93, Sti86, Str90, Stu90, Tim04, tWGB01, Dah81, Han79, Tho87]. Estimation [AC98a, AKLS10, BR97, BGG98, BS00b, Boy08, BH08, Bro93, CL07, CK01, DC90, Far90, HEJ11, HS03, Hen99, Kor90, LK99, Mee87, ME95, MS94, Mur86, OI07, PB92, PT91, PR94, Pul90, SB02, Wal00a, War97, AJ88, Bel87, CJT87, CH85, CH86, Cox84, GD89, Gol86, HTM89, JH83, Kah87, SR83, TK76, Tor78, VM89]. Estimator [Ang92, Fre06, Oak93, SD01, SI90, SJI05, Bly80, Kot86, PS89]. Estimators [Bas91, Buz97, CHS99, DN94b, FK97, Grö07, Kha92, Lee09, LT96, Lin93, Nay94, YT01, BvE88, DN94a, Kar85, Kar86, Kar87, Leh80, MA84, SC94, SI82]. Ethics [VM03]. Ethnicity [Saw98, Wal98]. Euler [Dav11]. Evaluating [GR04, KKO^{+}06, LC96, KHK89]. Evaluation [Coh97, Lac90, MJ04, Vic08, VH05, WY04]. Even [Tra07, Vid94]. Event [GC04, LHD00]. EVENTCHARTS [Gol92]. Events [Bar02, Gol92, WSS93, EG87]. Ever [Tra07].

Everyday [Ros00b, Sch06b]. Everyone [Tra07, Efr86]. Everything [AC05b]. Evidence [APT07, BB04, HB03, Luc97, She98]. Evidence-Based [APT07]. Evidential [SRC07]. Evolution [UJ08]. Evolutionary [Nie05]. Evolving [CY99]. Exact [AC98a, BK96, Che11, CL02, Hil02e, KR02, MW92, MH93, Nic94, Ost02, Ost03, OH08a, OH08b, Ros03, SM79, ZCM95]. Exam [BM10, Kva96]. Examination [LJLC05, Ost98, Ost02, Ost03, OH08a, OH08b]. Examining [SG01, WGHE09]. Example [AFV96, BS10b, BH08, ESL09, Gre10, Gur08, Hos92, JS03, LF00, LDM10, Sil97, TKM04, Tur97]. Examples [Beh90, Dob06b, KG95, LCD10, OS95, Pra08b, Pri05b, Ter05]. Excel [Cle99a, Sch99, LF00, Bon98, Wal02]. Excerpt [Dem06, SD04, SW05]. Excerpts [Pet99]. Exchange [CU92, Inm94]. Exchangeable [LS01]. Exclusion [Ham97, HR98]. Exemplar [Bul05]. Exercise [DLM01]. Exercises [Bel07, Cle99a, FZ00, KM92]. Existence [KC02, PC85]. Expect [PP98]. Expectation [Kon11, PP98]. Expected [Bar02, Bly80, CRM01, Hom97, SSC99]. Expenditure [CJ92]. Experience [Bra96a, CRV95, Kot10]. Experiences [BG08, JW04, SV03]. Experiment [AC98b, GL00, Pad08, PR98, PF95, Rie90, Sal99, SF00, SM79]. Experimental [ACA03, AC07a, DO04, Eas04, Gat95, GM96, Gun93, HMR^{+}01, HHB91, IM01, Mru07, Par98, Pie01, Tan02, Tim04, Hei81]. Experimenters [Ber06]. Experiments [Arn06, Ber96b, Czi99, Dan94, Koe09, Law96, Law03, LDL^{+}94, Len01a, Len01b, MLL10, NL95, Pra08a, PF95, Rob07a, Sul96, Ver99, Box80, Ank98, Cot98, Wul03]. expert [Tuk85a]. Explained [KS91, Moo07]. Explaining [CG92, CLR01, GC99, Lan06b, OW10]. Explanation [AP10, Few09, GKR94, Paw00, PG97]. Exploration [Cav07b, HBQ04, She97, YCF06]. Exploratory [Atk95, Dev07c, Fri02, Tim04, Tuk80]. Explore [DT00]. Exploring [HBC^{+}91, HM01, OH95, Rum99]. Explosion [GT95]. Exponential [BL94, CM92, ML09, MS94, She97, BS87, Gle89, McC88]. Exponentiated [Gre96]. Exposing [GM96]. Exposure [Dan92, WASB10]. Extend [Ano00g]. Extended [SJI05]. Extensions [JK90, LHD00, WS79, Far89]. Extent [SD04]. Extrapolation [Van96b]. extrapolations [MS87]. Extreme [Bur86, Che11, CT91, Hsi95, SB98]. Exuberance [SR05].

F [PS98b]. Face [PP10]. Faces [Str92]. Factor [Ber96b, Czi99, Dan94, GL00, Woo08, Mon79]. Factorial [Bec96, Dan94, LDL^{+}94, Nav97a, NL95, SVR11, Her86]. Factorials [LL97, SGK00]. Factors [JW04, LS99a, LDL^{+}94, Sea94, SS02, Sti95, WGHE09]. Faculty [BM10]. Fail [AS94]. Failure [MZ92]. Failures [AC00a]. Fall [LW98, VD05]. Fallacy [HH01]. Fallibility [FBG96, Paw00]. Families [CM92, Fer90, ML09, BS87, Gue87, McC88]. Family [MS94, She97, WWD03]. Fascination [Coc05b]. Fast [Buz97, LR01]. Fastest [Tib97]. Fate [Gun98]. Fathom [Col01]. Fathoms [Mil04b]. Faulty [Ste92]. Faulty-Inspection [Ste92]. Favor [Bet00]. Favorites [FDC09]. FDA [DuM99]. Fear [CW10]. Feared [Fox10, Men09, Stu09, vC10]. Features [JR92, SMC^{+}07]. Federal [Gon94, RWM^{+}72, Eld90, Tuk49b, Tuk49c, Wal98]. Feel [PF95]. Fences [DR07]. Few [MT11]. Fewer [SVR11]. Fiducial [Sch96b, Wan00]. Field [Ann06a]. Fields [Utt02]. Fieller [HL10, Zer78]. Fifth [Sha01]. Fifty [Mil04b].

Figure [EA11, WY04]. Figures [HQ08]. File [Hil96a]. Filter [CT91, EW02, MS87]. Fin [MCGM95]. Finance [Lun04b]. Financial [Dla05, KR99, vDK02]. Finding [EGCD91, HMR^{+}01, Hel99]. Finite [CLS04, DC90, Hou97, KT02, LaM08, Ter05, WS01, Bel87, Buo87, PG82]. Finite-Mixture [CLS04]. Finiteness [Muk10]. First [AC98b, BHM01, Cle03c, Dav95, Dav98a, Dev02, FVMU10, HM01, Igl07, KC02, Len01b, Mac01, Mar07, Mru07, Pil98, Rod04, Wul03, Her86, ME84, PC85, Rog84]. first-order [ME84]. First-Year [BHM01]. Fisher [Box80, Box81, Box05, Che11, Chr05, DO04, Hal10, Inm94, Lud05, McL99, Pri11, Ric10]. Fit [GKT91, HL99, KS91, PC02, Ros00a, Sim98, TC88]. Fitting [Gan95, HV03, HD94, Mad98, WG11]. Five [CS94, KDV94, Loc93, Lur95, Ost98]. Five-Pin [CS94]. Fixed [Fre10, HR10, Lee93, Sea94]. Fixed-Effects [Sea94]. Fixed-Width [Fre10]. Fixing [Sho03]. Flexible [PHSN08, Rub04a]. Flip [CS05, Gel06]. Flipper [Kon95]. Florent [FVMU10]. Florida [CRV95, WJM^{+}03]. Folded [Mon95]. Follow [LKNP95]. Follow-Up [LKNP95]. Following [ZRK^{+}11]. Football [AW06, Ano04h, Bil04, Col04, Gil00, Har04, Mas04, Mea03, Mea04, Ste91, Ste04]. Forecasting [Hyu01, IJW06, LS01, OL96, Bol86]. Foreign [Eng93]. Forensic [Wei07]. Forest [Grö09]. Forgotten [MF07]. Form [Wan10, Knü94]. Forms [DCWP93, Dri99, Eli00, Saw92]. Formula [Hay11, Hut93, LC99, Sko08, Wal95b]. Formulas [DCWP93, Jon99, Pei93, Har80]. formulated [Pei90]. Formulation [Olk92, Ric10, Smi95]. Formulations [LW90]. Forward [HS10, Sch03c]. Foundations [McC01, Sti99]. Four [BM94].

Fourth [DR99]. Fraction [Fre07b].

Fractional [Bec96, McL99]. Frailty [DJ05]. Frames [PHSN08]. Framework [HG97, KKO^{+}06, PO07]. Frank [NL95]. free [GS86, Kon87]. Freedom [AW91a, FMA^{+}09, Lev97, Sul96]. Frequencies [Nun92]. Frequency [DuM99, SMC^{+}07, SS08]. Frequentist [Ber96c, IBP05, MR09]. Fresh [LGH09]. Friendly [CG06, Ros00b]. Frontiers [Loh04]. FU [LKNP95]. FU-PLOT [LKNP95]. Fulcrum [DT00]. Function [AT99, Cai94, Cha95, Gro91, HT06, Inl10, Koo93, LGH09, McC94, Mon95, Pal08, Pos93, Sch91, Wal95a, vM93, CDFP81, KE83, McC88, Rup87]. Functions [DG07, DCWP93, EGCD91, Epp93, Jor94, LJ09, Luc97, PE05, Ren92, Ros00a, VE06, WTH95, Wal95b, BL87, Hel88]. Fundamental [Har08]. Fundamentals [Cle08a, Zha07b]. Funding [Der93]. Funnel [Gun93, SF00]. Further [Luc97]. Future [Ehr90, HHHZ99, Kot10, Moo01, RSW99, Tuk85a, ZH90, Bar90]. Fuzzy [AP10].

G [LDL^{+}94]. Galton [WWD03]. Gambling [Alb02, AC05b, CW93]. Game [BH08, Chu99, Coc05a, Coc05b, Gil00, GDL08, Shu01, Shu06, Ste91, Wes90, Kel94, Rum07]. Games [Pau92, ESL09]. Gamma [Hag09, Wie99, Gle89]. Gastwirth [MD97, PG97]. GAUSS [Yal07, Cle05a, Cle05b, Jes94, Sha91a]. Gaussian [Dav80, SS91, WS87]. GEE [HCW07, tWGB01]. Geisser [BSTM09, Zel85]. Gemmell [DM81]. Gendered [Ham97]. General [BM01, BM02b, BS10b, BHO98, CLR01, DS08, DM01, Eli00, EW02, FG00, Gad01, HH99, JP01, Kag01, KC02, LS01, LSO01, McC01, Mir01, Mor98, New01, Pes90, Ros01, Rub04a, Sal90, SD01, SC09, Shi01, Shu01, SS01, SB02, Stu90, TM95, Ver01, WBT01, WB02, Wil08, tWGB01, YT01, ZPM^{+}11, CJT87, DK95, Her80, Spi85, Zer78]. generalization [Har76, Kra88]. Generalized [AT99, BC92, Bir90, Boo92, Chu99, DD06, DJ05, Hil93, Hil94, HD94, HL99, Hu03, Hu08, Oks91, Pie92, Tju98, ZPH99, Som86]. Generate [MD96]. Generating [BS95, Cai94, CF07, EP91, Gan95, Inl10, Koo93, Lee93, McC94, PPS96, VE06, CDFP81, KE83, KP79]. Generation [DAF^{+}94, DL00, Jon02, LH72, PK82]. Generators [Tur07]. Genetic [HMR^{+}01]. Genetics [Sis07]. GENMOD [KDV94]. Genomics [MCM^{+}11]. GenStat [Hil07b]. Gentle [OF03]. Geometric [Bri96, HL10, Sam92, Sch93a, Ste96, TO93, Sch85, Sha82]. Geometrical [Bai92, Epp93, Hay11]. Geometrically [JB04]. Geometry [GM96, SK95, Her80]. George [AS86, Ord84]. Gertrude [Hal10, MM80]. Get [Sta01]. Getting [Sow03]. Gibbs [CG92, MB94]. GIS [Cra08]. Give [LF06, Wie99]. Given [ALTY00, MS94]. Glimpse [HHHZ99, Bol84]. Global [HV03, HBC^{+}91]. Go [VM07, Ver99]. Goalie [BS10a]. Goals [Saw98]. Gods [Jam99]. Goel [JP01]. Goel-Okumoto [JP01]. Going [MH93]. Gold [HLW09]. Golf [BS00a]. Good [Ver99]. Goodness [KS91, Ros00a, Sim98]. Goodness-of-Fit [Sim98]. Gosset [Bol84, Box81]. Gossip [Cle03d]. Got [VAD90]. Government [Ros95, Tuk49b, Tuk49c]. Grade [Dun04]. Graduate [BK93, GS08, Gel08, HR05, Moo05, Ter03]. Graph [FVMU10]. Graphic [Pic06, Ros00a]. Graphical [DGC97, FS01, Fra92, FW05, Gun98, Haz03, IM01, KAB^{+}97, Lee95, LKNP95, LBTM03, MC96, Nav97b, PC02, PS94, SW05, Sti95, Wes90, dML09, KD87]. Graphically [OH95]. Graphics [Alt02, AC07c, Atk95, Bel01, Bra08a, CC01, CF07, FF02, FS01, GEL01, HBC^{+}91, Hil02c, HL01, LR01, MMCsS96, MS02, Pra08b, SA00, Upt01, Wan02, WC07, Bea83, BR78]. Graphing [Bar98, Hyn96, McN92, NO98]. Graphs [CW99, GPD02, HKS06]. Graybill [St.90]. Greenberg [AS86]. Greenhouse [BSTM09]. Group [DS98, Eme96, LDM10, SS98, WB04, ZNY11]. Groups [Bir95, OW91, Par04]. Growth [Bel05, Lor90, Rum07, DBS^{+}85]. Guide [AC05a, AC05b, AC06b, Bra08b, Cle99a, Cle06b, Dev07c, Dix07, Dob06b, Gen03, Gil08, Gur08, Hil01, Hoe98, Lis08, Moo07, Mur97, Nel01, Par98, Pet99, Ros00b, Spe08, Tra07, Utt02, AC03]. Guided [KC09]. Guidelines [BGNP01, Fin06, Len01d]. GUM [Los04b].

H [TW91]. Haenszel

[Lev97, MW92, Som86, Upt00]. Hald

[PR98]. Half [LC99]. Hand [BD11, War95]. Handbook [Cle06c, Dav08b, Gup02, Lac00, Lun07b, Man03, Wie07]. Handicapping [BS00a]. Handling [Rub04a]. Hands [DPS04]. Hansen [TW91]. Happens [Dun03]. Harmonic [Ste96]. Harrison [Bol86]. Harrison-Stevens [Bol86]. Hartley [Edw86]. Harvard [Sti08]. Hasse [Loh95]. Hastings [CG95b, Hit03]. Having [Lee93]. Hawaiian [Sch81]. Hawkes [Hil08a]. Health [Dav08c, Fer04, Gel08, Gue06, JS11, SGCS04, Utt02]. Health-Care [JS11]. Heat [WF09]. Height [SWW02]. Helicopter [Ann05b]. Help [MFS^{+}03, Utt03a]. Helping [Bro08]. Here [SATV94]. Heredity [Nel98a]. Hesitate [Alb97a, Ber97a, Fre97, Lin97, Moo97a, Moo97b, Sch97, Sho97, Wit97]. Heterogeneity [DJ05, Ros05, VY85]. Heterogeneous [Bir95, JB04]. Heteroscedasticity [LE00]. heuristic [WS87]. Hierarchical [ALL10, CS91, JS11, SC99]. Higgins [Min07]. High [EP91, FHP^{+}11, HK92, SR04, TKM04].

High-Dimensional [EP91]. High-Tech [SR04]. Higher [Cle06a, ZCM95]. Highest [Hyn96]. Highlighting [Man07]. Hilfertys [Koe09]. Histogram [BH93, Wan97]. Histograms [SS08]. Historical [Dav09, Gor11, Jam01a]. History [HBF90, Hit03, Jam01b, Len02, Sti10, WF09, BR78, Her80, Her86, MMR90, WS80]. Hitter [Fre07c]. Hitters [Ann06a, Alb92]. HIV [BW08]. HLM [KDV94]. Hockey [BS10a, DM01, Gil00]. hold [Bar90]. Home [Ann06a, HS94, LP00, Alb92, BH98]. HOME-AWAY [BH98]. Home-Court [HS94]. Homogeneity [GGR05, MS92]. Honeycomb [Vid94]. Hope [Ehr90]. Horvitz [OS95]. Hot [DPS04, War95]. Hour [GS08]. Hours [Sch10b]. Huber [Fre06]. Human [EA11, SWW02]. Hurricanes [IJW06]. Hyper [NVB09]. Hyper-Efficient [NVB09]. hyperbolic [MN88]. hyperbolic-secant [MN88]. Hypergeometric [AC00b, LR01, MF07]. Hyperstat [Jer03]. Hypothese [Sal90]. Hypotheses [GKR97a, SBB01, Sil97, Hel88]. Hypothesis [All97, BH02, BH08, Cor00, FF02, GR96, KT02, Loo97, SK93, Nai87]. Hypothesis-Dependent [SK93]. Hypotheticals [GKR97a].

I-A [AW06]. Ideas [BUK^{+}08, Dob02, Pag00]. Identical [CF07]. Identification [JS11, Wom85]. Identifying [Sch10a]. Identities [Bai92, Roc88]. Identity [MS95, Wri92]. Idiots [Cle03d]. Ignoring [AFV96]. II [BS95, McC99, Ost03, Tuk49c, VD05]. iid [Jad00]. III [CH96]. I’ll [LF06]. Illustrate [Pau92]. Illustrates [JS03]. Illustrating [BR03, BP02, Jes94, SJI05, Tur97]. Illustration [Jol95, Nav97b]. Illustrations [Nas92]. Illustrative [Ter05]. Image [Yu94]. Imbalance [Kaz95]. Impact [Bor99, Hah95, HHHZ99, Ham97, HT90, Kaz95, LC96, SG91, SJI05].

Implementation [SR96, TK76]. Implementations [FF02]. Implementing [RKW98]. implicitly [BG89]. Importance [Grö07, Grö09, GM96, HB05a, Kot10, Tho04, Kru87]. Important [JR92]. Imposing [AR95]. Impossible [GKR94, GKR97b]. impressions [Noe84]. Improper [TL10]. Improve [YHZ08]. Improved [AS84, Car96, CLPC08, Dri99, Lev93, WSF10, Dah81]. Improvement [AC07b, HHHZ99, Hoe98, KM92]. Improvements [BS00b, Hog91, WY04]. Improving [Der93, Far90, Gon94, HQ08, SW90, YA09]. Imputation [HR06, HL01, HLP03, Sch05e, YHZ08, ZR10]. In-Sample [Fre05]. Inclusion [Sol93]. Inconsistent [Fre07a]. Incorporate [SA03]. Incorrect [Pei93]. Increased [WSF10]. Increasing [CF91, Mag96, Nie11]. Independence [CA97, DS08, Dav09, JT96, LaM94, Muk10, Pau08, SK95, Upt00, WSS93, Flu86, SD87]. Independent [EG87, LC99, Mea03, Zeh91, Bar76]. Indeterminacy [ALL10]. Index [Ano97b, Ano98d, Ano99c, Ano00c, Ano08e, Ano09c, Ano10c]. Indexe [Ley96]. Indices [dM07]. Induction [Gho96, SP85]. Inductive [SS91, FF80]. Industrial [Hah95, HHJL93, Hun98, MHT90, RKW98]. Industrial/Chemical [RKW98]. Industry [Cle08b, Dan92, Ket95, KM92, McD99, ST95]. Inequalities [Gho02, Nay02, SS97, Ste96]. Inequality [Cua93, Kag01, MN97, Shi01, Far89, Hil84, NW86, SYM84, SYC88, Spi85, WS79, WS80]. Inertia [Nel98b]. Infer.NET [WW11]. Inference [Bro11, Cra05, Epp93, Fre07b, Gad01, HM93, Hil02e, HHvdWZ06, HT90, LS97, LP76, LCD10, Mac01, NS97, NVB09, RS93, Rum99, ZR10, Edw86, GP89]. Inferences [CLS04, Gue87]. Infinite [Tuk47, AT86]. Inflation [AB03, Dun04, Sti95]. Influence [Fay02, Gro91, SC09, Rup87]. Information [BK93, Car96, Chu99, Day98, Don99, GT95, Igl98, McC88, McL99, Peñ97, PW91, SA03, TC88, TM99, AO89, Zel88, dCS86]. Information-theoretic [TC88]. Informative [DB90, WSF02]. Informed [Cle08b]. InfoTrac [Mil03]. Inning [RMY96]. Innovation [Ber06]. Innovations [Cle06a]. Input [Lar92]. Insight [Fra92, Gun93]. Insignificant [Chr03]. Inspection [Ste92, Sti86]. Institutional [Gib90]. Instruction [BA02, LG00]. Instructor [Kon95]. Instructors [LW98, SB09]. Instruments [NO98]. Insurance [Ver01]. integer [CDFP81, Dah81]. integer-parameter [Dah81]. Integral [Roc88]. Integrated [HG97, Owe01]. Integrating [BO92, SO94]. Integration [DG07]. intelligent [Tuk85a]. Intensified [BM10]. Intensive [Rob07b]. Interaction [SG91]. Interactions [SGK00, Mon79]. Interactive [AC07c, Atk95, Gol98]. Interactively [Dun99]. Intercept [GAR92]. Interdisciplinary [BV11, LRZG^{+}10]. Interesting [JS03, ME84]. Interim [Ber87]. Intermediate [Pri05b, Row02, Rum01, Fra00]. Internally [GW94]. Interpolation [Lau65]. Interpretation [Bai92, Ber96c, Dav08b, Hay11, HQ08, PS98b, PS98c, Ren92, Sti95, Str02, TO93, WB04]. Interpretations [Epp93]. Interpreting [Cua93, Pet99, Mon79]. Interrater [Mac93]. Interstates [LC96]. Interval [AC98a, And09, Ber96a, BS00b, Boy08, Bro93, CRM01, CL07, Ede90a, FM08, HM01, LT96, LK99, SL92, VW93, WBT01, War97, Zuo10]. Interval-Censored [FM08]. Intervals [Afs08, AC00a, Bar02, Bar09, BK96, BR03, BS03, BHO00, Fer90, Fre10, HL10, Juo93, LS08, LD94, LR08, LTN10, MW92, MG04, New01, Pes90, PE05, Pol99, Ros03, Rya09, SG01, SS08, TL93, Van96a, Var92, Ver99, VH05, Wan00, Wan10, WS01, WL99, YA09, ZG00, Ang89, Ang90, Ber81, Buo87, Khu93, Kon87]. intraclass [SP86]. Intro [McD04]. Introduce [KS08, KC09]. Introduction [Alt92, Ank98, CG06, Chi08, Cle08a, Cle08b, Coh02, Cow04b, Dav06, Dev07b, Fos08, Har06, Hay99, Jam03, II07, Law96, Len01a, Lun04b, Mäc07, McG05, Mil03, Min07, Nel07, OF03, Owe01, Pra08a, Rob07b, Rum99, Rum06, Sch03a, Sch03b, Sch04a, Sch04b, Sch04c, Sch04d, Sch05a, Sch05b, Sch05c, Sch05d, Sch06a, Sin08, SW05, Str02, SRC07, Stu09, Ver01, Wei07, Wue98, Cas85, Cox84, TK76, Cle00b, Ste01, Lun02a]. Introductory [AL07, BO92, Bra96a, CF01, CR01, DNG98, Fer04, FR04, HR00, Led95, Mag96, MGCR10, Moo07, SB09, Sol10, Ste90, SR96, Wat91, WC97, Bat04]. Intuitive [Rus08]. Invariance [PB92, Neu92]. Invariant [JRP11]. Inventions [SW05]. Inverse [Hu08, New01, SD01, SS91, Dav80, Rav85, WS87]. Inverse-Probability-of-Censoring [SD01]. Inversion [WTH95]. Inverted [LS88]. Inverting [Ros03]. Investigating [HJS06, Mor07]. Investigation [KR02]. Investigations [Cle01, GE11]. Invitation [Stu09]. Invited [HSF^{+}94, LDL^{+}94]. Involving [Fer90, HS03, Sil97]. Iowa [FDS05, LC96]. iPod [FDC09]. Iraq [Spi07]. Irrational [SR05]. Irregular [SS08]. Island [Meh09]. isn’t [Efr86]. Issue [Gra08]. Issues [AC05b]. Itching [MB06]. Iterative [Gan95, HD94, Jor94]. Itself [GS06].

J [Min07]. Jack [Bec82]. Jackknife [Buz97, EG83]. Jackson [Ben93]. Jacob [Wei82]. Jacobian [DD06]. Jai [CG95a]. James [Min07]. Japanese [Git94]. Java [Pra08b]. Jeffreys [LS97, TGM08, Zel82]. Jensen [Hil84, NW86, Spi85]. Jerzy [LR82]. Jim [Lea04]. JMP [Alt02]. Job [DWBQ98, Sta01]. Jobs [DWBQ98]. Joe [Ben93]. Joint [AS98, GC04, Mag90]. Jonckheere [PS81]. Jose [Gei84]. Journal [Pet99]. Journalists [Cle06c]. Journals [Gib90, TS03]. joy [GM86]. Judging [ESL09, SG01]. Jumping [Han92b]. Jury [Bar07]. Just [AC05b, Lar92, Lin92, MT11, Mil03]. Justification [Gre91, HD94].

Kalman [CT91, EW02, MS87]. Kalman-filter [MS87]. Kaplan [Bar09, Oak93, SD01, SJI05]. Kappa [dM07]. Kappa-Type [dM07]. Karl [Cle05a, Cle05b, Inm94]. Kempthorne [PS89]. Kendall [DF07, GB03, Ord84, VM90]. Kernel [Alt92, PW06]. Kernels [Zhu08]. Key [BHA02, GT95, Sch10a]. Key-Balanced [BHA02]. Khintchine [Jon02]. Kiefer [Bec82]. Kinetic [NQ96]. Know [BV11, Ter03, Utt03b]. Known [AH95, DS08, FVMU10, LC99, Tur97]. Kolmogorov [Har80]. Krieger [MD97, PG97]. Kronecker [MS98, Rog84]. Kullback [AO89]. Kurtosis [Gro98, JRP11, KS08, SI90, SB11, Rup87].

Lab [CF01]. Labor [SG98]. Laboratory [Ano98j, BN98, Bis98, Cle99a, Cry91, Czi98, Spu98, Var98]. Laboratory-Based [Ano98j, BN98, Bis98, Czi98, Spu98, Var98]. Labs [Dev01, KC09]. Lack [HS91]. Lagrange [Wri92]. Langren [FVMU10]. Language [Bro07, GEL01, Ham97]. Laplace [TGM08]. Laptops [BG08]. Large [BR03, BHO00, Bra08a, DuM99, FT93, HS94, JR92, Koo93, Mag96, MC96, OI07, Hid86, Tom84]. Larger [Bai95]. Last [SS06]. Late [Gil00]. Late-Game [Gil00]. Latent [HLW09, HLW09]. later [Hoe85]. Latin [Mur86]. Law [BR03, Koe09, LSE00, Tom84, AP10, CG07, Few09, Var72]. Laws [Gra08, Koo93]. Lead [Vid94]. Leaf [SMC^{+}07]. League [DM01, Gro90, HV95, RMY96, WC97]. Learn [Coc05a, Ros96, Tra07, Wat91]. Learning [Arn07, BGF^{+}08, Die93, GZK^{+}11, Gur08, Hil01, Hil08a, Kva00, MJ04, Moo93, RT01]. Least [Bas97, BK92, BC92, Dun94, HH99, HS92, Lee09, Lic90, ML92, Oks91, Rie90, Tar00, Van96b, deL06, BvE88, BM88b, MA84, PS89, SSM80, Wom85]. Least-Squares [Lic90]. left [Mul88]. left-truncated [Mul88]. Legal [Gas92]. Lego [HHvdWZ06]. Lehmann [Suk92]. Leibler [AO89]. leisurely [EG83]. Lemma [BP02]. Length [BCB00]. Lesson [BK92]. Lessons [Mos97]. Let [GPD02, Hog99a, MCDD91a]. Letter [Ano09d, Ano09e, Ano11b, Coh10, Goo65, HV09, HN10, SS09, Tuy09, Var72]. Letters [Ano97c, Ano98e, Ano99d, Ano00d, Ano00e, Ano00f, Ano01b, Ano02b, Ano02c, Ano02d, Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano04g, Ano05d, Ano05e, Ano06c, Ano06d, Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08f, Ano08g, Ano08h, Ano08i, Ano09f, Ano10d, Ano10e, Ano11c, Ano11d, Ano11e, Hig99a, HB05b]. Level [Ann06a, CF08, Gel05, HS10, LL97]. Levels [GL00, ME11]. Leverage [Sal90]. Leveraging [EA11]. Lévy [Inl10]. Li [vC10]. Licks [SS06]. Lies [AC05b]. Life [Cle05a, Cle05b, Cle06b, DB97, Kha08, Lud05, Moo07, Ros00b, Sch06b, SC99, Gol86, MS87]. Life-Testing [DB97]. Life-Time [SC99]. Lifelong [Coc05b]. Lifetime [OSC00]. Light [MD96]. lightly [Wio90]. Likelihood [Ass96, BC92, Bir90, Boy08, DN94b, HEJ11, Hen99, JP01, KR02, Kor90, LG05, LCD10, Mag90, Mea03, ME95, MV07, PB92, Pan02, Paw00, PT91, Sil97, Ste92, Tju98, WS01, YB10, Ber87, Cox84, Dah81, Dav80, DN94a, Han79, Hil87, Leh80, McC88, Tuk49a, VM89]. Likelihoods [BM03]. Likely [PP09]. LIMDEP [Hil06b]. Limit [Bel01, Inl10, Jol95, LC96, PG82, Yan88]. Limitations [MGLK99]. Limited [Rak06, TM99]. Limiting [DCWP93]. Limits [GE11, LDM10, Sch96b]. Lindeberg [Inl10]. Line [Wri92]. Linear [BR97, CR96, CBH09, Cow04a, DD06, DB97, Eli00, EGCD91, FW05, GAR92, Grö07, Grö09, Gur06, Hil94, HD94, Hu03, HT90, Ibr97, Jam03, Kot91, LaM94, LW90, Len01b, LE00, MN97, MS95, MV94, MSS91, MPB97, MB00, Nic94, PLC00, Pie92, PE05, Pri05a, Sal90, Sea97, SC09, SM91, Tar07, TM95, Tim04, Tju98, Van97, VSL01, WB91, WG11, XN03, ZPM^{+}11, ZPH99, vS96, BvE88, Bol86, CW88, CJT87, Duf86, Ham84, Her80, Kot86, MA84, PS89, SSM80, TC88, Thi87, Zer78]. Lines [MB00, Nel90, Nel98b]. Link [CG06]. Link-Probit-Normal [CG06]. Links [DHW11]. List [Dav98a]. Literacy [AC06a, Cle06a, Gal03, GNM^{+}98, Utt03a]. Literary [GGR05, PH02]. Literate [Sch90]. Literature [Ham97, Mur97]. Little [Dix07, Len01c, She07]. Liver [MB06]. Lives [She07]. Load [GN02]. Local [HV03, HBC^{+}91, Iri01]. Locally [Kuk09]. Locate [Hel99]. Locating [HBC^{+}91]. Location [HJP^{+}01, SB11, JH83, Spi85]. Log [Pie92, Wie99, vS96]. Log-Linear [vS96]. Log-Normal [Wie99]. Logical [Gre91]. Logistic [Bra96b, Fay02, HHB91, KR02, LM03, Men00, Men04, MFS^{+}03, OW91, PC02, Sim98, Str92, Tju09, WL99, CCW^{+}86, dCS86]. Logit [CK01, New01, Sap91]. Lognormal [Rea98, SA00, Sko08]. Logrank [Cal03]. LogXact [Har00]. Long [BH08, HR00]. Longitude [FVMU10]. Longitudinal [Dun94, GC04, Hor06, LKNP95, SMVD09, Zim00]. Look [Dan92, HV95, JL97, Kag01, Kot91, LGH09, Mir01, Nas06, RvE97, Sch03c, CJT87, EG83, Has87]. Looking [Sch03c]. Lord [Hil84]. loses [Sch84]. Lottery [Pau92, Ste08]. Love [AC05b, HR10]. Low [ME11]. Lower [DH11, GP09, Gre96, Lev97]. Lucullus [Mis06]. Lyell [MS01].

M [Gei84, SB02]. M-Estimation [SB02]. Machine [LH72]. Machine-generation [LH72]. Macintosh [Han05, SMP97]. MAD [Hil84]. Madness [FT10]. Magnitudes [Sch93a]. Main [SGK00, GS86]. Major [HV95, RMY96, WC97]. Majors [GH91, LRF11]. Make [Dav98b, MCDD91a, Wat91]. Making [CS05, Dev07c, Lon10, LH02, Mon97, Rum98, Sne08]. Man [Tib97]. Management [Git94, Obr98, Sha01, Tan02]. Managers [Bon98]. Managing [SB09]. Manipulating [Hel88]. Mann [BLS00, Ran01, Ahm96]. Mantel [Lev97, MW92, Som86, Upt00]. Manual [Kha91]. Manufacturing [SM09]. Many [ALTY00, MD96, Str92]. Map [SM97, WF09]. Maple [Tan99, Bag95]. Mar [SS09]. March [FT10]. Margin [SS99]. Marginal [JHFK04, Lee93, PLC00, Pan02, SA03, SSM80]. Marginals [JH96, GM86]. Margins [Mea03]. Markers [Vic08]. Market [AC05b, Ann06b, SG98]. Marketing [AR08]. Markov [AKLS10, Gho02, Jes94, KCGN98, Sha91a]. Mary [MM80]. Mass [DCWP93]. Match [HRS10]. Matched [HRS10, Lev97]. Matches [RvE97]. Matching [LGXB11, PMO08, RS91b, ZRK^{+}11]. Material [BGF^{+}08, FR04]. Materials [Ano91a, Ano91b, Ano92, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano97d, Ano97e, Ano97f, Ano97g, Ano98f, Ano02e, Ano03a, Ano04b, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, Ano10f, Ano10g, Ano10h, Ano11i, Ano11f, Ano11g, Ano11h, Cle01, Col01, Dev01, DL04, DL05, DAC05, DAC06, Hai01, Hil01, Hyd01, Hyu01, Jam01b, Jam01a, Kre01, Len01b, Len01c, Len02, Mac01, Mil01, Pou02, Ree02, RACK07, RD05, Rum01, Sha01, Str02, War02]. Math [Lan06a, Len01c]. Mathematica [Bag95, CB95b, Hil06a, Sti00]. Mathematical [Alb02, Coh02, Dav95, Dev01, Dob02, HBQ04, Lun04b, Mac99, Sta08]. Mathematics [AC03, Dav11, Dev06a, GH91, Hai01, HK92, Ter03]. MATLAB [Rus08]. Matrices [Fri02, KAB^{+}97, MS98, MC96, RM94]. Matrix [Dav06, McL99, Olk92, Rav85, dCS86]. Matters [HV03, Kah06, Muk10]. Maurice [DF07, Ord84]. Maximal [GK99, Pal08]. Maximization [LG05]. Maximum [BC92, DN94a, DN94b, GW94, HEJ11, Hen99, JP01, JB04, KR02, Kor90, LC96, Mea03, ME95, PB92, Shi88b, Tuk49a, YB10, Cox84, Dah81, VM89]. May [CW93, Dav98b]. MCLUST [HLW09]. Me [Bon02, LF06]. Mean [AH95, AS98, Bar09, BS99, Ber07, BS03, DK07, Gra08, GV09, GAR92, JB04, KP07, LF06, Lev93, MC91, Muk06, Nay02, O’C90, Per93, Sam91, SG00, She97, SP94, Ste96, WBT01, Zha07a, Aso80, Kot86, ME84, SYM84, SYC88, Shu82, Yan88]. mean-of-ratios [Kot86]. Meaning [Sen90]. Meaningful [Iri01]. Means [BC92, GR96, MS92, SK93, SF93, Tar07, ZG00, vM93, SSM80, Shi88a, SM79]. Measure [MN97, Ped85]. Measured [Cle03c, Bar76]. Measurement [LK99, VWB^{+}10, Spi86, Spi87]. Measures [DGC97, GKT98, Gro98, Hos92, HB03, JRP11, Mag90, SC09, Rav86, TC88]. Measuring [AG95, Fay02, HB05a, Sab02, Saw98, SGCS04, vS96]. Media [Dan92, GNM^{+}98]. Median [Bar09, Ber07, BH02, CLS07, FLP07, FKLL03, FG00, HJP^{+}01, HS92, HTM89, Koe09, LR08, Lee95, O’C90, SGC90, She97, SP94]. Medical [CJ92, Dav08c, SZM^{+}07, Wie06]. Medicare [GE11]. Medicine [APT07, Zam08, Abb02]. Medium [CW99]. Meet [MHT90]. Meeting [Saw98]. Meier [Bar09, Oak93, SD01, SJI05]. Melvin [Gei84]. Memorandum [Tuk49b, Tuk49c]. memoriam [AS86, DM81, FK87, Ord84]. Memorian [NL95]. Men [Han04]. Meng [vC10]. Mess [HR10]. Message [CF91, CW99, Koe09]. Meta [Nor95, Van96b]. Meta-Analysis [Nor95, Van96b]. Method [BK96, EP91, FMA^{+}09, Faw90, Fre05, GCGG95, HS92, LF06, LKNP95, McC01, McI05, Oeh92, PC02, PPS96, Pes90, Rid09, SD04, BG89, Cox84, Dah81, KP79, SM79]. methodology [Kra88, MH76]. Methods [Alb93, ACA03, And99, AT04, Bel07, Bon98, CHS04, Car00, CL02, Cle99b, Die93, Dob06a, DS02, Eme96, Gen03, Har06, Hyd01, Jam02, Joh99, Kva00, Lee08, Len02, Mar07, MCM^{+}11, MV94, Mor07, Mur97, Ost02, Ost03, OH08a, OH08b, PW06, Rie06, Rob07b, Rue04, Su08, VWB^{+}10, Wüt08, YA09, ZNY11, Han79, PK82, Len01c]. Metropolis [CG95b, Hit03]. Michael [FVMU10]. Microcomputer [HG97]. Microsoft [Cle99a, Wal02, Bon98, Bor02, LF00]. Mid [MW92]. Mid- [MW92]. Mighty [Kes00]. Milk [Pic06]. Milliken [St.90]. Milliken-Graybill [St.90]. Million [Bra08a]. Mind [Cav07a]. Minded [dML09]. Minimizes [SGC90]. Minimizing [Fre05]. Minimum [Ang92, BCB00, Cai94, HJP^{+}01, Eat78, GD89, Kar85, Kar86, Kar87, VM89]. Minimum-Deviation [HJP^{+}01]. Mining [DHN^{+}05, Dev07c, DuM99, Han98, HDE^{+}03, PW91, Su08]. Minitab [Cle99a, Alb93, AC07b, Arn06, Ste90]. MINITAB^{T M} [Cle01]. MINQUE [tWGB01]. Misleading [VW93]. Missed [Ter03]. Missing [HL01, PT91, Sne93, Spa08, YB10, vH04, HB96, Mur86]. Missingness [MGLK99]. Mississippi [BES98]. Missteps [Smi11]. Misuse [CR96].

Mixed [DJ05, Gur06, MSS91, MPB97, MB00, Nel07, PO07, Sch93c, Sea97, VSL01, Vos99, WG11, Wul07, HT83]. Mixed-Effects [MPB97, MB00]. Mixed-Model [Sch93c]. Mixture [Beh90, CLS04, CW11, FN92, LR06, MS02, PR98, VE06, Gle89, PK82]. Mixtures [Hou97, Rod04, Sap91]. ML [tWGB01]. ML3 [KDV94]. MLE [FM08]. MLE’s [SS91]. MM [HL04]. Mo [SYC88]. Mode [AG95, Sen90]. Model [AW01, AH02, AR95, And94, Bel87, Bri96, BCB00, CK01, Cor00, Eli00, Gur06, HQ08, HT90, JP01, JHFK04, JS11, Kot91, KB07, LRZG^{+}10, LR06, LE00, LDM10, MF07, MB00, PC02, PE05, PO07, SA00, Sap91, Sch93c, Stu90, TM95, Tim04, Van97, ZPM^{+}11, Alb92, Bol86, CJT87, Her80, HT83, Kot86, Mon79, ME84, Pau81, SSM80, SSH81, Zer78]. Model-Based [HQ08, Kot91, Bel87]. Model/Data [AW01]. Modeling [Alb02, Car96, Dan92, GPA97, GC04, Hai07, Hor06, KP07, NQ96, Peñ97, Rak06, RKW98, RMY96, Sch04e, Sim07, SB98, DBS^{+}85]. Modelling [Nel07]. Models [ALL10, AT04, BL94, BF91, CG06, CR96, CBH09, CLS04, CM92, CW04, Cow04a, CW11, DD06, DJ05, EGCD91, Fri95, GKT91, HV03, HT02, Hil93, Hil94, HD94, HL99, HL01, Hu03, Ibr97, JHFK04, LaM94, Len01b, Mam06, MN97, MS95, MV94, MSS91, McL99, MPB97, NS97, Nel98a, OF03, Pie92, PO07, Pri05a, Sch99, SMH91, SSH96, Sea94, Sea97, SMVD09, Sha91b, SC09, Su08, Tju98, Tju09, VSL01, Vic08, Vos99, Wal00a, WG11, Wie99, WGHE09, Wul07, YW09, ZPH99, Zho01, vS96, Cox84, GP89, Pei90, Fot02]. Modern [Abb02, AC08a, Fra00, KK99, Min07, Pil98, Rum99, SP98, Sta08]. Modernizing [Leh95]. Modes [DHW11]. Modification [Lev93]. Modified [Har80]. Molecular [Vic08]. Moment [Cai94, DR99, Inl10, KC02, Koo93, McC94, Nav97a, PN95, VE06, CDFP81, KE83, PC85].

Moment-Generating

[Cai94, McC94, CDFP81]. Moments [BSY92, Hos92, Kon11, LB00, MT11, Nel98b, Pau08, Sam92, Sko08, Smi95, WN07, BM86, CDFP81, NA85]. Monotone [RT03, Ste92, Rav86]. Monotonic [Bas91]. Monotonicity [FM84, Shi88a]. Monte [BS98, FF02, KCGN98, KBH09, LH72, Rie90]. Moore [Hu08]. Morris [TW91]. Most [Bir90, Ede90b, Kuk09, SS84]. Motivating [DNG98]. Motivational [FR04]. Mountain [Mon95]. Movies [ZL09]. Moving [Bel05, HS10, YW09, AT86]. mph [LC96]. MR [Kar86, Kar87, SYC88, Zel85]. MSOD [Ehr90]. Much [HK07, Len01c, LB00]. Mudgeon [Mon06]. Multi [Bor99]. Multi-Tiered [Bor99]. Multicenter [WKB^{+}95]. Multicollinearity [FW05]. Multidimensional [Dav08d, Par04]. multifactor [GS86]. Multilevel [Hor06, KDV94, ZPH99]. Multimedia [VM96]. Multinomial [Bro93, CK01, GGR05, KU96, Ang89, Ang90, Cox84, Tho87, Tor78]. Multiple [BHA02, Cua93, Fra92, FW05, HR06, HT83, HL01, HLP03, KC09, LD94, Men00, Nic94, PF95, Sav90, Sch05e, Smi11, TM99, WW97, ZR10, Pav88]. Multiple-Choice [BHA02]. Multiplicative [YW09]. Multiplicity [LC09]. multiprocess [Bol86]. Multistage [CF08]. Multivariable [Joh99, Lis08]. Multivariate [Ald00, AC06b, Ano00b, CY99, Dav08b, Dob06a, EP91, FN92, Gan95, GR04, Gon03, Hu07, Loo95, Sap91, Sko08, SA03, Wil02, ZPM^{+}11, Han79, Rav85, RC88, TC88, WC87]. Muncher [Mon97]. Murder [Gra08]. Mutual [SK95]. MVUE [Muk06]. Myths [BKM96].

Named [ML09]. Nano [BM10].

Nano-Project [BM10]. Napoleon [Gun98]. National [Bol03, You05, CJ92, DM01]. Natural [Jam01b, ML09, OS95, Rob07a]. Nature [BGG98, EA11, Inm94]. NBA [CY99]. NCAA [AW06, Car96, SMH91, SSH96, SS99]. Near [HMR^{+}01]. Near-Optimal [HMR^{+}01]. Nearest [Alt92, SM91]. Nearest-Neighbor [Alt92]. Nearly [JH96, Nel90, SA03]. Nearness [Kha92, Nay94, Ped85]. necessarist [Zel82]. Necessary [McL99, MA84]. Necessity [ST95]. Need [HG97, Hel99, Mon06, Ter03, Tuk80]. Needed [Hog91]. Needle [Rie90]. Needs [Bic95, Ket95, MHT90, McD99, Ros95]. Negative [KC02, MF07, MPR07, Nie11, VBB94, CDFP81, PC85]. Negligible [HS03]. Neighbor [Alt92, SM91]. Nested [LDL^{+}94, LS97, Sea94]. Network [SS01]. Networks [Pal99, WM96]. Neural [WM96]. Neutral [Ran01]. News [PR11, Utt02, DWBQ98]. Next [DAF^{+}94, HS10, Men09, Stu09, vC10]. Neyman [BP02, Chr05, LR82]. NFL [WB02]. NHL [BH08]. Nigerian [Oko99]. Nine [Bar02]. NLOGIT [Hil06b]. No [Lin92, Ang90, CH86, DN94a, GY90, Kar86, Kar87, SYC88, Spi87, WS94, Zel85]. Nominal [VW93, ZRK^{+}11]. Nominees [WASB10]. Non [Kha08]. Non-Simple-Random [Kha08]. nonapplicability [PG82]. Nonbipartite [LGXB11]. Noncalculus [SGC90]. Noncentral [Cha95, LR01, Pos93, WN07, Gue79]. Nonclinical [Kha04]. Noncollapsibility [Gre10]. Nonconstant [Ros03]. Nonestimable [BL87]. Nongaming [Bor97]. Nonhomogeneous [Hag09]. Nonidentifiable [NS97]. Nonlinear [Dav08d, SMVD09, Tju98]. Nonlinearity [WWD03]. Nonlocal [Mag98]. Nonmathematical [Hig99b]. nonnegative [Mul88]. Nonnormality [Ham84]. Nonoptimally [BM88b]. Nonparametric [Alt92, CHJ08, DS98, Jam02, Min07, OH08a, OH08b, Shu82, SS98, SA03, TK76]. Nonparametrics [Noe84]. Nonrandom [MGLK99]. Nonrandomized [TTT09]. Nonregular [NVB09]. Nonresponse [HH04, KH04, Rub04a]. Nonsingular [McL99]. Nonstandard [vS96]. Nonstatisticians [Bra96a, UJ08]. Nonuniform [Nun92]. Normal [AS98, BS99, BH93, CG06, Cai94, CL07, CLPC08, CP99, Dav05, GM91, Gre96, Hut93, JH96, LJ09, Pro08, Sap91, SF93, Van96a, Wal95b, Wie99, Wil08, YB10, Ham84, Han79]. Normal-Distribution-Based [YB10]. Normality [DB90, Haz03, Loo95, AO89]. Normally [tWGB01]. normals [Riv81]. North [Mon04, NOMLT95, Ste08]. Note [ALL10, Ang89, Ang90, BR97, BI09, CLPC08, Cha95, CK01, CH96, Dav93, Dav09, Don99, Ede90b, EM93, GM91, HL96, Hen99, HS92, Hil08b, Hsi95, Koo93, LaM94, LK99, ME11, Oak93, Oeh92, PLC00, Pan02, PT91, Pes90, Pul90, SI90, Sen08, Ste96, Tar00, TL93, Van97, Wes08, ZR10, AT86, All87, Aso80, BM86, BM88a, Buo87, CW88, Fre83, Kah87, KHK89, Khu93, KE83, MN88, Ped85, Tor78, dCS86]. Notes [Ano95e, Ano95f, Ano95g, Ano96d, Ano96e, Ano96f, Ano96g, Ano98g, Gol97a, Gol97b, HL98, Hil97a, Hil97b, Hil98, Hil99, Hil00, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil09, Ost09a, Ost09b, Ost11d, Ost11a, Ost11b, Ost11c, Gle86]. Nothing [HK07, Jam01b]. Notions [Muk10]. Novick [Gei84]. Null [SBB01]. Number [Ang92, Bar02, DL00, Fre05, WE08]. Numbers [AC05a, AC05b, BR03, Cle06c, Coc05b, Hel99, Koo93, She07, Utt02, Goo69, SSH81, Tom84]. Numeracy [Sow03]. Numerator [WSF02]. Numerical [Alt02, KD87, Luc97, Rid09, WTH95, Yal07]. Numerically [Wal95a]. Nursing [GS08].

Objection [MD97]. Observation [Muk06, Olk92]. Observations [SVR11, Han79, Tuk47]. obtain [Kap79]. Obtaining [Pes90, Smi95, TL93, WTH95]. Occupancy [WMAY09]. Occurrence [Dav95, Dav98a]. Odd [Ang92]. Odds [Han92b, LGH09, MW92, Rob03, WL99]. Office [Cle03d]. Official [Bra07, Hol07, Nor07]. Okumoto [JP01]. Old [Sen90, Smi95]. OLS [BF91]. Olympic [ESL09, Bra07, EA11, Hol07, NK96, Nor07]. Once [LC99]. One [AP10, Ano98j, BN98, BK96, Bis98, Cha04, CW04, Czi98, Czi99, Dan94, GS08, LC09, Law03, LT97, Lev97, LC99, LDM10, MS06, Muk06, O’C90, Sil97, Spu98, Var98, WBT01, ZG00, Bar76, Gue87, Hog79, SD87, Sch85]. One-Day [Cha04]. One-Dimensional [CW04, LT97]. One-Factor-at-a-Time [Czi99, Dan94]. One-Half [LC99]. One-Hour [GS08]. one-parameter [Gue87]. One-Semester [Ano98j, BN98, Bis98, Czi98, Spu98, Var98]. One-Sided [LC09, Sil97, ZG00]. One-Step [Law03]. One-Way [BK96, LDM10, SD87, Sch85]. Online [Hel99, HQ08, HJS06, LJLC05]. Operating [BK07, Hal08]. Operator [BSY92, Goo79]. Opportunities [KYA^{+}07]. Opposition [Dav98b]. Optimal [CB95a, HMR^{+}01, LGXB11, OF03, WC87, Zel88, Zuo10]. Optimality [Lee95, Shu82]. Optimization [Lan04, LG05, Law03, Pau81]. Optimizing [Nie05]. Oracle [MCM^{+}11]. Order [BR97, BP02, Dav93, DN94b, FP81, Mon06, SC97, BM86, DN94a, GY89, GY90, LH72, ME84]. order-statistics [GY89, GY90]. Ordered [LTN10, MFS^{+}03, BH86]. Ordering [Nie05, SRMP08]. orderings [Kru87]. Orders [KR99]. Ordinal [VW93]. ordinary [BvE88, MA84, PS89]. ordinary-least-squares [BvE88]. Oriented [Alb97b, Ano98j, BN98, Bis98, Czi98, Spu98, Var98]. Origins [BM02b]. orthant [KHK89]. Orthogonal [CR96, Lan06b, Sch85]. Orthogonality [deL06]. Oscar [PS89, WASB10]. Other [CF08, Cle03d, Gol92, Kir91, Lin92, Pic06, SG98, TGM08, Var92, Kon87]. Our [Hog99a, She07]. Outcome [GAR92, GM96, Mat99, MFS^{+}03, SRW95]. Outcomes [BLS00, CF08, LBTM03]. Outlier [Fin06]. outliers [Shi88b]. Output [OCE93, Pra08b, Sea94]. Outreach [KM94]. Overdispersion [Tju98]. Overlap [SG01]. overlapping [Tam84]. Overruled [MD97]. Overtime [BH08]. Overview [Hog92, WG11].

p [KP07]. Package

[DS98, Sea94, SS98, Spi90, Tho95].

Packages [BLS00, Cal03, Coh97, Dal90, DHN^{+}05, Eme96, GC04, HDE^{+}03, HKS06, HLW09, HL01, Hou97, HF96, Kel04, Loc93, Lur95, Mor98, OCE93, Ost98, Ost02, Ost03, OH08a, OH08b, Sch93c, ZPH99]. Pages [LP00]. Paired [DG97, Day98, McN92, MC91, OSC00]. Paired-Comparisons [Day98]. Paired-Data [OSC00]. Pairing [DG97]. Pairs [Dou93, EGS96, Gre96, HRS10]. Pairwise [SK95]. Paleontology [MS01]. Paper [Ann05b, Igl07, Sto98]. Papers [FGHK83, Sch10a]. Paradox [Ano00g, AFV96, CBH09, CU92, Fre65, KS05, Sen08, Shu06, SJD10, War95, Hil84, Bel05, CBH09, Gre10, PP09, RT03, SS01, SR04, SR05]. Paradoxes [BS00b, WB04]. Paradoxical [Hom97]. Parallel [Nel90]. Parameter [Bar02, Dou93, HEJ11, HM01, LaM08, LT96, MV07, MKV92, SGM07, AS84, Bly80, Dah81, GD89, Gue87, Kah87, Spi85]. Parameters [DN94b, Fer90, Iri01, RS93, SS91, TL93, War97, WB91, DN94a]. Parametric [CHJ08, Hil02e, OH08a, OH08b, Ros00a, SA03, BL87]. Pareto [Wil06a]. Part [Eng93, Hil04b, Hil05b, Sch93c, CH86, Kar86, Kar87, Hil03c, McC98, McC99, Ost03, Tuk49b, Tuk49c]. Partial [CW04, DK87, KR99, PP10, TO93, MC87]. Participation [Gel98, Mag96]. Partitioning [CS91]. partitions [Mee86]. Passing [Lan09]. Passion [Eas10, Lan09]. Passion-Driven [Eas10]. Past [RSW99, ZH90]. Pattern [Pri11]. Patterns [Rod04]. PC [Coh97]. PC-Based [Coh97]. PDF [HT06]. PDFs [Nad09]. Peace [You05]. Peak [KS08]. Pearson [BP02, Chr05, Cle05a, Cle05b, Dav81, Inm94]. Peculiar [Ley96]. Pedagogy [BM02b]. Peirce [BK07, LR06, RA04]. Penalized [Mea03, PW06]. Penn [HR05]. Penrose [Hu08]. People [Kir91]. Perceive [TS03]. Percent [BA06]. Percentage [MSL^{+}08]. Percentage-Based [MSL^{+}08]. Percentile [CL07, Pol99]. Percentiles [Dun91]. Perfect [CLR01, Mad91, DK87]. Performance [Ann06a, GAR92, MCM^{+}11, RMY96, Tob04]. Period [BH08, PS94]. Permutation [FF02, LD98, ZPM^{+}11]. Perot [Gun98]. Personal [BG08]. Persons [Moe05a, Moe05b]. Perspective [Bai95, Jam01a, OS95, SG92]. Perspectives [RKW98]. Perturbing [Wil08]. Pervasive [HH01]. Petersburg [Bel05, Rum07, SR04, SR05]. Petition [WE08]. Ph.D. [Leh95]. Pharmaceutical [Bra08b]. Phenomena [GPA97, NQ96]. Philosophical [Cle06b]. Phrases [GT95]. Physical [Old95, ZH90]. Pi [Jad00]. Picture [FS01]. Picturing [Cha08]. Pieces [Dob03b]. Pilot [MB06]. Pin [CS94]. Pinch [Dun91]. Pioneers [Hal10]. Pitching [RMY96]. Pitfalls [VM96, WASB10, Spi86, Spi87]. Pitman [FK97, Kha92, Nay94, Ped85]. Place [Jon02, Moo93]. Placement [RSW99]. Plan [SVR11]. Planning [IJW06]. Play [FDC09]. Player [MCDD91a, Shu06, WC97]. Players [Lac90, SG00]. Playfair [SW05]. Playing [Ann06a, Coc05a]. Playoff [AW06, BH98]. Playoffs [Bor99]. Plot [AH02, CC06, Cor00, FM92, HN98, LT97, Pad08, SMC^{+}07, LKNP95]. Plot-Density [HN98]. Plots [BM03, DR07, GKT91, GB03, HN98, KR99, Mon95, MB00, Nel90, Sal90, SW94, SRMP08, Wil99, MC87, MTL78, Mon79, Sti86]. Plotting [Bec96, Lan06b, WT92]. PLUS [Cow03, Hal92, KK99, Man03, Har06, Dla05, Pri05b, Lun02b]. Plutarch [Mis06]. Pocket [Gil08, Hoe98]. Point [Ano00g, Car96, CC01, GR96, Kes00, CH85, CH86]. Points [FRD07, FT93, Spi07]. Poisson [GY90, Alb92, Bar02, GY89, Ger79, Jol95, Kel94, Knü94, LJ09, NW86, Neu92, PA91, She84, Tie94, Wal95b, War97]. Poker [BD11]. poLCA [HLW09]. Policing [Men10]. Policy [Kaf98, Saw98]. Political [Gel08]. Polls [CF08]. Polychotomous [WB02]. Polynomial [BS79, Mag98, Pei90]. Polynomials [AC00b]. Polyplot [SB11]. Pool [DM01]. Pool-Adjacent-Violators [DM01]. Pooling [Cor00, Hin96]. Population [DC90, Kha92, LR08, MG04, Nay94, SSM80, SP94, Bel87, VY85]. Populations [KT02, RZR05, SGCS04, WS01, Buo87, PG82, Tor78]. Position [BTV90]. Positive [PN95, VAD90]. Positively [LSO01, ZG00]. Possible [Bly93, PG97, WY04, Mee86]. Post [Hog92, Zha00]. Post-Enumeration [Hog92]. Post-Stratification [Zha00]. Posterior [Ros96]. Posttest [YT01]. Potential [AW06, HLP03]. Power [Bro01, CRM01, CL02, HH01, Kor90, MSL^{+}08, MH93, ME11, Ott96, PA91, SG91, San06, She97, TM95, WB91, vDK02, Gue79, PS81, Shi88a]. Powerful [Ber96a, Bir90, DB90, Ede90b, Kuk09, SS84]. Powers [Suk92]. Pr [HL96]. Practical [Bra08b, DH11, Dev07c, Gen03, Hoe98, Huz99, Len01d, LF00, Lis08, NK96, Par98, Sav90, SB09]. Practice [BRRB01, Bro01, Can98, CRM01, CK01, DLM01, Dob03b, Dob04, Fay02, GPD02, HMR^{+}01, HL01, KCGN98, Len01d, Maa99, MS75, MS01, Nas06, Per98, Ran01, SG01, SB98, WSF02]. Practices [Fin06]. Pragmatic [GM96]. Pragmatics [Hin96]. Pre [CF08]. Pre-Election [CF08]. Preach [GPD02]. Precise [Dun91, SBB01]. Precision [BS11]. Predict [Len01c, SS99]. Predicting [BCJ97, CF08]. Prediction [Ald00, LS08, Nay02, SM91, Tar00, Vic08, Wan10, YA09, Knü94, Kon87]. predictions [Car86]. Predictive [AJ88, AT04, Ibr97, CH85, CH86]. Predictor [Eli00, GP09, Ver99]. Predictors [DB97, LBTM03, Sea97, Sim98]. Preliminary [Gol98, KR02, MM90]. Preparation [Wie08]. Prepared [BV11]. Preparing [HK92, Moo05]. Preposterous [Han92b]. Preschool [HGA01]. Presence [AB03, BH08, FKLL03, KP07]. Present [RSW99, ZH90]. Presentation [BKM96, HQ08, Wes90]. Presenting [FGHK83]. Preserving [FM84]. President [RWM^{+}72]. Presidential [CF08, WJM^{+}03]. Pretest [YT01]. Pretest-Posttest [YT01]. Pretesting [GAR92]. Prevalence [Kva96]. Previous [Kor90]. Previously [DHW11, WWD03]. Primer [For05]. Principal [BK99, FT93, HL98, JK99, NK96, Ren92]. Principle [BCB00, Hel95, Nel98a, Pri11, Ber87, Hil87]. Principles [Can98, Sch00, SGCS04]. Prior [LS97, Lon10, Mat99, Ros96, SS02, Gei84, Zel85]. priority [DN94a]. Priors [Ibr97, San06, TL10, TGM08, WSF02]. PRISM [Zim00]. Proactive [Hah02]. Probabilistic [GPA97, Gol86, Ros07, She98, Wüt08]. Probabilities [AC00b, BD11, Bro93, KU96, LTN10, PP10, PN95, Pul90, She07, Tie94]. Probability [Alb03a, Ang92, Ann05a, Bar99, BG08, BW08, BM94, Bor97, BSY92, Car08, Cle03d, Cle06b, Cle08a, Coc05a, Coh02, Cow06, Dav98a, Dem75, Dev06b, Dob04, DCWP93, FR04, Fot02, GKT91, Gho02, GEL01, Gor11, Hai07, Har08, Hut93, Kaz95, Kie06, Kir00, LS08, Lun04b, Mad91, Mil03, Nel01, OS95, Rob03, Rus08, SP98, SD01, Sch06b, SMH91, SSH96, Sol93, Ste01, Ste91, Utt03b, WT92, Wal95b, WSF10, EG87, FF80, Kel94, KHK89, Cle03d, Dev02, Dob03a]. Probit [CG06]. Problem [BES98, Chu99, Cle08b, Daw91, Day98, DO04, HH04, Hil96b, Hu92, Kuo00, Mal98, Meh09, Nun92, Pal99, RS91b, Rea98, Smi95, Sny05, SS09, WMAY09, Knü94, Pau81, Thi87, MS01, Rie90]. Problem-Solving [Hil96b]. Problems [Dan92, DBS^{+}85, Dob03a, Dob04, HV03, NVB09, Oko99, Sim94, UE90, WSF02, EW85]. PROC [Ann05c, Mon97]. Procedure [And09, CRM01, HD94, KH04, ML92, St.90]. Procedures [Ber96c, HS91, Hil02e, Sav90, UE90, WG11, AJ88, Pav88]. Process [AC08b, Ano00g, Ber96c, BM10, KM92, Law03, Sto98, GY89, GY90, LP76, Neu92]. Processes [Dev06b, Hai07, Lun02a, Lun04b, Lun04a, Per93, Rus08]. processing [Zel88]. productivity [HBF90]. Products [Hil03c, MS98, Rog84]. Profession [CF91, HS10, Igl98, Lan09]. Professional [Bor99, Gil00, Hil07a, Lac90]. Professor [Mon06]. Profiles [LRG90]. Profitable [Len01c]. Program [LRZG^{+}10, Zim07]. Program-Driven [Zim07]. Programmer [Gur08]. Programming [Cha04, GEL01, Spe08, Zim07]. Programs [Hog99a, KDV94, Leh95, Moo01, Tho95, Hei81]. Progression [Mad91]. Progressive [BS95]. Project [AC98b, BM10, DiP04, HR00, Yu94]. Projecting [Kor90]. Projection [BK99, Tuk85b]. Projections [Nic94]. Projects [AS99, GNM^{+}98, Led95, RT01, Smi11]. Promises [VM96]. Promoting [Bol03]. Proof [DS08, Gla04, Hwa09, Inl10, Ros07, SS91, SGC90, DK95, Far89, GY89, GY90, Har76, PG82, Tom84]. Proofs [Gla04, Ans85, FF80]. Propensity [PMO08]. Properties [DR01, Ibr97, JS03, Per93, WSS93, vDK02, Ber81, Len83, SP85]. Property [DB97, LSO01, Lee95, SC97, Tur97, Wil08, Ger79, ME84, Neu92, Pei90]. Proportion [Afs08, CRM01, CL02, Fre07b, Fre10, MG04, RvE97]. Proportional [Gan95]. Proportions [AC98a, AC00a, KT02, KMS10, WS01, WS93, Ang89, Ang90, Buo87, SR83, Tho87, WS94]. Proposal [Tju09]. Prospective [RSH01]. Prospects [Oko99]. Prosper [HHJL93]. Prosperity [Sny05]. Protect [KKO^{+}06]. Providers [JS11]. Proving [Zeh91]. proxy [Bar76]. Pseudo [YB10]. Pseudorandom [Dor96]. Psychiatry [Ham00]. Psychologists [Row02]. Psychology [Jam01a, Mru07]. Public [AC05b, CF91, Gel08, Kaf98, Saw98]. Publication [SRW95]. Publications [SATV94, Wio90]. Publish [SRW95]. Published [Igl07]. Pulling [BS10a]. Pursuit [BK99, Tuk85b]. Putting [Dob03b, SW90]. Puzzle [Fre65]. Puzzlers [Cle03d]. Puzzling [Sim94].

Q [GCGG95]. Q-TWiST [GCGG95]. Quadratic [Dri99, Saw92, Sch96b]. Quadratic-Type [Sch96b]. Qualifying [BM10]. Quality [AC07a, AC07b, Ano98j, BN98, Bis98, BO92, Czi98, GCGG95, Git94, Gon94, HHJL93, Hoe98, Hyd01, Imr94b, Kas98, Spu98, Str90, Var98, Zam08, HBF90]. Quality-Adjusted [GCGG95]. Quality-Oriented [Ano98j, BN98, Bis98, Czi98, Spu98, Var98]. Quantify [SR05]. Quantifying [BWM98]. Quantile [Gro98, He97, KR99, Ros00a]. Quantiles [HF96, Sho95]. Quantitative [BR78, CS94, Cle01, Cle06a, PH02]. Quantitatively [Sch90]. Quantities [JSM95]. Quantum [Tur07]. Quarter [GP09]. Quarterbacks [WB02]. Quasi [Sal99, Tju98]. Quasi-Experiment [Sal99]. Question [AS94, Tuk47, Tuk49a]. Questions [Mon06, TTT09]. Quick [AC08b, Ans85, MY05, Ang89, Ang90]. Quid [Fox10]. Quincunx [PR10]. Quintessential [PR10]. Quo [Fox10]. Quota [BWM98].

R [Ano91g, Ano91h, Cow03, Gei84, Mon06, Nas95, PS89, Bat04, Har06, HBQ04, Lun07b, Mäc07, Pri05a, Pri05b, Rob08, WC07]. R. [Box80, Box05, Inm94, Lud05]. R/S [Har06]. R/S-PLUS [Har06]. Race [Saw98, Wal98]. Racing [SB98]. Raindrop [BM03]. Random [ALL10, BRRB01, Beh90, Bel01, BK96, BH86, Cai94, CK01, DL00, FR04, GL00, Grö09, Hai07, HT06, HR98, JB04, Jon02, Kah06, Kha08, Lee93, LDM10, MB00, MTA08, Nav97a, Nel01, Pal08, Rus08, SSC99, Sea97, Sha91a, Sta04, Ter05, Tur07, WSS93, Wil08, YB10, Alb92, BG89, CS83, Flu86, Goo69, GP89, Ham84, HB96, KP79, PK82, Roc88]. Random-Effects [ALL10]. Randomization [Dav08a, Gad01, Moe05a, Moe05b, Old95]. Randomized [DiP04, GE11, Mur86]. Randomly [JK90]. Range [Bro01, Fay02]. Rank [Ald00, Kuk09, LW90, MV94, MLL10, Ros03, Suk92, VM90, WB91]. Rank-Based [MV94, MLL10]. Ranked [McI05]. Ranking [Fre05, Gro90, Mea03, WB02]. Rankings [GDL08, DK87]. Rao [Kag01, Nay02]. Raster [Bea83]. Rate [BGG98, ZL09]. Rater [LR06]. Raters [ZL09]. Rates [CLS07, Gra08, HS03, Lon10, Lor90, SD04, TM99, VD05, Pav88]. Rating [ZL09]. Ratings [Bas97, HQ08]. Ratio [Bir90, DC90, JSM95, Mag90, MW92, MV07, PE05, RS93, Sil97, Ste92, VW93, vM93, ABS81, Dav80, Han79]. Ratios [LGH09, WL99, AB92, Kot86]. Reaction [Ste90]. Read [Mur97]. Readmission [ZRK^{+}11]. Ready [SATV94]. Real [Ber97b, DLM01, Dor96, JS03, Kha08].

Real-Life [Kha08]. Realism [Ste07, WC97]. Really [Coc05a, FDC09, Gat95, Gla04, SS84]. Realm [Han92b]. Reasoning [AC06a, And99, Bra96a, Gas92, Hay99]. Reasons [Alb97a, Ber97a, Fre97, Lin97, Moo97a, Moo97b, Sch97, Sho97, Wit97]. Reassessing [CLS07]. Receiver [BK07, Hal08]. Recognition [Pri11]. recollections [Noe84]. recommendations [CGL83]. Reconstruction [Yu94]. Record [BTV90, Sol93, SS05]. Records [NK96]. Recreational [SB98]. Recurrence [DCWP93, WN07]. Recursive [Sko08, Smi95, Tie94, Tie98, CJT87]. Reduced [Ald00]. Reduced-Rank [Ald00]. Reducing [GR04, KH04]. Reduction [SM09]. redundant [Ham87]. refereeing [Gle86]. reference [Edw86]. Reflections [Hah95]. Reformulation [Lee09]. Regarding [Ter05]. Region [Nic94, CCW^{+}86]. Regional [SMH91, SSH96]. Regions [CC01, Dou93, Hyn96, LTN10, ME95, Car86, RC88]. Registration [SD04]. Regress [SG00]. Regression [Ald00, Alt92, AC04b, BCJ97, BS79, Bra96b, Bri94, Bri96, CR96, CBH09, CHS99, Coh91, CW04, Cua93, DC90, Dev07a, DR01, Dun94, EM93, Fay02, FLP07, Fra92, FW05, Gra08, GW94, GAR92, Grö07, Grö09, HL98, HHB91, Hol99, HL99, HL01, Hsi95, HT90, Iri01, Jam03, Joh99, JK99, Jor94, II07, KR02, KG90, Kot91, KP07, LaM94, LD94, Lev93, LM03, Lic90, LE00, MS75, MG97, MC91, Men00, Men04, Mil96, Mon06, MFS^{+}03, Nas92, Nel98b, Nel07, Nic94, OF03, OW91, OCE93, PLC00, PC02, Pei93, Pie92, Rak06, RT03, Sam91, Sen90, Sim07, Sim98, Smi98, Smi11, Str92, Tju98, Tju09, UE90, WWD03, WB02, WL99, XN03, Zho01, All87, Ans85, CW88, DH87, Fre83, Man82, Pei90, Spi86, Spi87, Thi87, Cow03]. Regression-Toward-the-Mean [Lev93]. Regressions [BF91, Mag98, Duf86, TC88].

Regressors [Sha91a]. Regrouping [RT03]. Reinforcing [CF91, Kha91]. Rejoinder [Ano03i, BK09a, Men10, MCDD91b, OH08b]. Relabeling [CW11]. Related [Dav05, Gho02, Sha91b, SC97, Utt02, XN03]. Relating [Dri99, JW04]. Relation [CBH09, JK99, WN07]. Relationship [Bro10, IBP05, Old95, SGK00, Sch93a, Sha91b, SR97]. Relationships [Ber96b, Jon99, Lee86, LM08, RS91b]. Relative [Far90, Grö07, HB05a, Kru87]. Relatives [ML09]. Release [Tan99]. Reliability [JP01, Mac93, McC98, McC99, SS01, WJM^{+}03, Yal07]. Remain [SR05]. Remaining [Str90]. Remarkable [Jam99, SG93]. remarks [BS87]. Reminder [FBG96, Paw00]. Reminiscence [Box05]. REML [Gur06, LDM10, tWGB01]. Removable [Sal99]. renewable [DBS^{+}85]. Repeated [DGC97]. Replacement [KH04, Aso80]. Replicated [LE93]. Replicating [Ros01]. Replications [BM02a, Ott96]. Reply [Alb97a, Ano98j, Ano99e, Ano04h, Ber97a, GKR97a, GKR97b, Hig99c, Hog99b, Imr94a, Moo97b, Efr86, Gei84, PS89, Zel88]. Report [Alb04, Alb05a, Alb05b, Ano98b, Ano99b, Dob06b, Lan97, Lan99, Lar94, Lar96, RWM^{+}72, Sch93b, Sto01, Wes07, Mon97]. Reported [Don99]. Reporting [DuM99, Utt02]. Reports [Alb03b, Sto02, SA02]. Representation [Gib90, YW09]. Representing [Bor02, FT93, Hid86]. Reproducibility [BS11]. Reproducing [PW06]. Republication [SW05]. Requirements [SL92]. Requiring [SVR11]. Resample [Ros96]. Resampling [Gen03, Har06, Sim94, SG92]. Research [Ann06b, AW91b, Gel04, HS04b, Hol99, Kha04, KC09, LRG90, LRZG^{+}10, Los04a, LD98, Rub04b, Sta04, SZM^{+}07, Str02, Woo08, Nag98]. Resequencing [MGCR10]. Residual [KS91, Nel90, Ste07, MC87].

Residuals [GW94]. Residue [Hag09]. Resolution [CU92, CH96]. Resolving [Vos99]. Resource [OS95, Won98]. resources [DBS^{+}85]. Respect [AG95]. Respondent [Mor95]. Response [CC01, DiP04, Fay02, HS04a, MKV92, Nel98a, PLC00, Rak06, Stu90, SR05, TTT09, Van96b, vC10, CCW^{+}86, MH76, MS87]. Response-Surface [Nel98a]. Responses [CW92, LDL^{+}94, Ran01]. Restrict [Bly93]. Restricted [Tar00]. Restriction [DN94b, DN94a]. Restrictions [Sea97]. Result [AC00a, Dri99, SC97]. Results [Bra96b, DS08, GK99, Wie99, CJT87, Kap79]. Retention [BUK^{+}08, Kva00]. Rethinking [Ann05b, GZK^{+}11]. Retired [FG00]. rev [McG05]. Revealing [Sch00]. Reveals [Wri92]. Revelations [Gun98]. Revenge [Lun07a]. Reversals [Gil00]. Reverse [Shu91]. Reversible [AKLS10]. Reversion [Sam91]. Review [Abb02, Alt02, AC02, ACA03, Ank98, Bar98, Bar99, Bat04, Bon02, Can98, Cle03c, Cle03b, Cle03a, Coh02, Cor00, Cot98, Cow03, Cow04b, Cow04a, Cow04c, DHN^{+}05, Dev02, Dob02, Dob03b, Dob03a, Fot02, Gen03, Gol98, Gon03, Gup02, Hai03, HH06, HDE^{+}03, HKS06, HLW09, Hay99, Hil03c, Hil04b, Hil05b, Hil05d, Hil05a, Hil05c, Hil06b, HL99, Hou97, Hu03, Jam99, Jam02, Jam03, Jer03, KK99, Kaf99, Kel04, KT94, KDV94, LJLC05, Led03, Los04b, Low98, Lun02a, Lun02b, Lur95, Maa99, Mac99, Man03, Mil04a, Mil04b, Mor98, Nag98, NO98, Nor95, Obr98, Par98, Pos90, RH02, Rob03, Row02, Rus98, SP98, Sho98, SMP97, Tan02, Ter03, Utt02, Wal02, Wil02, Won98, Wul03, AC03, Cle99a, Gol99, Min07, Sch99, Tan99]. Reviews [Alt02, Ano90d, Ano91e, Ano91f, Ano93b, Ano95h, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano96k, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98f, Ano98h, Ano98i, Ano02e, Ano03a, Ano04b, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, Ano10f, Ano10g, Ano10h, Ano11i, Ano11f, Ano11g, Ano11h, Bon98, Cle01, Cle03d, Col01, Dev01, DL04, DL05, DAC05, DAC06, Hai01, Hil01, Hil02c, Hil02d, Hil04f, Hil07g, Hol99, HL01, Hyd01, Hyu01, Jam01b, Jam01a, Kre01, Len01b, Len01c, Len02, Len04, Mac01, Mil01, Mil03, Pil98, Pou02, Ree02, RACK07, RD05, Rum01, Sha01, Smi98, Str02, War02, Zim00]. Revising [Wal98, Wil06a]. Revisited [Daw91, KK91, KC02, Meh09, SRW95, ZM94]. Revisiting [NK96, WMAY09]. Reviving [MF07]. Rho [RS91b]. Rhythm [EKG02]. Ridge [Hoe85, MS75, DH87]. ridge-regression [DH87]. Right [Ama99]. Risk [DNG98, GE11, Gue06, Hai03, KS91, KR99, MZ92, ABS81, Nai87, Jam99]. Risk-Benefit [Hai03]. Road [SB98]. Roads [Vid94]. robin [DK87]. Robust [Bas97, Fre06, MKV92, SI82]. Robustness [Sti10, CW88, Hog79]. Rock [BHM01]. Rocket [MZ92]. Role [Bil98, Boy08, HF04, LaM94, Old95, SK95, Wal05, WSF02, Ber87, JH83]. Roll [BHM01]. Rolled [Dun03]. ROM [Mil03, Sch04f, Sne98]. Rome [Vid94]. Ronald [Hal10]. Root [Fos11, Hut93]. Rosenbaum [MD97, PG97]. Ross [Wan91, Gun98]. round [DK87]. round-robin [DK87]. Rounding [HLP03, YHZ08]. Roundtable [KCGN98]. Rudolf [Rie90]. Ruin [Pal99]. Rule [Alb97b, Bet00, BP02, Fos11, JL97, Puk94, She07, WY04]. Rules [Cle03a, Sch06b]. run [Alb92]. Runners [SB98]. Runs [Ann06a, RMY96]. Rural [GKT98, LC96]. ruses [VY85].

S [Cow03, Dav81, KK99, Man03, Zel85, Dla05, Pri05b, Lun02b]. S-PLUS [Cow03, KK99, Man03, Har06, Dla05, Pri05b, Lun02b]. Saba [SS09]. Sacco [She98]. Saddlepoint [GC99, Huz99, Kuo00]. Salaries [HV95, WC97]. Salary [Lac90]. salvo [Tuk86]. Same [Sch10b, WS93, FK86, WS94]. Same-Day [Sch10b]. Sample [Afs08, Ama99, BM01, Bro93, Bro01, CRM01, CL02, CJ92, Die93, DK07, Ede90a, FMA^{+}09, Fre05, GJLW05, HH04, Hay11, HR98, HF96, IBP05, Jol95, Kha08, Kon11, KG95, LaM08, Len01d, LC99, LDM10, MS06, Mat99, MY05, MC91, MS92, NP97, PB03, Per93, Ros96, Rub04a, SL92, She97, SP94, SRC07, Suk92, TTT09, TM95, Tho87, War97, WSF10, Wri90, Zha07a, vM93, Ber81, CGL83, Eat78, Gue79, Knü94, ME84, Shu82, Thi87, Tor78]. Sampler [CG92, CLR01, MB94]. Samples [BS95, Kha08, Lev97, MFS^{+}03, Pes90, Wal00a, WBT01, Tam84]. Sampling [Alb93, AL07, Bel01, Bel07, Bul05, CLM92, DC90, Dor96, FKL^{+}96, FM84, HS03, LR01, Loh04, McI05, Mis06, OS95, PHSN08, Rya09, SS91, SG92, AJ88, Aso80, Bel87, SR83, WC87]. Samuel [DM06]. Samuelson [Far89, WS79, WS80]. Sandwich [Fre06]. SARS [YCF06]. SAS [Gup02, Ann05c, Ano00b, Bra08b, Cha04, Cot06, Cow04a, Dix07, Dob06b, DS98, Gur08, Hal08, Kas98, Mon97, Pra08b, Pri05b, SS98, Spe08, UE90, Wie06, Wie08, Wul07, Zim07]. SAS-Based [DS98, SS98]. Satisfying [LSE00]. Saw [CL02]. Saw-Toothed [CL02]. Say [Lin92]. Scale [WB91]. Scaled [FK97]. Scaling [Par04]. Scatter [Upt01]. Scatterplot [SG93]. Scatterplots [EGS96, KG98]. Schedule [FT10]. Scheffé [Khu93]. Schemes [Tie94, Tie98]. Scholarly [GT95]. Scholarship [Wil92]. School [Bol03, FHP^{+}11, HK92, Ter03, Tob04]. Schools [LH02]. Schucany [Ano91g, Ano91h]. Science [Ann05a, Arn07, BV11, Ber97b, BGNP01, Cle05a, Cle05b, CF91, Cow06, Dev06a, Gel08, Rie06, Wüt08, BGNP01]. Sciences [Dav08c, Fer04, Gon03, Kre01, Lee08, Rob07a, Wue98, ZH90, Bon98, Hu03, RH02, Tan02]. Scientific [Cle05a, Cle05b, SO94, HH06].

Scientists [Car00, Car08, Kie06, Moo07, SG98, Wei07, Ste01]. Score [Boo92, Fre07a, MV07, MPR07, PMO08, St.90, VM07]. Scores [Kva96, Shi88b]. Screening [Law03, Fre83]. Sealy [Bol84]. Searching [Ama99]. Searle [PS89]. Seasonal [YW09, Pie80]. secant [MN88]. Second [Dav98a, Dev07a, Jam01a]. Secondary [Bol03]. Secret [FVMU10]. Section [AG04, Hil04c, Hil04d, Hil04e, Hil05e, Hil05f, Hil05g, Hil05h, Hil06c, Hil06d, Hil06e, Hil06f, Hil07c, Hil07d, Hil07e, Hil08c, Hil08d, Hil08e, Hil08b, Hil09, Ost09a, Ost09b, Ost11d, Ost11a, Ost11b, Ost11c, Utt03a, ZH90, HBF90, Ano98g, Gol97a, Gol97b, Hil97a, Hil97b, Hil98, Hil99, Hil00, Hil02c, Hil02d, Hil02a, Hil02b, Hil03a, Hil03b, Hil04a]. Sectioned [CC06]. Seeding [SS99]. Seeing [AC07c, Cav07a]. Seek [BHA02]. Seemingly [BF91]. Segmented [BWM98]. Seldom [Don99]. Selected [Dob02, Igl98]. Selecting [Die93, Gur06, Par04]. Selection [Bar07, BWM98, BCB00, HCW07, HT90, JHFK04, MR09, Nel98a, OW91, Sal99, SA00, Sen08, VY85]. Selection/Quota [BWM98]. Selective [McI05]. self [Hid86]. self-representing [Hid86]. Semester [Ano98j, BN98, Bis98, Czi98, Spu98, Var98]. Semiconductor [ST95]. Semielasticity [HF04]. Seminar [GS08]. Semiparametric [BI09]. Sense [Dev07c]. Sensitive [TTT09]. Sensitivity [SGM07, SS02]. Separate [BM02b, GC04, PF95]. Separation [XN03]. Sequence [GGR05, HR00, Jad00]. Sequences [BHA02, BH98]. Sequencing [CR01]. Sequential [Eme96, Fre10, LS01, ZNY11]. serial [Sti86]. Series [BH98, Ben93, CHS99, Cry91, KT94, PT91, PA91, Pou02]. Service [AS99, BHM01, RT01]. Service-Learning [RT01]. Services [Gal03]. Sesquicentennial [Ano90a, Ano90b, Nor90, S^{+}90]. Set [MD96, Str90, WSS93]. Sets [AS98, FS93, McI05, Nai87]. Setting [DR01, Eng93, Gas92]. Seven [BH98]. Seventh [Hut93]. Seventh-Root [Hut93]. Several [MW92, ZRK^{+}11]. Sex [BHM01]. Shading [Jac08]. Shall [Gra08]. Shall-Issue [Gra08]. Shape [Hos92, RM94]. Shaped [LBTM03]. Shaping [McD99]. Shareware [Nas92]. sharing [AAF72]. Sharpening [Cot06]. Shaved [PP10]. Shayle [PS89]. Shewhart [SF00]. Shoeless [Ben93]. Shoes [AR95]. Short [Bir95, CRV95, Eng93, Kle95, Ped85]. Shortest [Fer90, Juo93, WL99]. Shot [RHCR06]. Should [Ano00g, FG00, Ros96, Utt03b]. Show [Chu99]. showcase [CB95a]. showdown [CB95a]. Shrunk [Ald00]. Shuttle [MZ92]. Sided [Ede90b, KB07, LC09, Sil97, ZG00, vDK02]. sidelight [DH87]. Sides [Atk94]. Sidney [DS81]. Siècle [MCGM95]. Sigma [AC07b, HHHZ99, Puk94]. SigmaPlot [Hil03c, Hil05a]. SigmaStat [Hil03c, Hil05i]. Sign [FKLL03, Ran01]. signal [Kra88]. signal-detection [Kra88]. Signatures [WE08]. Signed [Ros03]. Significance [Fos11, Mag90, ME11, SG01, ZCM95]. Significant [GS06, Rod04]. Significantly [Chr03]. Silvia [Eli00]. Similar [JR92]. Simple [AC00a, And09, BS95, BF91, Bel01, BK96, Eli00, EGCD91, Eub94, Few09, Fre07b, GS02, GAR92, GM96, HR98, Hel95, HD94, Hwa09, Jon99, Kha08, Kos05, Oks91, Pal08, PPS96, PO07, Ros07, SGK00, SGC90, Shi01, Sol00, Tur97, Yu94, Aso80, FF80, GP89, Nai87, PG82, Thi87]. Simplified [FS93, VM90]. Simplifying [HH99]. Simpson [AFV96, CBH09, Gre10, PP09, RT03, War95]. Simulated [Dun91]. Simulating [MS02]. Simulation [CLM92, KBH09, KP07, Lin93, MMCsS96, Rus91, Shu01, Sim94, dML09]. Simulation-Based [KBH09, dML09]. Simulational [BS95]. Simulations [AW01].

Simultaneous [Car86, KG90, LD94, PE05, Ang89, Ang90, Wom85]. simX [Zeh91]. sine [Bur86]. Single [Bon02, CL02, PF95]. singular [EW85, Man82]. singular-value [EW85]. Sinh [New01]. Siobhan [Daw91]. Sir [DF07]. Situations [Hom97]. Six [AC07b, HHHZ99, LJLC05, Men04]. Size [Afs08, Ama99, Bro93, Bro01, Bro10, CRM01, CL02, Dor96, Ede90a, FMA^{+}09, GW94, IBP05, Len01d, LC99, NP97, PB03, SL92, SP94, SRC07, TM95, WBT01, WSF10, vM93, Eat78, Gue79, Thi87, Tho87, Tor78]. Size-Biased [SP94]. Sizes [Jol95, Tim04]. Skating [EA11, WY04]. Skew [CLPC08]. Skew-Normal [CLPC08]. Skewed [ZG00]. Skewness [AG95, Gro91, JRP11, SB11]. Skewness-Invariant [JRP11]. Skills [Cot06, HHJL93]. Sleuth [Led03]. Sleuthing [EA11]. Slope [Nie11]. Slow [TKM04]. Slowly [Gun93]. Sluggers [Ann06a]. Small [FMA^{+}09, Lon10, LDM10, MFS^{+}03, Par04, War97, Ber81]. Small-Sample [FMA^{+}09, War97, Ber81]. SMD [RKW98]. Smoothed [SS08]. Smoothing [Eub94, Sol00, MS87]. So-Called [Fre06]. Social [Bon98, Gon03, Lee08, Nag98, Rak06, RH02, Wue98]. Society [Kot10]. Sociology [Gel08]. Softly [Ald00]. Software [Alt02, Ano90d, Ano91e, Ano91f, Ano93b, Ano95h, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano98h, Ano98i, Cal03, CL02, Coh97, Col01, DS98, HEJ11, HG97, HDE^{+}03, HKS06, Hil93, Hil96a, Hil02c, Hil02d, Hil04f, Hil07g, HL99, HL01, JP01, KP11, Kel04, KT94, LDM10, Mad98, McC98, McC99, Mor98, NQ96, Nor95, Ost98, Ost02, Ost03, OH08a, OH08b, SS98, Spi90, SMP97, WG11, ZPH99, ZNY11, Zim00, Tuk85a, Ano00b]. Sole [AR95]. Solid [MZ92]. Solution [Hu92, Nun92, Sav90]. Solutions [Bel07, CL02]. Solving [Cle08b, Hil96b, EW85]. Some [Alb97a, Ber81, Ber97a, BS87, Dan92, DS08, DHW11, Dun99, Fre97, GK99, Gel98, Gle86, God91, HL98, Hah95, HV95, JK90, Len83, Len01d, LJ09, Lin97, Mad98, Moo97a, Moo97b, NK96, Nay94, PR94, RS91b, Sch97, SSH81, Sho97, Sim94, Sta04, Ter05, VM03, Wit97, YA09, Ans85, BL87, CJT87, VY85]. Sometimes [Cor00, Hal10, Ham87, Sch84]. Sort [Ver99]. Sources [HB05a]. Space [MV07, PO07, EG87, MZ92]. Spaced [GL00]. Spaces [LaM08]. Sparse [ZRK^{+}11]. Spatial [HBC^{+}91, RHCR06]. Spatially [HR10]. Spatially-Correlated [HR10]. SPC [BA02]. Spearman [RS91b]. Special [AG04, Alt02, Dob04, Utt03a]. Specification [Mat99, Hei81]. Specified [Lee93, MS02]. Speed [LC96]. Spherical [TO93]. Spline [Eub94, Sol00, TH00]. Splines [BI09, PW06]. Split [AH02, Cor00, FM92, Pad08]. Split-Block [FM92]. Split-Plot [AH02, Cor00, FM92]. Splitting [GP09, PB90]. SPlus [Cow04b]. Spontaneous [DuM99]. Sports [Bas97, Coc05a, Fel08, Mos97]. Spread [Car96, SB11]. Spreadsheet [KP11]. Spreadsheets [Nas06]. SPSS [Dav07, Dav08b, Han05, Hil03c, Hil04b, Hil05b, Mac93, MJ04, Str02, vH04]. SQL [Ann05c]. Square [Abb02, Che11, DK07, Dri99, Fos11, WN07, Mur86]. Square-Root [Fos11]. Squared [Bai92, BK92, GAR92, Lev97, Mir01, Nay02, Shi01]. Squares [BK92, BC92, Dun94, HH99, HS92, Lee09, Lic90, ML92, MS98, Oks91, Rie90, Saw92, SBL97, SR97, Tar00, Van96b, deL06, BvE88, BM88b, MA84, PS89, SSM80, SP85, Wom85]. SSR [Sch93a]. St. [Bel05, Rum07, SR04, SR05]. Stability [Sto98]. Stabilizing [KMS10]. Stable [LR01]. stage [WC87]. stagewise [All87]. Standard [Fre06, Gad01, GC04, LE00, O’C90, Wal98, vDK02, vS96, Spi86, Spi87]. Standardize [Bri94]. Standardized [GW94, Men04]. Stanford [Sti99]. Stanley [DM06]. Stat [Dev01]. Stata [Gol99, Hor06, Hil05c, Lac00, McG05]. StatConcepts [Pag00]. State [ALTY00, CF08, FDS05, HR05, Mon04, PO07, SS04]. State-Level [CF08]. State-Space [PO07]. statements [AS84]. States [Gon94]. Stationary [CC01, Per93]. Statist [Ang90, CH86, DN94a, GY90, Kar86, Kar87, SYC88, Spi87, WS94, Zel85]. Statistic [BSTM09, FBG96, Lev97, MPR07, OH95, Paw00, PS98b, PS98c, SC97, Tar00, Har80, Mee87, Som86]. STATISTICA [Hil07f]. Statistical [Abb02, Alt02, AC02, ACA03, AC06a, AC08b, And99, Ano90c, Ano90d, Ano91c, Ano91d, Ano91e, Ano91f, Ano93b, Ano95h, Ano95i, Ano95j, Ano96h, Ano96i, Ano96j, Ano96k, Ano97h, Ano98h, Ano98i, BRRB01, BB04, Bel01, BUK^{+}08, BWM98, Bol03, Bon98, Bra96a, Bro11, Bro01, BGNP01, Cal03, CHS04, Car00, CC01, CRM01, CK01, Cle99a, Cle99b, Cle03a, Cle05a, Cle05b, Cle08a, Cle08b, CW99, Cra05, Dal90, Dan92, DLM01, Dla05, DHW11, DS02, Eld90, Eme96, EA11, ES93, Eng93, Epp93, FF02, Fay02, Fel08, FS01, FGHK83, FVMU10, Gal03, Gas92, Gel04, GNM^{+}98, Gen04, GEL01, Gon94, GKT98, Gue87, GB06, GM96, Gup02, HBF90, HM93, Hal10, HMR^{+}01, HS04b, Ham97, HEJ11, HS91, Hay99, HM01, Hil02c, Hil02d, Hil02e, Hil04f, Hil07g]. Statistical [Hog79, Hog91, HL01, HB03, Hun05, Hyd01, HF96, IJW06, Imr94b, Inm94, Jam00, Jam02, JW04, JSM95, Kas98, KP11, KM94, Kha04, Kir91, KBH09, Kre01, KC09, Lac00, Led03, Len01d, Len02, LR01, Lis08, Loh95, Loh04, Los04a, LGXB11, Lun02b, Lun07b, Lur95, MHT90, Mam06, Man03, MMCsS96, Mar07, MMR90, MSL^{+}08, MS01, McC01, McC98, McC99, MS02, Mon04, Mon97, Mor07, Mur97, Nag98, Nas92, Nas06, OCE93, Ost98, Ost02, Ost03, OH08a, OH08b, Pag00, Pau92, Peñ97, Per98, Pri11, Pri05b, Ran01, RKW98, Rid09, Rie06, Rub04b, Rue04, Sam91, SA00, Sch00, Sch10a, SG01, SMC^{+}07, Sne93, SS09, SW05, Spi90, SJD10, SMP97, SRW95, SB98, SGCS04, Tho95, Upt01, Utt02, Utt03a, VWB^{+}10, WB04, Wal05, Wan02, WW11, WM96, WOR98, Wie06, Wie07]. Statistical [Wil06b, Wil92, WSF02, Wio90, Wüt08, Yu94, YCF06, Zam08, ZH90, Zim00, Hel88, Tuk85a, Cow04b]. Statistical-Power-Based [MSL^{+}08]. Statistically [GS06, SO94]. Statistician [AR95, Ano98b, Bil98, Hah02, Hoe98, Hun98, LDL^{+}94, Pos90, Shu91, WKB^{+}95, Boa94, Alb03c, Alb04, Alb05a, Alb05b]. Statisticians [HHJL93, LRG90, RSH01, SG98, Sow03, TS03, VM03, Bar90]. Statistics [Abb02, Ahm96, Alb93, AG04, AL07, AR08, AC04a, AC05b, AC07c, AC08a, AS99, Ann05a, Ann06a, Ann06b, Ano98j, Ano00b, Ano04h, Arn07, BR97, Bar98, BN98, Bat04, BG08, Ber06, BLS00, Ber97b, BM02b, Bil98, Bil04, Bis91, Bis98, BK93, BHM01, Bon98, BO92, Bra07, Bra96a, Bra08b, Bro08, BK09a, BK09b, CF01, Car08, Cav07b, Cav07a, CR01, Cha08, CF07, Chi08, CHJ08, Cle99a, Cle00a, Cle03b, Cle05c, Cle06a, Cle06c, Cle08a, Cle08b, Coc05b, CF91, Coh02, Col04, CB95b, Cow06, Cra05, Czi98, DNG98, DS08, Dav93, Dav95, Dav98a, Dav06, Dav11, Dav08c, Dev01, Dev06a, Dev06b, Dev06c, Dev07a, Dev07b, Dev08, Dob02, Dob03b, Dob04, Eas10, Edw07, Ehr90, FP81, FR04, For05, Fra00, FHP^{+}11, FDS05, GS08, Gal03, GZK^{+}11]. Statistics [GH91, Gel98, Gel05, Gen04, Gib90, GR09, Gol98, Gol99, Gon03, Hah95, HHHZ99, Ham00, Han98, Har06, HR05, Har04, HF04, HK92, HS09, Hel99, Hig99b, Hil96b, HGA01, HS10, Hog94, Hog99a, HR00, Hol07, HBQ04, Igl98, Jam01a, Joh09, Kaf98, KK99, Kaf99, Kah06, Kha08, Kie06, Kir00, KG90, Kot10, Kva00, LaM94, Lan06a, Lan04, LJLC05, Lar92, Law96, LRF11, Lea04, Led95, Leh95, Len02, LP00, LF00, Loc93, LW98, LH02, LG00, Mäc07, Mac99, Mad98, MG09, Mag96, MGCR10, Mas04, McD99, Mea04, Mil01, Mil03, Mil04a, Mil04b, Mil04c, Moo93, MCGM95, Moo01, Moo05, Moo07, Mor98, ME11, Mos97, Mru07, NS99, NL09, NL10, Nor07, Obr98, OCE93, Owe01, PS98a, Pet99, Pil98, PR11, RS91a, Ree02, RH02, Ric08, RSW99, RT01]. Statistics [Ros00b, Row02, Rum98, Rum99, Rum01, Rum06, Rus98, SP98, SO94, SM97, SS04, Sch91, Sch99, Sha01, SW90, Sin08, Skl07, SG92, Smi98, SB09, Sne98, Sol10, SATV94, ST95, Spu98, Sta08, Ste90, Ste01, Ste04, Sti99, Sti08, Sti00, SZM^{+}07, SR96, SV03, Tal07, TS03, Tra07, Utt03b, UJ08, Var98, VM03, VM96, Ver01, Wal02, Wan02, Wat91, Wei07, Wel98, Wil94, WC97, Wue98, You05, deL06, BM86, BR78, Bur86, FK86, FF80, GY89, GY90, KD87, Len83, LH72, Sch81, S^{+}90, Tuk49b, Tuk49c, Tuk65, Can98, Low98, SP98, Sho98, Can01, Cle00b, Min07, Pie01, RWM^{+}72, Wal00b, Maa99, Rob03]. Statistix [Spi90]. Stats [McD04]. Stature [WWD03]. Status [Eld90]. StatXact [Cal03]. Steadfast [SR05]. Stem [SMC^{+}07]. Stem-and-Leaf [SMC^{+}07]. Step [CRM01, Law03, Rid09, Rog84]. Stepwise [Coh91, Mil96, OW91]. Stevens [Bol86]. Sting [Shu91]. Stirling [Wal95b]. Stochastic [Dev06b, Lun02a, Lun04b, Lun04a, RZR05, Sch04e]. Stock [AC05b, Bel05, Ley96]. Stocks [SR04]. Stopping [Bet00, BP02, CB95a]. Story [Jam99]. Straight [Wri92]. Straight-Line [Wri92]. Straightforward [BGG98]. Strategic [SMC^{+}07]. Strategies [BS10a, CW93, GM96]. Strategy [Loo97]. Stratification [Kaz95, Zha00]. Stratified [MLL10]. stratum [Hid86]. Streakiness [AW01]. Streets [SP94]. Strength [Car96, FT10, Per98]. Strengths [MGLK99]. Strikeouts [Ann06a]. Strong [Bro07, Nel98a, Tom84]. Stronger [Dav98b]. Structural [JHFK04, Har76]. Structure [DGC97, HCW07, Van97, Zim00]. Structured [SVR11]. Student [Alb03a, BUK^{+}08, GZK^{+}11, GNM^{+}98, Kva00, Loc93, Mag96, RT01, Smi11, Ste90, Tan99, Ber81, HJM08]. Studentized [GW94]. Student’s [Eis79, AR08, Ano98j, BN98, BV11, Bis98, BM10, BK93, Bol03, Bro08, Czi98, GS08, Gel08, HR05, Hog94, KYA^{+}07, Mon06, Moo05, Mur97, Spu98, Var98]. Studies [LKNP95, LE93, PHSN08, Tim04, MS87]. Studio [Pra08b]. Study [BA02, Cor00, DWBQ98, EA11, HEJ11, Kor90, Lis08, MB06, Shu01, Ste90, Wie99, WASB10, YT01, ZRK^{+}11]. Style [GGR05]. Styles [Moo93, PH02]. subclass [SSH81]. Subjects [Cot98]. Subsampling [MB94]. Subsets [MG97]. Succeed [BV11]. Success [JW04, McD99]. Successes [AC00a, Ang92]. Successful [BM02a]. Sufficiency [KS05, LaM08, Mee86, PR94]. Sufficient [CBH09, McL99, Mee87, MA84]. Suggestion [DB90, SW94]. Suggestions [Bir95, LJ09]. Sukhatme [Ful80]. Sum [Hag09, SGC90, Shi01, Tar00, CS83, Riv81]. Summarize [SMC^{+}07]. Summary [Lar92, OCE93, KD87]. Sums [MS98, Saw92, SBL97, SR97, Ter05, Flu86, SP85]. Sunflower [SW94]. Sunset [Tuk86]. Superior [LD98]. Supplement [CF01, Sch04f]. Support [HR98, Ter05, GD89]. Supporting [Luc97]. Suppression [FW05, Lyn03, SR97]. Surface [CC01, MKV92, Nel98a, Van96b, CCW^{+}86, MH76]. Surgery [ZRK^{+}11]. Surprising [FM08, SS05, NW86, VY85]. Survey [BHM01, CLM92, Coh97, FKL^{+}96, HG97, Hog92, KG90, KG95, KG98, Kot91, Loh04, NOMLT95, WS93, Hid86, Pie80, WS94, CJ92]. Surveys [Dem06, HH04, Kha08, Rub04a, TTT09]. Survival [Atk95, Bar09, BCJ97, Fre08, GCGG95, Gol92, HG97, Kel04, LW90, LSE00, McG05, WASB10]. Survive [AS94, HHJL93]. Suspension [Per98]. Swan [Lun07a, WH07]. Swans [Bro07, Tal07]. SWEEP [Goo79]. Swivelling [BD11]. Symbolic [Ham97]. Symmetric [Pal08, SC97, CS83]. Symmetrized [BA06]. Symmetry [vS96]. Synergism [HN98]. Synthesis [Zha00]. Systat [Hil05d, Hil08f]. System [CY99, DuM99, HHvdWZ06, Kas98, Pra08b, Rub04a]. Systems [AW06, Bag95, BWM98, Col01, Hil08a, Nas95, AAF72, Tuk85a].

T [MS06]. Table

[Len02, MH93, SMC^{+}07, Gol86]. Tables [BM94, Che11, Dav05, DuM99, GK99, GPD02, Gre91, Gre10, Hin96, Kos05, Lin92, MW92, Mir01, Mon97, Nie05, ZCM95, Lau65, Sha82]. TABULATE [Mon97]. Tabulation [MSL^{+}08]. Taguchi [MKV92]. Tail [Jor94, LB00, PN95]. Tails [KS08]. Take [MD96]. Tale [AH02]. Tales [Gun98, Wel98]. Tape [Pos90]. Teach [Bon02, CR01, Gun93, HK92, Koo93, MMCsS96, Moo05, Sim94]. Teach/Me [Bon02]. Teacher [Ano98j, Bis98, BHM01, CF01, CR01, Czi98, HJP^{+}01, HM01, HGA01, IM01, RT01, Spu98, Var98, Zho01]. Teachers [FDS05, HK92]. Teaching [Alb93, Alb97b, AG04, AR08, Ano91a, Ano91b, Ano92, Ano93a, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano97d, Ano97e, Ano97f, Ano97g, Ano98f, Ano02e, Ano03a, Ano04b, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano09g, Ano09h, Ano09i, Ano09j, Ano10f, Ano10g, Ano10h, Ano11i, Ano11f, Ano11g, Ano11h, BG08, Ber97b, Bir95, Bis91, Bra96a, BH08, CLM92, Cle01, Cle05c, Cle06a, Col01, CRV95, Dal90, Dev01, DL04, DL05, DAC05, DAC06, Eas04, FKL^{+}96, Gal03, Gel05, Gel08, GPA97, Hai01, HR05, Hil01, Hil96b, Hun05, Hyd01, Hyu01, IM01, Jam01b, Jam01a, Kle95, Kre01, KB07, Lea04, Lee07, Len01b, Len01c, Len02, LP00, LJ09, LW98, Loh95, Loo97, Mac01, Mad98, ME95, Mil01, Mil04a, MJ04, Mon04, Moo93, NS99]. Teaching [Oks91, Pou02, Ree02, RACK07, RD05, Rum01, SM97, Sha01, SW90, Sol10, Spi90, SM09, Sto98, Str02, SGCS04, SV03, UJ08, VM96, War02, AAF72, AT86, MH76]. Team [Car96, HS94]. Teams [FT10, Gro90, Mea03]. Tears [SG92]. Tech [SR04, BM05]. technical [Tuk65]. Technique [Bor97, DD06, KAB^{+}97, SI82]. Techniques [Cle08a, DGC97, Kha91, KC09, Nas92, SM97, Str90]. Tell [LF06]. Ten [Bir95]. Term [HR00]. Term-Long [HR00]. Terms [Dav95, Dav98a, Nel98a]. Ternary [Upt01]. Test [AP10, BM01, BRRB01, BLS00, Bro10, Cal03, Che11, DG97, Dor96, FKLL03, Fre07a, FG00, GJLW05, HRS10, MS06, MM90, MY05, MC91, MLL10, MPR07, ME11, ML92, MS92, Ran01, RZR05, Ros03, Sho03, Sil97, St.90, Upt00, VM07, AO89, Dav80, Har80, Har76, PS81]. Testing [All97, BRRB01, BW08, BH08, CRM01, Chr05, DB97, HB03, KG90, KT02, LC09, Lin93, Loo97, Ric10, Ros00a, SBB01, SK93, SVR11, MS87, Nai87]. Tests [BHA02, Ber96a, Bir90, Boo92, Chr03, Cor00, DB90, Dav08a, Ede90b, FF02, FS01, Gre91, GR96, Inm94, Kuk09, LW90, LC09, Loo95, LD98, Mag90, MH93, MV07, Sab02, SG91, Sho95, SRW95, Suk92, Vic08, WB91, WW97, ZCM95, ZPM^{+}11, GS86, Hel88, SM79, SS84]. Texas [NO98]. Text [DHN^{+}05]. Text-Mining [DHN^{+}05]. Textbook [Dav08c, Sch04f]. Their [BHA02, Bor99, Dal90, Epp93, JW04, Lun04a, ML09, PE05, San06, SGK00, SBL97, Sow03]. Them [AC04a, ZL09]. Theme [DNG98]. Themes [Dob06a]. Theodore [FK87]. Theorem [Bel01, BHO98, Gho02, Hwa09, Jes94, Jol95, Jon02, OS95, Ros07, Sha91a, Cox84, DK95, NA85, Pau81, PG82, Yan88, Zel88, Zer78, Inl10]. theorems [Ans85]. Theoretic [Ric10, TC88]. Theory [Coh02, Dun99, Gel98, Git94, Hun05, Lan04, LG00, Lun04b, NS99, Oks91, Sim94, SB98, Zim07, TK76]. Think [Cle00a]. Thinking [AC06a, Bol03, Dob06a, Fel08, Men10, Ros00b]. Third [GP09]. Thompson [OS95]. Thoughts [BM02b, Kot10, Sta04]. Three [Dun99, FF80, HEJ11, HKS06, HLW09, Hu08, JL97, KB07, LCD10, Per93, Puk94, Mon79]. Three-Condition [Hu08]. three-factor [Mon79]. Three-Parameter [HEJ11]. Three-Sided [KB07]. Throughout [NOMLT95]. Throw [Ben93]. Thumb [Cle03a]. TI [NO98]. TI-83 [NO98]. Tiered [Bor99, WB02]. Ties [FKLL03, Ran01]. Tighter [PN95]. Time [AAF72, Ann06a, CHS99, Czi99, Dan94, GC04, Hom97, KT94, PT91, Pou02, SC99, SJI05, WASB10, Zho01]. Time-Change [Zho01]. Time-Dependent [WASB10]. Time-Series [CHS99]. Time-sharing [AAF72]. Time-Varying [SJI05]. Timed [Gol92]. Timed-Events [Gol92]. Times [Cry91, SB98]. TIMESLAB [Cry91]. Timing [LKNP95]. Titan [Cle05a, Cle05b]. Toast [Sti08]. today [Hog79]. Together [Dob03b]. Tolerance [BK96, RC88]. Tool [CC01, Haz03, Hol99, Kha91, AAF72, EW85]. Toolbox [HBQ04]. Tools [Dun99, IM01, LG05, WM96, WOR98, Tuk65]. Toothed [CL02]. Topics [CR01, Dob02, MGCR10]. Toroidal [Upt01]. Total [Git94]. Tour [Pag00]. Tournament [Car96, SS99]. Tournaments [Edw96, SMH91, SSH96]. Toxicologists [Pie01]. Toys [HS03]. Trace [HN98]. Track [NK96]. Traditional [DS02]. Traffic [Hu92]. Training [BM05, DAF^{+}94, FDS05, GS08, HR05, Men10, RSH01]. Transformation [FLP07, New01]. Transformations [MPB97, SG91, SA00, SW91, Tar07, RC88]. Transforming [Atk94, AT99]. transition [Eis79]. Transitive [LSO01]. Transmuting [Han04]. Trauma [BCJ97]. Treatment [BA06, YT01]. Treatments [Faw90, GCGG95, Han92a]. Tree [Atk94]. Trees [Bor02]. Trends [Len01c]. Trial [YT01]. Trials [CRM01, San06, VBB94, WKB^{+}95, Ber87, Gei84, NW86, Zel85].

Triangles [TO93]. Tricks [Cle05c]. Trigonometric [Jon99]. Trivariate [Hut93]. Trout [Pic06]. True [Pol99]. Truncated [BS99, Mul88]. Truncation [Fer90, Man07]. Trust [Ros00b]. Truth [Sch93c]. Tukey [HR06]. Tunable [JB04]. Turning [GPD02, Spi07]. Tutorial [Goo79, HL04]. Twice [LC99]. Twins [BSTM09]. TWiST [GCGG95]. Two [Abb02, AC00a, BM01, CRM01, DS08, DHN^{+}05, DGC97, DS98, Ede90b, GJLW05, Gre91, Gro90, Hal10, Han92a, Hos92, KT02, KMS10, LW90, LC09, LC99, LBTM03, LL97, MS06, Mat99, MY05, MV94, Mon79, MS92, OW91, PE05, PS94, RS93, RZR05, Shu06, SS98, Spi86, Spi87, SF93, Suk92, WB04, WBT01, WS93, WASB10, vDK02, BvE88, Bar76, FK86, Gue87, Knü94, WC87, WS94]. Two-by-Two [Gre91]. Two-dimensional [Mon79]. Two-Group [DS98, SS98]. Two-Level [LL97]. Two-Period [PS94]. Two-Player [Shu06]. Two-Sample [BM01, GJLW05, MS06, Mat99, MY05, MS92, Suk92, Knü94]. Two-Sided [Ede90b, vDK02]. two-stage [WC87]. Two-Step [CRM01]. Type [Ahm96, AB03, BSTM09, Kaz95, LDM10, Nay02, Pav88, Pol99, Sch96b, VD05, dM07, BS95]. Type-II [BS95]. Typologies [VW93].

U.S [Gib90]. U.S. [CHS04, Gib90, Ley96]. ubiquitous [JH83]. UCLA [Lan04]. Ultimate [Won98]. Unbalanced [GL00, Sea94, Her86]. Unbiased [Ang92, BR97, DB97, Eli00, McI05, PR94, Sea97, SM91, BvE88, BM88a, GD89, HTM89, Kar85, Kar86, Kar87, Kot86, MA84, PS89, SC94, SS84, VM89]. Unbiasedness [Lin93, MS95]. Uncertainty [Ann05a, Jac08, Lon10]. Uncertified [WJM^{+}03]. Unconditional [MH93, Kap79]. Uncorrelated [Beh90]. Uncovering [Old95]. Undergraduate [GH91, Hig99b, Kot10, LRF11, LRZG^{+}10, Moo01, SS04, Wil06b]. Underlying [SF00]. Understand [Bel05, WGHE09]. Understanding [Abb02, Bro08, Can98, CG95b, CHS99, Cle99a, DJ05, Han05, Low98, Lun02b, Mil04c, RT01, Sch06b, Skl07, Sul96, WM96, Wue98, Zho01, dML09, Mil04b]. Undesirable [HR98]. Undiscovered [WWD03]. unequal [SSH81]. unequal-subclass-numbers [SSH81]. Unequally [GL00]. Unexpected [PP98]. Unified [Lor90, MSS91]. Uniform [Rod04, EG87, GM86]. Uniformly [Bir90, Ede90b, GD89, NVB09, SS84]. Unifying [ML09, OS95]. Unimodal [Ede90a, SS97]. Unique [Wil08]. Unit [AW91b]. United [Gon94]. Units [NOMLT95, Aso80, Hid86]. Univariate [Dav08b, LM08, Loo95, TM95, Lee86, TC88]. Universal [Sam91]. Universities [CF91, LW98]. University [Gel05, GDL08, Kir91, CRV95, FDS05, HR05, KM94, Mon04]. Unknown [Ede90a, MS06, Per93]. Unlocking [Cra08]. Unordered [EGS96]. Unquestionably [GKR97b]. Unrecognized [DHW11]. Unrelated [BF91]. Unusual [JS11]. Unweighted [KG95, SM79]. Updated [Dav07]. Updating [EM93]. Upon [GAR92]. Upper [DH11, GP09, Pol99, Wri90]. Urban [GKT98]. Urgent [Cle06a]. urn [Pau81]. Use [BRRB01, BB04, CHS04, CR96, Dav80, FKLL03, FG00, HL98, Hog99a, HBQ04, KG90, LP00, Loo95, Man82, Mor98, Pan02, San06, SM97, Sch04f, Sim94, Ste90, SZM^{+}07, Str02, Bar76, Gue79, Her80, Hog79, Kap79, Rav85, RC88]. Used [KC09, LC99, SS99]. Useful [Dav11, DGC97, DB97]. Usefulness [Luc97]. User [Bon02, Bor97, CG06]. User-Friendly [CG06]. Users [Spe08]. Uses [Cle06b, Mad98]. Using [Alb93, AW01, Ald00, All97, AC00b, Ann05c, AT99, Beh90, Bon98, Bor02, BSY92, Bra08b, Bro01, BCB00, CLM92, CG95a, CL02, CF08, Cle99a, Cle08b, CB95b, DB90, DD06, DT00, Gan95, GCGG95, GE11, GKT98, GC04, Han05, HB05a, HRS10, HR00, Hor06, Hwa09, Kas98, Kha92, Koo93, Kva96, Lac00, LG05, LC09, Lea04, LF00, LE00, Lun02b, Lun07b, MMCsS96, McG05, McI05, MS98, Ost02, Ost03, OH08a, OH08b, Pau92, Pra08b, RT01, Sch10a, SK93, Shu06, Smi98, SF00, SM09, SB98, SGCS04, TO93, Tim04, VE06, WW11, Wie06, Wie08, YHZ08, CH85, CH86, Cle01, Cow04b, Gup02, KD87, Mäc07, Man03, Mee87, Rus08, SP85, Shi88a, SR83, Spi86, Spi87]. Usual [WL99, Bly80]. Utility [Hil07a, KKO^{+}06]. Utilizes [SI90].

V [Ful80, Tan99]. vadis [Fox10]. Valid [WE08]. Validation [MCM^{+}11, EG83]. validity [Hil87]. Valuation [Bel05]. Value [BH02, BS11, Bro10, Don99, Rus91, SB98, vH04, EW85, Man82, MC87]. Values [Ber96a, Bly93, Imr94b, KP07, MTA08, PT91, SSC99, Sch96a, SBB01, Har80]. Vanzetti [She98]. VARCL [KDV94]. Variable [GR96, Grö09, HT06, OW91, Wil08]. Variables [AT99, BRRB01, Beh90, Bri96, Cai94, CR96, Coh91, HL01, JB04, Lic90, MTA08, Nav97a, Nel01, Rak06, SSC99, Ter05, Upt01, ZRK^{+}11, vS96, Bar76, BG89, CS83, Flu86, Ham84, Ham87, KP79, PK82, Roc88]. Variance [Ang92, AS98, BS99, CM92, Cor00, DK07, FN92, Grö07, HB05a, JB04, Kha92, KMS10, Lar92, Lin93, LDM10, MS06, MM90, MS92, Nay94, Nel07, Per93, Sch91, Sea94, SI90, SK93, Sti95, VM90, Zha07a, CGL83, GD89, Kar85, Kar86, Kar87, Mon79, Mul88, SYM84, SYC88, Sch85, SSH81, Shu82, VM89]. Variances [BS98, Sho03, NW86]. Variate [Jon02]. Variates [Bai92, Dav93, EP91, Gan95, Hag09, PPS96, Ren92]. Variation [AH95, HB05a, KS91, SM09, Sto98, Van96a, vM93]. Variational [OW10]. Variations [MTL78, Pad08]. Vary [HS94]. Varying [SJI05]. vec [Hu07]. Vector [Wan02, YW09]. Venn [LL97]. Versa [Smi95, SRW95]. Versatile [LT97]. Version [Hal92, Tan99]. Versions [Dav07, Loc93]. Versus [CL02, Czi99, HV03, HB03, Moe05a, Moe05b, PF95, WSF02, GKT98, Grö09, KR02, MR09, MSL^{+}08, Nai87, Sch10b]. Very [Cle06b]. Via [Don99, Ans85, Car96, DCWP93, FLP07, Gho96, HMR^{+}01, Pal99]. Vice [Smi95, SRW95]. Victory [Mea03, SS99, deL06]. Video [Moo93]. View [Wil94, Hog79, LP76]. Viewing [Zim00]. Viewpoints [MR09]. Views [AR95, FW05]. Violating [SF00]. violations [SD87]. Violators [DM01]. Violin [HN98]. Virginia [BM05]. Virtual [SM09]. Visibility [GT95]. Visual [AC07c, Fri95, Gun98, Pag00, Pic06, dML09, Sti86]. Visual-Minded [dML09]. Visualization [HJS06]. Visualizing [Bra08a, Dun99, FK09, Gol92, HJP^{+}01, KS08, LKNP95, SB11]. vol [Dob02]. Volume [Alb04, Alb05a, Alb05b, Ano97b, Ano99c, Ano00c, Ano09c, Ano10c, Sab02, Ano08e, Ano98d]. Voodoo [Cle03d]. Vote [MSL^{+}08].

W. [Boa94, Sne08]. Waiting [Hom97].

Wald

[FBG96, LDM10, Mag90, MV07, Paw00]. Waldo [FK09]. Walk [Pal08]. Walks [FR04, BH86]. Wanna [CW93]. War [You05]. Warren [FS06]. Way [BM03, BK96, DH11, Kuk09, LDM10, RvE97, SD87, Sch85]. Ways [Mir01]. Weak [Nel98a]. Weak-Heredity [Nel98a]. Web [LG05, SV03, WOR98]. Web-Based [LG05, SV03]. Weibull [HEJ11]. Weighing [Rob03]. Weighted [DM01, JK90, Kha08, KG95, Lee09, ML92, SD01, BM88b, SP86]. Weights [Jor94]. Weldon [KK91]. Welfare [ES93]. Well [LC99, Tur97, Pei90]. well-formulated [Pei90]. Well-Known [LC99, Tur97]. We’ve [VAD90]. Whale [RS93]. Whence [BHA02]. Where [Ano00g, DWBQ98, FK09, HV03]. Which [BRRB01, BF91, JW04]. while [FM84]. Whiskers [DR07]. White [Lun07a]. Whitney [Ahm96, BLS00, Ran01]. Who [Sim98, Sol10, Tib97, Tra07, ZL09, Boa94]. whose [AB92]. Wide [WOR98]. Wider [Wil94]. Width [CRM01, Fre10, Wan97]. Wilcoxon [Ahm96, BLS00, Ran01]. Wilks [DM06]. Will [BV11]. William [Ano91g, Ano91h, Bol84, DM81, SW05]. Williamsburg [Per98]. Wilson [Koe09]. Win [Alb02, Dav98b]. WinBUGS [Cow04c]. Windows [Dav07, Mac93, Mon97, Han05, Har00, Hil96a, Lur95, Str02]. Winning [BD11, Ste91, Kel94]. Wisdom [Sow03]. Wishart [Gho96, GS02]. Within [Cot98, Moe05a, Moe05b, O’C90, Sch10b]. Within-Subjects [Cot98]. Without [Bar99, DG07, HB05a, He97, Hu07, Lan06a, SG92]. Wolf [Rie90]. Wolfowitz [Wei82]. Women [Han04, S^{+}90]. Words [GT95]. Work [CW93, LRG90, Pos90]. Workbench [Los04b]. Workbook [Pet99]. Working [Cle06c, HCW07]. WorkPlace [HH06]. Workshop [Bar98, Cle99a, Law96, Ree02, Ric08]. World [Ben93, Cha08, DLM01, Kah06, Mil04c, Tib97, Vid94, Boa94, WOR98]. Worth [FS01]. Would [Kon95]. Would-Be [Kon95]. Wrestler [Gel06]. Writing [AC05a, AC06b, Dob06b, Mon06, RS91a, SO94, SR96]. Wrong [VM07].

Xiao [vC10]. Xiao-Li [vC10]. XploRe [Hil01]. XPro [Hil02e].

Yarnold [Eat78]. Yates [NL95, Upt00]. Year [BHM01]. Years [Men09, Stu09, vC10, Her86, Hoe85, Noe84]. yield [DBS^{+}85]. Youden [BK07]. Young [VM03].

Z [AC03]. Zellner [Gei84]. Zero [Dav09, Jam01b, WSF02, Wri90]. Zero-Numerator [WSF02]. Zeros [Ran01]. Zeroth [Mal98].

References

Abranovic:1972:TSC

[AAF72] W. Abranovic, R. Ageloff, and D. Frederick. Time-sharing computer systems as a teaching tool. The American Statistician, 26(1):34–38, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnold:1992:DWC

[AB92] Barry C. Arnold and Patrick L. Brockett. On distributions whose component ratios are Cauchy. The American Statistician, 46(1):25–26, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Austin:2003:TEI

[AB03] Peter C. Austin and Lawrence J. Brunner. Type I error inflation in the presence of a ceiling effect. The American Statistician, 57(2):97–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                                    tronic). URL http://oberon.
                                                    ingentaselect.com/cgi-bin/
                                                    linker?ini=asa&reqidx=/cw/
                                                    asa/00031305/v57n2/s8/p97.

Abbott:2002:BRS

[Abb02] Robert D. Abbott. Book review: Statistics at Square Two:

Understanding Modern Statistical Applications in Medicine. The American Statistician, 56 (4):331–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s16/p331.

Aitchison:1981:BRR

[ABS81] John Aitchison and John Bacon-Shone. Bayesian risk ratio analysis. The American Statistician, 35(4):254–257, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Agresti:1998:ABT

[AC98a] Alan Agresti and Brent A.

Coull. Approximate is better than “exact” for interval estimation of binomial proportions. The American Statistician, 52 (2):119–126, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/agresti.
html.

Anderson-Cook:1998:DFE

[AC98b] C. M. Anderson-Cook. Designing a first experiment: A project for design of experiment courses. The American Statistician, 52(4):338–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts_98/anderson-
                                        cook.html.

Agresti:2000:SEC

[AC00a] Alan Agresti and Brian Caffo. Simple and effective confidence intervals for proportions and differences of proportions result from adding two successes and two failures. The American Statistician, 54(4):280–??, November 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Alvo:2000:CHP

[AC00b] M. Alvo and P. Cabilio. Calculation of hypergeometric probabilities using Chebyshev polynomials. The American Statistician, 54(2):141–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Alvo.htm.

Anderson-Cook:2002:BRS

[AC02] C. M. Anderson-Cook. Book review: Statistical Consulting. The American Statistician, 56 (4):329–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n4/s13/p329.

Anderson-Cook:2003:BRZ

[AC03] Christine M. Anderson-Cook. Book review: The A to Z of Mathematics: A Basic Guide. The American Statistician, 57 (3):218–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s16/p218.

Anderson-Cook:2004:CES

[AC04a] Christine M. Anderson-Cook.

Common errors in statistics (and how to avoid them). The American Statistician, 58(4): 359, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2004:RA

[AC04b] Christine M. Anderson-Cook.

Regression and ANOVA. The American Statistician, 58(2): 172–173, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2005:CGW

[AC05a] Christine M. Anderson-Cook.

The Chicago guide to writing about numbers. The American Statistician, 59(3):274, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2005:MDL

[AC05b] Christine M. Anderson-Cook.

More damned lies and statistics: How numbers confuse public issues and chance: A guide to gambling, love, the stock market, & just about everything else. The American Statistician, 59(3):274–275, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2006:CDS

[AC06a] Christine M. Anderson-Cook. The challenge of developing statistical literacy, reasoning and thinking. The American Statistician, 60(1):99, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2006:CGW

[AC06b] Christine M. Anderson-Cook. The Chicago guide to writing about multivariate analysis. The American Statistician, 60 (2):203–204, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:QED

[AC07a] Christine M. Anderson-Cook. Quality by experimental design. The American Statistician, 61 (2):183–184, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:SSQ

[AC07b] Christine M. Anderson-Cook. Six sigma quality improvement with MINITAB. The American Statistician, 61(2):185, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2007:VSS

[AC07c] Christine M. Anderson-Cook.

Visual statistics: Seeing data with dynamic interactive graphics. The American Statistician, 61(3):278, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2008:MES

[AC08a] Christine M. Anderson-Cook.

Modern engineering statistics. The American Statistician, 62 (3):270–271, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2008:QCS

[AC08b] Christine M. Anderson-Cook.

A quick course in statistical process control. The American Statistician, 62(2):183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Cook:2003:BRS

[ACA03] Christine M. Anderson-Cook and Marcus M. Alley. Book review: Statistical Methods in Agricultural and Experimental Biology (3rd ed.). The American Statistician, 57(4): 317–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s18/p317.

Afshartous:2008:SSD

[Afs08]  David Afshartous.  Sample

size determination for binomial proportion confidence intervals. The American Statistician, 62 (1):27–31, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Appleton:1996:ICE

[AFV96] David R. Appleton, Joyce M. French, and Mark P. J. Vanderpump. Ignoring a covariate: An example of Simpson’s paradox. The American Statistician, 50(4):340–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnold:1995:MSR

[AG95] Barry C. Arnold and Richard A. Groeneveld. Measuring skewness with respect to the mode. The American Statistician, 49 (1):34–38, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2004:SST

[AG04] James H. Albert and James E. Gentle. Special section: Teaching computational statistics. The American Statistician, 58 (1):1–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s1/p1.

Arnholt:1995:EMK

[AH95] Alan T. Arnholt and Jaimie L.

Hebert. Estimating the mean with known coefficient of variation. The American Statistician, 49(4):367–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aldworth:2002:SPM

[AH02] Jeremy Aldworth and Wherly P.

Hoffman. Split-plot model with covariate: A cautionary tale. The American Statistician, 56 (4):284–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s4/p284.

Ahmad:1996:CMW

[Ahm96] Ibrahim A. Ahmad. A class of Mann–Whitney–Wilcoxon type statistics. The American Statistician, 50(4):324–327, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Agrawal:1988:PED

[AJ88] M. C. Agrawal and Nirmal Jain.

Predictive estimation in double sampling procedures. The American Statistician, 42(3): 184–186, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2010:ERM

[AKLS10] David H. Annis, Peter C. Kiessler, Robert Lund, and Tara L. Steuber. Estimation in reversible Markov chains. The American Statistician, 64 (2):116–120, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Alf:2007:SAI

[AL07] Cherié Alf and Sharon Lohr. Sampling assumptions in introductory statistics classes. The American Statistician, 61(1): 71–77, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:1992:BAP

[Alb92] James Albert. A Bayesian analysis of a Poisson random effects model for home run hitters. The American Statistician, 46(4):246–253, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:1993:TBS

[Alb93] James H. Albert. Teaching Bayesian statistics using sampling methods and MINITAB. The American Statistician, 47 (3):182–??, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:1997:RBB

[Alb97a] Jim Albert. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):271–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:1997:TBR

[Alb97b] Jim Albert. Teaching Bayes’ rule: A data-oriented approach. The American Statistician, 51 (3):247–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/albert.html.

Albert:2002:CBC

[Alb02] James H. Albert. Calculated bets: Computers, gambling, and mathematical modeling to win. The American Statistician, 56(4):329–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s14/p329.

Albert:2003:CSC

[Alb03a] James H. Albert. College students’ conceptions of probability. The American Statistician, 57(1):37–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s6/p37.

Albert:2003:ER

[Alb03b]  James H. Albert.  Edi-

tor’s reports. The American Statistician, 57(3):153– ??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n3/s1/p153.

Albert:2003:NCA

[Alb03c] James H. Albert. A new column in The American Statistician. The American Statistician, 57(2):94–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n2/s6/p94.

Albert:2004:ERA

[Alb04] James H. Albert. Editor’s report for The American Statistician, volume 57 (2003). The American Statistician, 58(2): 91–92, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2005:ERAa

[Alb05a] James H. Albert. Editor’s report for The American Statistician, volume 58 (2004). The American Statistician, 59(3): 205–206, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2005:ERAb

[Alb05b] James H. Albert. Editor’s report for The American Statistician, volume 59 (2005). The American Statistician, 59(4): 281, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aldrin:2000:MPU

[Ald00] M. Aldrin. Multivariate prediction using softly shrunk reduced-rank regression. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Alley:1987:NSR

[All87] William M. Alley. A note on stagewise regression. The American Statistician, 41(2): 132–134, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Allen:1997:HTU

[All97]  David L. Allen.  Hypoth-

esis testing using an L1distance bootstrap. The American Statistician, 51(2): 145–150, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/allen.html.

Alonso:2010:NIR

[ALL10] Ariel Alonso, Saskia Litière, and Annouschka Laenen. A note on the indeterminacy of the random-effects distribution in hierarchical models. The American Statistician, 64(4): 318–324, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Altman:1992:IKN

[Alt92] N. S. Altman. An introduction to kernel and nearest-neighbor nonparametric regression. The American Statistician, 46(3): 175–185, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Altman:2002:SCG

[Alt02] Micah Altman. Statistical computing and graphics — statistical computing software reviews — A review of JMP 4.03 with special attention to its numerical accuracy. The American Statistician, 56(1):72–75, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Almond:2000:DCG

[ALTY00] Russell G. Almond, Charles Lewis, John W. Tukey, and Duanli Yan. Displays for comparing a given state to many others. The American Statistician, 54(2):89–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Almond.htm.

Amaratunga:1999:SRS

[Ama99] Dhammika Amaratunga. Searching for the right sample size. The American Statistician, 53 (1):52–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anderson-Sprecher:1994:MCR

[And94] Richard Anderson-Sprecher.

Model comparisons and R^{2}. The American Statistician, 48 (2):113–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Andrews:1999:SRM

[And99] Douglas M. Andrews. Statistical reasoning and methods. The American Statistician, 53 (3):291–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Andersson:2009:SCA

[And09] Per Gösta Andersson. A simple correlation adjustment procedure applied to confidence interval construction. The American Statistician, 63(3):258–262, August 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Angers:1989:NQS

[Ang89] Claude Angers. Note on quick simultaneous confidence intervals for multinomial proportions. The American Statistician, 43(2):91, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [Ang90].

Angers:1990:CNQ

[Ang90] Claude Angers. Correction: “Note on quick simultaneous confidence intervals for multinomial proportions” [Amer. Statist. 43 (1989), no. 2, 91]. The American Statistician, 44 (1):65, ???? 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Ang89].

Angus:1992:MVU

[Ang92] John E. Angus. Minimum variance unbiased estimator of the probability of an odd number of successes. The American Statistician, 46(2):115–117, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ankenman:1998:BRI

[Ank98] Bruce E. Ankenman. Book review: Introduction to the Design and Analysis of Experiments. The American Statistician, 52(4):368–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:PSS

[Ann05a] Dave H. Annis. Probability and statistics: The science of uncertainty. The American Statistician, 59(3):276–277, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:RPH

[Ann05b] David H. Annis. Rethinking the paper helicopter. The American Statistician, 59(4):320–326, November 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2005:PSB

[Ann05c] Izabela E. Annis. PROC SQL:

Beyond the basics using SAS. The American Statistician, 59 (4):350, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2006:BAT

[Ann06a] David H. Annis. Baseball’s all-time best hitters: How statistics can level the playing field and baseball’s all-time best sluggers: Adjusted batting performance from strikeouts to home runs. The American Statistician, 60(1):97, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2006:SMR

[Ann06b] Izabela E. Annis. Statistics in market research. The American Statistician, 60(4):345, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Annis:2007:DDA

[Ann07]  David H. Annis.  Dyadic

data analysis. The American Statistician, 61(4):371, November 2007. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:ASCa

[Ano90a] Anonymous. ASA’s sesquicentennial celebration. The American Statistician, 44(2):113–??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:ASCb

[Ano90b] Anonymous. ASA’s sesquicentennial celebration. The American Statistician, 44(3):200–??, August 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:NDS

[Ano90c] Anonymous. New developments in statistical computing. The American Statistician, 44 (3):242–??, August 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1990:SCS

[Ano90d] Anonymous. Statistical computing software reviews. The American Statistician, 44(3): 236–??, August 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:ATMa

[Ano91a] Anonymous. Accent on teaching materials. The American Statistician, 45(1):64–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:ATMb

[Ano91b] Anonymous. Accent on teaching materials. The American Statistician, 45(2):141–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:NDSa

[Ano91c] Anonymous. New developments in statistical computing. The American Statistician, 45(1): 74–??, February 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:NDSb

[Ano91d] Anonymous. New developments in statistical computing. The American Statistician, 45(2): 154–??, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:SCSa

[Ano91e] Anonymous. Statistical computing software reviews. The American Statistician, 45(1): 68–??, February 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:SCSb

[Ano91f] Anonymous. Statistical computing software reviews. The American Statistician, 45(2): 146–??, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:WRSa

[Ano91g] Anonymous. William R. Schucany. The American Statistician, 45(1):1–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1991:WRSb

[Ano91h] Anonymous. William R. Schucany. The American Statistician, 45(4):263–??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1992:ATM

[Ano92] Anonymous. Accent on teaching materials. The American Statistician, 46(2):129–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1993:ATM

[Ano93a] Anonymous. Accent on teaching materials. The American Statistician, 47(4):277–??, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1993:SCS

[Ano93b] Anonymous. Statistical computing software reviews. The American Statistician, 47(4): 297–??, November 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ATMa

[Ano95a] Anonymous. Accent on teaching materials. The American Statistician, 49(1):62–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ATMb

[Ano95b] Anonymous. Accent on teaching materials. The American Statistician, 49(2):219–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ATMc

[Ano95c] Anonymous. Accent on teaching materials. The American Statistician, 49(4):372–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:EC

[Ano95d] Anonymous. Editorial collaborators. The American Statistician, 49(4):404–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ENa

[Ano95e] Anonymous. Editor’s notes.

The American Statistician, 49 (1):84–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ENb

[Ano95f] Anonymous. Editor’s notes.

The American Statistician, 49 (2):235–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:ENc

[Ano95g] Anonymous. Editor’s notes. The American Statistician, 49 (4):395–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:SCSa

[Ano95h] Anonymous. Statistical computing software reviews. The American Statistician, 49(1): 84–??, ???? 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:SCSb

[Ano95i] Anonymous. Statistical computing software reviews. The American Statistician, 49(2): 235–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1995:SCSc

[Ano95j] Anonymous. Statistical computing software reviews. The American Statistician, 49(4): 395–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ATMa

[Ano96a] Anonymous. Accent on teaching materials. The American Statistician, 50(2):168–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ATMb

[Ano96b] Anonymous. Accent on teaching materials. The American Statistician, 50(3):257–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:EC

[Ano96c] Anonymous. Editorial collaborators. The American Statistician, 50(4):390–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENa

[Ano96d] Anonymous. Editor’s notes.

The American Statistician, 50 (1):86–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENb

[Ano96e] Anonymous. Editor’s notes.

The American Statistician, 50 (2):182–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENc

[Ano96f] Anonymous. Editor’s notes.

The American Statistician, 50 (3):266–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:ENd

[Ano96g] Anonymous. Editor’s notes.

The American Statistician, 50 (4):366–??, November 1996.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSa

[Ano96h] Anonymous. Statistical computing software reviews. The American Statistician, 50(1): 86–??, ???? 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSb

[Ano96i] Anonymous. Statistical computing software reviews. The American Statistician, 50(2): 182–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSc

[Ano96j] Anonymous. Statistical computing software reviews. The American Statistician, 50(3): 266–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1996:SCSd

[Ano96k] Anonymous. Statistical computing software reviews. The American Statistician, 50(4): 366–??, November 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:EC

[Ano97a] Anonymous. 1997 editorial collaborators. The American Statistician, 51(4):376–??, November 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:IV

[Ano97b] Anonymous. Index to volume 51. The American Statistician, 51(4):378–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:LE

[Ano97c] Anonymous. Letters to the editor. The American Statistician, 51(2):210–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTa

[Ano97d] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(1): 93–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTb

[Ano97e] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(2): 206–??, May 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTc

[Ano97f] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(3): 293–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:RBTd

[Ano97g] Anonymous. Reviews of books and teaching materials. The American Statistician, 51(4): 374–??, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1997:SCS

[Ano97h] Anonymous. Statistical computing software reviews. The American Statistician, 51(4):??, November 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:EC

[Ano98a] Anonymous. 1998 editorial collaborators. The American Statistician, 52(4):372– ??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/1998%20referees.html.

Anonymous:1998:ASE

[Ano98b] Anonymous. The American Statistician Editor’s report for 1997. The American Statistician, 52(3):288–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:C

[Ano98c] Anonymous. Correction. The American Statistician, 52(4): 369–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:IV

[Ano98d] Anonymous. Index to Volume 52. The American Statistician, 52(4):374–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:LE

[Ano98e] Anonymous. Letters to the editor. The American Statistician, 52(4):370–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:RBT

[Ano98f] Anonymous. Reviews of books and teaching materials. The American Statistician, 52(1): 80–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SEN

[Ano98g] Anonymous. Section Editor’s notes. The American Statistician, 52(3):266–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SCSa

[Ano98h] Anonymous. Statistical computing software reviews. The American Statistician, 52(1): 68–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:SCSb

[Ano98i] Anonymous. Statistical computing software reviews. The American Statistician, 52(1): 70–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1998:TCR

[Ano98j] Anonymous. Teacher’s corner — reply — A onesemester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):243–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:EC

[Ano99a] Anonymous. 1999 editorial collaborators. The American Statistician, 53(4):394–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:ER

[Ano99b] Anonymous. Editor’s report for 1998. The American Statistician, 53(2):175–??, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:IV

[Ano99c] Anonymous. Index of volume 53. The American Statistician, 53(4):396–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:LE

[Ano99d] Anonymous. Letters to the editor. The American Statistician, 53(4):393–??, November 1999.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1999:R

[Ano99e] Anonymous. Reply. The American Statistician, 53(3):201–??, August 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:EC

[Ano00a] Anonymous. 2000 editorial collaborators. The American Statistician, 54(4):329–??, November 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:AMS

[Ano00b] Anonymous. Applied Multivariate Statistics with SAS Software. The American Statistician, 54(2):154–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:IV

[Ano00c] Anonymous. Index of volume 54. The American Statistician, 54(4):331–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEa

[Ano00d] Anonymous. Letters to the editor. The American Statistician, 54(2):156–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEb

[Ano00e] Anonymous. Letters to the editor. The American Statistician, 54(3):227–??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:LEc

[Ano00f] Anonymous. Letters to the editor. The American Statistician, 54(4):325–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:PPP

[Ano00g] Anonymous. The point process paradox: Where should we extend the conversation? The American Statistician, 54 (2):119–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Singpurwalla.htm.

Anonymous:2001:EC

[Ano01a] Anonymous. Editorial collaborators. The American Statistician, 55(4):377, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2001:LE

[Ano01b] Anonymous. Letters to the editor. The American Statistician, 55(1):85–??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2002:C

[Ano02a] Anonymous. Corrections. The American Statistician, 56(3): 258–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s22/p258.

Anonymous:2002:LEa

[Ano02b] Anonymous. Letters to the Editor. The American Statistician, 56(2):159–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s14/p159.

Anonymous:2002:LEb

[Ano02c] Anonymous. Letters to the Editor. The American Statistician, 56(3):252–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s21/p252.

Anonymous:2002:LEc

[Ano02d] Anonymous. Letters to the Editor. The American Statistician, 56(4):337–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s24/p337.

Anonymous:2002:RBT

[Ano02e] Anonymous. Reviews of books and teaching materials. The American Statistician, 56(2): 156–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n2/s13/p156.

Anonymous:2003:BRT

[Ano03a]  Anonymous.  Brief reviews

of teaching materials. The American Statistician, 57(4): 319–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s20/p318.

Anonymous:2003:Ca

[Ano03b] Anonymous. Correction. The American Statistician, 57(2): 151–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n2/s21/p151.

Anonymous:2003:Cb

[Ano03c] Anonymous. Correction. The American Statistician, 57(3): 227–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.

ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s21/p227.

Anonymous:2003:Cc

[Ano03d] Anonymous. Corrections. The American Statistician, 57(4): 322–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s21/p322.

Anonymous:2003:LEa

[Ano03e] Anonymous. Letters to the Editor. The American Statistician, 57(1):71–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s19/p71.

Anonymous:2003:LEb

[Ano03f] Anonymous. Letters to the Editor. The American Statistician, 57(2):147–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s20/p147.

Anonymous:2003:LEc

[Ano03g] Anonymous. Letters to the Editor. The American Statistician, 57(3):221–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n3/s20/p221.

Anonymous:2003:LEd

[Ano03h] Anonymous. Letters to the Editor. The American Statistician, 57(4):319–??, February 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s20/p319.

Anonymous:2003:R

[Ano03i] Anonymous. Rejoinder. The American Statistician, 57(3): 189–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n3/s10/p181.

Anonymous:2004:EC

[Ano04a] Anonymous. 2004 editorial collaborators. The American Statistician, 58(4):365–368, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:BRT

[Ano04b] Anonymous. Brief reviews of teaching materials. The American Statistician, 58(3):262–263, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:Cb

[Ano04c] Anonymous. Correction. The American Statistician, 58(4): 364, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:Ca

[Ano04d] Anonymous. Corrections. The American Statistician, 58(2): 178, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:LEa

[Ano04e] Anonymous. Letters to the Editor. The American Statistician, 58(1):90–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s23/p90.

Anonymous:2004:LEb

[Ano04f] Anonymous. Letters to the Editor. The American Statistician, 58(2):176–177, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:LEc

[Ano04g] Anonymous. Letters to the Editor. The American Statistician, 58(4):362–363, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2004:RSC

[Ano04h] Anonymous. Reply: Statistics and the college football championship. The American Statistician, 58(3):194–195, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:EC

[Ano05a] Anonymous. 2005 editorial collaborators. The American Statistician, 59(4):355–358, November 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:Ca

[Ano05b] Anonymous. Correction. The American Statistician, 59(1): 119, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:Cb

[Ano05c] Anonymous. Correction. The American Statistician, 59(3): 280, August 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:LEa

[Ano05d] Anonymous. Letters to the Editor. The American Statistician, 59(1):116–118, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2005:LEb

[Ano05e] Anonymous. Letters to the Editor. The American Statistician, 59(3):279, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:EC

[Ano06a] Anonymous. 2006 editorial collaborators. The American Statistician, 60(4):350–353, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:C

[Ano06b] Anonymous. Correction. The American Statistician, 60(2): 212, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEa

[Ano06c] Anonymous. Letters to the Editor. The American Statistician, 60(1):101–103, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEb

[Ano06d] Anonymous. Letters to the Editor. The American Statistician, 60(2):208–211, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEc

[Ano06e] Anonymous. Letters to the Editor. The American Statistician, 60(3):295–297, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2006:LEd

[Ano06f] Anonymous. Letters to the Editor. The American Statistician, 60(4):347–349, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:EC

[Ano07a] Anonymous. 2007 editorial collaborators. The American Statistician, 61(4):374–379, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTa

[Ano07b] Anonymous. Brief reviews of teaching materials. The American Statistician, 61(2):186, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTb

[Ano07c] Anonymous. Brief reviews of teaching materials. The American Statistician, 61(3):279, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:BRTc

[Ano07d]  Anonymous.  Brief reviews

of teaching materials. The American Statistician, 61(4): 372, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEa

[Ano07e] Anonymous. Letters to the Editor. The American Statistician, 61(2):187–188, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEb

[Ano07f] Anonymous. Letters to the Editor. The American Statistician, 61(3):280–284, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2007:LEc

[Ano07g] Anonymous. Letters to the Editor. The American Statistician, 61(4):373, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:EC

[Ano08a] Anonymous. 2008 editorial collaborators. The American Statistician, 62(4):366–367, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:BRTa

[Ano08b]  Anonymous.  Brief reviews

of teaching materials. The American Statistician, 62(1): 94, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:BRTb

[Ano08c] Anonymous. Brief reviews of teaching materials. The American Statistician, 62(2):185, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:C

[Ano08d] Anonymous. Corrections. The American Statistician, 62(2): 188, May 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:IV

[Ano08e] Anonymous. Index to volume 62. The American Statistician, 62(4):368–370, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEa

[Ano08f] Anonymous. Letters to the Editor. The American Statistician, 62(1):95–96, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEb

[Ano08g] Anonymous. Letters to the Editor. The American Statistician, 62(2):186–187, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEc

[Ano08h] Anonymous. Letters to the Editor. The American Statistician, 62(3):274–278, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2008:LEd

[Ano08i] Anonymous. Letters to the Editor. The American Statistician, 62(4):364–365, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:EC

[Ano09a] Anonymous. 2009 editorial collaborators. The American Statistician, 63(4):407–408, November 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:E

[Ano09b] Anonymous. Errata. The American Statistician, 63(2): 200, May 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:IV

[Ano09c] Anonymous. Index to volume 63. The American Statistician, 63(4):409–411, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEa

[Ano09d] Anonymous. Letter to the Editor. The American Statistician, 63(1):102–103, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEc

[Ano09e] Anonymous. Letter to the Editor. The American Statistician, 63(4):406, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:LEb

[Ano09f] Anonymous. Letters to the Editor. The American Statistician, 63(2):197–199, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTa

[Ano09g] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(1): 92–101, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTb

[Ano09h] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(2): 189–196, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTc

[Ano09i] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(3): 281–293, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2009:RBTd

[Ano09j] Anonymous. Reviews of books and teaching materials. The American Statistician, 63(4): 401–405, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:EC

[Ano10a] Anonymous. 2010 editorial collaborators. The American Statistician, 64(4):364–365, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:C

[Ano10b] Anonymous. Correction. The American Statistician, 64(2): 195, May 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:IV

[Ano10c] Anonymous. Index to volume 64. The American Statistician, 64(4):366–368, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:LEa

[Ano10d] Anonymous. Letters to the Editor. The American Statistician, 64(3):276, August 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:LEb

[Ano10e] Anonymous. Letters to the Editor. The American Statistician, 64(4):362–363, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTa

[Ano10f] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(1):

88–96, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTb

[Ano10g] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(3): 268–275, August 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2010:RBTc

[Ano10h] Anonymous. Reviews of books and teaching materials. The American Statistician, 64(4): 357–361, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:EC

[Ano11a] Anonymous. 2011 editorial collaborators. The American Statistician, 65(4):298–299, November 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEc

[Ano11b] Anonymous. Letter to the Editors. The American Statistician, 65(2):141, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEa

[Ano11c] Anonymous. Letters to the Editors. The American Statistician, 65(1):69, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEb

[Ano11d] Anonymous. Letters to the Editors. The American Statistician, 65(2):140, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:LEd

[Ano11e] Anonymous. Letters to the Editors. The American Statistician, 65(3):209–211, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTa

[Ano11f] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(1): 61–68, February 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTb

[Ano11g] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(2): 136–139, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBTc

[Ano11h] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(3): 201–208, August 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2011:RBT

[Ano11i] Anonymous. Reviews of books and teaching materials. The American Statistician, 65(4): 283–297, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ansley:1985:QPS

[Ans85] Craig F. Ansley. Quick proofs of some regression theorems via the qr algorithm. The American Statistician, 39(1):55–59, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arizono:1989:TNB

[AO89] Ikuo Arizono and Hiroshi Ohta. A test for normality based on Kullback–Leibler information. The American Statistician, 43(1):20–22, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aldous:2010:WCO

[AP10] David Aldous and Tung Phan. When can one test an explanation? Compare and contrast Benford’s Law and the fuzzy CLT. The American Statistician, 64(3):221–227, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Athanasios:2007:EEB

[APT07] Diamandopoulos A. Athanasios, Goudas C. Pavlos, and Kassimatis I. Theodoros. Early evidence-based medicine. The American Statistician, 61(2): 154–158, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Alper:1995:IVI

[AR95] Paul Alper and Robert L. Raymond. Imposing views, imposing shoes: A statistician as a sole model. The American Statistician, 49(3):316–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Allenby:2008:TBS

[AR08] Greg M. Allenby and Peter E.

Rossi. Teaching Bayesian statistics to marketing and business students. The American Statistician, 62(3):195–198, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnold:2006:DEM

[Arn06] Steven F. Arnold. Design of experiments with MINITAB. The American Statistician, 60(2): 205, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnholt:2007:SAS

[Arn07] Alan T. Arnholt. Statistics:

The art and science of learning from data. The American Statistician, 61(2):181–182, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Angus:1984:ICS

[AS84] John E. Angus and Ray E.

Schafer. Improved confidence statements for the binomial parameter. The American Statistician, 38(3):189–191, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Abernathy:1986:MBG

[AS86] James R. Abernathy and Pranab K. Sen. In memoriam: Bernard George Greenberg, 1919–1985. The American Statistician, 40(3):183–184, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Abel:1994:SFQ

[AS94] Patricia S. Abel and Nozer D. Singpurwalla. To survive or to fail: That is the question. The American Statistician, 48 (1):18–??, February 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arnold:1998:JCS

[AS98] Barry C. Arnold and Robert M. Shavelle. Joint confidence sets for the mean and variance of a normal distribution. The American Statistician, 52(2): 133–140, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/arnold.
html.

Anderson:1999:CSS

[AS99] Jon E. Anderson and Engin A. Sungur. Community service statistics projects. The American Statistician, 53(2):132–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/anderson.pdf.

Asok:1980:NCB

[Aso80] C. Asok. A note on the comparison between simple mean and mean based on distinct units in sampling with replacement. The American Statistician, 34(3):158, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Assuncao:1996:LEE

[Ass96] Renato Assuncao. Likelihood elasticity and error bounds. The American Statistician, 50(2): 165–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Aguirre-Torres:1986:NTI

[AT86] Victor Aguirre-Torres. A note on teaching infinite moving averages. The American Statistician, 40(1):40–41, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Au:1999:TVU

[AT99] Chi Au and Judy Tam. Transforming variables using the Dirac generalized function. The American Statistician, 53(3): 270–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Austin:2004:BMD

[AT04] Peter C. Austin and Jack V. Tu. Bootstrap methods for developing predictive models. The American Statistician, 58 (2):131–137, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Atkinson:1994:TBS

[Atk94] A. C. Atkinson. Transforming both sides of a tree. The American Statistician, 48(4):307–??, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Atkinson:1995:IDG

[Atk95] E. Neely Atkinson. Interactive dynamic graphics for exploratory survival analysis. The American Statistician, 49(1): 77–??, ???? 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ames:1991:EAD

[AW91a] Michael H. Ames and John T. Webster. On estimating approximate degrees of freedom. The American Statistician, 45 (1):45–??, February 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Arndt:1991:EBC

[AW91b] Stephan Arndt and Robert F. Woolson. Establishing a biostatistical care unit in a clinical research center. The American Statistician, 45(1):22–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Albert:2001:UMD

[AW01] Jim Albert and Patricia Williamson.

Using model/data simulations to detect streakiness. The American Statistician, 55(1): 41–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Albert.htm.

Annis:2006:CPP

[AW06] David H. Annis and Samuel S.

Wu. A comparison of potential playoff systems for NCAA I-A football. The American Statistician, 60(2):151–157, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brady:2002:CSB

[BA02] James E. Brady and Theodore T.

Allen. Case study based instruction of DOE and SPC. The American Statistician, 56(4): 312–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s9/p312.

Berry:2006:SPC

[BA06] Donald A. Berry and Gregory D. Ayers. Symmetrized percent change for treatment comparisons. The American Statistician, 60(1):27–31, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baglivo:1995:CAS

[Bag95] Jenny Baglivo. Computer algebra systems: Maple and Mathematica. The American Statistician, 49(1):86–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bailey:1992:DIB

[Bai92] Ralph W. Bailey. Distributional identities of beta and chisquared variates: A geometrical interpretation. The American Statistician, 46(2):117–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bailar:1995:LP

[Bai95] John C. Bailar III. A larger perspective. The American Statistician, 49(1):10–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barnow:1976:UPV

[Bar76] Burt S. Barnow. The use of proxy variables when one or two independent variables are measured with error. The American Statistician, 30(??):119– 121, 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barabba:1990:WDF

[Bar90] Vincent P. Barabba. What does the future hold for statisticians. The American Statistician, 44 (2):125–128, ???? 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bartelson:1998:BRW

[Bar98] Becki Bucher Bartelson. Book review: Workshop Statistics: Discovery with Data and the Graphing Calculator. The American Statistician, 52(4): 367–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bartelson:1999:BRP

[Bar99] Becki Bucher Bartelson. Book review: Probability Without Equations: Concepts for Clinicians. The American Statistician, 53(2):170–??, May 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2002:CNC

[Bar02] Lawrence Barker. A comparison of nine confidence intervals for a Poisson parameter when the expected number of events is 5. The American Statistician, 56(2):85–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s2/p85.

Barrett:2007:DBJ

[Bar07] Bruce E. Barrett. Detecting bias in jury selection. The American Statistician, 61(4): 296–301, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2009:MMC

[Bar09]  Chris Barker.  The mean,

median, and confidence intervals of the Kaplan–Meier survival estimate — computations and applications. The American Statistician, 63(1):78–80, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bassett:1991:EMB

[Bas91] Gilbert W. Bassett. Equivariant, monotonic, 50% breakdown estimators. The American Statistician, 45(2):135–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bassett:1997:RSR

[Bas97] Gilbert W. Bassett, Jr. Robust sports ratings based on least absolute errors. The American Statistician, 51(2):99–105, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts/bassett.html.

Bates:2004:BRI

[Bat04] Samantha C. Bates. Book review: Introductory Statistics with R. The American Statistician, 58(1):88– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s20/p88.

Barry:2004:DUS

[BB04] Daniel Barry and Philip J.

Boland. Debating the use of statistical evidence in allegations of age discrimination. The American Statistician, 58 (2):102–109, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berger:1992:DGM

[BC92] Roger L. Berger and George Casella. Deriving generalized means as least squares and maximum likelihood estimates. The American Statistician, 46(4):279–??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bryant:2000:MSU

[BCB00] Peter G. Bryant and Olga I.

Cordero-Brana. Model selection using the minimum description length principle. The American Statistician, 54(4): 257–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Bryant.htm.

Bedrick:1997:BBR

[BCJ97] Edward J. Bedrick, Ronald Christensen, and Wesley Johnson. Bayesian binomial regression: Predicting survival at a trauma center. The American Statistician, 51(3): 211–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/bedrick.html.

Barr:2011:SPW

[BD11] G. D. I. Barr and K. Dowie. Swivelling probabilities — the winning hand at poker. The American Statistician, 65(3): 170–176, August 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Beatty:1983:RGC

[Bea83] J. C. Beatty. Raster graphics and color. The American Statistician, 37(1):60–75, February 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bechhofer:1982:JCK

[Bec82] Robert Bechhofer. Jack Carl Kiefer, 1924–1981. The American Statistician, 36(4):356–357, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Beckman:1996:PFF

[Bec96] Richard J. Beckman. Plotting pk factorial or pn − k fractional factorial data. The American Statistician, 50(2):170–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Behboodian:1990:EUD

[Beh90] Javad Behboodian. Examples of uncorrelated dependent random variables using a bivariate mixture. The American Statistician, 44(3):218, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Belsley:1984:DCD

[Bel84] D. A. Belsley. Demeaning conditioning diagnostics through centering. The American Statistician, 38(??):73–77, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bellhouse:1987:MBE

[Bel87] D. R. Bellhouse. Model-based estimation in finite population sampling. The American Statistician, 41(4):260–262, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bellhouse:2001:SCG

[Bel01]  D. R. Bellhouse.  Statisti-

cal computing and graphics — the central limit theorem under simple random sampling. The American Statistician, 55 (4):352–357, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Belden:2005:DMB

[Bel05]  Susan Belden.  Discussion:

Moving beyond the St. Petersburg Paradox to understand growth stock valuation. The American Statistician, 59(3): 233–234, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bellhouse:2007:SME

[Bel07] David R. Bellhouse. Sampling methods: Exercises and solutions. The American Statistician, 61(3):276, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bennett:1993:DSJ

[Ben93] Jay Bennett. Did Shoeless Joe Jackson throw the 1919 World Series? The American Statistician, 47(4):241–??, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berengut:1981:SSS

[Ber81] David Berengut. Some smallsample properties of the Student t confidence intervals. The American Statistician, 35(3): 144–147, ???? 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berry:1987:IAC

[Ber87] Donald A. Berry. Interim analysis in clinical trials. the role of the likelihood principle. The American Statistician, 41(2):

117–122, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berger:1996:MPT

[Ber96a] Roger L. Berger. More powerful tests from confidence interval p values. The American Statistician, 50(4):314–??, November 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bergerud:1996:DFR

[Ber96b] Wendy A. Bergerud. Displaying factor relationships in experiments. The American Statistician, 50(3):228–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bernard:1996:BIF

[Ber96c] Jean-Marc Bernard. Bayesian interpretation of frequentist procedures for a Bernoulli process. The American Statistician, 50(1):7–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berry:1997:RBB

[Ber97a] Donald A. Berry. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):270–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berry:1997:TEB

[Ber97b] Donald A. Berry. Teaching elementary Bayesian statistics with real applications in science. The American Statistician, 51(3):241–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/berry.htm.

Berk:2003:D

[Ber03] Kenneth N. Berk. Discussion. The American Statistician, 57 (3):178–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s6/p178.

Berengut:2006:SED

[Ber06] David Berengut. Statistics for experimenters: Design, innovation, and discovery. The American Statistician, 60(4):341–342, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berrendero:2007:BMB

[Ber07] José R. Berrendero. The bagged median and the bragged mean. The American Statistician, 61 (4):325–330, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blom:1998:MP

[BES98] Gunnar Blom, Jan-Eric Englund, and Dennis Sandell. The Mississippi problem. The American Statistician, 52(1): 47–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/blom.html.

Betensky:2000:ADR

[Bet00] Rebecca A. Betensky. Alternative derivations of a rule for early stopping in favor of H_{0}. The American Statistician, 54 (1):??, February 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bartels:1991:SCS

[BF91] Robert Bartels and Denzil G.

Fiebig. A simple characterization of seemingly unrelated regressions models in which OLS is BLUE. The American Statistician, 45(2):137–140, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Benichou:1989:DMI

[BG89] Jacques Benichou and Mitchell H.

Gail. A delta method for implicitly defined random variables. The American Statistician, 43 (1):41–44, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bell:2008:ETP

[BG08] Richard C. Bell, Jr. and Andrew G. Glen. Experiences teaching probability and statistics with personal laptops in the classroom daily. The American Statistician, 62(2):155–160, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Busstra:2008:DDD

[BGF^{+}08] M. C. Busstra, A. Geelen, E. J. Feskens, R. J. M. Hartog, and P. van ’t Veer. Design and development of digital learning material for applied data analysis. The American Statistician, 62(4):329–339, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barr:1998:SNA

[BGG98] Donald R. Barr, Maj. Andrew G. Glen, and Harvey F. Graf. The “straightforward” nature of arrival rate estimation? The American Statistician, 52(4):346–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/barr.html.

Bryce:2001:CGB

[BGNP01] G. Rex Bryce, Robert Gould, William I. Notz, and Roxy L. Peck. Curriculum guidelines for Bachelor of Science degrees in statistical science. The American Statistician, 55(1):7– ??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blom:1986:RWO

[BH86] Gunnar Blom and Lars Holst.

Random walks of ordered elements with applications. The American Statistician, 40(4): 271–274, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:1993:HAH

[BH93] Lawrence D. Brown and J. T. Gene Hwang. How to approximate a histogram by a normal density. The American Statistician, 47(4):251–255, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bassett:1998:EAH

[BH98] G. W. Bassett and W. J. Hurley. The effects of alternative HOME-AWAY sequences in a best-of-seven playoff series. The American Statistician, 52 (1):49–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/bassett.
html.

Bhattacharya:2002:MVU

[BH02] Bhaskar Bhattacharya and Desale Habtzghi. Median of the p value under the alternative hypothesis. The American Statistician, 56(3): 202–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s7/p202.

Brimberg:2008:OPN

[BH08] Jack Brimberg and W. J. Hurley. Is the overtime period in an NHL game long enough? an example for teaching estimation and hypothesis testing in the presence of censored data. The American Statistician, 62 (2):151–154, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bar-Hillel:2002:SWA

[BHA02] Maya Bar-Hillel and Yigal Attali. Seek whence: Answer sequences and their consequences in key-balanced multiple-choice tests. The American Statistician, 56(4): 299–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s6/p299.

Bolstad:2001:TCS

[BHM01] W. M. Bolstad, Lyn A. Hunt, and Judith L. McWhirter. Teacher’s corner — sex, drugs, and rock & roll survey in a firstyear service course in statistics. The American Statistician, 55 (2):145–149, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boos:1998:GAB

[BHO98] Dennis D. Boos and Jacqueline M. Hughes-Oliver. General — applications of Basu’s theorem. The American Statistician, 52(3):218–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/boos.html.

Boos:2000:HLD

[BHO00] Dennis D. Boos and Jacqueline M. Hughes-Oliver. How large does n have to be for Z and t intervals? The American Statistician, 54(2):121–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Boos.htm.

Bornkamp:2009:NBS

[BI09] Björn Bornkamp and Katja Ickstadt. A note on B-splines for semiparametric elicitation. The American Statistician, 63 (4):373–377, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bickel:1995:WAN

[Bic95] Peter J. Bickel. What academia needs. The American Statistician, 49(1):5–??, ???? 1995.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Billard:1998:RSS

[Bil98] Lynne Billard. The role of statistics and the statistician. The American Statistician, 52 (4):319–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/billard.
html.

Billingsley:2004:DSC

[Bil04] Richard Billingsley. Discussion: Statistics and the college football championship. The American Statistician, 58(3):190, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Birkes:1990:GLR

[Bir90]  David Birkes.  Generalized

likelihood ratio tests and uniformly most powerful tests. The American Statistician, 44(2): 163–166, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Birch:1995:TSE

[Bir95] Jeffrey B. Birch. Ten suggestions for effectively teaching short courses to heterogeneous groups. The American Statistician, 49(2):190–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bisgaard:1991:TSE

[Bis91]  Søren Bisgaard.  Teaching

statistics to engineers. The American Statistician, 45(4): 274–??, November 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bisgaard:1998:TCD

[Bis98]  Søren Bisgaard.  Teacher’s

corner — discussion — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):238–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Becker:1992:LLS

[BK92] William Becker and Peter Kennedy. A lesson in least squares and R squared. The American Statistician, 46(4): 282–??, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boen:1993:CIS

[BK93] James R. Boen and Marcus O.

Kjelsberg. Career information for statistics graduate students. The American Statistician, 47 (2):148–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bhaumik:1996:SEM

[BK96] D. K. Bhaumik and P. M.

Kulkarni. A simple and exact method of constructing tolerance intervals for the oneway ANOVA with random effects. The American Statistician, 50(4):319–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bolton:1999:CPC

[BK99] Richard J. Bolton and Wojtek J. Krzanowski. A characterization of principal components for projection pursuit. The American Statistician, 53 (2):108–??, 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baker:2007:PYR

[BK07] Stuart G. Baker and Barnett S. Kramer. Peirce, Youden, and receiver operating characteristic curves. The American Statistician, 61(4):343–346, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:2009:RWS

[BK09a] Emery N. Brown and Robert E. Kass. Rejoinder: What is statistics? The American Statistician, 63(2):122–123, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:2009:WS

[BK09b] Emery N. Brown and Robert E. Kass. What is statistics? The American Statistician, 63 (2):105–110, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Becker:1996:PM

[BKM96] Richard A. Becker and Sallie Keller-McNulty. Presentation myths. The American Statistician, 50(2):112–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bar-Lev:1987:NPF

[BL87] Shaul K. Bar-Lev. Nonestimable parametric functions for some discrete distributions. The American Statistician, 41 (3):199–200, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bar-Lev:1994:DCC

[BL94] Shaul K. Bar-Lev. A derivation of conditional cumulants in exponential models. The American Statistician, 48(2):126–129, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bergmann:2000:DOW

[BLS00] Reinhard Bergmann, John Ludbrook, and Will P. J. M. Spooren. Different outcomes of the Wilcoxon–Mann–Whitney test from different statistics packages. The American Statistician, 54(1):??, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blyth:1980:EAE

[Bly80] Colin R. Blyth. Expected absolute error of the usual estimator of the binomial parameter. The American Statistician, 34 (3):155–157, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blyth:1993:REP

[Bly93] Colin R. Blyth. Restrict estimates to the possible values? The American Statistician, 47 (1):73–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Balakrishnan:1986:NMO

[BM86] N. Balakrishnan and Henrick J. Malik. A note on moments of order statistics. The American Statistician, 40(2):147–148, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bickel:1988:NUB

[BM88a] P. J. Bickel and C. L. Mallows. A note on unbiased Bayes estimates. The American Statistician, 42(2):132–134, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bloch:1988:NWL

[BM88b] Daniel A. Bloch and Lincoln E. Moses. Nonoptimally weighted least squares. The American Statistician, 42(1):50–53, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boomsma:1994:FET

[BM94] Anne Boomsma and Ivo W.

Molenaar. Four electronic tables for probability distributions. The American Statistician, 48(2):153–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Balkin:2001:GAA

[BM01] Sandy D. Balkin and Colin L.

Mallows. General — an adjusted, asymmetric two-sample t test. The American Statistician, 55(3):203–206, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bayarri:2002:BDS

[BM02a] M. J. Bayarri and A. M.

Mayoral. Bayesian design of “successful” replications. The American Statistician, 56(3): 207–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s8/p207.

Bessant:2002:GTO

[BM02b] Kenneth C. Bessant and Eric D.

MacPherson. General — thoughts on the origins, concepts, and pedagogy of statistics as a “separate discipline”. The American Statistician, 56 (1):22–28, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barrowman:2003:RPN

[BM03] Nicholas J. Barrowman and Ransom A. Myers. Raindrop plots: A new way to display collections of likelihoods and distributions. The American Statistician, 57(4):268– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s10/p268.

Birch:2005:TVT

[BM05] Jeffrey B. Birch and J. P. Morgan. TA training at Virginia Tech. The American Statistician, 59(1):14–18, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blitzstein:2010:NPQ

[BM10] Joseph Blitzstein and Xiao-Li Meng. Nano-project qualifying exam process: An intensified dialogue between students and faculty. The American Statistician, 64(4):282–290, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barton:1998:OSL

[BN98] Russell R. Barton and Craig A. Nowack. A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):

233–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/barton.
html.

Bowerman:1992:IQC

[BO92] Bruce L. Bowerman and Richard T. O’Connell. Integrating quality control concepts into the introductory business statistics course. The American Statistician, 46(2):124–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boardman:1994:SWC

[Boa94] Thomas J. Boardman. The statistician who changed the world: W. Edwards Deming, 1900–1993. The American Statistician, 48(3):179–187, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boland:1984:BGW

[Bol84] Philip J. Boland. A biographical glimpse of William Sealy Gosset. The American Statistician, 38(??):179–183, 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bolstad:1986:HSF

[Bol86] William M. Bolstad. Harrison-Stevens forecasting and the multiprocess dynamic linear model. The American Statistician, 40(2):129–135, ???? 1986.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boland:2003:PST

[Bol03] Philip J. Boland. Promoting statistical thinking among secondary school students in the national context. The American Statistician, 57(2): 85–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s4/p85.

Bond:1998:BRA

[Bon98] Marjorie E. Bond. Book reviews: Applied Statistical Methods for Business, Economics, and the Social Sciences and Statistics for Managers: Using Microsoft Excel. The American Statistician, 52(2):201–?? (or 199–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bond:2002:BRT

[Bon02] Marjorie E. Bond. Book review: Teach/Me Data Analysis CD: Single User Edition. The American Statistician, 56(4): 335–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s23/p335.

Boos:1992:GST

[Boo92] Dennis D. Boos. On generalized score tests. The American Statistician, 46(4):327–??, November 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Borack:1997:TEP

[Bor97] Jules I. Borack. A technique for estimating the probability of detecting a nongaming drug user. The American Statistician, 51(2):134–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/borack.html.

Boronico:1999:MTP

[Bor99] Jess S. Boronico. Multi-tiered playoffs and their impact on professional baseball. The American Statistician, 53(1): 56–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bordley:2002:RTU

[Bor02]  Robert F. Bordley.  Rep-

resenting trees using Microsoft doughnut charts. The American Statistician, 56(2): 139–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s10/p139.

Box:1980:RFD

[Box80]  Joan Fisher Box.  R. A.

Fisher and the design of experiments, 1922–1926. The American Statistician, 34(1):1– 7, ???? 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Box:1981:GFD

[Box81]  Joan Fisher Box.  Gosset,

Fisher, and the t distribution. The American Statistician, 35 (2):61–66, ???? 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Box:2005:RRF

[Box05] Joan Fisher Box. A reminiscence of R. A. Fisher. The American Statistician, 59(4): 312–314, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boyles:2008:RLI

[Boy08] Russell A. Boyles. The role of likelihood in interval estimation. The American Statistician, 62(1):22–26, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2002:INP

[BP02] Thomas E. Bradstreet and Milton N. Parnes. Illustrating the Neyman–Pearson lemma with a stopping rule of order k. The American Statistician, 56 (2):113–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s7/p113.

Beniger:1978:QGS

[BR78] James R. Beniger and Dorothy L.

Robyn. Quantitative graphics in statistics: A brief history. The American Statistician, 32 (??):1–11, 1978. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Balakrishnan:1997:NBL

[BR97] N. Balakrishnan and C. R. Rao.

A note on the best linear unbiased estimation based on order statistics. The American Statistician, 51(2):181–185, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/balakrishnan.
html.

Blume:2003:ILL

[BR03] Jeffrey D. Blume and Richard M.

Royall. Illustrating the law of large numbers (and confidence intervals). The American Statistician, 57(1):51– ??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s8/p51.

Bradstreet:1996:TIS

[Bra96a] Thomas E. Bradstreet. Teaching introductory statistics courses so that nonstatisticians experience statistical reasoning. The American Statistician, 50 (1):69–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brant:1996:DLR

[Bra96b] Rollin Brant. Digesting logistic regression results. The American Statistician, 50(2):117–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brackstone:2007:DOS

[Bra07] Gordon Brackstone. Discussion: The official statistics Olympic challenge. The American Statistician, 61(1):9–12, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2008:GLD

[Bra08a] Thomas E. Bradstreet. Graphics of large datasets: Visualizing a million. The American Statistician, 62(2):180–181, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradstreet:2008:PSU

[Bra08b] Thomas E. Bradstreet. Pharmaceutical statistics using SAS: A practical guide. The American Statistician, 62(2):183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bring:1994:HSR

[Bri94] Johan Bring. How to standardize regression coefficients. The American Statistician, 48 (3):209–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bring:1996:GAC

[Bri96] Johan Bring. A geometric approach to compare variables in a regression model. The American Statistician, 50(1):57–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bromaghin:1993:SSD

[Bro93] Jeffrey F. Bromaghin. Sample size determination for interval estimation of multinomial probabilities. The American Statistician, 47(3):203–??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Browne:2001:SPU

[Bro01] Richard H. Browne. Statistical practice — using the sample range as a basis for calculating sample size in power calculations. The American Statistician, 55(4):293–298, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Brown:2007:SLB

[Bro07] Aaron Brown. Strong language on black swans. The American Statistician, 61(3):195–197, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Broers:2008:HSB

[Bro08] Nick J. Broers. Helping students to build a conceptual understanding of elementary statistics. The American Statistician, 62(2):161–166, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Browne:2010:TVR

[Bro10] Richard H. Browne. The t-test p value and its relationship to the effect size and P (X > Y). The American Statistician, 64 (1):30–33, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Broemeling:2011:AES

[Bro11] Lyle D. Broemeling. An account of early statistical inference in Arab cryptology. The American Statistician, 65(4): 255–257, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Barker:2001:SPE

[BRRB01] Lawrence Barker, Henry Rolka, Deborah Rolka, and Cedric Brown. Statistical practice — equivalence testing for binomial random variables: Which test to use? The American Statistician, 55(4):279–287, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bryce:1999:D

[Bry99] G. Rex Bryce. Discussion.

The American Statistician, 53 (1):21–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bradley:1979:CPR

[BS79] Ralph A. Bradley and Sushil S.

Srivastava. Correlation in polynomial regression. The American Statistician, 33(??):11–14, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Buhler:1987:SRE

[BS87] Wolfgang J. Bühler and Joachim Sehr. Some remarks on exponential families. The American Statistician, 41(4):279–280, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Balakrishnan:1995:SSA

[BS95] N. Balakrishnan and R. A.

Sandhu. A simple simulational algorithm for generating progressive Type-II censored samples. The American Statistician, 49(2):229–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Booth:1998:MCA

[BS98] James G. Booth and Somnath Sarkar. Monte Carlo approximation of bootstrap variances. The American Statistician, 52(4):354–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/booth.html.

Barr:1999:MVT

[BS99] Donald R. Barr and E. Todd Sherrill. Mean and variance of truncated normal distributions. The American Statistician, 53(4):357–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bingham:2000:EHG

[BS00a] Derek R. Bingham and Tim B. Swartz. Equitable handicapping in golf. The American Statistician, 54(3):170– ??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Bingham.htm.

Blaker:2000:PII

[BS00b] Helge Blaker and Emil Spjotvoll. Paradoxes and improvements in interval estimation. The American Statistician, 54(4): 242–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Blaker.htm.

Bonett:2003:CIM

[BS03] Douglas G. Bonett and Edith Seier. Confidence intervals for mean absolute deviations. The American Statistician, 57(4): 233–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s2/p233.

Beaudoin:2010:SPG

[BS10a] David Beaudoin and Tim B.

Swartz. Strategies for pulling the goalie in hockey. The American Statistician, 64(3): 197–204, August 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Block:2010:GEB

[BS10b] Henry W. Block and Thomas H.

Savits. A general example for Benford data. The American Statistician, 64(4):335–339, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boos:2011:VPR

[BS11] Dennis D. Boos and Leonard A.

Stefanski. P -value precision and reproducibility. The American Statistician, 65(4):213– 221, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bathke:2009:GGA

[BSTM09] Arne C. Bathke, Oliver Schabenberger, Randall D. Tobias, and Laurence V. Madden. Greenhouse–Geisser adjustment and the ANOVA-type statistic: Cousins or twins? The American Statistician, 63 (3):239–246, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Boullion:1992:MDP

[BSY92] Thomas L. Boullion, John W. Seaman, Jr., and Dean M. Young. Moments of discrete probability distributions derived using a differential operator. The American Statistician, 46(1):22–24, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Blom:1990:DRP

[BTV90] Gunnar Blom, Daniel Thorburn, and Theodore A. Vessey. The distribution of the record position and its applications. The American Statistician, 44 (2):151–153, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berenson:2008:ASR

[BUK^{+}08] Mark L. Berenson, Jessica Utts, Karen A. Kinard, Deborah J. Rumsey, Albyn Jones, and Leonard M. Gaines. Assessing student retention of essential statistical ideas. The American Statistician, 62(1):54–61, February 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bull:2005:ES

[Bul05] Brian Bull. Exemplar sampling. The American Statistician, 59 (2):166–172, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Buonaccorsi:1987:NCI

[Buo87] John P. Buonaccorsi. A note on confidence intervals for proportions in finite populations. The American Statistician, 41 (3):215–218, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Burrows:1986:ESS

[Bur86] Peter M. Burrows. Extreme statistics from the sine distribution. The American Statistician, 40(3):216–217, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Buzas:1997:FEJ

[Buz97] J. S. Buzas. Fast estimators of the jackknife. The American Statistician, 51(3): 235–240, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/buzas.html.

Begg:2011:BSP

[BV11] Melissa D. Begg and Roger D.

Vaughan. Are biostatistics students prepared to succeed in the era of interdisciplinary science? (and how will we know ?). The American Statistician, 65(2):71–79, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Baksalary:1988:CTC

[BvE88] Jerzy K. Baksalary and Adrian C. van Eijnsbergen. A comparison of two criteria for ordinaryleast-squares estimators to be best linear unbiased estimators. The American Statistician, 42 (3):205–208, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bhatti:2008:CPH

[BW08] Chad R. Bhatti and Jennifer L. Wightman. Conditional probability and HIV testing. The American Statistician, 62(3): 238–241, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Berger:1998:QSA

[BWM98] Paul D. Berger, Chen Wang, and James P. Monahan. Quantifying a statistical aspect of segmented selection/quota systems. The American Statistician, 52(3):228–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts_98/berger.
                                        html.

Chen:1997:ID

[CA97] Gemai Chen and A. Adatia. Independence and t distribution. The American Statistician, 51(2):176–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.

amstat.org/publications/
tas/abstracts/chen.html.

Cain:1994:MGF

[Cai94] Michael Cain. The momentgenerating function of the minimum of bivariate normal random variables. The American Statistician, 48(2):124–125, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Callaert:2003:CSS

[Cal03] Herman Callaert. Comparing statistical software packages: The case of the logrank test in StatXact. The American Statistician, 57(3): 214–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s15/p214.

Cannon:1998:BRU

[Can98] Ann R. Cannon. Book review:

Understanding Data: Principles and Practice of Statistics. The American Statistician, 52(4):367–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cannon:2001:ESE

[Can01] Ann Cannon. Essential Statistics (3rd ed.). The American Statistician, 55(1):83–??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carlstein:1986:SCR

[Car86] E. Carlstein. Simultaneous confidence regions for predictions. The American Statistician, 40 (4):277–279, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carlin:1996:INB

[Car96] Bradley P. Carlin. Improved NCAA basketball tournament modeling via point spread and team strength information. The American Statistician, 50(1): 39–??, ???? 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carlton:2000:DAS

[Car00] Matthew A. Carlton. Data analysis: Statistical and computational methods for scientists and engineers. The American Statistician, 54(2):155–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carlton:2003:D

[Car03] Matthew A. Carlton. Discussion. The American Statistician, 57(3):179–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

                                        oberon.ingentaselect.com/
                                        cgi-bin/linker?ini=asa&reqidx=
                                        /cw/asa/00031305/v57n3/s7/
                                        p179.

Carlton:2008:PSC

[Car08] Matthew A. Carlton. Probability and statistics for computer scientists. The American Statistician, 62(3):271–272, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Casella:1985:IEB

[Cas85] George Casella. An introduction to empirical Bayes data analysis. The American Statistician, 39(2):83–87, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cavanaugh:2007:STS

[Cav07a]  Joseph Cavanaugh.  Seeing

through statistics and mind on statistics. The American Statistician, 61(1):95, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cavanaugh:2007:SEA

[Cav07b] Joseph Cavanaugh. Statistics:

The exploration and analysis of data. The American Statistician, 61(1):97–98, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Coe:1995:OSS

[CB95a] Paul R. Coe and William Butterworth. Optimal stopping in “The showcase showdown”. The American Statistician, 49 (3):271–275, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cook:1995:BSU

[CB95b] Peyton Cook and Lyle D. Broemeling. Bayesian statistics using Mathematica. The American Statistician, 49(1):70–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:2009:RPL

[CBH09] Aiyou Chen, Thomas Bengtsson, and Tin Kam Ho. A regression paradox for linear models: Sufficient conditions and relation to Simpson’s Paradox. The American Statistician, 63 (3):218–225, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Castillo:2001:SCG

[CC01] Enrique Del Castillo and Suntara Cahya. Statistical computing and graphics — A tool for computing confidence regions on the stationary point of a response surface. The American Statistician, 55(4): 358–365, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:2006:SDP

[CC06] Dale J. Cohen and Jon Cohen. The sectioned density plot. The American Statistician, 60 (2):167–174, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carter:1986:ACR

[CCW^{+}86] Walter H. Carter, Jr., Vernon M. Chinchilli, John D. Wilson, Eleanor D. Campbell, Fay K. Kessler, and Richard A. Carchman. An asymptotic confidence region for the ED_{100p} from the logistic response surface for a combination of agents. The American Statistician, 40 (??):124–128, 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cressie:1981:MGF

[CDFP81] Noel Cressie, Anne S. Davis, J. Leroy Folks, and George E. Policello II. The momentgenerating function and negative integer moments. The American Statistician, 35(3): 148–150, ???? 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cockerill:1991:IPA

[CF91] Rhonda Cockerill and Bruce Fried. Increasing public awareness of statistics as a science and a profession — reinforcing the message in universities. The American Statistician, 45 (3):174–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cabilio:2001:TCC

[CF01] P. Cabilio and P. J. Farrell. Teacher’s corner — A computer-based lab supplement to courses in introductory statistics. The American Statistician, 55(3):228–232, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chatterjee:2007:GDI

[CF07] Sangit Chatterjee and Aykut Firat. Generating data with identical statistics but dissimilar graphics. The American Statistician, 61(3):248–254, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2008:PPO

[CF08] William F. Christensen and Lindsay W. Florence. Predicting presidential and other multistage election outcomes using state-level pre-election polls. The American Statistician, 62 (1):1–10, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Casella:1992:EGS

[CG92] George Casella and Edward I. George. Explaining the Gibbs sampler. The American Statistician, 46(3):167–174, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Charnes:1995:UCC

[CG95a] John M. Charnes and Howard S. Gitlow. Using control charts to corroborate bribery in jai alai. The American Statistician, 49 (4):386–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chib:1995:UMH

[CG95b] Siddartha Chib and Edward Greenberg. Understanding the Metropolis–Hastings algorithm. The American Statistician, 49 (4):327–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Caffo:2006:UFI

[CG06] Brian Caffo and Michael Griswold. A user-friendly introduction to link-probit-normal models. The American Statistician, 60(2):139–145, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cho:2007:BBL

[CG07] Wendy K. Tam Cho and Brian J. Gaines. Breaking the (Benford) Law. The American Statistician, 61(3):218–223, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chan:1983:ACS

[CGL83] Tony F. Chan, Gene H. Golub, and Randall J. LeVeque. Algorithms for computing the sample variance: Analysis and recommendations. The American Statistician, 37(3):242–247, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:1985:BPE

[CH85] Ronald Christensen and Michael D.

Huffman. Bayesian point estimation using the predictive distribution. The American Statistician, 39(4):319–321, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [CH86].

Christensen:1986:CBP

[CH86] Ronald Christensen and Michael D. Huffman. Correction to: “Bayesian point estimation using the predictive distribution” [amer. statist. 39 (1985), no. 4, part 1, 319–321]. The American Statistician, 40(3):258, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [CH85].

Christensen:1996:NAR

[CH96] Ronald Christensen and Aparna V. Huzurbazar. A note on augmenting resolution III designs. The American Statistician, 50 (2):175–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chattamvelli:1995:NNB

[Cha95] R. Chattamvelli. A note on the noncentral beta distribution function. The American Statistician, 49(2):231–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chambers:1999:CDC

[Cha99] John Chambers. Computing with data: Concepts and challenges. The American Statistician, 53(1):73–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chan:2004:SPO

[Cha04] Grace Chan. SAS programming: The one-day course. The American Statistician, 58(3): 260, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chapman:2008:ESP

[Cha08] Jessica Chapman. Elementary statistics: Picturing the world. The American Statistician, 62 (4):357–358, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:2011:DCS

[Che11] Yung-Pin Chen. Do the chisquare test and Fisher’s exact test agree in determining extreme for 2 × 2 tables? The American Statistician, 65(4): 239–245, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chitturi:2008:IBS

[Chi08] Pallavi Chitturi. Introduction to business statistics. The American Statistician, 62(3): 269, August 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2008:PNS

[CHJ08] Ronald Christensen, Timothy Hanson, and Alejandro Jara. Parametric nonparametric statistics. The American Statistician, 62(4):296–306, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:2003:SIT

[Chr03] Ronald Christensen. Significantly insignificant F tests. The American Statistician, 57(1): 27–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s4/p27.

Christensen:2005:TFN

[Chr05] Ronald Christensen. Testing Fisher, Neyman, Pearson, and Bayes. The American Statistician, 59(2):121–126, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Choudhury:1999:UTS

[CHS99] Askar H. Choudhury, Robert Hubata, and Robert D. St. Louis. Understanding timeseries regression estimators. The American Statistician, 53 (4):342–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cantwell:2004:USM

[CHS04] Patrick J. Cantwell, Howard Hogan, and Kathleen M. Styles. The use of statistical methods in the U.S. Census. The American Statistician, 58(3):203–212, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chun:1999:IEA

[Chu99] Young H. Chun. On the information economics approach to the generalized game show problem. The American Statistician, 53(1):43–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:1992:ASE

[CJ92] Steven B. Cohen and Ayah E.

Johnson. Address sample effects in the National Medical Expenditure Survey. The American Statistician, 46(2): 100–??, May 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chib:1987:ALS

[CJT87] Siddhartha Chib, S. Rao Jammalamadaka, and Ram C. Tiwari. Another look at some results on the recursive estimation in the general linear model. The American Statistician, 41 (1):56–58, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:2001:SPN

[CK01] Zhen Chen and Lynn Kuo. Statistical practice — A note on the estimation of the multinomial logit model with random effects. The American Statistician, 55(2):89–95, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chernick:2002:STB

[CL02] Michael R. Chernick and Christine Y. Liu. The saw-toothed behavior of power versus sample size and software solutions: Single binomial proportion using exact methods. The American Statistician, 56 (2):149–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s12/p149.

Chakraborti:2007:CIE

[CL07] S. Chakraborti and J. Li. Confidence interval estimation of a normal percentile. The American Statistician, 61(4):331–336, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:1999:BRW

[Cle99a] Richard Cleary. Book review: Workshop Statistics: Discovery with Data and Minitab, Data Analysis with Microsoft Excel, and Statistical Laboratory Exercises Using Excel: A Guide to Understanding Data. The American Statistician, 53(4): 388–??, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:1999:SME

[Cle99b] Richard Cleary. Statistical methods for engineers. The American Statistician, 53(3): 292–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2000:HTA

[Cle00a] Richard Cleary. How to think about statistics. The American Statistician, 54(4):322–??, November 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2000:SIE

[Cle00b] Richard Cleary. Statistics: An Introduction, 5th edition. The American Statistician, 54(4): 324–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2001:RBT

[Cle01] Richard J. Cleary. Reviews of books and teaching materials — quantitative investigations in the biosciences using MINITAB^{T M}. The American Statistician, 55(2):167, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2003:BRSb

[Cle03a] Richard J. Cleary. Book review:

Statistical Rules of Thumb. The American Statistician, 57 (2):142–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s14/p142.

Cleary:2003:BRSa

[Cle03b] Richard J. Cleary. Book review: Statistics with Confidence (2nd ed.). The American Statistician, 57(1):66– ??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s11/p66.

Cleary:2003:BRF

[Cle03c] Richard J. Cleary. Book review: The First Measured Century. The American Statistician, 57(2):143–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s16/p143.

Cleary:2003:BRD

[Cle03d] Richard J. Cleary. Book reviews: Duelling Idiots and Other Probability Puzzlers and What are the Chances? Voodoo Deaths and Office Gossip and Other Adventures in Probability. The American Statistician, 57(3):218–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/

linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s18/p218.

Cleary:2005:GTSa

[Cle05a] Richard J. Cleary. Gauss: Titan of science and Karl Pearson: The scientific life in a statistical age. The American Statistician, 59(4):349, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2005:GTSb

[Cle05b] Richard J. Cleary. Gauss: Titan of science and Karl Pearson: The scientific life in a statistical age. The American Statistician, 59(4):349, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2005:TSB

[Cle05c] Richard J. Cleary. Teaching statistics: A bag of tricks. The American Statistician, 59(3): 275, August 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:ITS

[Cle06a]  Richard J. Cleary.  Inno-

vations in teaching statistics and achieving quantitative literacy: An urgent challenge for higher education. The American Statistician, 60(1):99–100, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:PVG

[Cle06b] Richard J. Cleary. Probability is the very guide of life: The philosophical uses of chance. The American Statistician, 60 (3):290, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2006:WNS

[Cle06c] Richard J. Cleary. Working with numbers and statistics: A handbook for journalists. The American Statistician, 60(2): 203, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2008:FPS

[Cle08a] Richard J. Cleary. Fundamentals of probability and statistics for engineers, statistical techniques for data analysis and statistics: An introduction. The American Statistician, 62(1): 90–91, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2008:SID

[Cle08b] Richard J. Cleary. Statistics: Informed decisions using data and an introduction to statistical analysis for business and industry: A problem solving approach. The American Statistician, 62(2):183–184, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chang:1992:TSS

[CLM92] Ted C. Chang, Sharon L. Lohr, and C. Graham McLaren. Teaching survey sampling using simulation. The American Statistician, 46(3):232–??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chang:2008:NIA

[CLPC08] Ching-Hui Chang, Jyh-Jiuan Lin, Nabendu Pal, and Miao-Chen Chiang. A note on improved approximation of the binomial distribution by the skew-normal distribution. The American Statistician, 62(2): 167–170, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Casella:2001:GEP

[CLR01] George Casella, Michael Lavine, and Christian P. Robert. General — explaining the perfect sampler. The American Statistician, 55(4):299–305, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chung:2004:DDI

[CLS04] Hwan Chung, Eric Loken, and Joseph L. Schafer. Difficulties in drawing inferences with finite-mixture models. The American Statistician, 58(2): 152–158, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Capotorti:2007:RAR

[CLS07] Andrea Capotorti, Frank Lad, and Giuseppe Sanfilippo. Reassessing accuracy rates of median decisions. The American Statistician, 61(2):132–138, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Clason:1992:EFV

[CM92] Dennis L. Clason and Leigh W. Murray. Exponential families and variance component models. The American Statistician, 46(1):29–??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cobb:1999:D

[Cob99] George W. Cobb. Discussion. The American Statistician, 53 (1):16–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cochran:2005:CYR

[Coc05a] James J. Cochran. Can you really learn basic probability by playing a sports board game? The American Statistician, 59 (3):266–272, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cochran:2005:NGB

[Coc05b] James J. Cochran. The numbers game: Baseball’s lifelong fascination with statistics. The American Statistician, 59(2): 203, May 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:1991:DVS

[Coh91] Ayala Cohen. Dummy variables in stepwise regression. The American Statistician, 45 (3):226–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cohen:1997:EAP

[Coh97] Steven B. Cohen. An evaluation of alternative PC-Based software packages developed for the analysis of complex survey data. The American Statistician, 51(3):285–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/cohen.html.

Cohen:2002:BRI

[Coh02] Arthur Cohen. Book review:

An Introduction to Probability Theory and Mathematical Statistics, (2nd ed.). The American Statistician, 56(4): 331–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s17/p331.

Cohen:2010:LE

[Coh10] Michael P. Cohen. Letter to the Editor. The American Statistician, 64(2):192, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Collings:2001:RBT

[Col01] Patti B. Collings. Reviews of books and teaching materials — Fathom: Dynamic systems software. The American Statistician, 55(3):258, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Colleys:2004:DSC

[Col04]  Wes Colleys.  Discussion:

Statistics and the college football championship. The American Statistician, 58(3):191–192, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Corner:2000:SPA

[Cor00] Denis G. Janky Teacher’s Corner. Sometimes pooling for analysis of variance hypothesis tests: A review and study of a split-plot model. The American Statistician, 54(4): 269–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Agresti.htm;   http://
                                        www.amstat.org/publications/
                                        tas/Janky.htm.

Cotton:1998:BRA

[Cot98]  John W. Cotton.  Book

review: Analyzing Within-Subjects Experiments. The American Statistician, 52(4): 368–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cotsonis:2006:SYS

[Cot06] George Cotsonis. Sharpening your SAS skills. The American Statistician, 60(4):346, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cowles:2003:BRR

[Cow03] Mary Kathryn Cowles. Book review: An R and S-PLUS Companion to Applied Regression. The American Statistician, 57 (4):316–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s16/p316.

Cowles:2004:BRSb

[Cow04a] Mary Kathryn Cowles. Book review: SAS for Linear Models (4th ed.). The American Statistician, 58(1):89– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s22/p89.

Cowles:2004:BRSa

[Cow04b] Mary Kathryn Cowles. Book review: Statistical Computing: An Introduction to Data Analysis using SPlus. The American Statistician, 58(1): 89–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/

                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s21/p89.

Cowles:2004:RW

[Cow04c] Mary Kathryn Cowles. Review of WinBUGS 1.4. The American Statistician, 58(4):330–336, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cowles:2006:PSC

[Cow06] Mary Kathryn Cowles. Probability and statistics for computer science. The American Statistician, 60(1):98, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cox:1984:EIM

[Cox84] Christopher Cox. An elementary introduction to maximum likelihood estimation for multinomial models: Birch’s theorem and the delta method. The American Statistician, 38 (4):283–287, ???? 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Clemons:1999:BN

[CP99] Traci Clemons and Marcello Pagano. Are babies normal? The American Statistician, 53(4):298–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cameron:1995:CE

[CR95] Joseph M. Cameron and Joan R. Rosenblatt. Churchill Eisenhart, 1913-1994. The American Statistician, 49(3): 243–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carroll:1996:UMO

[CR96] R. J. Carroll and David Ruppert. The use and misuse of orthogonal regression in linear errors-in-variables models. The American Statistician, 50(1):1– ??, ???? 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chance:2001:TCS

[CR01] Beth L. Chance and Allan J.

Rossman. Teacher’s corner — sequencing topics in introductory statistics: A debate on what to teach when. The American Statistician, 55(2): 140–144, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Craigmile:2005:ASC

[Cra05] Peter F. Craigmile. All of statistics: A concise course in statistical inference. The American Statistician, 59(2):203–204, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Crawford:2008:UCG

[Cra08] Carol A. Gotway Crawford. Unlocking the census with GIS. The American Statistician, 62 (4):363, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cesana:2001:SPS

[CRM01] B. M. Cesana, G. Reina, and E. Marubini. Statistical practice — sample size for testing a proportion in clinical trials: A “two-step” procedure combining power and confidence interval expected width. The American Statistician, 55(4): 288–292, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cornell:1995:CTS

[CRV95] John A. Cornell, Ron H. Randles, and G. Geoff Vining. Challenges in teaching short courses by an academic department: The University of Florida experience. The American Statistician, 49(2):195–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cryer:1991:TTS

[Cry91] Jonathan D. Cryer. TIMES-LAB: A times series laboratory. The American Statistician, 45 (3):243–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:1983:SSR

[CS83] Robert Chen and Larry A. Shepp. On the sum of symmetric random variables. The American Statistician, 37(3): 237, ???? 1983. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chevan:1991:HP

[CS91] Albert Chevan and Michael Sutherland. Hierarchical partitioning. The American Statistician, 45(2):90–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chen:1994:QAF

[CS94] Wenjun Chen and Tim Swartz.

Quantitative aspects of five-pin bowling. The American Statistician, 48(2):92–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carlton:2005:MBF

[CS05] Matthew A. Carlton and William D. Stansfield. Making babies by the flip of a coin? The American Statistician, 59 (2):180–182, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chib:1991:EBA

[CT91] Siddhartha Chib and Ram C.

Tiwari. Extreme bounds analysis in the Kalman filter. The American Statistician, 45(2): 113–114, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Christensen:1992:BRE

[CU92] Ronald Christensen and Jessica Utts. Bayesian resolution of the “exchange paradox”. The American Statistician, 46(4):

274–276, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cuadras:1993:IIM

[Cua93] Carles M. Cuadras. Interpreting an inequality in multiple regression. The American Statistician, 47(4):256–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Carroll:1988:NAR

[CW88] Raymond J. Carroll and A. H. Welsh. A note on asymmetry and robustness in linear regression. The American Statistician, 42(4):285–287, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cox:1992:CCD

[CW92] D. R. Cox and Nanny Wermuth. A comment on the coefficient of determination for binary responses. The American Statistician, 46(1):1–??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Coyle:1993:WBG

[CW93] Cynthia A. Coyle and Chamont Wang. Wanna bet? on gambling strategies that may or may not work in a casino. The American Statistician, 47 (2):108–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cook:1999:GSA

[CW99] R. Dennis Cook and Sanford Weisberg. Graphs in statistical analysis: Is the medium the message? The American Statistician, 53(1):29–??, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cook:2004:POD

[CW04] R. Dennis Cook and Sanford Weisberg. Partial onedimensional regression models. The American Statistician, 58 (2):110–116, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cleary:2010:BDF

[CW10] Rick Cleary and Sam Woolford.

The business of desire and fear. The American Statistician, 64 (1):21–22, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Cron:2011:ECB

[CW11] Andrew J. Cron and Mike West.

Efficient classification-based relabeling in mixture models. The American Statistician, 65(1): 16–20, February 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chatterjee:1999:NEM

[CY99] Sangit Chatterjee and Mustafa R.

Yilmaz. The NBA as an evolving multivariate system. The American Statistician, 53(3):

257–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Czitrom:1998:TCD

[Czi98] Veronica Czitrom. Teacher’s corner — discussion — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):240–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Czitrom:1999:OFT

[Czi99]  Veronica Czitrom.  One-

factor-at-a-time versus designed experiments. The American Statistician, 53(2):126–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/czitrom.pdf.

Dobler:2005:BRTb

[DAC05] Carolyn Pillers Dobler and Christine M. Anderson-Cook. Brief reviews of teaching materials. The American Statistician, 59(3):278, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2006:BRT

[DAC06] Carolyn Pillers Dobler and Christine M. Anderson-Cook. Brief reviews of teaching materials. The American Statistician, 60(2):207, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DeMets:1994:TNG

[DAF^{+}94] David L. DeMets, Dan Anbar, William Fairweather, Thomas A. Louis, and Robert T. O’Neill. Training the next generation of biostatisticians. The American Statistician, 48(4): 280–??, November 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dahiya:1981:IME

[Dah81] Ram C. Dahiya. An improved method of estimating an integer-parameter by maximum likelihood. The American Statistician, 35(1):34–37, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dallal:1990:SCP

[Dal90] Gerard E. Dallal. Statistical computing packages: Dare we abandon their teaching to others? The American Statistician, 44(4):265–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Danaher:1992:SSM

[Dan92] Peter J. Danaher. Some statistical modeling problems in the advertising industry: A look at media exposure distributions. The American Statistician, 46(4):254–??, November 1992. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Daniel:1994:FOF

[Dan94] Cuthbert Daniel. Factorial onefactor-at-a-time experiments. The American Statistician, 48 (2):132–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:1980:ULR

[Dav80] Anne S. Davis. Use of the likelihood ratio test on the inverse Gaussian distribution. The American Statistician, 34 (2):108–110, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:1981:EP

[Dav81] Herbert A. David. Egon S. Pearson, 1895–1980. The American Statistician, 35(2):94–95, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:1993:NOS

[Dav93] H. A. David. A note on order statistics for dependent variates. The American Statistician, 47(3):198–199, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:1995:FOC

[Dav95] H. A. David. First (?) occurrence of common terms in mathematical statistics. The American Statistician, 49(2):

121–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:1998:FOC

[Dav98a] H. A. David. First (?) occurrence of common terms in probability and statistics — A second list, with corrections. The American Statistician, 52 (1):36–40, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/david.html.

David:1998:SOM

[Dav98b] H. A. David. Stronger opposition may make it easier to win. The American Statistician, 52(4):351–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/david2.
html.

David:2005:TRN

[Dav05] Herbert A. David. Tables related to the normal distribution. The American Statistician, 59(4):309–311, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2006:IMA

[Dav06] H. A. David. The introduction of matrix algebra into statistics. The American Statistician, 60(2):162, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2007:CCS

[Dav07] J. Wade Davis. A crash course in SPSS for windows: Updated for versions 10, 11, 12 and 13. The American Statistician, 61 (1):99, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2008:BRT

[Dav08a] Herbert A. David. The beginnings of randomization tests. The American Statistician, 62 (1):70–72, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:HUM

[Dav08b] J. Wade Davis. Handbook of univariate and multivariate data analysis and interpretation with SPSS. The American Statistician, 62(3):268, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:MST

[Dav08c]  J. Wade Davis.  Medical

statistics: A textbook for the health sciences. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davis:2008:MND

[Dav08d]  J. Wade Davis.  Multidi-

mensional nonlinear descriptive analysis. The American Statistician, 62(3):271, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2009:HNZ

[Dav09] Herbert A. David. A historical note on zero correlation and independence. The American Statistician, 63(2):185–186, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2011:ECM

[Dav11] Herbert A. David. Euler’s contributions to mathematics useful in statistics. The American Statistician, 65(1):37–42, February 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dawkins:1991:SPC

[Daw91] Brian Dawkins. Siobhan’s problem: The coupon collector revisited. The American Statistician, 45(1):76–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dayton:1998:ICP

[Day98] C. Mitchell Dayton. Information criteria for the pairedcomparisons problem. The American Statistician, 52(2): 143–?? (or 144–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/

                                        tas/abstracts_98/dayton.
                                        html.

DAgostino:1990:SUP

[DB90] Ralph B. D’Agostino and Albert Belanger. A suggestion for using powerful and informative tests of normality. The American Statistician, 44(4):316–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Doganaksoy:1997:UPB

[DB97] Necip Doganaksoy and N. Balakrishnan. A useful property of best linear unbiased predictors with applications to lifetesting. The American Statistician, 51(1):22–28, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts/doganaksoy.
                                        html.

Dennis:1985:PMG

[DBS^{+}85] B. Dennis, B. E. Brown, A. R. Stage, H. E. Burkhart, and S. Clark. Problems of modeling growth and yield of renewable resources. The American Statistician, 39(4):374–383, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deng:1990:RRE

[DC90] Lih-Yuan Deng and Raj S. Chhikara. On the ratio and regression estimation in finite population sampling. The American Statistician, 44(4): 282–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deCani:1986:NDI

[dCS86] John S. de Cani and Robert A.

Stine. A note on deriving the information matrix for a logistic distribution. The American Statistician, 40(3):220–222, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Drane:1993:LFP

[DCWP93] J. Wanzer Drane, Su Hua Cao, Lixia Wang, and T. Postelnicu. Limiting forms of probability mass functions via recurrence formulas. The American Statistician, 47(4):269–274, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Das:2006:BAG

[DD06] Sourish Das and Dipak K. Dey.

On Bayesian analysis of generalized linear models using the Jacobian technique. The American Statistician, 60(3): 264–268, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deLaubenfels:2006:VLS

[deL06] Ralph deLaubenfels. The victory of least squares and orthogonality in statistics. The American Statistician, 60(4):315–321, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deming:1975:PBA

[Dem75] W. Edwards Deming. On probability as a basis for action. The American Statistician, 29 (4):146–152, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deming:2006:ESE

[Dem06] W. Edwards Deming. On errors in surveys (an excerpt). The American Statistician, 60 (1):34–38, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Derr:1993:BCI

[Der93] Janice Derr. Biostatistics cores: Improving the chances for funding. The American Statistician, 47(2):99–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2001:RBT

[Dev01] Jay Devore. Reviews of books and teaching materials — stat labs: Mathematical statistics through applications. The American Statistician, 55(3): 257, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2002:BRF

[Dev02] Jay Devore. Book review: A First Course in Probability. The American Statistician, 56(3):

247–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s14/p247.

Devore:2006:EMS

[Dev06a] Jay Devore. Essential mathematics and statistics for science. The American Statistician, 60(4):344–345, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2006:PSS

[Dev06b] Jay Devore. Probability, statistics, and stochastic processes. The American Statistician, 60 (3):290, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2006:SBE

[Dev06c] Jay Devore. Statistics for business and economics. The American Statistician, 60(4): 342–343, November 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:ARA

[Dev07a] Jay Devore. Applied regression analysis: A second course in business and economic statistics. The American Statistician, 61(4):367, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:IBS

[Dev07b]  Jay Devore.  Introduction

to business statistics. The American Statistician, 61(1): 97, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2007:MSD

[Dev07c] Jay Devore. Making sense of data: A practical guide to exploratory data analysis and data mining. The American Statistician, 61(4):370–371, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Devore:2008:S

[Dev08] Jay Devore. Statistics. The American Statistician, 62(4): 362–363, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2007:SMK

[DF07] Herbert A. David and Wayne A. Fuller. Sir Maurice Kendall (1907–1983). The American Statistician, 61(1):41–46, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:1997:PTU

[DG97] H. A. David and Jason L. Gunnink. The paired t test under artificial pairing. The American Statistician, 51(1): 9–12, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/david.html.

Datta:2007:CFC

[DG07] Gauri Sankar Datta and Malay Ghosh. Characteristic functions without contour integration. The American Statistician, 61(1):67–70, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dawson:1997:TGT

[DGC97] Kathryn S. Dawson, Chris Gennings, and Walter H. Carter. Two graphical techniques useful in detecting correlation structure in repeated measures data. The American Statistician, 51 (3):275–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/dawson.html.

Draper:1987:RRS

[DH87] Norman R. Draper and Agnes M.

Herzberg. A ridge-regression sidelight. The American Statistician, 41(4):282–283, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Danaher:2005:BYE

[DH05] Peter J. Danaher and Bruce G. S. Hardie. Bacon with your eggs? Applications of a new bivariate beta-binomial distribution. The American Statistician, 59(4):282–286, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Demirtas:2011:PWC

[DH11] Hakan Demirtas and Donald Hedeker. A practical way for computing approximate lower and upper correlation bounds. The American Statistician, 65 (2):104–109, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Davi:2005:RTT

[DHN^{+}05] Angelique Davi, Dominique Haughton, Nada Nasr, Gaurav Shah, Maria Skaletsky, and Ruth Spack. A review of two text-mining packages. The American Statistician, 59(1): 89–103, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Du:2011:PUL

[DHW11] Yeting Du, Christopher Hundt, and David B. Wolfson. Previously unrecognized links between statistical dependence and some common modes of convergence. The American Statistician, 65(1):55–60, February 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dietz:1993:CLA

[Die93] E. Jacquelin Dietz. A cooperative learning activity on methods of selecting a sample. The American Statistician, 47 (2):104–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DiPietro:2004:BRR

[DiP04] Michele DiPietro. Bayesian randomized response as a class project. The American Statistician, 58(4):303–309, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dixon:2007:LSB

[Dix07] Philip Dixon. The little SAS book for enterprise guide 3.0. The American Statistician, 61 (3):277, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Duchateau:2005:UHG

[DJ05] Luc Duchateau and Paul Janssen. Understanding heterogeneity in generalized mixed and frailty models. The American Statistician, 59(2):143–146, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dawes:1987:PRR

[DK87] Robyn M. Dawes and Joseph B.

Kadane. Partial round-robin comparisons with perfect rankings. The American Statistician, 41(3):204–205, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Driscoll:1995:APC

[DK95] Michael F. Driscoll and Barbara Krasnicka. An accessible proof of Craig’s theorem in the general case. The American Statistician, 49(1):59–62, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dodd:2007:BVS

[DK07] Lori E. Dodd and Edward L. Korn. The bootstrap variance of the square of a sample mean. The American Statistician, 61 (2):127–131, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Deng:2000:RNG

[DL00] Lih-Yuan Deng and Dennis K. J. Lin. Random number generation for the new century. The American Statistician, 54 (2):145–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Deng.htm.

Dobler:2004:BRT

[DL04] Carolyn Pillers Dobler and Russell V. Lenth. Brief reviews of teaching materials. The American Statistician, 58(2): 175, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2005:BRTa

[DL05] Carolyn Pillers Dobler and Russell V. Lenth. Brief reviews of teaching materials. The American Statistician, 59(1):114–115, February 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dlas:2005:SAF

[Dla05] Alexandra Dlas. Statistical analysis of financial data in S-Plus. The American Statistician, 59(4):351, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Denby:2001:SPE

[DLM01] L. Denby, J. M. Landwehr, and C. L. Mallows. Statistical practice — an exercise in the real world of design and analysis. The American Statistician, 55 (4):263–271, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dempster:1981:MWG

[DM81] A. P. Dempster and Frederick Mosteller. In memoriam: William Gemmell Cochran, 1909–1980. The American Statistician, 35(1):38, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dawson:2001:GNH

[DM01] Don Dawson and Lonnie Magee. General — the National Hockey League Entry Draft, 1969–1995: An application of a weighted pool-adjacent-violators algorithm. The American Statistician, 55(3):194– 199, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

David:2006:SSW

[DM06] Herbert A. David and Donald F. Morrison. Samuel Stanley Wilks (1906–1964). The American Statistician, 60(1): 46–49, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deMast:2007:AKT

[dM07] Jeroen de Mast. Agreement and kappa-type indices. The American Statistician, 61(2):148–153, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

deMicheaux:2009:UCC

[dML09] Pierre Lafaye de Micheaux and Benoit Liquet. Understanding convergence concepts: A visual-minded and graphical simulation-based approach. The American Statistician, 63 (2):173–178, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dinh:1994:APM

[DN94a] K. T. Dinh and Truc T. Nguyen. Acknowledgement of priority: “Maximum likelihood estimators of binomial parameters under an order restriction” [Amer. Statist. 48 (1994), no. 1, 29–30]. The American Statistician, 48(4):354, ???? 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [DN94b].

Dinh:1994:MLE

[DN94b] Khoan T. Dinh and Truc T.

Nguyen. Maximum likelihood estimators of binomial parameters under an order restriction. The American Statistician, 48 (1):29–30, February 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See acknowledgement of priority [DN94a].

Dargahi-Noubary:1998:RMT

[DNG98] G. R. Dargahi-Noubary and Jo Anne S. Growney. Risk — A motivating theme for an introductory statistics course. The American Statistician, 52 (1):42–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/noubary.
html.

Dette:2004:EED

[DO04] Holger Dette and Timothy E.

O’Brien. Efficient experimental design for the Behrens– Fisher problem with application to bioassay. The American Statistician, 58(2):138–143, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2002:BRM

[Dob02] Carolyn Pillers Dobler. Book review: Mathematical Statistics: Basic Ideas and Selected Topics (vol. 1, 2nd ed.). The American Statistician, 56(4): 332–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s18/p332.

Dobler:2003:BRPb

[Dob03a] Carolyn Pillers Dobler. Book review: Problems in Probability. The American Statistician, 57 (3):218–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s17/p218.

Dobler:2003:BRPa

[Dob03b] Carolyn Pillers Dobler. Book review: The Practice of Statistics: Putting the Pieces Together. The American Statistician, 57(2):142–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s15/p142.

Dobler:2004:SPP

[Dob04] Carolyn Pillers Dobler. 101 special practice problems in probability and statistics. The American Statistician, 58(2): 173, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2006:EMT

[Dob06a] Carolyn Pillers Dobler. The essence of multivariate thinking: Basic themes and methods. The American Statistician, 60(3):292, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2006:SGR

[Dob06b] Carolyn Pillers Dobler. SAS guide to report writing: Examples. The American Statistician, 60(3):294, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Donahue:1999:NIS

[Don99] Rafe M. J. Donahue. A note on information seldom reported via the P value. The American Statistician, 53(4):303–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dorfman:1996:PSR

[Dor96] Jeffrey H. Dorfman. Pseudorandom sampling has a real effect on test size. The American Statistician, 50(2):151–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Douglas:1993:CRP

[Dou93] J. B. Douglas. Confidence regions for parameter pairs. The American Statistician, 47(1): 43–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dorsey-Palmateer:2004:BHH

[DPS04] Reid Dorsey-Palmateer and Gary Smith. Bowlers’ hot hands. The American Statistician, 58(1):38–??, February 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

oberon.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=
/cw/asa/00031305/v58n1/s10/
p38.

Dodge:1999:CFC

[DR99] Yadolah Dodge and Valentin Rousson. The complications of the fourth central moment. The American Statistician, 53 (3):267–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dodge:2001:APC

[DR01] Yadolah Dodge and Valentin Rousson. On asymmetric properties of the correlation coefficient in the regression setting. The American Statistician, 55 (1):51–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Dodge.htm.

Dumbgen:2007:FAB

[DR07] Lutz Dümbgen and Hans Riedwyl. On fences and asymmetry in box-and-whiskers plots. The American Statistician, 61 (4):356–359, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Driscoll:1999:IRR

[Dri99] Michael F. Driscoll. An improved result relating quadratic forms and chi-square distributions. The American Statistician, 53(3):273–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Desu:1981:SA

[DS81] M. M. Desu and Norman C.

Severo. Sidney Addelman (1932–1979). The American Statistician, 35(??):39, 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DuarteSilva:1998:NTG

[DS98] A. Pedro Duarte Silva and Antonie Stam. Nonparametric two-group classification: Concepts and a SAS-based software package. The American Statistician, 52(2):183–?? (or 185–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/silva.html.

Duckworth:2002:BTS

[DS02] William M. Duckworth and W. Robert Stephenson. Beyond traditional statistical methods. The American Statistician, 56 (3):230–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s11/p230.

Datta:2008:GPS

[DS08] Gauri S. Datta and Sanat K. Sarkar. A general proof of some known results of independence between two statistics. The American Statistician, 62 (2):141–143, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Doane:2000:UBF

[DT00] David P. Doane and Ronald L. Tracy. Using beam and fulcrum displays to explore data. The American Statistician, 54(4): 289–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Doane.htm.

Dufour:1986:BLR

[Duf86] Jean-Marie Dufour. Bias of S^{2} in linear regressions with dependent errors. The American Statistician, 40(4):284–285, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DuMouchel:1999:BDM

[DuM99] William DuMouchel. Bayesian data mining in large frequency tables, with an application to the FDA spontaneous reporting system. The American Statistician, 53(3):177–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dunn:1991:PSP

[Dun91] Charles L. Dunn. Precise simulated percentiles in a pinch. The American Statistician, 45 (3):207–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dunlop:1994:RLD

[Dun94] Dorothy D. Dunlop. Regression for longitudinal data: A bridge from least squares regression. The American Statistician, 48(4):299–303, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dunn:1999:TTI

[Dun99] Peter K. Dunn. Three tools for interactively visualizing some distribution theory concepts. The American Statistician, 53(2):137–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/dunn.pdf.

Dunn:2003:WHW

[Dun03] Peter K. Dunn. What happens when a 1 × 1 × r die is rolled? The American Statistician, 57 (4):258–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.

                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s8/p258.

Duncan:2004:GIC

[Dun04] Kristin A. Duncan. Grade inflation: A crisis in college education. The American Statistician, 58(3):259, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

DeMets:1998:WJS

[DWBQ98] David L. DeMets, Robert Woolson, Curtis Brooks, and Roger Qu. Where the jobs are: A study of amstat news job advertisements. The American Statistician, 52(4):303–??, November 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Emerson:2011:SSL

[EA11] John W. Emerson and Taylor B. Arnold. Statistical sleuthing by leveraging human nature: A study of Olympic figure skating. The American Statistician, 65(3):143–148, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Easterling:2004:TED

[Eas04] Robert G. Easterling. Teaching experimental design. The American Statistician, 58(3): 244–252, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Easterling:2010:PDS

[Eas10] Robert G. Easterling. Passiondriven statistics. The American Statistician, 64(1):1–5, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eaton:1978:YCM

[Eat78] P. W. Eaton. Yarnold’s criterion and minimum sample size. The American Statistician, 32 (3):102–103, ???? 1978. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Edelman:1990:CIC

[Ede90a] David Edelman. A confidence interval for the center of an unknown unimodal distribution based on a sample of size 1. The American Statistician, 44(4): 285–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Edelman:1990:NUM

[Ede90b] David Edelman. A note on uniformly most powerful twosided tests. The American Statistician, 44(3):219–220, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Edwards:1986:RHB

[Edw86] A. W. F. Edwards. Is the reference in Hartley (1749) to Bayesian inference. The American Statistician, 40(2):109–110, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Edwards:1996:DET

[Edw96] Christopher T. Edwards. Doubleelimination tournaments: Counting and calculating. The American Statistician, 50(1):27–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Edwards:2007:ESB

[Edw07] Phillip Edwards. Essentials of statistics for business and economics. The American Statistician, 61(2):182–183, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Efron:1986:WIE

[Efr86] B. Efron. Why isn’t everyone a Bayesian? with discussion and a reply by the author. The American Statistician, 40(1):1– 11, ???? 1986. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Efron:1983:LLB

[EG83] Bradley Efron and Gail Gong. A leisurely look at the bootstrap, the jackknife, and crossvalidation. The American Statistician, 37(1):36–48, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eisenberg:1987:IED

[EG87] Bennett Eisenberg and B. K. Ghosh. Independent events in a discrete uniform probability space. The American Statistician, 41(1):52–56, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Elswick:1991:SAF

[EGCD91] R. K. Elswick, Chris Gennings, Vernon M. Chinchilli, and Kathryn S. Dawson. A simple approach for finding estimable functions in linear models. The American Statistician, 45(1):51–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ernst:1996:SUP

[EGS96] Michael D. Ernst, Rudy Guerra, and William R. Schucany. Scatterplots for unordered pairs. The American Statistician, 50(3):260–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ehrenberg:1990:HFS

[Ehr90] A. S. C. Ehrenberg. A hope for the future of statistics: MSOD. The American Statistician, 44 (3):195–??, August 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eisenhart:1979:TSS

[Eis79] Churchill Eisenhart. On the transition from “Student’s” z to “Student’s” t. The American Statistician, 33(1):6–10, ???? 1979. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Erosheva:2002:CDR

[EKG02] Elena A. Erosheva, Patricia D. Kroboth, and Joel B. Greenhouse. Characterizing the diurnal rhythm of DHEA. The American Statistician, 56(4): 273–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s3/p273.

Eldridge:1990:SAC

[Eld90] Marie D. Eldridge. The status of advisory committees to the Federal Statistical Agencies. The American Statistician, 44(2):154–??, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Elian:2000:SFB

[Eli00] N. Elian. Simple forms of the best linear unbiased predictor in the general linear model silvia. The American Statistician, 54(1):??, February 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Escobar:1993:NUR

[EM93] Luis A. Escobar and E. Barry Moser. A note on the updating of regression estimates. The American Statistician, 47 (3):192–194, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Emerson:1996:SPG

[Eme96] Scott S. Emerson. Statistical packages for group sequential methods. The American Statistician, 50(2):183–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Engeman:1993:SCP

[Eng93] Richard M. Engeman. Short courses as part of statistical consulting in a foreign setting. The American Statistician, 47 (2):122–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Emrich:1991:MGH

[EP91] Lawrence J. Emrich and Marion R. Piedmonte. A method for generating high-dimensional multivariate binary variates. The American Statistician, 45 (4):302–??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Epps:1993:CFT

[Epp93] T. W. Epps. Characteristic functions and their empirical counterparts: Geometrical interpretations and applications to statistical inference. The American Statistician, 47(1): 33–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Engeman:1993:AWS

[ES93] Richard M. Engeman and Stephen A. Shumake. Animal welfare and the statistical consultant. The American Statistician, 47(3):229–??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Emerson:2009:AJB

[ESL09] John W. Emerson, Miki Seltzer, and David Lin. Assessing judging bias: An example from the 2000 Olympic Games. The American Statistician, 63(2): 124–131, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eubank:1994:SSS

[Eub94]  R. L. Eubank.  A simple

smoothing spline. The American Statistician, 48(2):103–106, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eubank:1985:SVD

[EW85] R. L. Eubank and J. T. Webster. The singular-value decomposition as a tool for solving estimability problems. The American Statistician, 39(1):64–66, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Eubank:2002:GEB

[EW02] R. L. Eubank and Suojin Wang. General — the equivalence between the Cholesky decomposition and the Kalman filter. The American Statistician, 56 (1):39–43, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Farnum:1989:APS

[Far89] Nicholas R. Farnum. An alternate proof of Samuelson’s inequality and its extensions. The American Statistician, 43 (1):46–47, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Farnum:1990:IRE

[Far90] Nicholas R. Farnum. Improving the relative error of estimation. The American Statistician, 44(4):288–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fawcett:1990:AMD

[Faw90] Richard F. Fawcett. The alignment method for displaying and analyzing treatments in blocking designs. The American Statistician, 44(3):204–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fay:2002:SPM

[Fay02] Michael P. Fay. Statistical practice — measuring a binary response’s range of influence in logistic regression. The American Statistician, 56(1):5–9, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fears:1996:RFW

[FBG96] Thomas R. Fears, Jacques Benichou, and Mitchell H. Gail. A reminder of the fallibility of the Wald statistic. The American Statistician, 50(3):226–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Froelich:2009:DYI

[FDC09] Amy G. Froelich, William M. Duckworth, and Jessica Culhane. Does your iPod really play favorites? The American Statistician, 63(3):263–268, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Froelich:2005:TST

[FDS05] Amy G. Froelich, William M. Duckworth, and W. Robert Stephenson. Training statistics teachers at Iowa State University. The American Statistician, 59(1):8–10, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fellingham:2008:STS

[Fel08] Gilbert W. Fellingham. Statistical thinking in sports. The American Statistician, 62(4): 360–361, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ferentinos:1990:SCI

[Fer90] Kosmas K. Ferentinos. Shortest confidence intervals for families of distributions involving truncation parameters. The American Statistician, 44(2):167–168, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ferketich:2004:IBH

[Fer04] Amy K. Ferketich. Introductory biostatistics for the health sciences. The American Statistician, 58(2):172, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fewster:2009:SEB

[Few09] R. M. Fewster. A simple explanation of Benford’s Law. The American Statistician, 63 (1):26–32, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Feldman:1980:TSI

[FF80] Dorian Feldman and Martin Fox. Three simple inductive proofs in probability and statistics. The American Statistician, 34(1):50–51, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fay:2002:SCG

[FF02] Michael P. Fay and Dean A.

Follmann. Statistical computing and graphics — designing Monte Carlo implementations of permutation or bootstrap hypothesis tests. The American Statistician, 56(1):63–70, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freidlin:2000:SMT

[FG00] Boris Freidlin and Joseph L. Gastwirth. Should the median test be retired from general use? The American Statistician, 54 (3):161–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Freidlin.htm.

Freeman:1983:PSP

[FGHK83] Daniel H. Freeman, Jr., María Elena González, David C. Hoaglin, and Beth A. Kilss. Presenting statistical papers. The American Statistician, 37(2):106–110, 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Franklin:2011:ASB

[FHP^{+}11] Christine Franklin, Brad Hartlaub, Roxy Peck, Richard Scheaffer, David Thiel, and Katherine Tranbarger Freier. AP statistics: Building bridges between high school and college statistics education. The American Statistician, 65(3): 177–182, August 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Finney:2006:CGC

[Fin06] David J. Finney. Calibration guidelines challenge outlier practices. The American Statistician, 60(4):309–314, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Feingold:1986:CDB

[FK86] Marcia Feingold and P. E. Korsog. The correlation and dependence between two F statistics with the same denominator. The American Statistician, 40 (3):218–219, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fuller:1987:MTA

[FK87] Wayne A. Fuller and Oscar Kempthorne. In memoriam: Theodore Alfonso Bancroft, 1907–1986. The American Statistician, 41(3):167–168, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:1997:PCC

[FK97] Michael R. Frey and John C.

Kern. The Pitman closeness of a class of scaled estimators. The American Statistician, 51 (2):151–154, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/frey.html.

Friendly:2009:WWV

[FK09] Michael Friendly and Ernest Kwan. Where’s Waldo? visualizing collinearity diagnostics. The American Statistician, 63 (1):56–65, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fesco:1996:TSS

[FKL^{+}96] Ronald S. Fesco, William D. Kalsbeek, Sharon L. Lohr, Richard L. Scheaffer, Fritz J. Scheuren, and Elizabeth A. Stasny. Teaching survey sampling. The American Statistician, 50(4):328–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fong:2003:UST

[FKLL03] Daniel Y. T. Fong, C. W. Kwan, K. F. Lam, and Karen S. L. Lam. Use of the sign test for the median in the presence of ties. The American Statistician, 57(4):237–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s3/p237.

Fitzmaurice:2007:AMR

[FLP07] Garrett M. Fitzmaurice, Stuart R. Lipsitz, and Michael Parzen. Approximate median regression via the Box–Cox transformation. The American Statistician, 61(3):233–238, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Flury:1986:SRV

[Flu86] Bernhard K. Flury. On sums of random variables and independence. The American Statistician, 40(3):214–215, ???? 1986.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fishman:1984:SDD

[FM84] G. S. Fishman and L. R.

Moore. Sampling from a discrete distribution while preserving monotonicity. The American Statistician, 38(??):219– 223, ???? 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Federer:1992:CAS

[FM92] Walter T. Federer and Michael P.

Meredith. Covariance analysis for split-plot and split-block designs. The American Statistician, 46(2):155–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2008:SMI

[FM08] Jesse Frey and Osvaldo Marrero. A surprising MLE for interval-censored binomial data. The American Statistician, 62(2):135–137, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Faes:2009:ESS

[FMA^{+}09] Christel Faes, Geert Molenberghs, Marc Aerts, Geert Verbeke, and Michael G. Kenward. The effective sample size and an alternative small-sample degrees-of-freedom method. The American Statistician, 63 (4):389–399, November 2009.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Flury:1992:MAM

[FN92] Bernard D. Flury and A. Narayanan. A mixture approach to multivariate analysis of variance. The American Statistician, 46 (1):31–??, February 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ford:2005:PES

[For05] E. David Ford. A primer of ecological statistics. The American Statistician, 59(4):350, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Foster:2008:IB

[Fos08] Eric D. Foster. Introduction to bioinformatics. The American Statistician, 62(3):268–269, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Foster:2011:SSR

[Fos11] Colin Foster. The significance of a square-root rule. The American Statistician, 65(4): 222, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fotopoulos:2002:BRA

[Fot02] Stergios B. Fotopoulos. Book review: Applied Probability Models. The American Statistician, 56(3):248– ??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s15/p248.

Fox:2010:DFQ

[Fox10] David R. Fox. Desired and feared— quo vadis or quid agis? The American Statistician, 64 (1):6–9, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fahmy:1981:BDO

[FP81] Salwa Fahmy and Frank Proschan. Bounds on differences of order statistics. The American Statistician, 35(1):??, February 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fisher:2004:RWM

[FR04] Lynn A. Fisher and Donald St. P. Richards. Random walks as motivational material in introductory statistics and probability courses. The American Statistician, 58(4):310–316, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Franklin:1992:GIM

[Fra92] LeRoy A. Franklin. Graphical insight into multiple regression concepts. The American Statistician, 46(4):284–??, November 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Franklin:2000:ISM

[Fra00] Christine Franklin. Intermediate Statistics: A Modern Approach (2nd ed.). The American Statistician, 54(3):225–??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Finner:2007:CDC

[FRD07] Helmut Finner, Markus Roters, and Thorsten Dickhaus. Characterizing density crossing points. The American Statistician, 61(1):28–33, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freund:1965:PP

[Fre65] J. E. Freund. Puzzle or paradox? The American Statistician, 19(4):29–44, ???? 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:1983:NSR

[Fre83] David A. Freedman. A note on screening regression equations. The American Statistician, 37 (2):152–155, ???? 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:1997:DBB

[Fre97] David A. Freedman. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):266–??, August 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2005:RMB

[Fre05] Jesse Frey. A ranking method based on minimizing the number of in-sample errors. The American Statistician, 59(3): 207–216, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2006:CHS

[Fre06] David A. Freedman. On the so-called “Huber Sandwich Estimator” and “Robust Standard Errors”. The American Statistician, 60(4):299–302, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2007:HCS

[Fre07a] David A. Freedman. How can the score test be inconsistent? The American Statistician, 61 (4):291–295, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2007:BIP

[Fre07b] Jesse Frey. Bayesian inference on a proportion believed to be a simple fraction. The American Statistician, 61(3):201–206, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2007:HBT

[Fre07c] Jesse Frey. Is an .833 hitter better than a .338 hitter? The American Statistician, 61 (2):105–111, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Freedman:2008:SA

[Fre08] David A. Freedman. Survival analysis. The American Statistician, 62(2):110–119, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Frey:2010:FWS

[Fre10] Jesse Frey. Fixed-width sequential confidence intervals for a proportion. The American Statistician, 64(3):242–249, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friendly:1995:CVM

[Fri95] Michael Friendly. Conceptual and visual models for categorical data. The American Statistician, 49(2):153–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friendly:2002:CED

[Fri02] Michael Friendly. Corrgrams: Exploratory displays for correlation matrices. The American Statistician, 56(4):316– ??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s10/p316.

Fisch:1993:SAC

[FS93] Roland D. Fisch and Gunther A. Strehlau. A simplified approach to calibration confidence sets. The American Statistician, 47(3):168–??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fisher:2001:SCG

[FS01] N. I. Fisher and P. Switzer. Statistical computing and graphics — graphical assessment of dependence: Is a picture worth 100 tests? The American Statistician, 55(3):233–239, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Foster:2006:BWB

[FS06] Dean P. Foster and Robert A.

Stine. Being Warren Buffett. The American Statistician, 60 (1):53–60, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Flury:1993:RLC

[FT93] Bernard D. Flury and Thaddeus Tarpey. Representing a large collection of curves: A case for principal points. The American Statistician, 47(4): 304–??, November 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fearnhead:2010:CSS

[FT10] Paul Fearnhead and Benjamin M. Taylor. Calculating strength of schedule, and choosing teams for march madness. The American Statistician, 64 (2):108–115, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Fuller:1980:BVS

[Ful80] Wayne A. Fuller. Balkrishna V. Sukhatme (1924–1979). The American Statistician, 34(1): 49, ???? 1980. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friendly:2010:FKS

[FVMU10] Michael Friendly, Pedro Valero-Mora, and Joaquín Ibáñez Ulargui. The first (known) statistical graph: Michael Florent van Langren and the “Secret” of Longitude. The American Statistician, 64(2):185–191, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Friedman:2005:GVS

[FW05] Lynn Friedman and Melanie Wall. Graphical views of suppression and multicollinearity in multiple linear regression. The American Statistician, 59 (2):127–136, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Flury:2000:EE

[FZ00] Bernard Flury and Alice Zoppe. Exercises in EM. The American Statistician, 54(3):207– ??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Flury.htm.

Gadbury:2001:GRI

[Gad01] Gary L. Gadbury. General — randomization inference and bias of standard errors. The American Statistician, 55(4): 310–313, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gal:2003:TSL

[Gal03] Iddo Gal. Teaching for statistical literacy and services of statistics agencies. The American Statistician, 57(2): 80–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s3/p80.

Gambino:2009:DEC

[Gam09] Jack G. Gambino. Design effect caveats. The American Statistician, 63(2):141–146, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gange:1995:GMC

[Gan95] Stephen J. Gange. Generating multivariate categorical variates using the iterative proportional fitting algorithm. The American Statistician, 49(2): 134–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gregoire:1992:MSE

[GAR92] Timothy G. Gregoire, Alexandros A. Arabatzis, and Marion R. Reynolds. Mean squared error performance of simple linear regression conditional upon the outcome of pretesting the intercept. The American Statistician, 46(2):89–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gastwirth:1992:SRL

[Gas92] Joseph L. Gastwirth. Statistical reasoning in the legal setting. The American Statistician, 46 (1):55–??, February 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gates:1995:WRE

[Gat95] Charles E. Gates. What really is experimental error in block designs? The American Statistician, 49(4):362–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Genest:2003:DDK

[GB03] Christian Genest and Jean-Claude Boies. Detecting dependence with Kendall plots. The American Statistician, 57(4): 275–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/

linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s11/p275.

Gullion:2006:WSC

[GB06] Christina M. Gullion and Nancy Berman. What statistical consultants do. The American Statistician, 60(2):130–138, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goutis:1999:ESA

[GC99] Constantino Goutis and George Casella. Explaining the saddlepoint approximation. The American Statistician, 53(3): 216–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guo:2004:SJM

[GC04] Xu Guo and Bradley P. Carlin. Separate and joint modeling of longitudinal and event time data using standard computer packages. The American Statistician, 58(1):16– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s7/p16.

Gelber:1995:CTU

[GCGG95] Richard D. Gelber, Bernard F.

Cole, Shari Gelber, and Aron Goldhirsch. Comparing treatments using qualityadjusted survival: The Q-TWiST method. The American Statistician, 49(2):161–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:1989:UMV

[GD89] Malay Ghosh and Gauri Sankar Datta. Uniformly minimum variance unbiased estimation for discrete distributions with support depending on the parameter. The American Statistician, 43(1):17–19, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Grewal:2008:URG

[GDL08] Rajdeep Grewal, James A. Dearden, and Gary L. Llilien. The university rankings game. The American Statistician, 62 (3):232–237, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gilliland:2011:URC

[GE11] Dennis Gilliland and Don Edwards. Using randomized confidence limits to balance risk: An application to medicare investigations. The American Statistician, 65(3):149–153, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Geisser:1984:PDB

[Gei84] Seymour Geisser. On prior distributions for binary trials. with comments by Jose M. Bernardo, Melvin R. Novick and Arnold Zellner and a reply by the author. The American Statistician, 38(4):244–251, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [Zel85].

Gelman:1998:SCP

[Gel98]  Andrew Gelman.  Some

class-participation demonstrations for decision theory and Bayesian statistics. The American Statistician, 52(2):165–?? (or 167–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/gelman2.
html.

Glen:2001:SCG

[GEL01] Andrew G. Glen, Diane L.

Evans, and Lawrence M. Leemis. Statistical computing and graphics — APPL: A probability programming language. The American Statistician, 55 (2):156–166, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelfand:2004:DSR

[Gel04] Alan E. Gelfand. Discussion of “statistical research. The American Statistician, 58(3): 197–198, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2005:CTS

[Gel05] Andrew Gelman. A course on teaching statistics at the university level. The American Statistician, 59(1):4–7, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2006:BWC

[Gel06] Andrew Gelman. The boxer, the wrestler, and the coin flip. The American Statistician, 60 (2):146–150, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2008:TBG

[Gel08] Andrew Gelman. Teaching Bayes to graduate students in political science, sociology, public health, education, economics, …. The American Statistician, 62(3):202–205, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gentleman:2003:BRR

[Gen03] Robert Gentleman. Book review: Resampling Methods: A Practical Guide to Data Analysis (2nd ed.). The American Statistician, 57(1): 70–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n1/s18/p70.

Gentle:2004:CSC

[Gen04] James E. Gentle. Courses in statistical computing and computational statistics. The American Statistician, 58(1):

2–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s2/p2.

Gerber:1979:CPP

[Ger79] Hans U. Gerber. A characteristic property of the Poisson distribution. The American Statistician, 33(2):85–86, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Giron:2005:BAM

[GGR05] Javier Girón, Josep Ginebra, and Alex Riba. Bayesian analysis of a multinomial sequence and homogeneity of literary style. The American Statistician, 59(1):19–30, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gaudard:1991:UCA

[GH91] Marie Gaudard and Gerald J.

Hahn. An undergraduate concentration in applied statistics for mathematics majors. The American Statistician, 45(2): 115–??, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:1996:WDI

[Gho96] Malay Ghosh. Wishart distribution via induction. The American Statistician, 50(3):

243–246, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:2002:PIR

[Gho02]  B. K. Ghosh.  Probabil-

ity inequalities related to Markov’s theorem. The American Statistician, 56(3):186– ??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s4/p186.

Gibbons:1990:UIR

[Gib90] Jean D. Gibbons. U.S. institutional representation on editorial boards of U.S statistics journals. The American Statistician, 44(3):210–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gill:2000:LGR

[Gil00] Paramjit S. Gill. Late-game reversals in professional basketball, football, and hockey. The American Statistician, 54 (2):94–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Gill.htm.

Gilbert:2008:PGE

[Gil08] Gregory E. Gilbert. A pocket guide to epidemiology. The American Statistician, 62(1):

92–93, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gitlow:1994:CJT

[Git94] Howard S. Gitlow. A comparison of Japanese total quality control and Dr Deming’s theory of management. The American Statistician, 48(3):197–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gonen:2005:BTS

[GJLW05] Mithat Gönen, Wesley O. Johnson, Yonggang Lu, and Peter H. Westfall. The Bayesian twosample t test. The American Statistician, 59(3):252–257, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gautam:1999:SRM

[GK99] Shiva Gautam and George Kimeldorf. Some results on the maximal correlation in 2 × k contingency tables. The American Statistician, 53(4):336–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gastwirth:1994:HCA

[GKR94] Joseph Gastwirth, Abba Krieger, and Paul Rosenbaum. How a court accepted an impossible explanation. The American Statistician, 48(4):313–??, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gastwirth:1997:RHH

[GKR97a] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Reply: Hypotheticals and hypotheses. The American Statistician, 51(2):120–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gastwirth:1997:RUI

[GKR97b] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Reply: Unquestionably impossible. The American Statistician, 51(2):115–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gan:1991:PPD

[GKT91] F. F. Gan, Kenneth J. Koehler, and John C. Thompson. Probability plots and distribution curves for assessing the fit of probability models. The American Statistician, 45(1):14–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goodall:1998:CUR

[GKT98] Colin R. Goodall, Karen Kafadar, and John W. Tukey. Computing and using rural versus urban measures in statistical applications. The American Statistician, 52(2): 101–111, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/goodall.
html;   http://www.jstor.
org/stable/2685467.

Ganju:2000:AUD

[GL00] Jitendra Ganju and James M.

Lucas. Analysis of unbalanced data from an experiment with random block effects and unequally spaced factor levels. The American Statistician, 54 (1):??, February 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Glass:2004:PRC

[Gla04] D. Glass. Proofs that really count: The art of combinatorial proof. The American Statistician, 58(4):360, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gleser:1986:SNR

[Gle86] Leon J. Gleser. Some notes on refereeing. The American Statistician, 40(4):310–312, 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gleser:1989:GDM

[Gle89] Leon Jay Gleser. The gamma distribution as a mixture of exponential distributions. The American Statistician, 43(2): 115–117, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Genest:1986:JCB

[GM86] Christian Genest and Jock MacKay. The joy of copulas: bivariate distributions with uniform marginals. The American Statistician, 40(4):280–283, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:1991:NBD

[GM91] Andrew Gelman and Xiao-Li Meng. A note on bivariate distributions that are conditionally normal. The American Statistician, 45(2):125–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gunst:1996:IOD

[GM96] Richard F. Gunst and Gary C. McDonald. The importance of outcome dynamics, simple geometry, and pragmatic statistical arguments in exposing deficiencies of experimental design strategies. The American Statistician, 50(1):44–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2002:YCL

[GN02] Andrew Gelman and Deborah Nolan. You can load a die, but you can’t bias a coin. The American Statistician, 56(4): 308–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.

ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s8/p308.

Gelman:1998:SPS

[GNM^{+}98] Andrew Gelman, Deborah Nolan, Anna Men, Steve Warmerdam, and Michelle Bautista. Student projects on statistical literacy and the media. The American Statistician, 52(2):160–??, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goddard:1991:CSC

[God91] M. J. Goddard. Constructing some categorical anomalies. The American Statistician, 45 (2):129–134, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Golbeck:1986:PAC

[Gol86] Amanda L. Golbeck. Probabilistic approaches to current life table estimation. The American Statistician, 40(3): 185–190, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldman:1992:EVS

[Gol92] Anne I. Goldman. EVENTCHARTS:

Visualizing survival and other timed-events data. The American Statistician, 46(1):13–??, February 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldstein:1997:SENa

[Gol97a] Richard Goldstein. Section Editor’s notes. The American Statistician, 51(1):65–??, February 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldstein:1997:SENb

[Gol97b] Richard Goldstein. Section Editor’s notes. The American Statistician, 51(2):193–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldman:1998:BRI

[Gol98] Robert N. Goldman. Book review: Interactive Statistics: Preliminary Edition. The American Statistician, 52(3): 283–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goldstein:1999:BRS

[Gol99] Richard Goldstein. Book review: Statistics with Stata 5 . The American Statistician, 53(4):392–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gonzalez:1994:IDQ

[Gon94] Maria Elena Gonzalez. Improving data quality awareness in the United States Federal Statistical Agencies. The American Statistician, 48(1):12–??, February 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gonzalez:2003:BRA

[Gon03] Richard Gonzalez. Book review:

Applied Multivariate Statistics for the Social Sciences (4th ed.). The American Statistician, 57 (1):68–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s15/p68.

Good:1965:LE

[Goo65] I. J. Good. Letter to the editor. The American Statistician, 19(3):43, June 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2681423.

Good:1969:HRR

[Goo69] I. J. Good. How random are random numbers? The American Statistician, 23(4): 42–45, October 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2681742.

Goodnight:1979:TSO

[Goo79] James H. Goodnight. A tutorial on the SWEEP operator. The American Statistician, 33 (??):149–158, ???? 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gorroochurn:2011:EPH

[Gor11] Prakash Gorroochurn. Errors of probability in historical context. The American Statistician, 65(4):246–254, November 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gumpertz:1989:SAI

[GP89] Marcia Gumpertz and Sastry G. Pantula. A simple approach to inference in random coefficient models. The American Statistician, 43(4):203–210, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gelman:2009:SPU

[GP09] Andrew Gelman and David K. Park. Splitting a predictor at the upper quarter or third and the lower quarter or third. The American Statistician, 63(1):1– 8, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Goel:1997:CAT

[GPA97] Prem K. Goel, Mario Peruggia, and Baoshe An. Computeraided teaching of probabilistic modeling for biological phenomena. The American Statistician, 51(2):164–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/goel.html.

Gelman:2002:LPW

[GPD02] Andrew Gelman, Cristian Pasarica, and Rahul Dodhia. Let’s practice what we preach:

Turning tables into graphs. The American Statistician, 56 (2):121–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s8/p121.

Grimmett:1996:EVD

[GR96] D. R. Grimmett and J. R. Ridenhour. The effect of a variable data point on hypothesis tests for means. The American Statistician, 50(2):145–150, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gervini:2004:CED

[GR04] Daniel Gervini and Valentin Rousson. Criteria for evaluating dimension-reducing components for multivariate data. The American Statistician, 58 (1):72–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s14/p72.

Gibbs:2009:CWS

[GR09] Alison Gibbs and Nancy Reid.

Comment: What is statistics? The American Statistician, 63 (2):112–113, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Grambsch:2008:RMM

[Gra08] Patricia Grambsch. Regression to the mean, murder rates, and shall-issue laws. The American Statistician, 62(4):289–295, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greenland:1991:LJC

[Gre91] Sander Greenland. On the logical justification of conditional tests for two-by-two contingency tables. The American Statistician, 45(3):248–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greenland:1996:LBC

[Gre96] Sander Greenland. A lower bound for the correlation of exponentiated bivariate normal pairs. The American Statistician, 50(2):163–164, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Greenland:2010:SPA

[Gre10] Sander Greenland. Simpson’s Paradox from adding constants in contingency tables as an example of Bayesian noncollapsibility. The American Statistician, 64(4):340–344, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Groeneveld:1990:RTL

[Gro90] Richard A. Groeneveld. Ranking teams in a league with two divisions of t teams. The American Statistician, 44(4):277–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Groeneveld:1991:IFA

[Gro91] Richard A. Groeneveld. An influence function approach to describing the skewness of a distribution. The American Statistician, 45(2):97–102, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Groeneveld:1998:CQM

[Gro98] Richard A. Groeneveld. A class of quantile measures for kurtosis. The American Statistician, 52(4):325–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/groen.html.

Groeneveld:1999:D

[Gro99] Richard A. Groeneveld. Discussion. The American Statistician, 53(1):23–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gromping:2007:ERI

[Grö07] Ulrike Grömping. Estimators of relative importance in linear regression based on variance decomposition. The American Statistician, 61(2):139–147, May 2007. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Gromping:2009:VIA

[Grö09] Ulrike Grömping. Variable importance assessment in regression: Linear regression versus random forest. The American Statistician, 63(4):308–319, November 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Groggel:1986:DFT

[GS86] David J. Groggel and John H. Skillings. Distribution-free tests for main effects in multifactor designs. The American Statistician, 40(2):99–102, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ghosh:2002:SDW

[GS02] Malay Ghosh and Bimal K. Sinha. A simple derivation of the Wishart distribution. The American Statistician, 56 (2):100–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s4/p100.

Gelman:2006:DBS

[GS06] Andrew Gelman and Hal Stern. The difference between “significant” and “not significant” is not itself statistically significant. The American Statistician, 60(4):328–331, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gajewski:2008:OHT

[GS08] Byron J. Gajewski and Stephen D.

Simon. A one-hour training seminar on Bayesian statistics for nursing graduate students. The American Statistician, 62 (3):190–194, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gbur:1995:KWP

[GT95] Edward E. Gbur, Jr. and Bruce E. Trumbo. Key words and phrases — the key to scholarly visibility and efficiency in an information explosion. The American Statistician, 49(1): 29–??, ???? 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guenther:1979:UNA

[Gue79] William C. Guenther. The use of noncentral F approximations for calculation of power and sample size. The American Statistician, 33(??):209– 210, 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guenther:1987:SIT

[Gue87] William C. Guenther. Statistical inferences for two oneparameter families. The American Statistician, 41(1):50–52, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Guerra:2006:DDC

[Gue06] Rudy Guerra. Dicing with death: Chance, risk, and health. The American Statistician, 60(2):204, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gunter:1993:TFS

[Gun93] Bert Gunter. Through a funnel slowly with ball bearing and insight to teach experimental design. The American Statistician, 47(4):265–??, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gunter:1998:VRG

[Gun98] Berton Gunter. Visual revelations: Graphical tales of fate and deception from Napoleon Bonaparte to Ross Perot. The American Statistician, 52(1): 83–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gupta:2002:BRH

[Gup02] Ghanshyam Gupta. Book review: A Handbook of Statistical Analyses using SAS, 2nd ed. The American Statistician, 56(4):334–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n4/s22/p334.

Gurka:2006:SBL

[Gur06] Matthew J. Gurka. Selecting the best linear mixed model under REML. The American Statistician, 60(1):19–26, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gurwit:2008:LSE

[Gur08] Margie R. Gurwit. Learning SAS by example: A programmer’s guide. The American Statistician, 62(2):182–183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Graziani:2009:HCM

[GV09] Rebecca Graziani and Piero Veronese. How to compute a mean? the Chisini approach and its applications. The American Statistician, 63(1):33–36, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gray:1994:MSS

[GW94] J. Brian Gray and William H.

Woodall. The maximum size of standardized and internally studentized residuals in regression analysis. The American Statistician, 48(2):111–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Gan:1989:EPO

[GY89] Gaoxiong Gan and Shie-Shien Yang. An elementary proof of the order-statistics characterization of the Poisson process. The American Statistician, 43 (1):45–46, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [GY90].

Gan:1990:CEP

[GY90] Gaoxiong Gan and Shie-Shien Yang. Correction: “An elementary proof of the order-statistics characterization of the Poisson process” [Amer. Statist. 43 (1989), no. 1, 45–46]. The American Statistician, 44(1): 65, ???? 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [GY89].

Garfield:2011:RAS

[GZK^{+}11] Joan Garfield, Andrew Zieffler, Daniel Kaplan, George W. Cobb, Beth L. Chance, and John P. Holcomb. Rethinking assessment of student learning in statistics courses. The American Statistician, 65(1): 1–10, February 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hagwood:2009:ARC

[Hag09] Charles Hagwood. An application of the residue calculus: The distribution of the sum of nonhomogeneous gamma variates. The American Statistician, 63 (1):37–39, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:1995:DII

[Hah95] Gerald J. Hahn. Deming’s impact on industrial statistics: Some reflections. The American Statistician, 49(4):336–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:2002:DPS

[Hah02] Gerald J. Hahn. Deming and the proactive statistician. The American Statistician, 56(4): 290–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s5/p290.

Haigh:2001:RBT

[Hai01]  John Haigh.  Reviews of

books and teaching materials — mathematics of chance. The American Statistician, 55(4): 372–??, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haigh:2003:BRR

[Hai03] John Haigh. Book review: Risk-Benefit Analysis (2nd ed.). The American Statistician, 57(1): 69–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s16/p69.

Haigh:2007:PMA

[Hai07] John Haigh. Probability: Modeling and applications to random processes. The American Statistician, 61(3):276–277, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hallahan:1992:DCD

[Hal92] Charles Hallahan. DBMS/ COPY and DBMS/COPY Plus — (Version 2.0). The American Statistician, 46(1):49–??, February 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hall:2008:ARO

[Hal08] Daniel B. Hall. Analyzing receiver operating characteristic curves with SAS. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hall:2010:RFG

[Hal10] Nancy S. Hall. Ronald Fisher and Gertrude Cox: Two statistical pioneers sometimes cooperate and sometimes collide. The American Statistician, 64 (3):212–220, August 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamedani:1984:NLC

[Ham84] G. G. Hamedani. Nonnormality of linear combinations of normal random variables. The American Statistician, 38(4):

295–296, ???? 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamilton:1987:SCV

[Ham87] David Hamilton. Sometimes R^{2} > r_{yx_{1}}^{2} + r_{yx_{2}}^{2}. correlated variables are not always redundant. The American Statistician, 41(2):129–132, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hammer:1997:SES

[Ham97] Heather Hammer. Symbolic exclusion in statistical literature: The impact of gendered language. The American Statistician, 51(1):13–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/hammer.html.

Hamer:2000:SP

[Ham00] Robert M. Hamer. Statistics in psychiatry. The American Statistician, 54(4):322–??, November 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Han:1979:AME

[Han79] Chien Pai Han. Alternative methods of estimating the likelihood ratio in classification of multivariate normal observations. The American Statistician, 33(4):204–206, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hand:1992:CTT

[Han92a] David J. Hand. On comparing two treatments. The American Statistician, 46(3):190–??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hanley:1992:JCD

[Han92b] James A. Hanley. Jumping to coincidences: Defying odds in the realm of the preposterous. The American Statistician, 46 (3):197–??, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hand:1998:DMS

[Han98] David J. Hand. Data mining: Statistics and more? The American Statistician, 52 (2):112–??, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/hand.html.

Hanley:2004:TWM

[Han04] James A. Hanley. “transmuting” women into men. The American Statistician, 58(3): 237–243, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hansen:2005:USW

[Han05] John. Hansen. Using SPSS for Windows and Macintosh: Analyzing and understanding data.

The American Statistician, 59 (1):113, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harvey:1976:APG

[Har76] A. C. Harvey. An alternative proof and generalization of a test for structural change. The American Statistician, 30(??): 122–123, 1976. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harter:1980:MAF

[Har80] H. Leon Harter. Modified asymptotic formulas for critical values of the Kolmogorov test statistic. The American Statistician, 34(2):110–111, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardin:2000:LW

[Har00] James W. Hardin. LogXact 4.1 for Windows. The American Statistician, 54(4):320–??, November 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harville:2004:DSC

[Har04] David A. Harville. Discussion:

Statistics and the college football championship. The American Statistician, 58(3):187–189, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardin:2006:IST

[Har06] Jo Hardin. Introduction to statistics through resampling methods and R/S-PLUS. The American Statistician, 60(4): 343–344, November 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harvill:2008:FPC

[Har08] Jane L. Harvill. Fundamental probability: A computational approach. The American Statistician, 62(2):179–180, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hastie:1987:CLD

[Has87] Trevor Hastie. A closer look at the deviance. The American Statistician, 41(1):16–20, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayden:1999:BRI

[Hay99] Robert W. Hayden. Book review: Introduction to Statistical Reasoning. The American Statistician, 53(1):86, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayes:2011:GIA

[Hay11] Kevin Hayes. A geometrical interpretation of an alternative formula for the sample covariance. The American Statistician, 65(2):110–112, May 2011.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hazelton:2003:GTA

[Haz03] Martin L. Hazelton. A graphical tool for assessing normality. The American Statistician, 57(4):285–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s12/p285.

Heitjan:1996:DMR

[HB96] Daniel F. Heitjan and Srabashi Basu. Distinguishing “missing at random” and “missing completely at random”. The American Statistician, 50(3):207–213, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hubbard:2003:CME

[HB03] Raymond Hubbard and M. J.

Bayarri. Confusion over measures of evidence (p’s) versus errors (α’s) in classical statistical testing. The American Statistician, 57(3): 171–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s5/p171.

Harris:2005:MRI

[HB05a] Ian R. Harris and Brent D. Burch. Measuring relative importance of sources of variation without using variance. The American Statistician, 59 (3):217–222, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hubbard:2005:LE

[HB05b] Raymond. Hubbard and M. J. Bayarri. Letters to the Editor. The American Statistician, 59(4):353–354, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haslett:1991:DGE

[HBC^{+}91] John Haslett, Ronan Bradley, Peter Craig, Antony Unwin, and Graham Wills. Dynamic graphics for exploring spatial data with application to locating global and local anomalies. The American Statistician, 45 (3):234–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:1990:HSQ

[HBF90] Gerald J. Hahn, Thomas J. Boardman, and David Fluharty. History of the section on quality and productivity of the American Statistical Association. The American Statistician, 44(3): 197–199, ???? 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2004:URT

[HBQ04] Nicholas J. Horton, Elizabeth R. Brown, and Linjuan Qian. Use of R as a toolbox for mathematical statistics exploration. The American Statistician, 58(4):343–357, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hin:2007:CWC

[HCW07] Lin-Yee Hin, Vincent J. Carey, and You-Gan Wang. Criteria for working–correlation– structure selection in GEE. The American Statistician, 61(4): 360–364, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hillis:1994:SJI

[HD94] Stephen L. Hillis and Charles S.

Davis. A simple justification of the iterative fitting procedure for generalized linear models. The American Statistician, 48(4):288–289, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haughton:2003:RSP

[HDE^{+}03] Dominique Haughton, Joel Deichmann, Abdolreza Eshghi, Selin Sayek, Nicholas Teebagy, and Heikki Topi. A review of software packages for data mining. The American Statistician, 57(4):290–??, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s14/p290.

He:1997:QCC

[He97] Xuming He. Quantile curves without crossing. The American Statistician, 51(2):186–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

                                        www.amstat.org/publications/
                                        tas/abstracts/he.html.

Heiberger:1981:SED

[Hei81] Richard M. Heiberger. The specification of experimental designs to ANOVA programs. The American Statistician, 35 (??):98–108, 1981. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harper:2011:CAM

[HEJ11] William V. Harper, Ted G. Eschenbach, and Thomas R. James. Concerns about maximum likelihood estimation for the three-parameter Weibull distribution: Case study of statistical software. The American Statistician, 65(1):44–54, February 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Helms:1988:MSH

[Hel88] Ronald W. Helms. Manipulating statistical hypotheses and tests as Boolean functions. The American Statistician, 42(4):

253–256, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Helland:1995:SCA

[Hel95] Inge S. Helland. Simple counterexamples against the conditionality principle. The American Statistician, 49(4):351–356, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Helberg:1999:FSO

[Hel99] Clay Helberg. Finding statistics online: How to locate the elusive numbers you need. The American Statistician, 53 (3):293–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hengartner:1999:NML

[Hen99] Nicolas W. Hengartner. A note on maximum likelihood estimation. The American Statistician, 53(2):123–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Herr:1980:HUG

[Her80] David G. Herr. On the history of the use of geometry in the general linear model. The American Statistician, 34 (1):43–47, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Herr:1986:HAU

[Her86] David G. Herr. On the history of ANOVA in unbalanced, factorial designs: the first 30 years.

The American Statistician, 40 (4):265–270, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyndman:1996:SQS

[HF96] Rob J. Hyndman and Yanan Fan. Sample quantiles in statistical packages. The American Statistician, 50(4):361–??, November 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hatzinikolaou:2004:RSS

[HF04] Dimitris Hatzinikolaou and Kosmas Ferentinos. On the role of semielasticity in statistics. The American Statistician, 58 (2):150–151, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Heller:2008:BD

[HFDJ08] Gillian Z. Heller, Andrew B. Forbes, Keith B. G. Dear, and Erica Jobling. Biostatistics @ distance. The American Statistician, 62(4):321–328, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harrell:1997:SMS

[HG97] Frank E. Harrell, Jr. and Richard Goldstein. A survey of microcomputer survival analysis software: The need for an integrated framework. The American Statistician, 51(4): 360–??, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/harrell.html.

Hilton:2001:TCS

[HGA01] Sterling C. Hilton, Scott D.

Grimshaw, and Genan T. Anderson. Teacher’s corner — statistics in preschool. The American Statistician, 55(4): 332–336, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayes:1999:SGL

[HH99] Kevin Hayes and John Haslett.

Simplifying general least squares. The American Statistician, 53 (4):376–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoenig:2001:APP

[HH01] John M. Hoenig and Dennis M.

Heisey. The abuse of power: The pervasive fallacy of power calculations for data analysis. The American Statistician, 55 (1):19–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Hoenig.htm.

Hansen:2004:PNS

[HH04] Morris H. Hansen and William N.

Hurwitz. The problem of nonresponse in sample surveys. The American Statistician, 58(4): 292–294, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardin:2006:RSW

[HH06] James W. Hardin and Joseph M. Hilbe. A review of Scientific WorkPlace 5.5. The American Statistician, 60(1):93–96, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hill:1991:LRA

[HHB91] L. Robert Hill, William G. Hammond, and John R. Benfield. Logistic regression analysis in experimental bronchial carcinogenesis. The American Statistician, 45(3):184–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:1999:ISS

[HHHZ99] Gerald J. Hahn, William J. Hill, Roger W. Hoerl, and Stephen A. Zinkgraf. The impact of six sigma improvement— A glimpse into the future of statistics. The American Statistician, 53(3):208–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:1993:SIS

[HHJL93] Roger W. Hoerl, Jeffrey H. Hooper, Peter J. Jacobs, and James M. Lucas. Skills for industrial statisticians to survive and prosper in the emerging quality environment. The American Statistician, 47(4):

280–??, November 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hothorn:2006:LSC

[HHvdWZ06] Torsten Hothorn, Kurt Hornik, Mark A. van de Wiel, and Achim Zeileis. A Lego system for conditional inference. The American Statistician, 60 (3):257–263, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hidiroglou:1986:CSR

[Hid86] M. A. Hidiroglou. The construction of a self-representing stratum of large units in survey design. The American Statistician, 40(1):27–31, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Higgins:1999:LE

[Hig99a] James J. Higgins. Letters to the editor. The American Statistician, 53(2):170–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Higgins:1999:NSN

[Hig99b] James J. Higgins. Nonmathematical statistics: A new direction for the undergraduate discipline. The American Statistician, 53(1):1–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Higgins:1999:R

[Hig99c] James J. Higgins. Reply. The American Statistician, 53(1): 26–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hildebrand:1984:LMP

[Hil84] David K. Hildebrand. Lord’s MAD paradox and Jensen’s inequality. The American Statistician, 38(4):296–297, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hill:1987:VLP

[Hil87] Bruce M. Hill. The validity of the likelihood principle. The American Statistician, 41 (2):95–100, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1993:GAM

[Hil93] Joseph M. Hilbe. Generalized additive models software. The American Statistician, 47 (1):59–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1994:GLM

[Hil94] Joseph M. Hilbe. Generalized linear models. The American Statistician, 48(3):255–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1996:WFC

[Hil96a] Joseph Hilbe. Windows file conversion software. The American Statistician, 50(3):268–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hillmer:1996:PSA

[Hil96b] Steven C. Hillmer. A problemsolving approach to teaching business statistics. The American Statistician, 50(3):249–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1997:SENa

[Hil97a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 51(3):284–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1997:SENb

[Hil97b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 51(4):359–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/hilbe.html.

Hilbe:1998:SEN

[Hil98] Joseph Hilbe. Section Editor’s notes. The American Statistician, 52(1):70–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:1999:SEN

[Hil99] Joseph Hilbe. Section Editor’s notes. The American Statistician, 53(2):147–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2000:SEN

[Hil00] Joseph Hilbe. Section Editor’s notes. The American Statistician, 54(4):319–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2001:RBT

[Hil01]  Joseph Hilbe.  Reviews of

books and teaching materials — XploRe: Learning guide. The American Statistician, 55 (2):169–170, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2002:SENa

[Hil02a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 56(2):148–??, May 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n2/s11/p148.

Hilbe:2002:SENb

[Hil02b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 56(3):234– ??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s12/p234.

Hilbe:2002:SCG

[Hil02c] Joseph Hilbe. Statistical computing and graphics — statistical computing software reviews — Section Editor’s notes. The American Statistician, 56 (1):71, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2002:SCS

[Hil02d] Joseph Hilbe. Statistical computing software reviews: Section Editor’s notes. The American Statistician, 56(4):325– ??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s11/p325.

Hilbe:2002:XEP

[Hil02e] Joseph M. Hilbe. XPro 5.1:

Exact procedures for parametric statistical inference. The American Statistician, 56(4): 326–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s12/p326.

Hilbe:2003:SENa

[Hil03a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 57(3):200– ??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n3/s13/p200.

Hilbe:2003:SENb

[Hil03b] Joseph Hilbe. Section Editor’s notes. The American Statistician, 57(4):289– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s13/p289.

Hilbe:2003:RCS

[Hil03c] Joseph M. Hilbe. A review of current SPSS products: SPSS 12, SigmaPlot 8.02, SigmaStat 3.0, Part 1. The American Statistician, 57(4):310– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n4/s15/p310.

Hilbe:2004:SENa

[Hil04a] Joseph Hilbe. Section Editor’s notes. The American Statistician, 58(1):85– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s16/p85.

Hilbe:2004:RSP

[Hil04b] Joseph M. Hilbe. A review of SPSS 12.01, part 2. The American Statistician, 58(2):168–171, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2004:SENb

[Hil04c]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 58(2):159, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2004:SENc

[Hil04d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 58(3):253, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2004:SENd

[Hil04e]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 58(4):328–329, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2004:SCS

[Hil04f] Joseph M. Hilbe. Statistical computing software reviews.

The American Statistician, 58 (2):92, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSb

[Hil05a] Joseph M. Hilbe. Review of SigmaPlot 9.0. The American Statistician, 59(1):111–112, February 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSP

[Hil05b] Joseph M. Hilbe. A review of SPSS, part 3. The American Statistician, 59(2):185–186, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSc

[Hil05c] Joseph M. Hilbe. A review of Stata 9.0. The American Statistician, 59(4):335–348, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:RSa

[Hil05d] Joseph M. Hilbe. A review of Systat 11. The American Statistician, 59(1):104–110, February 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENa

[Hil05e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(1):88, February 2005. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENb

[Hil05f]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 59(2):183–184, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENc

[Hil05g] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(3):273, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:SENd

[Hil05h] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 59(4):334, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2005:S

[Hil05i] Joseph M. Hilbe. SigmaStat 3.1. The American Statistician, 59(2):187–191, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:M

[Hil06a] Joseph M. Hilbe. Mathematica 5.2. The American Statistician, 60(2):176–186, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:RLN

[Hil06b] Joseph M. Hilbe. A review of LIMDEP 9.0 and NLOGIT 4.0. The American Statistician, 60 (2):187–202, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENa

[Hil06c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(1):92, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENb

[Hil06d]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 60(2):175, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENc

[Hil06e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(3):271, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:SENd

[Hil06f] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 60(4):340, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:EPC

[Hil07a] Joseph M. Hilbe. ePrint 5 professional conversion utility. The American Statistician, 61 (2):179–180, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:G

[Hil07b] Joseph M. Hilbe. GenStat 9.

The American Statistician, 61 (3):269–273, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENa

[Hil07c]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 61(2):178, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENb

[Hil07d] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 61(3):260–261, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SENc

[Hil07e]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 61(4):365–366, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:S

[Hil07f] Joseph M. Hilbe. STATIS-TICA 7. The American Statistician, 61(1):91–94, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2007:SCS

[Hil07g] Joseph M. Hilbe. Statistical computing software reviews. The American Statistician, 61 (1):78, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:HLS

[Hil08a] Joseph M. Hilbe. Hawkes learning systems. The American Statistician, 62(1):85–89, February 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENd

[Hil08b] Joseph M. Hilbe. Section Editor’s note. The American Statistician, 62(4):355, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENa

[Hil08c] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 62(1):73, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENb

[Hil08d]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 62(2):171–172, May 2008. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:SENc

[Hil08e] Joseph M. Hilbe. Section Editor’s notes. The American Statistician, 62(3):267, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2008:S

[Hil08f] Joseph M. Hilbe. Systat 12.2.

The American Statistician, 62 (2):177–178, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2009:SEN

[Hil09]  Joseph M. Hilbe.  Section

Editor’s notes. The American Statistician, 63(1):75–77, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hines:1996:PPA

[Hin96] W. G. S. Hines. Pragmatics of pooling in ANOVA tables. The American Statistician, 50 (2):127–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hitchcock:2003:HMH

[Hit03]  David B. Hitchcock.  A

history of the Metropolis– Hastings algorithm. The American Statistician, 57(4):254– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.

ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s7/p254.

Hanley:2008:S

[HJM08] James A. Hanley, Marilyse Julien, and Erica E. M. Moodie. Student’s z, t, and s. The American Statistician, 62(1): 64–69, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hanley:2001:TCV

[HJP^{+}01] James A. Hanley, Lawrence Joseph, Platt, Robert, et al. Teacher’s corner — visualizing the median as the minimumdeviation location. The American Statistician, 55(2):150–152, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyde:2006:ICO

[HJS06] Valerie Hyde, Wolfgang Jank, and Galit Shmueli. Investigating concurrency in online auctions through visualization. The American Statistician, 60 (3):241–250, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayden:1992:PHS

[HK92] Robert W. Hayden and Farid Kianifard. Preparing high school mathematics teachers to teach statistics. The American Statistician, 46(4):290–??, November 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2007:MAA

[HK07] Nicholas J. Horton and Ken P.

Kleinman. Much ado about nothing. The American Statistician, 61(1):79–90, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haughton:2006:RTD

[HKS06] Dominique Haughton, Arnold Kamis, and Patrick Scholten. A review of three directed acyclic graphs software packages. The American Statistician, 60(3): 272–286, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hayter:1996:NCP

[HL96] A. J. Hayter and W. Liu. A note on the calculation of Pr(X_{1} < X_{2} ⋯ < X_{k}). The American Statistician, 50(4):365–??, November 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hadi:1998:SCN

[HL98] Ali S. Hadi and Robert F. Ling.

Some cautionary notes on the use of principal components regression. The American Statistician, 52(1):15–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/hadi.html.

Horton:1999:RSF

[HL99] Nicholas J. Horton and Stuart R. Lipsitz. Review of software to fit generalized estimating equation regression models. The American Statistician, 53(2):160–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/horton.pdf.

Horton:2001:SCG

[HL01] Nicholas J. Horton and Stuart R. Lipsitz. Statistical computing and graphics — statistical computing software reviews — multiple imputation in practice: Comparison of software packages for regression models with missing variables. The American Statistician, 55(3): 244–254, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hunter:2004:TMA

[HL04] David R. Hunter and Kenneth Lange. A tutorial on MM algorithms. The American Statistician, 58(1):30– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s9/p30.

Hirschberg:2010:GCD

[HL10] Joe Hirschberg and Jenny Lye. A geometric comparison of the delta and Fieller confidence intervals. The American Statistician, 64(3):234–241, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Horton:2003:PBW

[HLP03] Nicholas J. Horton, Stuart R.

Lipsitz, and Michael Parzen. A potential for bias when rounding in multiple imputation. The American Statistician, 57(4): 229–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s1/p229.

Haughton:2009:RTL

[HLW09] Dominique Haughton, Pascal Legrand, and Sam Woolford. Review of three latent class cluster analysis packages: Latent Gold, poLCA, and MCLUST. The American Statistician, 63(1):81–91, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hahn:1993:ASI

[HM93] Gerald J. Hahn and William Q.

Meeker. Assumptions for statistical inference. The American Statistician, 47(1):1–??, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Henderson:2001:TCE

[HM01] Michael Henderson and Mary C.

Meyer. Teacher’s corner — exploring the confidence interval for a binomial parameter in a first course in statistical computing. The American Statistician, 55(4):337–344, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamada:2001:SPF

[HMR^{+}01] M. Hamada, H. F. Martz, C. S. Reese, et al. Statistical practice — finding nearoptimal Bayesian experimental designs via genetic algorithms. The American Statistician, 55 (3):175–181, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hintze:1998:VPB

[HN98] Jerry L. Hintze and Ray D. Nelson. Violin plots: A box plotdensity trace synergism. The American Statistician, 52(2): 179–?? (or 181–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/hintze.
                                        html.

Hogbin:2010:LE

[HN10] Martin Hogbin and W. Nijdam. Letter to the Editor. The American Statistician, 64 (2):193–194, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:1985:RAY

[Hoe85] Roger W. Hoerl. Ridge analysis 25 years later. The American Statistician, 39(3):186–192, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:1998:PSP

[Hoe98]  Roger Hoerl.  The pocket

statistician: A practical guide to quality improvement. The American Statistician, 52(1): 84–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogg:1979:SRO

[Hog79] Robert V. Hogg. Statistical robustness: one view of its use in applications today. The American Statistician, 33(3):108–115, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogg:1991:SEI

[Hog91] Robert V. Hogg. Statistical education: Improvements are badly needed. The American Statistician, 45(4):342–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogan:1992:PES

[Hog92] Howard Hogan. The 1990 post-enumeration survey: An overview. The American Statistician, 46(4):261–??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogg:1994:CSE

[Hog94] Robert V. Hogg. A core in statistics for engineering students. The American Statistician, 48(4):285–??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogg:1999:LUC

[Hog99a] Robert V. Hogg. Let’s use CQI in our statistics programs. The American Statistician, 53(1):7– ??, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hogg:1999:R

[Hog99b] Robert V. Hogg. Reply. The American Statistician, 53(1): 27–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Holcomb:1999:BRA

[Hol99]  John P. Holcomb.  Book

reviews: Applied Regression Analysis, 3rd edition and Applied Regression Analysis: A Research Tool, 2nd edition. The American Statistician, 53(2): 170–??, May 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Holt:2007:OSO

[Hol07] D. Tim Holt. The official statistics Olympic challenge. The American Statistician, 61(1):1– 8, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hombas:1997:WTE

[Hom97] V. C. Hombas. Waiting time and expected waiting time:

Paradoxical situations. The American Statistician, 51(2): 130–133, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/hombas.html.

Horton:2006:MLM

[Hor06] Nicholas J. Horton. Multilevel and longitudinal modeling using stata. The American Statistician, 60(3):293–294, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hosking:1992:MME

[Hos92] J. R. M. Hosking. Moments or L moments? an example comparing two measures of distributional shape. The American Statistician, 46(3):186–??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Houghton:1997:PEF

[Hou97] Dominique Houghton. Packages for estimating finite mixtures: A review. The American Statistician, 51(2):194–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/haughton.html.

Ho:2008:IPI

[HQ08] Daniel E. Ho and Kevin M.

Quinn. Improving the presentation and interpretation of online ratings data with model-based figures. The American Statistician, 62(4):279–288, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hedayat:1998:EUS

[HR98] A. S. Hedayat and Weining Z. Robieson. Exclusion of an undesirable sample from the support of a simple random sample. The American Statistician, 52 (1):41–43, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/hedayat.
html.

Holcomb:2000:UTL

[HR00] John P. Holcomb, Jr. and Rochelle L. Ruffer. Using a term-long project sequence in introductory statistics. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harkness:2005:TGS

[HR05] W. L. Harkness and J. L. Rosenberger. Training graduate students at Penn State University in teaching statistics. The American Statistician, 59 (1):11–13, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

He:2006:TDM

[HR06] Yulei He and Trivellore E. Raghunathan. Tukey’s gh distribution for multiple imputation. The American Statistician, 60(3):251–256, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hodges:2010:ASC

[HR10] James S. Hodges and Brian J.

Reich. Adding spatiallycorrelated errors can mess up the fixed effect you love. The American Statistician, 64(4): 325–334, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Heller:2010:UCM

[HRS10] Ruth Heller, Paul R. Rosenbaum, and Dylan S. Small. Using the cross-match test to appraise covariate balance in matched pairs. The American Statistician, 64(4):299–309, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Haunsperger:1991:LCS

[HS91] Deanna B. Haunsperger and Donald G. Saari. The lack of consistency for statistical decision procedures. The American Statistician, 45(3):252–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hettmansperger:1992:CNM

[HS92] Thomas P. Hettmansperger and Simon J. Sheather. A cautionary note on the method of least median squares. The American Statistician, 46(2):79–83, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harville:1994:HCA

[HS94] David A. Harville and Michael H. Smith. The home-court advantage: How large is it and does it vary from team to team? The American Statistician, 48 (1):22–??, February 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hedayat:2003:SDE

[HS03] A. S. Hedayat and Bikas K. Sinha. On a sampling design for estimation of negligible accident rates involving electronic toys. The American Statistician, 57(4):249– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s5/p249.

Hamada:2004:R

[HS04a] Michael Hamada and Randy Sitter. Response. The American Statistician, 58(3):202, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hamada:2004:SR

[HS04b] Michael Hamada and Randy Sitter. Statistical research. The American Statistician, 58 (2):93–101, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hedayat:2009:CWS

[HS09] Samad Hedayat and John Stufken. Comment: What is statistics? The American Statistician, 63(2):115–116, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoerl:2010:MSP

[HS10] Roger W. Hoerl and Ronald D.

Snee. Moving the statistics profession forward to the next level. The American Statistician, 64(1):10–14, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hunter:1994:IC

[HSF^{+}94] J. Stuart Hunter, Stephen M.

Stigler, Stephen E. Fienberg, James Boen, and Joseph B. Kadane. Invited comments. The American Statistician, 48 (2):70–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hsieh:1995:CNA

[Hsi95] Frank Y. Hsieh. A cautionary note on the analysis of extreme data with Cox regression. The American Statistician, 49 (2):226–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hochberg:1983:MCM

[HT83] Yosef Hochberg and Ajit C. Tamhane. Multiple comparisons in a mixed model. The American Statistician, 37(4): 305–307, ???? 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hurvich:1990:IMS

[HT90] Clifford M. Hurvich and Chih-Ling Tsai. The impact of model selection on inference in linear regression. The American Statistician, 44(3):214–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Heiberger:2002:DDC

[HT02] Richard M. Heiberger and Paulo Teles. Displays for direct comparison of ARIMA models. The American Statistician, 56(2):131–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n2/s9/p131.

Hanley:2006:PFR

[HT06] James A. Hanley and Dana Teltsch. The PDF of a function of a random variable. The American Statistician, 60(1): 61–67, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hirji:1989:MUE

[HTM89] Karim F. Hirji, Anastasios A.

Tsiatis, and Cyrus R. Mehta. Median unbiased estimation for binary data. The American Statistician, 43(1):7–11, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hu:1992:EST

[Hu92] Inchi Hu. An elementary solution to a traffic problem. The American Statistician, 46(2): 120–??, May 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hu:2003:BRG

[Hu03] X. Joan Hu. Book review:

Generalized Linear Models: With Applications in Engineering and the Sciences. The American Statistician, 57(1): 67–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s14/p67.

Hu:2007:MAV

[Hu07] Xiaomi Hu. Multivariate analysis without “vec” and “⊗”. The American Statistician, 61 (1):64–66, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hu:2008:TCC

[Hu08] Xiaomi Hu. A three-condition characterization of the Moore– Penrose generalized inverse. The American Statistician, 62 (3):216–218, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hunter:1998:CDI

[Hun98] J. Stuart Hunter. Cuthbert Daniel: Industrial statistician. The American Statistician, 52 (2):89–92, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hunter:2005:TCS

[Hun05] David R. Hunter. Teaching computing in statistical theory courses. The American Statistician, 59(4):327–333, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hutchinson:1993:SRF

[Hut93] T. P. Hutchinson. The seventhroot formula for a trivariate normal probability. The American Statistician, 47(2):102–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Huzurbazar:1999:PSA

[Huz99] S. Huzurbazar. Practical saddlepoint approximations. The American Statistician, 53(3): 225–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hoaglin:1995:CLS

[HV95] David C. Hoaglin and Paul F. Velleman. A critical look at some analyses of major league baseball salaries. The American Statistician, 49(3):277–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hand:2003:LVG

[HV03] David J. Hand and Veronica Vinciotti. Local versus global models for classification problems: Fitting models where it matters. The American Statistician, 57(2):124–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s11/p124.

Hamedani:2009:LE

[HV09] G. G. Hamedani and H. W.

Volkmer. Letter to the Editor. The American Statistician, 63(3):295, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hwang:2009:SPB

[Hwa09] Leng-Cheng Hwang. A simple proof of the binomial theorem using differential calculus. The American Statistician, 63 (1):43–44, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyden:2001:RBT

[Hyd01] Paul Hyden. Reviews of books and teaching materials — statistical quality assurance methods for engineers. The American Statistician, 55(2):167–??, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyndman:1996:CGH

[Hyn96] Rob J. Hyndman. Computing and graphing highest density regions. The American Statistician, 50(2):120–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyung:2001:RBT

[Hyu01] Namwon Hyung. Reviews of books and teaching materials — elements of forecasting, 2nd edition. The American Statistician, 55(4):371, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Inoue:2005:RBB

[IBP05] Lurdes Y. T. Inoue, Donald A. Berry, and Giovanni Parmigiani. Relationship between Bayesian and frequentist sample size determination. The American Statistician, 59(1): 79–87, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ibrahim:1997:PPP

[Ibr97] Joseph G. Ibrahim. On properties of predictive priors in linear models. The American Statistician, 51(4):333–337, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/ibrahim.html.

Iglewicz:1998:SIS

[Igl98] Boris Iglewicz. Selected information on the statistics profession. The American Statistician, 52(4):289–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/iglewicz.
html.

Iglewicz:2007:EFP

[Igl07] Boris Iglewicz. Einstein’s first published paper. The American Statistician, 61(4):339–342, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kern:2007:IRA

[II07]  John C. Kern II.  Intro-

duction to regression analysis. The American Statistician, 61 (1):101, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Iman:2006:SAF

[IJW06] Ronald L. Iman, Mark E. Johnson, and Charles C. Watson, Jr. Statistical aspects of forecasting and planning for hurricanes. The American Statistician, 60 (2):105–121, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Iversen:2001:TCD

[IM01] Philip Iversen and Mervyn Marasinghe. Teacher’s corner — dynamic graphical tools for teaching experimental design and analysis concepts. The American Statistician, 55(4): 345–351, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Imrey:1994:R

[Imr94a] Peter B. Imrey. Reply. The American Statistician, 48(2): 82–??, May 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Imrey:1994:SVQ

[Imr94b] Peter B. Imrey. Statistical values, quality, and certification. The American Statistician, 48 (2):65–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Inlow:2010:MGF

[Inl10] Mark Inlow. A moment generating function proof of the Lindeberg–Lévy Central Limit Theorem. The American Statistician, 64(3):228–230, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Inman:1994:KPR

[Inm94] Henry F. Inman. Karl Pearson and R. A. Fisher on statistical tests: A 1935 exchange from Nature. The American Statistician, 48(1):2–11, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Irizarry:2001:LRM

[Iri01] Rafael A. Irizarry. Local regression with meaningful parameters. The American Statistician, 55(1):72–??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Irizarry.htm.

Jackson:2008:DUS

[Jac08] Christopher H. Jackson. Displaying uncertainty with shading. The American Statistician, 62(4):340–347, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jaditz:2000:DPI

[Jad00] Ted Jaditz. Are the digits of pi an iid sequence? The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:1999:BRA

[Jam99] Sreenivasa Rao Jammalamadaka.

Book review: Against the Gods: The Remarkable Story of Risk. The American Statistician, 53 (2):170–??, May 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2000:SDE

[Jam00] Sreenivasa Rao Jammalamadaka.

Statistical distributions in engineering. The American Statistician, 54(3):226–??, August 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2001:RBTb

[Jam01a] Sreenivasa Rao Jammalamadaka. Reviews of books and teaching materials — statistics in psychology: An historical perspective, second edition. The American Statistician, 55(4):372–??, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2001:RBTa

[Jam01b] Sreenivasa Rao Jammalamadaka. Reviews of books and teaching materials — the nothing that is: A natural history of zero. The American Statistician, 55 (4):372, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jammalamadaka:2002:BRA

[Jam02] Sreenivasa Rao Jammalamadaka. Book review: Applied Nonparametric Statistical Methods, 3rd Edition. The American Statistician, 56(3):251– ??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s20/p251.

Jammalamadaka:2003:BRI

[Jam03] Sreenivasa Rao Jammalamadaka. Book review: Introduction to Linear Regression Analysis (3rd ed.). The American Statistician, 57(1):67– ??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s13/p67.

Jeske:2004:TAM

[JB04] Daniel R. Jeske and Todd Blessinger. Tunable approximations for the mean and variance of the maximum of heterogeneous geometrically distributed random variables. The American Statistician, 58(4):322–327, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jersky:2003:BRH

[Jer03] Brian Jersky. Book review:

Hyperstat (2nd ed.). The American Statistician, 57(4): 316–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s17/p316.

Jeske:1994:IGM

[Jes94] Daniel R. Jeske. Illustrating the Gauss–Markov theorem. The American Statistician, 48 (3):237–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Joiner:1983:URE

[JH83] Brian L. Joiner and David L. Hall. The ubiquitous role of f′/f in efficient estimation of location. The American Statistician, 37(2):128–133, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jolliffe:1996:BBD

[JH96] Ian T. Jolliffe and Peter B. Hope. Bounded bivariate distributions with nearly normal marginals. The American Statistician, 50(1):17–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Joffe:2004:MSC

[JHFK04] Marshall M. Joffe, Thomas R. Ten Have, Harold I. Feldman, and Stephen E. Kimmel. Model selection, confounder control, and marginal structural models. The American Statistician, 58(4):272–279, November 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnson:1990:RWA

[JK90] Norman L. Johnson and Samuel Kotz. Randomly weighted averages: Some aspects and extensions. The American Statistician, 44(3):245–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jong:1999:RBP

[JK99] Jyh-Cherng Jong and Samuel Kotz. On a relation between principal components and regression analysis. The American Statistician, 53(4):349–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jovanovic:1997:LRT

[JL97] B. D. Jovanovic and P. S. Levy.

A look at the rule of three. The American Statistician, 51 (2):137–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/jovanovic.
html.

Johnson:1999:ARA

[Joh99] Roger W. Johnson. Applied regression analysis and multivariable methods, 3rd edition. The American Statistician, 53 (3):292–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnstone:2009:CWS

[Joh09] Iain Johnstone. Comment:

What is statistics? The American Statistician, 63(2):111, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jolliffe:1995:SSC

[Jol95] Ian T. Jolliffe. Sample sizes and the central limit theorem: The Poisson distribution as an illustration. The American Statistician, 49(3):269–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:1999:DRA

[Jon99] M. C. Jones. Distributional relationships arising from simple trigonometric formulas. The American Statistician, 53(2): 99–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:2002:KTP

[Jon02] M. C. Jones. On Khintchine’s theorem and its place in random variate generation. The American Statistician, 56(4): 304–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s7/p304.

Jorgensen:1994:TFI

[Jor94] Murray A. Jorgensen. Tail functions and iterative weights in binary regression. The American Statistician, 48(3):230–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jeske:2001:GML

[JP01] Daniel R. Jeske and Hoang Pham. General — on the maximum likelihood estimates for the Goel-Okumoto software reliability model. The American Statistician, 55(3):219–222, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:1992:DIF

[JR92] M. C. Jones and John A. Rice.

Displaying the important features of large collections of similar curves. The American Statistician, 46(2):140–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jones:2011:SIM

[JRP11] M. C. Jones, J. F. Rosco, and Arthur Pewsey. Skewnessinvariant measures of kurtosis. The American Statistician, 65 (2):89–95, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Jeske:2003:REI

[JS03] Daniel R. Jeske and Ashwin Sampath. A real example that illustrates interesting properties of bootstrap bias correction. The American Statistician, 57 (1):62–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s10/p62.

Jones:2011:IUH

[JS11] Hayley E. Jones and David J.

Spiegelhalter. The identification of “unusual” health-care providers from a hierarchical model. The American Statistician, 65(3):154–163, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnston:1995:SAR

[JSM95] Roger G. Johnston, Shayla D. Schroder, and A. Rajika Mallawaaratchy. Statistical artifacts in the ratio of discrete quantities. The American Statistician, 49(3):285–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnson:1996:CIC

[JT96] Bruce R. Johnson and Benjamin J. Tilly. On the construction of independence counterexamples. The American Statistician, 50(1):14–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Juola:1993:MSC

[Juo93] R. C. Juola. More on shortest confidence intervals. The American Statistician, 47(2): 117–119, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Johnson:2004:FRD

[JW04] H. Dean Johnson and Dennis A. Warner. Factors relating to the degree to which statistical consulting clients deem their consulting experiences to be a success. The American Statistician, 58(4):280–289, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koziol:1997:GTD

[KAB^{+}97] James A. Koziol, Joel E.

Alexander, Lance O. Bauer, Samuel Kuperman, Sandra Morzorati, Sean J. O’Connor, John Rohrbaugh, Bernice Porjesz, Henri Begleiter, and John Polich. A graphical technique for displaying correlation matrices. The American Statistician, 51(4):301–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kafadar:1998:SPP

[Kaf98] Karen Kafadar. Statistics and public policy. The American Statistician, 52(1):86–??, February 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kafadar:1999:BRS

[Kaf99] Karen Kafadar. Book review:

Statistics: Concepts and Controversies, 4th edition. The American Statistician, 53(2): 170–??, May 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kagan:2001:GAL

[Kag01] Abram Kagan. General — another look at the Cramer-Rao inequality. The American Statistician, 55(3):211–212, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kahn:1987:CNB

[Kah87] William D. Kahn. A cautionary note for Bayesian estimation of the binomial parameter n. The American Statistician, 41(1):38–40, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kahn:2006:DMC

[Kah06] Michael Kahn. Data matters: Conceptual statistics for a random world. The American Statistician, 60(3):292–293, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kappenman:1979:UCC

[Kap79] Russell F. Kappenman. On the use of a certain conditional distribution to obtain unconditional distributional results. The American Statistician, 33 (1):23–24, ???? 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Karakostas:1985:MVU

[Kar85] K. X. Karakostas. On minimum variance unbiased estimators. The American Statistician, 39(4):303–305, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See corrections [Kar86, Kar87].

Karakostas:1986:CMV

[Kar86] K. X. Karakostas. Correction to: “On minimum variance unbiased estimators” [Amer. Statist. 39 (1985), no. 4, part 1, 303–305, MR 87b:62031]. The American Statistician, 40 (3):258, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Kar85, Kar87].

Karakostas:1987:CMV

[Kar87] K. X. Karakostas. Correction: “On minimum variance unbiased estimators” [Amer. Statist. 39 (1985), no. 4, part 1, 303–305, MR 87b:62031]. The American Statistician, 41 (1):92, ???? 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Kar85, Kar86].

Kastner:1998:SQC

[Kas98] William Kastner. Statistical quality control using the SAS system. The American Statistician, 52(1):85–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kazempour:1995:ISI

[Kaz95] Kazem Kazempour. Impact of stratification imbalance on probability of type I error. The American Statistician, 49(2): 170–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuindersma:2007:TBM

[KB07] Scott R. Kuindersma and Brian S. Blais. Teaching Bayesian model comparison with the three-sided coin. The American Statistician, 61(3): 239–244, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koehler:2009:AMC

[KBH09] Elizabeth Koehler, Elizabeth Brown, and Sebastien J.-P. A. Haneuse. On the assessment of Monte Carlo error in simulation-based statistical analyses. The American Statistician, 63(2):155–162, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khuri:2002:GEF

[KC02] Andre Khuri and George Casella. General — the existence of the first negative moment revisited. The American Statistician, 56(1):44–47, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuiper:2009:GLI

[KC09] Shonda Kuiper and Linda Collins. Guided labs that introduce statistical techniques used in research from multiple disciplines. The American Statistician, 63(4):343–347, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kass:1998:MCM

[KCGN98] Robert E. Kass, Bradley P.

Carlin, Andrew Gelman, and Radford M. Neal. Markov chain Monte Carlo in practice: A roundtable discussion. The American Statistician, 52 (2):93–100, May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/kass.html.

Kaigh:1987:NGD

[KD87] W. D. Kaigh and Michael F.

Driscoll. Numerical and graphical data summary using Ostatistics. The American Statistician, 41(1):25–32, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kreft:1994:RFM

[KDV94] Ita G. G. Kreft, Jan De Leeuw, and Rien Van Der Leeden. Review of five multilevel analysis programs: BMDP-5V, GENMOD, HLM, ML3, and VARCL. The American Statistician, 48(4):324–??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kirmani:1983:NMG

[KE83] S. N. U. A. Kirmani and E. Mirhakkak Esfahani. A note on the moment generating function. The American Statistician, 37(2):161, ???? 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Keller:1994:CPD

[Kel94] Joseph B. Keller. A characterization of the Poisson distribution and the probability of winning a game. The American Statistician, 48(4):294–298, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kelly:2004:RSP

[Kel04] Patrick J. Kelly. A review of software packages for analyzing correlated survival data. The American Statistician, 58(4): 337–342, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kesling:2000:MBP

[Kes00]  George D. Kesling.  The

mighty bonus point. The American Statistician, 54(3): 215–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Kesling.htm;   http://
                                        www.amstat.org/publications/
                                        tas/Long.htm.

Kettenring:1995:WIN

[Ket95] Jon R. Kettenring. What industry needs. The American Statistician, 49(1):2–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Korn:1990:STR

[KG90] Edward L. Korn and Barry I.

Graubard. Simultaneous testing of regression coefficients with complex survey data: Use of Bonferroni t statistics. The American Statistician, 44(4): 270–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Korn:1995:EDW

[KG95] Edward L. Korn and Barry I.

Graubard. Examples of differing weighted and unweighted estimates from a sample survey. The American Statistician, 49 (3):291–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Korn:1998:SSD

[KG98] Edward L. Korn and Barry I.

Graubard. Scatterplots with survey data. The American Statistician, 52(1):56– ??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/korn.html.

Kish:2004:RPR

[KH04] Leslie Kish and Irene Hess. A “replacement” procedure for reducing the bias of nonresponse. The American Statistician, 58 (4):295–297, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khamis:1991:MCT

[Kha91] H. J. Khamis. Manual computations — A tool for reinforcing concepts and techniques. The American Statistician, 45(4): 294–??, November 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khattree:1992:CEP

[Kha92] Ravindra Khattree. Comparing estimators for population variance using Pitman nearness. The American Statistician, 46 (3):214–??, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khatry:2004:SCB

[Kha04] Deepak B. Khatry. Statistical contributions in biotechnology and nonclinical drug-discovery research. The American Statistician, 58(3):199, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khan:2008:SRL

[Kha08]  M. G. M. Khan.  Statis-

tics for real-life sample surveys: Non-simple-random samples and weighted data. The American Statistician, 62(4): 361, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kepner:1989:NEC

[KHK89] James L. Kepner, James D. Harper, and Sandra Z. Keith. A note on evaluating a certain orthant probability. The American Statistician, 43(1):48–49, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khuri:1993:NSC

[Khu93] André I. Khuri. A note on Scheffé’s confidence intervals. The American Statistician, 47 (3):176–178, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kiessler:2006:PSE

[Kie06] Peter C. Kiessler. Probability and statistics for engineers and scientists. The American Statistician, 60(4):343, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kirk:1991:SCU

[Kir91] Roger E. Kirk. Statistical consulting in a university: Dealing with people and other challenges. The American Statistician, 45(1):28–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kirmani:2000:ESA

[Kir00] Syed N. U. A. Kirmani. Engineering Statistics and Applied Statistics and Probability for Engineers (2nd ed.). The American Statistician, 54(3):226–??, August 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kemp:1991:WDD

[KK91] Adrienne W. Kemp and C. David Kemp. Weldon’s dice data revisited. The American Statistician, 45(3):216–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kafadar:1999:BRM

[KK99] Karen Kafadar and James R. Koehler. Book review: Modern Applied Statistics with S-Plus, 2nd edition. The American Statistician, 53(1):86–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Karr:2006:FEU

[KKO^{+}06] A. F. Karr, C. N. Kohnen, A. Oganian, J. P. Reiter, and A. P. Sanil. A framework for evaluating the utility of data altered to protect confidentiality. The American Statistician, 60 (3):224–232, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kleinbaum:1995:TSC

[Kle95] David G. Kleinbaum. Teaching short courses in biostatistics and epidemiology. The American Statistician, 49(2):200–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kopas:1992:PIE

[KM92] Dale A. Kopas and Paul R. McAllister. Process improvement exercises for the chemical industry. The American Statistician, 46(1):34–??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khamis:1994:OUS

[KM94] Harry J. Khamis and Barbara L. Mann. Outreach at a University Statistical Consulting Center. The American Statistician, 48(3):204–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kulinskaya:2010:VSD

[KMS10] Elena Kulinskaya, Stephan Morgenthaler, and Robert G. Staudte. Variance stabilizing the difference of two binomial proportions. The American Statistician, 64(4):350–356, November 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Knusel:1994:PPD

[Knü94] Leo Knüsel. The prediction problem as the dual form of the two-sample problem with applications to the Poisson and the binomial distribution. The American Statistician, 48(3): 214–219, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koenker:2009:MMW

[Koe09] Roger Koenker. The median is the message: Wilson and Hilfertys experiments on the law of errors. The American Statistician, 63(1):20–25, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Konijn:1987:DFO

[Kon87] Hendrik S. Konijn. Distributionfree and other prediction intervals. The American Statistician, 41(1):11–15, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Konold:1995:CCF

[Kon95] Clifford Konold. Confessions of a coin flipper and would-be instructor. The American Statistician, 49(2):203–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kong:2011:ESC

[Kon11] Yong Kong. The expectation of sample central moments. The American Statistician, 65 (3):198–199, August 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koopmans:1993:NUM

[Koo93] Lambert H. Koopmans. A note on using the moment generating function to teach the laws of large numbers. The American Statistician, 47(3):199–202, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Korn:1990:PPP

[Kor90] Edward L. Korn. Projecting power from a previous study: Maximum likelihood estimation. The American Statistician, 44(4):290–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Koschat:2005:CST

[Kos05] Martin A. Koschat. A case for simple tables. The American Statistician, 59(1):31–40, February 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kott:1986:WMR

[Kot86] Phillip S. Kott. When a meanof-ratios is the best linear unbiased estimator under a model. The American Statistician, 40 (3):202–204, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kott:1991:MBL

[Kot91] Phillip S. Kott. A modelbased look at linear regression with survey data. The American Statistician, 45(2):107–112, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kotz:2010:TIU

[Kot10] Brian C. Kotz. Thoughts on the importance of the undergraduate statistics experience to the discipline’s (and society’s) future. The American Statistician, 64(1):15–18, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kronmal:1979:AMG

[KP79] Richard A. Kronmal and Arthur V. Peterson, Jr. On the alias method for generating random variables from a discrete distribution. The American Statistician, 33(4):214–218, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Krause:2007:MSA

[KP07] Andreas Krause and José Pinheiro. Modeling and simulation to adjust p values in presence of a regression to the mean effect. The American Statistician, 61(4):302–307, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Keeling:2011:SAS

[KP11] Kellie B. Keeling and Robert J. Pavur. Statistical accuracy of spreadsheet software. The American Statistician, 65(4): 265–273, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuczmarski:1999:QPP

[KR99] James G. Kuczmarski and Paul R. Rosenbaum. Quantile plots, partial orders, and financial risk. The American Statistician, 53(3):239–??, August 1999. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

King:2002:PIM

[KR02] Elizabeth N. King and Thomas P.

Ryan. A preliminary investigation of maximum likelihood logistic regression versus exact logistic regression. The American Statistician, 56(3): 163–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s1/p163.

Kraemer:1988:AAG

[Kra88] Helena Chmura Kraemer. Assessment of 2 × 2 associations: generalization of signaldetection methodology. The American Statistician, 42(1): 37–49, ???? 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kreft:2001:RBT

[Kre01] Ita G. G. Kreft. Reviews of books and teaching materials — statistical concepts: Education and behavioral sciences (2nd ed.). The American Statistician, 55(3):259, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kruskal:1987:RIA

[Kru87] William Kruskal. Relative importance by averaging over orderings. The American Statistician, 41(??):6–10, 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Korn:1991:ERV

[KS91] Edward L. Korn and Richard Simon. Explained residual variation, explained risk, and goodness of fit. The American Statistician, 45(3):201–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kagan:2005:SP

[KS05] Abram Kagan and Lawrence A. Shepp. A sufficiency paradox. The American Statistician, 59 (1):54–56, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kotz:2008:VPT

[KS08] Samuel Kotz and Edith Seier. Visualizing peak and tails to introduce kurtosis. The American Statistician, 62(4):348–354, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kim:1994:ETS

[KT94] Junghun Kim and Pravin K. Trivedi. Econometric time series analysis software: A review. The American Statistician, 48(4):336–??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Krishnamoorthy:2002:HTA

[KT02] K. Krishnamoorthy and Jessica Thomson. Hypothesis testing about proportions in two finite populations. The American Statistician, 56(3): 215–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s9/p215.

Kunte:1996:EMP

[KU96] S. Kunte and K. S. Upadhya.

Estimating multinomial probabilities. The American Statistician, 50(3):214–216, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuk:2009:NWD

[Kuk09] Anthony Y. C. Kuk. A new way to derive locally most powerful rank tests. The American Statistician, 63(3):278–280, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kuonen:2000:SAC

[Kuo00]  Diego Kuonen.  A saddle-

point approximation for the collector’s problem. The American Statistician, 54(3): 165–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Kuonen.htm.

Kvam:1996:UES

[Kva96] Paul H. Kvam. Using exam scores to estimate the prevalence of classroom cheating. The American Statistician, 50 (3):238–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kvam:2000:EAL

[Kva00] Paul H. Kvam. The effect of active learning methods on student retention in engineering statistics. The American Statistician, 54(2):136–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Kvam.htm.

Kennedy:2007:OBS

[KYA^{+}07] Richard E. Kennedy, Sharon D. Yeatts, Kellie J. Archer, Chris Gennings, and Viswanathan Ramakrishnan. Opportunities for biostatistics students. The American Statistician, 61(2): 120–126, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lackritz:1990:SEP

[Lac90] James R. Lackritz. Salary evaluation for professional baseball players. The American Statistician, 44(1):4–??, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lachenbruch:2000:HSA

[Lac00] Peter A. Lachenbruch. A handbook of statistical analyses using Stata. The American Statistician, 54(2):153–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

LaMotte:1994:NRI

[LaM94] Lynn Roy LaMotte. A note on the role of independence in t statistics constructed from linear statistics in regression models. The American Statistician, 48(3):238–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

LaMotte:2008:SFP

[LaM08] Lynn Roy LaMotte. Sufficiency in finite parameter and sample spaces. The American Statistician, 62(3):211–215, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Landwehr:1997:ER

[Lan97] James M. Landwehr. Editor’s report. The American Statistician, 51(1):2–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/ed_rept.html.

Landwehr:1999:ER

[Lan99] James M. Landwehr. Editor’s report. The American Statistician, 53(4):297–??, November 1999. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Lange:2004:CSO

[Lan04] Kenneth Lange. Computational statistics and optimization theory at UCLA. The American Statistician, 58(1): 9–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s4/p9.

Lane:2006:SM

[Lan06a] David Lane. Statistics without math. The American Statistician, 60(2):205–206, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Langsrud:2006:ECP

[Lan06b] Øyvind Langsrud. Explaining correlations by plotting orthogonal contrasts. The American Statistician, 60(4):335–339, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Landes:2009:PPP

[Lan09] Reid D. Landes. Passing on the passion for the profession. The American Statistician, 63 (2):163–172, May 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larson:1992:AVJ

[Lar92] David A. Larson. Analysis of variance with just summary statistics as input. The American Statistician, 46(2):151–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larntz:1994:ER

[Lar94] Kinley Larntz. Editor’s report.

The American Statistician, 48 (1):1–??, February 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larntz:1996:ER

[Lar96] Kinley Larntz. Editor’s report. The American Statistician, 50(4):283–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Laubscher:1965:IT

[Lau65] N. F. Laubscher. Interpolation in F -tables. The American Statistician, 19(1):28, 40, 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lawrance:1996:DEW

[Law96] A. J. Lawrance. A design of experiments workshop as an introduction to statistics. The American Statistician, 50(2):156–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lawson:2003:OSS

[Law03] John Lawson. One-step screening and process optimization experiments. The American Statistician, 57(1):15– ??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s2/p15.

Lindsay:2000:MDT

[LB00] Bruce G. Lindsay and Prasanta Basak. Moments determine the tail of a distribution (but not much else). The American Statistician, 54(4): 248–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Lindsay.htm.

Li:2003:DSE

[LBTM03] Xiuhong Li, Jennifer M. Buechner, Patrick M. Tarwater, and Alvaro Muñoz. A diamondshaped equiponderant graphical display of the effects of two categorical predictors on continuous outcomes. The American Statistician, 57(3): 193–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s12/p193.

Ledolter:1996:EIM

[LC96] Johannes Ledolter and K. S.

Chan. Evaluating the impact of the 65 mph maximum speed limit on Iowa rural interstates. The American Statistician, 50 (1):79–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Levin:1999:OHC

[LC99] Bruce Levin and Xun Chen. Is the one-half continuity correction used once or twice to derive a well-known approximate sample size formula to compare two independent binomial distributions? The American Statistician, 53(1):62–??, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lauzon:2009:EMC

[LC09] Carolyn Lauzon and Brian Caffo. Easy multiplicity control in equivalence testing using two one-sided tests. The American Statistician, 63(2):147–154, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lozada-Can:2010:TEA

[LCD10] C. Lozada-Can and A. C. Davison. Three examples of accurate likelihood inference. The American Statistician, 64(2): 131–139, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lane:1994:SCI

[LD94] Thomas P. Lane and William H. DuMouchel. Simultaneous confidence intervals in multiple regression. The American Statistician, 48(4):315–??, November 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ludbrook:1998:WPT

[LD98] John Ludbrook and Hugh Dudley. Why permutation tests are superior to t and F tests in biomedical research. The American Statistician, 52(2):127–?? (or 126–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/ludbrook.
html.

Lazaro:1994:IRD

[LDL^{+}94] Carlos G. Lazaro, David DeLong, Ming-Long Lam, William H. Rogers, and Patrick J. Fleury. Invited responses to Dallal, G. E. (1992), “The Computer Analysis of Factorial Experiments With Nested Factors,” The American Statistician, 46, 240. The American Statistician, 48(2):140–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lynn:2010:CSA

[LDM10] Henry S. Lynn, Zhanjian Dong, and Zhe Mu. Comparison of software algorithms for calculating REML Wald type confidence limits for the betweengroup variance component in a small sample one-way random effects model example. The American Statistician, 64(1): 83–87, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lindsay:1993:DRS

[LE93] R. Murray Lindsay and A. S. C.

Ehrenberg. The design of replicated studies. The American Statistician, 47(3):217–??, August 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Long:2000:UHC

[LE00] J. Scott Long and Laurie H.

Ervin. Using heteroscedasticity consistent standard errors in the linear regression model. The American Statistician, 54 (3):217–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leary:2004:TSU

[Lea04] Richard J. C. Leary. Teaching statistics using baseball. Jim Albert. The American Statistician, 58(4):358, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ledolter:1995:PIS

[Led95] Johannes Ledolter. Projects in introductory statistics courses. The American Statistician, 49 (4):364–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ledolter:2003:BRS

[Led03] Johannes Ledolter. Book review: The Statistical Sleuth (2nd ed.). The American Statistician, 57(2):145–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n2/s18/p145.

Leemis:1986:RAC

[Lee86] Lawrence M. Leemis. Relationships among common univariate distributions. The American Statistician, 40(2):143–146, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:1993:GRB

[Lee93] A. J. Lee. Generating random binary deviates having fixed marginal distributions and specified degrees of association. The American Statistician, 47 (3):209–??, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:1995:GDO

[Lee95]  Yoong-Sin Lee.  Graphical

demonstration of an optimality property of the median. The American Statistician, 49 (4):369–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2007:CCC

[Lee07] Herbert K. H. Lee. Chocolate chip cookies as a teaching aid. The American Statistician, 61(4):351–355, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2008:BMS

[Lee08] Herbert K. H. Lee. Bayesian methods: A social and behavioral sciences approach. The American Statistician, 62(4): 356, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:2009:RWL

[Lee09] Jaechoul Lee. A reformulation of weighted least squares estimators. The American Statistician, 63(1):49–55, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lehmann:1980:ELE

[Leh80] E. L. Lehmann. Efficient likelihood estimators. The American Statistician, 34(4):233–235, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lehoczky:1995:MSP

[Leh95] John Lehoczky. Modernizing statistics Ph.D. programs. The American Statistician, 49(1): 12–??, ???? 1995. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:1983:SPS

[Len83] Russell V. Lenth. Some properties of U statistics. The American Statistician, 37(4):311–313, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:IDA

[Len01a] Russell V. Lenth. Introduction to design and analysis of experiments. The American Statistician, 55(1):80–??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:RBTa

[Len01b] Russell V. Lenth. Reviews of books and teaching materials — A first course in the design of experiments: A linear models approach. The American Statistician, 55(4):370–??, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:RBTb

[Len01c] Russell V. Lenth. Reviews of books and teaching materials — So Much Data So Little Math: How to Predict Data Trends, 5 Easy Profitable Methods. The American Statistician, 55(3):256, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2001:SPS

[Len01d] Russell V. Lenth. Statistical practice — some practical guidelines for effective sample size determination. The American Statistician, 55(3): 187–193, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2002:RBT

[Len02] Russell V. Lenth. Reviews of books and teaching materials — statistics on the table: The history of statistical concepts and methods. The American Statistician, 56(1):76–??, 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lenth:2004:BR

[Len04] Russell V. Lenth. Book reviews.

The American Statistician, 58 (2):92, May 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Levin:1993:IMR

[Lev93] Joel R. Levin. An improved modification of a regressiontoward-the-mean demonstration. The American Statistician, 47(1):24–??, February 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Levin:1997:LBM

[Lev97] Bruce Levin. A lower bound for the Mantel–Haenszel one degree of freedom chi-squared statistic in 1: m matched samples. The American Statistician, 51(4):318–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/levin.html.

Ley:1996:PDU

[Ley96] Eduardo Ley. On the peculiar distribution of the U.S. stock indexes’ digits. The American Statistician, 50(4): 311–313, November 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2684926.

Levine:2000:PSE

[LF00] Richard A. Levine and Juanjuan Fan. Practical statistics by example using Microsoft Excel. The American Statistician, 54(2):151–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lann:2006:TMM

[LF06] Avital Lann and Ruma Falk. Tell me the method, I’ll give you the mean. The American Statistician, 60(4):322–327, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lovett:2000:ACT

[LG00] Marsha C. Lovett and Joel B. Greenhouse. Applying cognitive theory to statistics instruction. The American Statistician, 54 (3):196–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Lovett.htm.

Langohr:2005:LMU

[LG05] Klaus Langohr and Guadalupe Gómez. Likelihood maximization using Web-based optimization tools. The American Statistician, 59(2):192–202, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lyles:2009:FLD

[LGH09] Robert H. Lyles, Ying Guo, and Andrew N. Hill. A fresh look at the discriminant function approach for estimating crude or adjusted odds ratios. The American Statistician, 63 (4):320–327, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lu:2011:ONM

[LGXB11] Bo Lu, Robert Greevy, Xinyi Xu, and Cole Beck. Optimal nonbipartite matching and its statistical applications. The American Statistician, 65(1): 21–30, February 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lurie:1972:MGO

[LH72] D. Lurie and H. O. Hartley. Machine-generation of order statistics for Monte Carlo computations. The American Statistician, 26(1):26–27, February 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Love:2002:SEM

[LH02] Thomas E. Love and David K. Hildebrand. Statistics education and the making statistics more effective in schools of business conferences. The American Statistician, 56(2): 107–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s6/p107.

Lee:2000:EAE

[LHD00] J. Jack Lee, Kenneth R. Hess, and Joel A. Dubin. Extensions and applications of event charts. The American Statistician, 54(1):??, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lichtenberg:1990:AVL

[Lic90] Frank R. Lichtenberg. Aggregation of variables in leastsquares regression. The American Statistician, 44(2):169–??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ling:1992:JSN

[Lin92] Robert F. Ling. Just say no to binomial (and other discrete distributions) tables. The American Statistician, 46(1): 53–??, February 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Link:1993:STU

[Lin93] William A. Link. Simulation testing of unbiasedness of variance estimators. The American Statistician, 47(2):132–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lindley:1997:DBB

[Lin97] D. V. Lindley. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):265–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liss:2008:MAP

[Lis08] Charles L. Liss. Multivariable analysis: A practical guide for clinicians and study design and statistical analysis: A practical guide for clinicians. The American Statistician, 62 (1):92, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Little:2006:CB

[Lit06] Roderick J. Little. Calibrated Bayes. The American Statistician, 60(3):213–223, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lesch:2009:SST

[LJ09] Scott M. Lesch and Daniel R.

Jeske. Some suggestions for teaching about normal approximations to Poisson and binomial distribution functions.

The American Statistician, 63 (3):274–277, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larreamendy-Joerns:2005:SOS

[LJLC05] Jorge Larreamendy-Joerns, Gaea Leinhardt, and Javier Corredor. Six online statistics courses: Examination and review. The American Statistician, 59(3): 240–251, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lyles:1999:NCI

[LK99] Robert H. Lyles and Lawrence L. Kupper. A note on confidence interval estimation in measurement error adjustment. The American Statistician, 53(3): 247–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lesser:1995:FPG

[LKNP95] Martin L. Lesser, Nina E. Kohn, Barbara A. Napolitano, and Savita Pahwa. The FU-PLOT: A graphical method for visualizing the timing of followup in longitudinal studies. The American Statistician, 49(2): 139–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lin:1997:CBT

[LL97] Dennis K. J. Lin and Amy W. Lam. Connections between twolevel factorials and Venn diagrams. The American Statistician, 51(1):49–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/lin.html.

Liao:2003:ACD

[LM03] J. G. Liao and Dan McGee.

Adjusted coefficients of determination for logistic regression. The American Statistician, 57 (3):161–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s3/p161.

Leemis:2008:UDR

[LM08] Lawrence M. Leemis and Jacquelyn T. McQueston. Univariate distribution relationships. The American Statistician, 62(1):45–53, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lock:1993:CFS

[Loc93] Robin H. Lock. A comparison of five student versions of statistics packages. The American Statistician, 47(2):136–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lohr:1995:HDS

[Loh95] Sharon L. Lohr. Hasse diagrams in statistical consulting and teaching. The American Statistician, 49(4):376–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lohr:2004:SFS

[Loh04] Sharon L. Lohr. Statistical frontiers in survey sampling. The American Statistician, 58 (2):145–149, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Longford:2010:BDM

[Lon10] Nicholas T. Longford. Bayesian decision making about small binomial rates with uncertainty about the prior. The American Statistician, 64(2):164–169, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Looney:1995:HUT

[Loo95] Stephen W. Looney. How to use tests for univariate normality to assess multivariate normality. The American Statistician, 49(1):64–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Loosen:1997:CST

[Loo97] Franz Loosen. A concrete strategy for teaching hypothesis testing. The American Statistician, 51(2):158–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/loosen.html.

Lorenzen:1990:UAC

[Lor90] Gunter Lorenzen. A unified approach to the calculation of growth rates. The American Statistician, 44(2):148–150, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Losinger:2004:HDS

[Los04a] Willard C. Losinger. How to do statistical research. The American Statistician, 58(3):200, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Losinger:2004:RGW

[Los04b] Willard C. Losinger. A review of the GUM Workbench. The American Statistician, 58 (2):165–167, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Low:1998:BRU

[Low98] Elizabeth Low. Book review:

Understanding Basic Statistics. The American Statistician, 52 (2):198–?? (or 196–196??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lindley:1976:IBP

[LP76] D. V. Lindley and L. D.

Phillips. Inference for a Bernoulli process (a Bayesian view). The American Statistician, 30(3):112–119, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leon:2000:UCH

[LP00] Ramón V. León and William C. Parr. Use of course home pages in teaching statistics. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lehmann:1982:JN

[LR82] E. L. Lehmann and Constance Reid. Jerzy Neyman, 1894– 1981. The American Statistician, 36(3):161–162, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Liao:2001:SCG

[LR01] J. G. Liao and Ori Rosen. Statistical computing and graphics — fast and stable algorithms for computing and sampling from the noncentral hypergeometric distribution. The American Statistician, 55(4): 366–369, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Loken:2006:PCM

[LR06] Eric Loken and Michael J. Rovine. Peirce’s 19th Century mixture model approach to rater agreement. The American Statistician, 60(2):158–161, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Larocque:2008:CID

[LR08] Denis Larocque and Ronald H. Randles. Confidence intervals for a discrete population median. The American Statistician, 62(1):32–39, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lazar:2011:CCU

[LRF11] Nicole A. Lazar, Jaxk Reeves, and Christine Franklin. A capstone course for undergraduate statistics majors. The American Statistician, 65(3):183–189, August 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lane:1990:WPR

[LRG90] Julia Lane, Russ Ray, and Dennis Glennon. Work profiles of research statisticians. The American Statistician, 44(1):9– ??, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Legler:2010:MIU

[LRZG^{+}10] Julie Legler, Paul Roback, Kathryn Ziegler-Graham, James Scott, Sharon Lane-Getaz, and Matthew Richey. A model for an interdisciplinary undergraduate research program. The American Statistician, 64(1): 59–69, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lehmann:1988:ID

[LS88] E. L. Lehmann and Juliet Popper Shaffer. Inverted distributions. The American Statistician, 42(3):191–194, ???? 1988.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Li:1997:BIN

[LS97] Heng Li and Hal S. Stern. Bayesian inference for nested designs based on Jeffreys’ prior. The American Statistician, 51 (3):219–224, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/li.html.

Lavine:1999:BFW

[LS99a] Michael Lavine and Mark J. Schervish. Bayes factors: What they are and what they are not. The American Statistician, 53 (2):119–??, 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Louis:1999:D

[LS99b] Thomas A. Louis and Wei Shen. Discussion. The American Statistician, 53(3):196–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lad:2001:GDA

[LS01] Frank Lad and Romano Scozzafava. General — distributions agreeing with exchangeable sequential forecasting. The American Statistician, 55(2): 131–139, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Landon:2008:CCP

[LS08] Joshua Landon and Nozer D.

Singpurwalla. Choosing a coverage probability for prediction intervals. The American Statistician, 62(2):120–124, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leemis:2000:SDS

[LSE00] Lawrence M. Leemis, Bruce W.

Schmeiser, and Diane L. Evans. Survival distributions satisfying Benford’s law. The American Statistician, 54(4):236– ??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Leemis.htm.

Langford:2001:GPB

[LSO01] Eric Langford, Neil Schwertman, and Margaret Owens. General — is the property of being positively correlated transitive? The American Statistician, 55(4):322–325, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Leemis:1996:CAI

[LT96] Lawrence M. Leemis and Kishor S. Trivedi. A comparison of approximate interval estimators for the Bernoulli parameter. The American Statistician, 50(1):63–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lee:1997:VOD

[LT97] J. Jack Lee and Z. Nora Tu. A versatile one-dimensional distribution plot: The BLiP plot. The American Statistician, 51(4):353–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Li:2010:CCI

[LTN10] Zhiguo Li, Jeremy M. G. Taylor, and Bin Nan. Construction of confidence intervals and regions for ordered binomial probabilities. The American Statistician, 64(4):291–298, November 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Luceno:1997:FES

[Luc97] Alberto Luceno. Further evidence supporting the numerical usefulness of characteristic functions. The American Statistician, 51(3):233– ??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/luceno.html.

Ludbrook:2005:RFL

[Lud05]  John Ludbrook.  R. A.

Fisher’s life and death in Australia, 1959–1962. The American Statistician, 59(2):164–165, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2002:BRS

[Lun02a]  Robert Lund.  Book re-

view: Stochastic Processes. An Introduction. The American Statistician, 56(4):332– ??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s19/p332.

Lund:2002:BRU

[Lun02b]  Ulric J. Lund.  Book re-

view: Understanding Statistical Concepts Using S-Plus. The American Statistician, 56(3): 249–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s18/p249.

Lund:2004:SPT

[Lun04a] Robert Lund. Stochastic processes and their applications. The American Statistician, 58 (3):261, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2004:EPT

[Lun04b] Robert B. Lund. Elementary probability theory with stochastic processes and an introduction to mathematical finance. The American Statistician, 58 (2):173–174, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2007:RWS

[Lun07a] Robert Lund. Revenge of the white swan. The American Statistician, 61(3):189–192, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lund:2007:HSA

[Lun07b] Ulric J. Lund. A handbook of statistical analyses using R. The American Statistician, 61 (3):277–278, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lurie:1995:RFS

[Lur95] Philip M. Lurie. A review of five statistical packages for Windows. The American Statistician, 49(1):99–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lan:1990:LRT

[LW90] Kuang-Kuo Gordon Lan and Janet Wittes. Linear rank tests for survival data: Equivalence of two formulations. The American Statistician, 44(1):23–??, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Loftsgaarden:1998:STC

[LW98] Don O. Loftsgaarden and Ann E. Watkins. Statistics teaching in colleges and universities: Courses, instructors, and degrees in fall 1995. The American Statistician, 52(4):

308–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/loft.html.

Lynn:2003:SCA

[Lyn03] Henry S. Lynn. Suppression and confounding in action. The American Statistician, 57(1): 58–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s9/p58.

Milliken:1984:NSC

[MA84] George A. Milliken and Mohammed Albohali. On necessary and sufficient conditions for ordinary least squares estimators to be best linear unbiased estimators. The American Statistician, 38(4):298–299, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Maatta:1999:BRA

[Maa99] Jon Maatta. Book review:

ActivStats 2.0 and The Active Practice of Statistics. The American Statistician, 53(1): 85–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

MacLennan:1993:IRS

[Mac93] Richard N. MacLennan. Interrater reliability with SPSS for Windows 5.0. The American Statistician, 47(4):292–??, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Macky:1999:BRC

[Mac99] David W. Macky. Book review: A Course in Mathematical Statistics, 2nd edition. The American Statistician, 53 (1):87–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mackisack:2001:RBT

[Mac01] Margaret Mackisack. Reviews of books and teaching materials — chance encounters: A first course in data analysis and inference. The American Statistician, 55(3):255, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Machler:2007:SIU

[Mäc07] Martin Mächler. Statistics: An introduction using R. The American Statistician, 61(1): 100–101, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Madsen:1991:PPP

[Mad91] Richard Madsen. On the probability of a perfect progression. The American Statistician, 45 (3):214–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Madgett:1998:SUD

[Mad98] Alan Madgett. Some uses for distribution-fitting software in teaching statistics. The American Statistician, 52(3): 253–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/madgett.
html.

Madigan:1999:D

[Mad99] David Madigan. Discussion.

The American Statistician, 53 (3):198–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Magee:1990:MBW

[Mag90] Lonnie Magee. R^{2} measures based on Wald and likelihood ratio joint significance tests. The American Statistician, 44 (3):250–??, August 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Magel:1996:ISP

[Mag96] Rhonda C. Magel. Increasing student participation in large introductory statistics classes. The American Statistician, 50 (1):51–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Magee:1998:NBP

[Mag98] Lonnie Magee. Nonlocal behavior in polynomial regressions. The American Statistician, 52 (1):20–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/magee.html.

Mallows:1998:ZP

[Mal98] Colin Mallows. The zeroth problem. The American Statistician, 52(1):1–9, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/mallows.
html.

Mammen:2006:SM

[Mam06] Enno Mammen. Statistical models. The American Statistician, 60(2):204–205, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mandel:1982:USV

[Man82] J. Mandel. Use of the singular value decomposition in regression analysis. The American Statistician, 36(1):15–24, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manola:2003:BRH

[Man03] Judith Manola. Book review: A Handbook of Statistical Analyses using S-Plus (2nd ed.). The American Statistician, 57 (2):146–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s19/p146.

Mandel:2007:CTH

[Man07]  Micha Mandel.  Censoring

and truncation — highlighting the differences. The American Statistician, 61(4):321–324, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Marasinghe:2007:FCS

[Mar07] Mervyn Marasinghe. A first course in statistical methods. The American Statistician, 61 (1):95–97, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Massey:2004:DSC

[Mas04] Kenneth Massey. Discussion:

Statistics and the college football championship. The American Statistician, 58(3):185–187, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Matthews:1999:EPS

[Mat99] J. N. S. Matthews. Effect of prior specification on Bayesian design for two-sample comparison of a binary outcome. The American Statistician, 53(3): 254–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

MacEachern:1994:SGS

[MB94] Steven N. MacEachern and L. Mark Berliner. Subsampling the Gibbs sampler. The American Statistician, 48(3):188–??, August 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morrell:2000:LRE

[MB00] Christopher H. Morrell and Larry J. Brant. Lines in random effects plots from the linear mixed-effects model. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McGee:2006:APS

[MB06] Monnie McGee and Nora V. Bergasa. Analysis of a pilot study for amelioration of itching in liver disease. The American Statistician, 60(4):303–308, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mansfield:1987:DVR

[MC87] Edward R. Mansfield and Michael D. Conerly. Diagnostic value of residual and partial residual plots. The American Statistician, 41(??):107– 116, 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mee:1991:RTM

[MC91] Robert W. Mee and Tin Chiu Chua. Regression toward the mean and the paired sample t test. The American Statistician, 45(1):39–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murdoch:1996:GDL

[MC96] D. J. Murdoch and E. D. Chow.

A graphical display of large correlation matrices. The American Statistician, 50(2):178–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCulloch:1988:ILF

[McC88] Robert E. McCulloch. Information and the likelihood function in exponential families. The American Statistician, 42(1): 73–75, ???? 1988. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCullagh:1994:DMG

[McC94] Peter McCullagh. Does the moment-generating function characterize a distribution? The American Statistician, 48 (3):208, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCulloch:1998:ARS

[McC98] B. D. McCulloch. Assessing the reliability of statistical software: Part I. The American Statistician, 52(4): 358–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/mccullogh.
html.

McCullough:1999:ARS

[McC99] B. D. McCullough. Assessing the reliability of statistical software: Part II. The American Statistician, 53(2):149–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/mccull.pdf.

McCue:2001:GSF

[McC01] Kenneth E. McCue. General — the statistical foundations of the EI method. The American Statistician, 55(2):106–110, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McDonald:1999:SSS

[McD99] Gary C. McDonald. Shaping statistics for success in the 21st Century: The needs of industry. The American Statistician, 53 (3):203–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McDonald:2004:IS

[McD04]  Elaine T. McDonald.  In-

tro stats. The American Statistician, 58(1):86– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s18/p86.

Morgan:1991:LMD

[MCDD91a] J. P. Morgan, N. R. Chaganty, R. C. Dahiya, and M. J. Doviak. Let’s make a deal: The player’s dilemma. The American Statistician, 45(4):284–287, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). With comment by R. G. Seymann (pp. 287–288) and rejoinder (p. 289).

Morgan:1991:R

[MCDD91b] J. P. Morgan, N. R. Chaganty, R. C. Dahiya, and M. J. Doviak. Rejoinder. The American Statistician, 45(4):289–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McGee:2005:ISA

[McG05] Daniel L. McGee. An introduction to survival analysis using Stata (rev.). The American Statistician, 59(1):113, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:1995:SEF

[MCGM95] David S. Moore, George W.

Cobb, Joan Garfield, and William Q. Meeker. Statistics education fin de siècle. The American Statistician, 49 (3):250–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McIntyre:2005:MUS

[McI05] G. A. McIntyre. A method for unbiased selective sampling, using ranked sets. The American Statistician, 59(3):230–232, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McLeod:1999:NSC

[McL99] A. Ian McLeod. Necessary and sufficient condition for nonsingular Fisher information matrix in ARMA and fractional ARIMA models. The American Statistician, 53(1):71–??, February 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Martinez:2011:EPC

[MCM^{+}11] Josue G. Martinez, Raymond J. Carroll, Samuel Müller, Joshua N. Sampson, and Nilanjan Chatterjee. Empirical performance of crossvalidation with oracle methods in a genomics context. The American Statistician, 65(4): 223–228, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McNeil:1992:GPD

[McN92] Don McNeil. On graphing paired data. The American Statistician, 46(4):307–??, November 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mac:1996:HML

[MD96] Robert J. Mac and G. Dawson. How many light bulbs does it take to generate a data set? The American Statistician, 50 (3):247–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mintz:1997:OOC

[MD97] Jim Mintz and Wilfrid Dixon.

Objection overruled: A comment on Gastwirth, Krieger, and Rosenbaum. The American Statistician, 51(2):117–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/mintz.html.

Morris:1984:IPS

[ME84] Max D. Morris and Sherwood F. Ebey. An interesting property of the sample mean under a first-order autoregressive model. The American Statistician, 38(2):127–129, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meeker:1995:TAA

[ME95] William Q. Meeker and Luis A.

Escobar. Teaching about approximate confidence regions based on maximum likelihood estimation. The American Statistician, 49(1):48–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morris:2011:NCP

[ME11] Nathan Morris and Robert Elston. A note on comparing the power of test statistics at low significance levels. The American Statistician, 65(3):164–166, August 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mease:2003:PML

[Mea03] David Mease. A penalized maximum likelihood approach for the ranking of college football teams independent of victory margins. The American Statistician, 57(4):241– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s4/p241.

Mease:2004:DSC

[Mea04]  David Mease.  Discussion:

Statistics and the college football championship. The American Statistician, 58(3):192–194, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meeden:1986:SPC

[Mee86] Glen Meeden. Sufficiency and partitions of the class of all possible discrete distributions. The American Statistician, 40 (1):42–44, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meeden:1987:EWU

[Mee87] Glen Meeden. Estimation when using a statistic that is not sufficient. The American Statistician, 41(2):135–136, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehlum:2009:IPR

[Meh09] Halvor Mehlum. The island problem revisited. The American Statistician, 63(3):269–273, August 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Menard:2000:CDM

[Men00] Scott Menard. Coefficients of determination for multiple logistic regression analysis. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Menard:2004:SAC

[Men04] Scott Menard. Six approaches to calculating standardized logistic regression coefficients. The American Statistician, 58 (3):218–223, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meng:2009:DFW

[Men09] Xiao-Li Meng. Desired and feared — what do we do now and over the next 50 years? The American Statistician, 63 (3):202–210, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meng:2010:RBT

[Men10] Xiao-Li Meng. Rejoinder: Better training, deeper thinking, and more policing. The American Statistician, 64(1):26–29, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:2007:FDD

[MF07] Gregory K. Miller and Stephanie L. Fridell. A forgotten discrete distribution? reviving the negative hypergeometric model. The American Statistician, 61(4): 347–350, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murad:2003:SSO

[MFS^{+}03] Havi Murad, Anat Fleischman, Siegal Sadetzki, Orna Geyer, and Laurence S. Freedman. Small samples and ordered logistic regression: Does it help to collapse categories of outcome? The American Statistician, 57 (3):155–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s2/p155.

Mayo:1997:ESB

[MG97] Matthew S. Mayo and J. Brian Gray. Elemental subsets: The building blocks of regression. The American Statistician, 51 (2):122–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/mayo.html.

Miao:2004:EDC

[MG04] Weiwen Miao and Joseph L.

Gastwirth. The effect of dependence on confidence intervals for a population proportion. The American Statistician, 58(2):124–130, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Madigan:2009:CWS

[MG09] David Madigan and Andrew Gelman. Comment: What is statistics? The American Statistician, 63(2):114, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Malone:2010:RTI

[MGCR10] Christopher J. Malone, John Gabrosek, Phyllis Curtiss, and Matt Race. Resequencing topics in an introductory applied statistics course. The American Statistician, 64(1):52–58, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Molenberghs:1999:NMC

[MGLK99] Geert Molenberghs, Els J. T.

Goetghebeur, Stuart R. Lipsitz, and Michael G. Kenward. Nonrandom missingness in categorical data: Strengths and limitations. The American Statistician, 53(2):110–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Maghsoodloo:1976:RSM

[MH76] S. Maghsoodloo and J. N. Hool. On response surface methodology and its computer-assisted teaching. The American Statistician, 30(3):140–144, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehta:1993:EPC

[MH93] Cyrus R. Mehta and Joan F. Hilton. Exact power of conditional and unconditional tests: Going beyond the 2 × 2 contingency table. The American Statistician, 47(2):91–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Makuch:1990:SCC

[MHT90] William M. Makuch, Gerald J. Hahn, and William T. Tucker. A statistical computing curriculum to meet industrial needs. The American Statistician, 44 (1):42–??, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:1996:CES

[Mil96] Alan J. Miller. The convergence of Efroymson’s stepwise regression algorithm. The American Statistician, 50(2):180–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:2001:RBT

[Mil01] Jacqueline B. Miller. Reviews of books and teaching materials — electronic companions to statistics, biostatistics, and business statistics. The American Statistician, 55(2):168, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Miller:2003:BRB

[Mil03] Stephen J. Miller. Book reviews: Basic Statistics and Data Analysis (with InfoTrac and CD-ROM) (1st ed.), Introduction to Probability and Statistics (11th ed.) and Just the Essentials of Elementary Statistics (3rd ed.). The American Statistician, 57(3): 219–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s19/p219.

Miller:2004:BRC

[Mil04a]  Stephen J. Miller.  Book

review: CAST (Computer-Assisted Statistics Teaching). The American Statistician, 58 (1):87–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s19/p87.

Miller:2004:BRF

[Mil04b] Stephen J. Miller. Book review: Fifty Fathoms: Statistics Demonstrations for Deeper Understanding. The American Statistician, 58(1):86– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s17/p86.

Miller:2004:SAU

[Mil04c] Stephen J. Miller. Statistics in action: Understanding a world of data. The American Statistician, 58(3):259–260, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Minnotte:2007:BRI

[Min07] Michael Minnotte. Book review: Introduction to Modern Nonparametric Statistics. James J. Higgins. The American Statistician, 61(2):184, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mirkin:2001:GEW

[Mir01] Boris Mirkin. General — eleven ways to look at the chi-squared coefficient for contingency tables. The American Statistician, 55(2):111–120, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Missiakoulis:2006:PLA

[Mis06] Spyros Missiakoulis. Plutarch and Lucullus’ applied sampling. The American Statistician, 60 (3):269–270, August 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mills:2004:EAS

[MJ04] Jamie D. Mills and Elisa L.

Johnson. An evaluation of activstats for SPSS for teaching and learning. The American Statistician, 58(3):254–258, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Myers:1992:RSA

[MKV92] Raymond H. Myers, Andre I.

Khuri, and Geoffrey Vining. Response surface alternatives to the Taguchi Robust Parameter Design Approach. The American Statistician, 46(2):131–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moser:1992:ECT

[ML92] Barry Kurt Moser and Yuh-Ren Lin. Equivalence of the corrected F test and the weighted least squares procedure. The American Statistician, 46(2): 122–124, May 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morris:2009:UNN

[ML09] Carl N. Morris and Kari F.

Lock. Unifying the named natural exponential families and their relatives. The American Statistician, 63(3):247–253, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehrotra:2010:RBA

[MLL10] Devan V. Mehrotra, Xiaomin Lu, and Xiaoming Li. Rankbased analyses of stratified experiments: Alternatives to the van Elteren test. The American Statistician, 64(2):121–130, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Monroe:1980:GMC

[MM80] Robert J. Monroe and Francis E. McVay. Gertrude Mary Cox (1900–1978). The American Statistician, 34(1):48, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Markowski:1990:CEP

[MM90] Carol A. Markowski and Edward P. Markowski. Conditions for the effectiveness of a preliminary test of variance. The American Statistician, 44(4): 322–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Marasinghe:1996:UGS

[MMCsS96] Mervyn G. Marasinghe, William Q. Meeker, Dianne Cook, and Tae sung Shin. Using graphics and simulation to teach statistical concepts. The American Statistician, 50(4):342–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mason:1990:BHA

[MMR90] R. L. Mason, J. D. McKenzie, Jr., and S. J. Ruberg. A brief history of the American Statistical Association, 1839–1989. The American Statistician, 44 (2):68–73, ???? 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Manoukian:1988:NHS

[MN88] Edward B. Manoukian and Pierre Nadeau. A note on the hyperbolic-secant distribution. The American Statistician, 42(1):77–79, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mathew:1997:IMD

[MN97] Thomas Mathew and Kenneth Nordström. An inequality for a measure of deviation in linear models. The American Statistician, 51(4):344–349, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moerbeek:2005:RCVa

[Moe05a] Mirjam Moerbeek. Randomization of clusters versus randomization of persons within clusters. The American Statistician, 59(1):72–78, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moerbeek:2005:RCVb

[Moe05b] Mirjam Moerbeek. Randomization of clusters versus randomization of persons within clusters. The American Statistician, 59(2):173–179, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Monlezun:1979:TDP

[Mon79] Charles J. Monlezun. Twodimensional plots for interpreting interactions in the threefactor analysis of variance model. The American Statistician, 33(2):63–69, ???? 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Monti:1995:FED

[Mon95] Katherine L. Monti. Folded empirical distribution function curves — mountain plots. The American Statistician, 49(4): 342–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Monti:1997:MST

[Mon97] Katherine L. Monti. Making statistical tables in windows: A comparison of Data Muncher and SAS (PROC TABULATE and PROC REPORT). The American Statistician, 51(1): 83–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/monti.html.

Monahan:2004:TSC

[Mon04] John Monahan. Teaching statistical computing at North Carolina State University. The American Statistician, 58(1): 6–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s3/p6.

Monahan:2006:PCR

[Mon06] John F. Monahan. Professor C. R. Mudgeon and the “order,” or writing regression questions so that students do not need a calculator. The American Statistician, 60(1):50–52, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:1993:PVN

[Moo93] David S. Moore. The place of video in new styles of teaching and learning statistics. The American Statistician, 47(3): 172–??, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:1997:BBS

[Moo97a] David S. Moore. Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):254–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/moore.html.

Moore:1997:RBB

[Moo97b] David S. Moore. Reply — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):272–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:2001:UPF

[Moo01] David S. Moore. Undergraduate programs and the future of academic statistics. The American Statistician, 55(1):1– ??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:2005:PGS

[Moo05]  David S. Moore.  Prepar-

ing graduate students to teach statistics. The American Statistician, 59(1):1–3, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moore:2007:SEI

[Moo07] Page C. Moore. Statistics explained: An introductory guide for life scientists. The American Statistician, 61(3):274, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morris:1995:R

[Mor95] Carl N. Morris. Respondent. The American Statistician, 49 (1):21–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morgan:1998:RES

[Mor98] Walter T. Morgan. A review of eight statistics software packages for general use. The American Statistician, 52(1):70–??, February 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moreno:2007:ISC

[Mor07] Jerry L. Moreno. Investigating statistical concepts, applications, and methods. The American Statistician, 61(1):98–99, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mosteller:1997:LSS

[Mos97] Frederick Mosteller. Lessons from sports statistics. The American Statistician, 51(4): 305–310, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Morrell:1997:LTL

[MPB97] Christopher H. Morrell, Jay D.

Pearson, and Larry J. Brant. Linear transformations of linear mixed-effects models. The American Statistician, 51(4): 338–??, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/morrell.html.

Morgan:2007:NST

[MPR07] B. J. T. Morgan, K. J. Palmer, and M. S. Ridout. Negative score test statistic. The American Statistician, 61(4):285–288, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mandel:2009:SBC

[MR09] Micha Mandel and Yosef Rinott. A selection bias conflict and frequentist versus Bayesian viewpoints. The American Statistician, 63(3):211–217, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mrug:2007:EDS

[Mru07] Sylvie Mrug. Experimental design and statistics for psychology: A first course. The American Statistician, 61(3):275–276, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Marquardt:1975:RRP

[MS75] Donald W. Marquardt and Ronald D. Snee. Ridge regression in practice. The American Statistician, 29(??):3–20, ???? 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Meinhold:1987:KFS

[MS87] Richard J. Meinhold and Nozer D. Singpurwalla. A Kalman-filter smoothing approach for extrapolations in certain dose-response, damageassessment, and acceleratedlife-testing studies. The American Statistician, 41(2):101–106, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moser:1992:HVT

[MS92] Barry K. Moser and Gary R.

Stevens. Homogeneity of variance in the two-sample means test. The American Statistician, 46(1):19–??, February 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moschopoulos:1994:EGC

[MS94] Panagis Moschopoulos and Joan G. Staniswalis. Estimation given conditionals from an exponential family. The American Statistician, 48(4):271–275, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCulloch:1995:IDU

[MS95] C. E. McCulloch and S. R.

Searle. On an identity derived from unbiasedness in linear models. The American Statistician, 49(1):39–42, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Moser:1998:ASS

[MS98] Barry Kurt Moser and Julia K.

Sawyer. Algorithms for sums of squares and covariance matrices using Kronecker products. The American Statistician, 52 (1):54–57, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/moser.html.

McCready:2001:SPS

[MS01] Thomas A. McCready and Neil C. Schwertman. Statistical practice — the statistical paleontology of Charles Lyell and the Coupon Problem. The American Statistician, 55(4): 272–278, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Michael:2002:SCG

[MS02] John R. Michael and William R. Schucany. Statistical computing and graphics — the mixture approach for simulating bivariate distributions with specified correlations. The American Statistician, 56(1):48–54, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Maity:2006:TST

[MS06] Arnab Maity and Michael Sherman. The two-sample T test with one variance unknown. The American Statistician, 60 (2):163–166, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McCarthy:2008:PBV

[MSL^{+}08] John McCarthy, Howard Stanislevic, Mark Lindeman, Arlene S. Ash, Vittorio Addona, and Mary Batcher. Percentagebased versus statistical-power-based vote tabulation audits.

The American Statistician, 62 (1):11–16, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McLean:1991:UAM

[MSS91] Robert A. McLean, William L.

Sanders, and Walter W. Stroup. A unified approach to mixed linear models. The American Statistician, 45(1):54–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McDonald:2011:DCJ

[MT11] James B. McDonald and Patrick Turley. Distributional characteristics: Just a few more moments. The American Statistician, 65(2):96–103, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murdoch:2008:VRV

[MTA08] Duncan J. Murdoch, Yu-Ling Tsai, and James Adcock. P - values are random variables. The American Statistician, 62 (3):242–245, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McGill:1978:VBP

[MTL78] R. McGill, John W. Tukey, and W. A. Larsen. Variations of Box plots. The American Statistician, 32(??):12–16, ???? 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mukhopadhyay:2006:MMO

[Muk06] Nitis Mukhopadhyay. MVUE for the mean with one observation. The American Statistician, 60(1):71–74, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mukhopadhyay:2010:WFM

[Muk10] Nitis Mukhopadhyay. When finiteness matters: Counterexamples to notions of covariance, correlation, and independence. The American Statistician, 64(3):231–233, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mullooly:1988:VLT

[Mul88] John P. Mullooly. The variance of left-truncated continuous nonnegative distributions. The American Statistician, 42 (3):208–210, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murray:1986:EMC

[Mur86] Leigh W. Murray. Estimation of missing cells in randomized block and Latin square designs. The American Statistician, 40 (4):289–293, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Murphy:1997:HRS

[Mur97] James R. Murphy. How to read the statistical methods literature: A guide for students.

The American Statistician, 51 (2):155–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/murphy.html.

McKean:1994:CTR

[MV94] Joseph W. McKean and Thomas J.

Vidmar. A comparison of two rank-based methods for the analysis of linear models. The American Statistician, 48(3): 220–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Molenberghs:2007:LRS

[MV07] Geert Molenberghs and Geert Verbeke. Likelihood ratio, score, and Wald tests in a constrained parameter space. The American Statistician, 61(1): 22–27, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mehta:1992:CEM

[MW92] Cyrus R. Mehta and Stephen J.

Walsh. Comparison of exact, mid-p, and Mantel–Haenszel confidence intervals for the common odds ratio across several 2 × 2 contingency tables. The American Statistician, 46 (2):146–??, May 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

McGrath:2005:QCT

[MY05] Richard N. McGrath and Arthur B. Yeh. A quick, compact, two-sample dispersion test. The American Statistician, 59(1):47–53, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Martz:1992:RCF

[MZ92] H. F. Martz and W. J. Zimmer. The risk of catastrophic failure of the solid rocket boosters on the Space Shuttle. The American Statistician, 46(1):42–??, February 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Norton:1985:TM

[NA85] Robert M. Norton and Steven Arnold. A theorem on moments. The American Statistician, 39(2):106, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nadarajah:2009:PDP

[Nad09] Saralees Nadarajah. PDFs and dual PDFs. The American Statistician, 63(1):45–48, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nagler:1998:BRS

[Nag98] Jonathan Nagler. Book review: Statistical Computing Environments for Social Research. The American Statistician, 52(2): 198–?? (or 196–196??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Naiman:1987:CRS

[Nai87] Daniel Q. Naiman. Characterization of risk sets for simple versus simple hypothesis testing. The American Statistician, 41(3):218–220, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nash:1992:SSI

[Nas92] John C. Nash. Statistical shareware: Illustrations from regression techniques. The American Statistician, 46(4):312–??, November 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nash:1995:CAS

[Nas95] John C. Nash. Computer algebra systems: DERIVE(R). The American Statistician, 49 (1):93–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nash:2006:SSP

[Nas06] J. C. Nash. Spreadsheets in statistical practice — another look. The American Statistician, 60(3):287–289, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Naveau:1997:CBC

[Nav97a] Philippe Naveau. Comparison between the Chernoff and factorial moment bounds for discrete random variables. The American Statistician, 51(1):

40–41, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/naveau.html.

Navidi:1997:GIE

[Nav97b] William Navidi. A graphical illustration of the EM algorithm. The American Statistician, 51 (1):29–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/navidi.html.

Nayak:1994:PNC

[Nay94] Tapan K. Nayak. Pitman nearness comparison of some estimators of population variance. The American Statistician, 48 (2):99–102, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nayak:2002:RCT

[Nay02] Tapan K. Nayak. Rao–Cramer type inequalities for mean squared error of prediction. The American Statistician, 56 (2):102–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s5/p102.

Nelder:1990:NPL

[Nel90] J. A. Nelder. Nearly parallel lines in residual plots. The American Statistician, 44(3): 221–??, August 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nelder:1998:STR

[Nel98a] John A. Nelder. The selection of terms in response-surface models: How strong is the weak-heredity principle? The American Statistician, 52(4): 315–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/nelder.
html.

Nelsen:1998:CRL

[Nel98b] Roger B. Nelsen. Correlation, regression lines, and moments of inertia. The American Statistician, 52(4):343–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/nelsen.
html.

Nelsen:2001:PRV

[Nel01] Roger Nelsen. Probability and random variables: A beginner’s guide. The American Statistician, 55(1):82–??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nelson:2007:IMM

[Nel07] Kerrie Nelson. Introduction to mixed modelling: Beyond regression and analysis of variance. The American Statistician, 61(4):368–369, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Neuts:1992:IPP

[Neu92] Marcel F. Neuts. An invariance property of the Poisson process. The American Statistician, 46(4):276–278, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Newton:1999:D

[New99] H. Joseph Newton. Discussion. The American Statistician, 53 (1):15–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Newcombe:2001:GLC

[New01] Robert G. Newcombe. General — logit confidence intervals and the inverse sinh transformation. The American Statistician, 55 (3):200–202, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nickerson:1994:CCC

[Nic94] David M. Nickerson. Construction of a conservative confidence region from projections of an exact confidence region in multiple linear regression. The American Statistician, 48 (2):120–124, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Niermann:2005:OOT

[Nie05] Stefan Niermann. Optimizing the ordering of tables with evolutionary computation. The American Statistician, 59(1): 41–46, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nievergelt:2011:IDN

[Nie11] Yves Nievergelt. Increasing data with a negative slope. The American Statistician, 65 (4):262, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Naik:1996:ROT

[NK96] Dayanand N. Naik and Ravindra Khattree. Revisiting Olympic track records: Some practical considerations in the principal component analysis. The American Statistician, 50 (2):140–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nelder:1995:CAF

[NL95] John A. Nelder and Peter W.

Lane. The computer analysis of factorial experiments: In memorian — Frank Yates. The American Statistician, 49 (4):382–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2009:CWS

[NL09] Deborah Nolan and Duncan Temple Lang. Comment:

What is statistics? The American Statistician, 63(2):117–121, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nolan:2010:CSC

[NL10] Deborah Nolan and Duncan Temple Lang. Computing in the statistics curricula. The American Statistician, 64 (2):97–107, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nash:1998:BRT

[NO98] John C. Nash and Chris Olsen. Book review: The Texas Instruments TI-83 Graphing Calculator. The American Statistician, 52(3):285–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/ti-83.html.

Noether:1984:NEY

[Noe84] Gottfried E. Noether. Nonparametrics: The early years — impressions and recollections. The American Statistician, 38(??): 173–178, 1984. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Niland:1995:SBC

[NOMLT95] Joyce C. Niland, Tamara L. Odom-Maryon, Jennifer Lee, and Barbara C. Tilley. A survey of biostatistical consulting units throughout North America. The American Statistician, 49(2):183–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Norwood:1990:AS

[Nor90]  Janet L. Norwood.  ASA

sesquicentennial. The American Statistician, 44(2):67–??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Normand:1995:MAS

[Nor95] Sharon-Lise T. Normand. Metaanalysis software: A comparative review. The American Statistician, 49(3):297–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Norwood:2007:DOS

[Nor07] Janet L. Norwood. Discussion: The official statistics Olympic challenge. The American Statistician, 61(1):13–15, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nicewander:1997:CCC

[NP97] W. Alan Nicewander and James M. Price. A consonance criterion for choosing sample size. The American Statistician, 51(4):311–??, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/nicewander.
html.

Nash:1996:SMK

[NQ96] John C. Nash and Tony K. Quon. Software for modeling kinetic phenomena. The American Statistician, 50(4):368–??, November 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Neath:1997:EBI

[NS97] Andrew A. Neath and Francisco J. Samaniego. On the efficacy of Bayesian inference for nonidentifiable models. The American Statistician, 51(3): 225–232, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/neath.html.

Nolan:1999:TST

[NS99] D. Nolan and T. P. Speed. Teaching statistics theory through applications. The American Statistician, 53(4):370–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nunnikhoven:1992:BPS

[Nun92] Thomas S. Nunnikhoven. A birthday problem solution for nonuniform birth frequencies. The American Statistician, 46 (4):270–??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nordman:2009:UHE

[NVB09] Daniel J. Nordman, Stephen B.

Vardeman, and Melissa A. Bingham. Uniformly hyperefficient Bayes inference in a class of nonregular problems. The American Statistician, 63 (3):234–238, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nedelman:1986:BTP

[NW86] Jerry Nedelman and Ted Wallenius. Bernoulli trials, Poisson trials, surprising variances, and Jensen’s inequality. The American Statistician, 40(4):286–289, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oakes:1993:NKM

[Oak93] David Oakes. A note on the Kaplan–Meier estimator. The American Statistician, 47(1): 39–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Obremski:1998:BRS

[Obr98] Tom Obremski. Book review:

Statistics for Management, 7th ed. The American Statistician, 52(2):198–?? (or 196–196??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

OCinneide:1990:MWO

[O’C90] Colm Art O’Cinneide. The mean is within one standard deviation of any median. The American Statistician, 44(4): 292–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Okunade:1993:CAR

[OCE93] Albert K. Okunade, Cyril F. Chang, and Richard D. Evans. Comparative analysis of regression output summary statistics in common statistical packages. The American Statistician, 47(4):298–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oehlert:1992:NDM

[Oeh92] Gary W. Oehlert. A note on the delta method. The American Statistician, 46(1):27–29, February 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

OBrien:2003:GIO

[OF03] Timothy E. O’Brien and Gerald M. Funk. A gentle introduction to optimal design for regression models. The American Statistician, 57(4): 265–??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s9/p265.

OHaraHines:1995:ECS

[OH95] R. J. O’Hara Hines and W. G. S. Hines. Exploring Cook’s statistic graphically.

The American Statistician, 49 (4):389–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2008:ESS

[OH08a] Robert A. Oster and Joseph M.

Hilbe. An examination of statistical software packages for parametric and nonparametric data analyses using exact methods. The American Statistician, 62 (1):74–84, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2008:RES

[OH08b] Robert A. Oster and Joseph M.

Hilbe. Rejoinder to “An Examination of Statistical Software Packages for Parametric and Nonparametric Data Analyses Using Exact Methods”. The American Statistician, 62 (2):173–176, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ord:2007:ECD

[OI07] J. Keith Ord and Harvey J.

Iglarsh. The estimation of conditional distributions from large databases. The American Statistician, 61(4):308–314, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Okolo:1999:NCP

[Oko99] Abraham Okolo. The Nigerian census: Problems and prospects. The American Statistician, 53(4):321–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oksanen:1991:SAT

[Oks91] E. H. Oksanen. A simple approach to teaching generalized least squares theory. The American Statistician, 45(3):229–??, August 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ord:1996:AF

[OL96] Keith Ord and Sam Lowe. Automatic forecasting. The American Statistician, 50(1):88–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oldford:1995:PDD

[Old95] R. W. Oldford. A physical device for demonstrating confounding, blocking, and the role of randomization in uncovering a causal relationship. The American Statistician, 49 (2):210–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Olkin:1992:MFH

[Olk92] Ingram Olkin. A matrix formulation on how deviant an observation can be. The American Statistician, 46(3):205–209, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ord:1984:MMG

[Ord84] Keith Ord. In memoriam:

Maurice George Kendall, 1907– 1983. The American Statistician, 38(1):36–37, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Overton:1995:HTT

[OS95] W. Scott Overton and Stephen V.

Stehman. The Horvitz– Thompson theorem as a unifying perspective for probability sampling: With examples from natural resource sampling. The American Statistician, 49 (3):261–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

ONeill:1999:D

[OS99] Robert T. O’Neill and Ana Szarfman. Discussion. The American Statistician, 53(3): 190–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Owen:2000:PDA

[OSC00] W. J. Owen, D. Sinha, and M. H. Capozzoli. A paired-data analysis for a lifetime distribution. The American Statistician, 54(4):252–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Owen.htm.

Oster:1998:EFS

[Ost98] Robert A. Oster. An examination of five statistical software packages for epidemiology. The American Statistician, 52 (3):267–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/oster.html.

Oster:2002:ESS

[Ost02] Robert A. Oster. An examination of statistical software packages for categorical data analysis using exact methods. The American Statistician, 56 (3):235–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s13/p235.

Oster:2003:ESS

[Ost03] Robert A. Oster. An examination of statistical software packages for categorical data analysis using exact methods — Part II. The American Statistician, 57(3):201–??, August 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s14/p201.

Oster:2009:SENa

[Ost09a]  Robert A. Oster.  Section

Editor’s notes. The American Statistician, 63(2):187–188, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2009:SENb

[Ost09b] Robert A. Oster. Section Editor’s notes. The American Statistician, 63(4):400, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENa

[Ost11a] Robert A. Oster. Section editor’s notes. The American Statistician, 65(1):43, February 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENb

[Ost11b]  Robert A. Oster.  Section

editor’s notes. The American Statistician, 65(2):113–114, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SENc

[Ost11c] Robert A. Oster. Section editor’s notes. The American Statistician, 65(3):200, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Oster:2011:SEN

[Ost11d]  Robert A. Oster.  Section

editor’s notes. The American Statistician, 65(4):263–264, November 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ottenbacher:1996:PRR

[Ott96] Kenneth J. Ottenbacher. The power of replications and replications of power. The American Statistician, 50(3):271–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

OGorman:1991:VSD

[OW91] Thomas W. O’Gorman and Robert F. Woolson. Variable selection to discriminate between two groups: Stepwise logistic regression or stepwise discriminant analysis? The American Statistician, 45(3):187–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ormerod:2010:EVA

[OW10] J. T. Ormerod and M. P. Wand. Explaining variational approximations. The American Statistician, 64(2):140–153, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Owen:2001:IBS

[Owe01] William B. Owen. Introduction to business statistics: A computer integrated, data analysis approach. The American Statistician, 55(1):83–??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Perez-Abreu:1991:PAP

[PA91] Víctor Pérez-Abreu. Poisson approximation to power series distributions. The American Statistician, 45(1):42–45, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Padgett:2008:VSP

[Pad08] Lakshmi Padgett. Variations on split plot and split block experiment designs. The American Statistician, 62(2):184, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pagano:2000:SVT

[Pag00] Marcello Pagano. StatConcepts: A visual tour of statistical ideas. The American Statistician, 54(4):323–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Palacios:1999:RPE

[Pal99] Jose Luis Palacios. The ruin problem via electric networks. The American Statistician, 53 (1):67–??, February 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Palacios:2008:SSR

[Pal08] José Luis Palacios. On the simple symmetric random walk and its maximal function. The American Statistician, 62(2): 138–140, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pan:2002:NUM

[Pan02] Wei Pan. A note on the use of marginal likelihood and conditional likelihood in analyzing clustered data. The American Statistician, 56(3): 171–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s2/p171.

Parr:1998:BRP

[Par98] William C. Parr. Book review: Practical Guide to Experimental Design. The American Statistician, 52(3):283–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pardoe:2004:MSS

[Par04]  Iain Pardoe.  Multidimen-

sional scaling for selecting small groups in college courses. The American Statistician, 58(4): 317–321, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Paul:1981:BTU

[Pau81] Jerome L. Paul. A Bayes theorem urn model optimization problem. The American Statistician, 35(4):247–249, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Paulson:1992:ULG

[Pau92] Richard A. Paulson. Using lottery games to illustrate statistical concepts and abuses. The American Statistician, 46 (3):202–??, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Paula:2008:CMI

[Pau08]  Áureo De Paula.  Condi-

tional moments and independence. The American Statistician, 62(3):219–221, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pavur:1988:TER

[Pav88] Robert Pavur. Type i error rates for multiple comparison procedures with dependent data. The American Statistician, 42(3):171–173, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pawitan:2000:RFW

[Paw00] Yudi Pawitan. A reminder of the fallibility of the Wald statistic: Likelihood explanation. The American Statistician, 54(1):??, February 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Picard:1990:DS

[PB90] Richard R. Picard and Kenneth N. Berk. Data splitting. The American Statistician, 44 (2):140–??, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pal:1992:IML

[PB92] Nabendu Pal and J. Calvin Berry. On invariance and maximum likelihood estimation. The American Statistician, 46(3): 209–212, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Parker:2003:SSM

[PB03] Robert A. Parker and Nancy G. Berman. Sample size: More than calculations. The American Statistician, 57(3):166– ??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s4/p166.

Piegorsch:1985:EFN

[PC85] Walter W. Piegorsch and George Casella. The existence of the first negative moment. The American Statistician, 39 (1):60–62, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pardoe:2002:GMA

[PC02] Iain Pardoe and R. Dennis Cook. A graphical method for assessing the fit of a logistic regression model. The American Statistician, 56(4): 263–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s2/p263.

Piepho:2005:SCI

[PE05] Hans-Peter Piepho and Katharina Emrich. Simultaneous confidence intervals for two estimable functions and their ratio under a linear model. The American Statistician, 59(4): 292–300, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peddada:1985:SNP

[Ped85] Shyamal Das Peddada. A short note on Pitman’s measure of nearness. The American Statistician, 39(4):298–299, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peixoto:1990:PWF

[Pei90] Julio L. Peixoto. A property of well-formulated polynomial regression models. The American Statistician, 44(1):26–30, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peixoto:1993:IFR

[Pei93] Julio L. Peixoto. Incorrect F formulas in regression analysis. The American Statistician, 47 (3):194–??, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pena:1997:CIS

[Peñ97] Daniel Peña. Combining information in statistical modeling. The American Statistician, 51(4):326–332, November 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/pena.html.

Percival:1993:TCP

[Per93] Donald B. Percival. Three curious properties of the sample variance and autocovariance for stationary processes with unknown mean. The American Statistician, 47(4):274–276, November 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Perry:1998:SPE

[Per98]  Robert J. Perry.  Statis-

tical practice — estimating strength of the Williamsburg Bridge suspension cables. The American Statistician, 52(3): 211–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/perry.html.

Peskun:1990:NGM

[Pes90] Peter H. Peskun. A note on a general method for obtaining confidence intervals from samples from discrete distributions. The American Statistician, 44 (1):31–35, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peterson:1999:IBS

[Pet99] William P. Peterson. Interpreting basic statistics: A guide and workbook based on excerpts from journal articles. The American Statistician, 53 (3):294–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Proschan:1995:MCC

[PF95] Michael A. Proschan and Dean A. Follmann. Multiple comparisons with control in a single experiment versus separate experiments: Why do we feel differently? The American Statistician, 49(2):144–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Plane:1982:SPN

[PG82] Donald R. Plane and Kenneth R. Gordon. A simple proof of the nonapplicability of the central limit theorem to finite populations. The American Statistician, 36(3):175–176, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Poole:1997:HCA

[PG97] Charles Poole and Sander Greenland. How a court accepted a possible explanation: A comment on Gastwirth, Krieger, and Rosenbaum. The American Statistician, 51(2):112–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/poole.html.

Peng:2002:QAL

[PH02] Roger D. Peng and Nicolas W. Hengartner. Quantitative analysis of literary styles. The American Statistician, 56(3): 175–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s3/p175.

Potthoff:2008:FFC

[PHSN08] Richard F. Potthoff, Susan Halabi, Joellen M. Schildkraut, and Beth Newman. Flexible frames and control sampling in casecontrol studies. The American Statistician, 62(4):307–313, November 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pickle:2006:GDT

[Pic06] Linda Pickle. Graphic discovery: A trout in the milk and other visual adventures. The American Statistician, 60(2): 203, May 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pierce:1980:SRD

[Pie80] David A. Pierce. A survey of recent developments in seasonal adjustment. The American Statistician, 34(3):125–134, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piegorsch:1992:CLR

[Pie92] Walter W. Piegorsch. Complementary log regression for generalized linear models. The American Statistician, 46(2): 94–??, May 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piegorsch:2001:SED

[Pie01] Walter W. Piegorsch. Statistics and Experimental Design for Toxicologists (3rd ed.). The American Statistician, 55(1): 81–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

PillersDobler:1998:BRM

[Pil98] Carolyn Pillers Dobler. Book reviews: Modern Elementary Statistics (9th ed.), A First Course in Statistics (6th ed.). The American Statistician, 52 (3):284–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Peterson:1982:MMC

[PK82] Arthur V. Peterson, Jr. and Richard A. Kronmal. On mixture methods for the computer generation of random variables. The American Statistician, 36 (3):184–191, August 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pan:2000:NML

[PLC00] Wei Pan, Thomas A. Louis, and John E. Connett. A note on marginal linear regression with correlated response data. The American Statistician, 54 (3):191–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Pan.htm.

Peikes:2008:PSM

[PMO08] Deborah N. Peikes, Lorenzo Moreno, and Sean Michael Orzol. Propensity score matching. The American Statistician, 62 (3):222–231, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Philips:1995:MBT

[PN95] Thomas K. Philips and Randolph Nelson. The moment bound is tighter than Chernoff’s bound for positive tail probabilities. The American Statistician, 49(2):175–178, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piepho:2007:SSS

[PO07] Hans-Peter Piepho and Joseph Ochieng Ogutu. Simple state-space models in a mixed model framework. The American Statistician, 61(3):224–232, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Polansky:1999:UBT

[Pol99] Alan M. Polansky. Upper bounds on the true coverage of bootstrap percentile type confidence intervals. The American Statistician, 53(4):362–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Posten:1990:RAT

[Pos90] Harry O. Posten. Review of audio tape: The Statistician at Work. The American Statistician, 44(4):302–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Posten:1993:EAN

[Pos93] Harry O. Posten. An effective algorithm for the noncentral beta distribution function. The American Statistician, 47 (2):129–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pourahmadi:2002:RBT

[Pou02] Mohsen Pourahmadi. Reviews of books and teaching materials — A course in time series analysis. The American Statistician, 56(1):77–??, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Proschan:1998:EUC

[PP98] Michael A. Proschan and Brett Presnell. Expect the unexpected from conditional expectation. The American Statistician, 52(3):248–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/proschan.
html.

Pavlides:2009:HLS

[PP09] Marios G. Pavlides and Michael D. Perlman. How likely is Simpson’s Paradox? The American Statistician, 63(3):226–233, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pavlides:2010:EFP

[PP10] Marios G. Pavlides and Michael D. Perlman. On estimating the face probabilities of shaved dice with partial data. The American Statistician, 64(1):37–45, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Park:1996:SMG

[PPS96] Chul Gyu Park, Taesung Park, and Dong Wan Shin. A simple method for generating correlated binary variates. The American Statistician, 50(4): 306–310, November 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pena:1994:SCA

[PR94] Edsel A. Peña and Vijay Rohatgi. Some comments about sufficiency and unbiased estimation. The American Statistician, 48(3):242–243, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Piepel:1998:MEA

[PR98] Greg Piepel and Trish Redgate.

A mixture experiment analysis of the Hald cement data. The American Statistician, 52 (1):23–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/piepel.
html.

Proschan:2010:BQQ

[PR10] Michael A. Proschan and Jeffrey S. Rosenthal. Beyond the quintessential quincunx. The American Statistician, 64(1): 78–82, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Plante:2011:SN

[PR11] Jean-François Plante and Nancy Reid. Statistics in the news. The American Statistician, 65 (2):80–88, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prado:2008:IDE

[Pra08a] Raquel Prado. Introduction to design of experiments. The American Statistician, 62(4): 362, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pratter:2008:SGJ

[Pra08b] Frederick M. Pratter. SAS graphics for java: Examples using SAS appdev studio and the output delivery system. The American Statistician, 62(4): 359, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:LMR

[Pri05a] Samantha C. Bates Prins. Linear models with R. The American Statistician, 59(3):276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:SAD

[Pri05b] Samantha C. Bates Prins. Statistical analysis and data display: An intermediate course with examples in S-Plus, R, and SAS. The American Statistician, 59(3):275–276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Priebe:2011:FCP

[Pri11] Carey E. Priebe. Fisher’s conditionality principle in statistical pattern recognition. The American Statistician, 65(3): 167–169, August 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Proschan:2008:NAB

[Pro08] Michael A. Proschan. The normal approximation to the binomial. The American Statistician, 62(1):62–63, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Potter:1981:PJT

[PS81] R. W. Potter and G. W. Sturm.

The power of Jonckheere’s test. The American Statistician, 35 (4):249–250, ???? 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Puntanen:1989:EOL

[PS89] Simo Puntanen and George P. H. Styan. The equality of the ordinary least squares estimator and the best linear unbiased estimator. with comments by Oscar Kempthorne and Shayle R. Searle and a reply by the authors. The American Statistician, 43(3):153–164, ???? 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pontius:1994:GAT

[PS94] Jeffrey S. Pontius and Rhoda M.

Schantz. Graphical analyses of a two-period crossover design. The American Statistician, 48 (3):249–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Parr:1998:DCB

[PS98a] William C. Parr and Marlene A.

Smith. Developing case-based business statistics courses. The American Statistician, 52(4): 330–??, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/parr.html.

Pratt:1998:NIF

[PS98b] John W. Pratt and Robert Schlaifer. A new interpretation of the F statistic. The American Statistician, 52(2):140–??, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts_98/pratt.html.

Pratt:1998:NIS

[PS98c] John W. Pratt and Robert Schlaifer. A new interpretation of the F statistic. The American Statistician, 52(2):141–143, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pena:1991:NLE

[PT91] Daniel Peña and George C. Tiao. A note on likelihood estimation of missing values in time series. The American Statistician, 45(3):212–213, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pukelsheim:1994:TSR

[Puk94]  Friedrich Pukelsheim.  The

three sigma rule. The American Statistician, 48(2):88–91, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pulskamp:1990:NEB

[Pul90] Richard J. Pulskamp. A note on the estimation of binomial probabilities. The American Statistician, 44(4):293–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Penner:1991:MI

[PW91] Ralph Penner and Donald G.

Watts. Mining information. The American Statistician, 45 (1):4–??, February 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pearce:2006:PSR

[PW06] N. D. Pearce and M. P. Wand.

Penalized splines and reproducing kernel methods. The American Statistician, 60(3):233–240, August 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rovine:2004:PB

[RA04] Michael J. Rovine and Douglas R. Anderson. Peirce and Bowditch. The American Statistician, 58(3):232–236, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Richards:2007:BRT

[RACK07] Donald Richards, Christine M.

Anderson-Cook, and Alina Kline. Brief reviews of teaching materials. The American Statistician, 61(1):102, February 2007. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Rakovski:2006:RSD

[Rak06] Carter Cleveland Rakovski. Regression with social data: Modeling continuous and limited response variables. The American Statistician, 60(2):206, May 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Randles:2001:SPN

[Ran01] Ronald H. Randles. Statistical practice — on neutral responses (zeros) in the sign test and ties in the Wilcoxon–Mann– Whitney test. The American Statistician, 55(2):96–101, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Raveh:1985:UIC

[Rav85] Adi Raveh. On the use of the inverse of the correlation matrix in multivariate data analysis. The American Statistician, 39(1):39–42, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Raveh:1986:MMA

[Rav86] Adi Raveh. On measures of monotone association. The American Statistician, 40(2): 117–123, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rode:1988:UBC

[RC88] Richard A. Rode and Vernon M. Chinchilli. The use of Box–Cox transformations in the development of multivariate tolerance regions with applications to clinical chemistry. The American Statistician, 42 (1):23–30, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosoff:2005:BRT

[RD05] Jeffrey Rosoff and Carolyn Pillers Dobler. Brief reviews of teaching materials. The American Statistician, 59(4):352, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Read:1998:LAC

[Rea98] K. L. Q. Read. A lognormal approximation for the collector’s problem. The American Statistician, 52(2):173–?? (or 175–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/read.html.

Rees:2002:RBT

[Ree02] Kevin D. Rees. Reviews of books and teaching materials — workshop statistics: Discovery with data, A Bayesian approach. The American Statistician, 56(1):77–78, 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rencher:1992:ICD

[Ren92] Alvin C. Rencher. Interpretation of canonical discriminant functions, canonical variates, and principal components. The American Statistician, 46 (3):217–??, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Reiss:2002:BRE

[RH02] Elayne Reiss and Lorrie L. Hoffman. Book review: The Elements of Statistics: With Applications to Economics and the Social Sciences. The American Statistician, 56(4):333– ??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s20/p333.

Reich:2006:SAB

[RHCR06] Brian J. Reich, James S. Hodges, Bradley P. Carlin, and Adam M. Reich. A spatial analysis of basketball shot chart data. The American Statistician, 60(1):3–12, February 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Richardson:2008:WSD

[Ric08] Mary Richardson. Workshop statistics: Discovery with data. The American Statistician, 62 (4):363, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rice:2010:DTF

[Ric10] Kenneth Rice. A decisiontheoretic formulation of Fisher’s approach to testing. The American Statistician, 64(4):345– 349, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ridout:2009:SAC

[Rid09] Martin S. Ridout. Statistical applications of the complexstep method of numerical differentiation. The American Statistician, 63(1):66–74, February 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Riedwyl:1990:RWC

[Rie90] Hans Riedwyl. Rudolf Wolf’s contribution to the Buffon Needle Problem (an early Monte Carlo experiment) and application of least squares. The American Statistician, 44(2): 138–139, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ries:2006:SMC

[Rie06] Lawrence D. Ries. Statistical methods for communication science. The American Statistician, 60(3):293, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rivest:1981:SCN

[Riv81] Louis-Paul Rivest. On the sum of contaminated normals.

The American Statistician, 35 (3):155–156, ???? 1981. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rappaport:1998:PIS

[RKW98] Karen D. Rappaport, Nouna Kettaneh, and Svante Wold. Perspectives on implementing statistical modeling and design (SMD) in an industrial/ chemical environment. The American Statistician, 52(2): 151–?? (or 152–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/rappaport.
html.

Rousseeuw:1994:SCM

[RM94] Peter J. Rousseeuw and Geert Molenberghs. The shape of correlation matrices. The American Statistician, 48(4):276–279, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosner:1996:MPP

[RMY96] Bernard Rosner, Frederick Mosteller, and Cleo Youtz. Modeling pitching performance and the distribution of runs per inning in major league baseball. The American Statistician, 50(4):352–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Roback:2003:BRW

[Rob03] Paul Roback. Book review:

Weighing the Odds: A Course in Probability and Statistics. The American Statistician, 57 (2):144–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s17/p144.

Robinson:2007:DEA

[Rob07a] Timothy J. Robinson. Design of experiments for agriculture and the natural sciences. The American Statistician, 61(3):275, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Robinson:2007:ICI

[Rob07b] Timothy J. Robinson. Introduction to computer-intensive methods of data analysis in biology. The American Statistician, 61(4):371, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Robinson:2008:RB

[Rob08] Timothy J. Robinson. The R book. The American Statistician, 62(3):272–273, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rockower:1988:IIR

[Roc88] Edward B. Rockower. Integral identities for random variables.

The American Statistician, 42 (1):68–72, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rodriguez:2004:FSD

[Rod04] Ricardo J. Rodriguez. First significant digit patterns from mixtures of uniform distributions. The American Statistician, 58(1):64– 71, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s13/p64.

Rogers:1984:KPA

[Rog84] Gerald S. Rogers. Kronecker products in ANOVA—a first step. The American Statistician, 38(3):197–202, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ross:1995:WGN

[Ros95] N. Phillip Ross. What the government needs. The American Statistician, 49(1):7–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ross:1996:BSR

[Ros96] Sheldon M. Ross. Bayesians should not resample a prior sample to learn about the posterior. The American Statistician, 50(2):116, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosenkrantz:2000:CBQ

[Ros00a] Walter A. Rosenkrantz. Confidence bands for quantile functions: A parametric and graphic alternative for testing goodness of fit. The American Statistician, 54(3): 185–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Rosenkrantz.htm.

Rossman:2000:SYC

[Ros00b] Allan Rossman. Statistics you can’t trust: A friendly guide to clear thinking about statistics in everyday life. The American Statistician, 54(2):151–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosenbaum:2001:GRE

[Ros01] Paul R. Rosenbaum. General — replicating effects and biases. The American Statistician, 55 (3):223–227, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosenbaum:2003:ECI

[Ros03] Paul R. Rosenbaum. Exact confidence intervals for nonconstant effects by inverting the signed rank test. The American Statistician, 57(2): 132–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n2/s12/p132.

Rosenbaum:2005:HC

[Ros05] Paul R. Rosenbaum. Heterogeneity and causality. The American Statistician, 59(2): 147–152, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rosalsky:2007:SPP

[Ros07] Andrew Rosalsky. A simple and probabilistic proof of the binomial theorem. The American Statistician, 61(2):161–162, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rowell:2002:BRS

[Row02] Ginger Holmes Rowell. Book review: Statistics for Psychologists: An Intermediate Course. The American Statistician, 56 (3):248–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s16/p248.

Rousseeuw:1999:BBB

[RRT99] Peter J. Rousseeuw, Ida Ruts, and John W. Tukey. The bagplot: A bivariate boxplot. The American Statistician, 53(4): 382–??, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Radke-Sharpe:1991:WCS

[RS91a] Norean Radke-Sharpe. Writing as a component of statistics education. The American Statistician, 45(4):292–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rae:1991:ASR

[RS91b] Gordon Rae and John E.

Spencer. Average Spearman’s rho, concordance, and the matching problem: Some relationships. The American Statistician, 45(2):161–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Raftery:1993:IAR

[RS93] Adrian E. Raftery and Tore Schweder. Inference about the ratio of two parameters, with application to whale censusing. The American Statistician, 47(4):259–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ritter:2001:APE

[RSH01] Mary Ann Ritter, Robert R.

Starbuck, and Robert V. Hogg. Advice from prospective employers on training BS statisticians. The American Statistician, 55(1):14–??, February 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Roberts:1999:APS

[RSW99] Rosemary Roberts, Richard Scheaffer, and Ann Watkins. Advanced placement statistics— past, present, and future. The American Statistician, 53(4): 307–??, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Root:2001:TCC

[RT01] Rob Root and Trisha Thorme. Teacher’s corner — communitybased projects in applied statistics: Using service-learning to enhance student understanding. The American Statistician, 55 (4):326–331, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rinott:2003:MRR

[RT03] Yosef Rinott and Michael Tam. Monotone regrouping, regression, and Simpson’s Paradox. The American Statistician, 57 (2):139–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s13/p139.

Rubin:2004:DGF

[Rub04a] Donald B. Rubin. The design of a general and flexible system for handling nonresponse in sample surveys. The American Statistician, 58(4):298–302, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rubin:2004:ABS

[Rub04b] Donald B. Rubin. On advice for beginners in statistical research. The American Statistician, 58 (3):196–197, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rue:2004:SMA

[Rue04] Montserrat Rue. Statistical methods for the analysis of biomedical data. The American Statistician, 58(2):174, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rumsey:1998:BSD

[Rum98] Deborah J. Rumsey. Business statistics: Decision making with data. The American Statistician, 52(1):85–??, February 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rumsey:1999:ESM

[Rum99]  Deborah J. Rumsey.  Ex-

ploring statistics: A modern introduction to data analysis and inference. The American Statistician, 53(3):291–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rumsey:2001:RBT

[Rum01] Deborah J. Rumsey. Reviews of books and teaching materials — statistics and data analysis: From elementary to intermediate. The American Statistician, 55(3):255–??, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rumsey:2006:IBS

[Rum06] Deborah J. Rumsey. Introduction to Bayesian statistics. The American Statistician, 60 (1):98–99, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rump:2007:CGS

[Rum07] Christopher M. Rump. Capital growth and the St. Petersburg Game. The American Statistician, 61(3):213–217, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Ruppert:1987:WKI

[Rup87] David Ruppert. What is kurtosis? an influence function approach. The American Statistician, 41(1):1–5, ???? 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Russell:1991:EVS

[Rus91] K. G. Russell. Estimating the value of e by simulation. The American Statistician, 45 (1):66–??, February 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Russek:1998:BRA

[Rus98] Bernadette Russek. Book review: The Assessment Challenge in Statistics Education. The American Statistician, 52 (3):281–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Russo:2008:IPR

[Rus08] Ralph P. Russo. Intuitive probability and random processes using MATLAB. The American Statistician, 62(2):181–182, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Rovine:1997:WLC

[RvE97] Michael J. Rovine and Alexander von Eye. A 14th way to look at a correlation coefficient: Correlation as the proportion of matches. The American Statistician, 51(1):42–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/rovine.html.

Rathbun:1972:RPC

[RWM^{+}72] Daniel B. Rathbun, W. Allen Wallis, Frederick Mosteller, Ansley J. Coale, Paul M. Densen, Solomon Fabricant, Robert D. Fisher, W. Braddock Hickman, William Kruskal, Stanley Lebergott, Richard M. Scammon, William H. Shaw, Frank D. Stella, James A. Suffridge, and John W. Tukey. The report of the President’s Commission on Federal Statistics. The American Statistician, 26 (1):15–18, February 1972. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2682840.

Ryan:2009:ACI

[Rya09] Kenneth J. Ryan. Approximate confidence intervals for p when double sampling. The American Statistician, 63(2):132–140, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Reiczigel:2005:BTS

[RZR05] Jenő Reiczigel, Ildikó Zakariás, and Lajos Rózsa. A bootstrap test of stochastic equality of two populations. The American Statistician, 59(2):156–161, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stinnett:1990:WSS

[S^{+}90] Sandra Stinnett et al. Women in statistics: sesquicentennial activities. The American Statistician, 44(2):74–80, ???? 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sansgiry:2000:SCG

[SA00] Prashant Sansgiry and Olcay Akman. Statistical computing and graphics transformations of the lognormal distribution as a selection model. The American Statistician, 54(4): 307–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Sansgiry.htm.

Stokes:2002:ERb

[SA02] Lynne Stokes and James H. Albert. Editor’s reports. The American Statistician, 56(4): 261–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s1/p261.

Spiegelman:2003:NNM

[SA03] Clifford Spiegelman and Eun Sug Ark. Nearly nonparametric multivariate density estimates that incorporate marginal parametric density information. The American Statistician, 57(3): 183–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s9/p183.

Sabatti:2002:MDV

[Sab02] Chiara Sabatti. Measuring dependency with volume tests. The American Statistician, 56 (3):191–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n3/s5/p191.

Sall:1990:LPG

[Sal90] John Sall. Leverage plots for general linear hypothese. The American Statistician, 44(4): 308–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Salzberg:1999:RSB

[Sal99] Alan J. Salzberg. Removable selection bias in quasiexperiment. The American Statistician, 53(2):103–??, 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samuels:1991:SRT

[Sam91] Myra L. Samuels. Statistical reversion toward the mean: More universal than regression toward the mean. The American Statistician, 45(4):344–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samaniego:1992:EDG

[Sam92] Francisco J. Samaniego. Elementary derivations of geometric moments. The American Statistician, 46(2):108–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Santis:2006:PPT

[San06] Fulvio De Santis. Power priors and their use in clinical trials. The American Statistician, 60(2):122–129, May 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sapra:1991:CBL

[Sap91] Sunil K. Sapra. A connection between the logit model, normal discriminant analysis, and multivariate normal mixtures. The American Statistician, 45(4):265–??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solomon:1994:EPS

[SATV94] Daniel L. Solomon, J. Tim Arnold, Bruce E. Trumbo, and Paul F. Velleman. Electronic publications in statistics — ready or not, here they come. The American Statistician, 48 (3):191–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Saville:1990:MCP

[Sav90] D. J. Saville. Multiple comparison procedures: The practical solution. The American Statistician, 44(2):174–??, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Saw:1992:ASS

[Saw92] Sutaip L. C. Saw. ANOVA sums of squares as quadratic forms. The American Statistician, 46(4):288–??, November 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sawyer:1998:MRE

[Saw98] Hon. Thomas C. Sawyer. Measuring race and ethnicity:

Meeting public policy goals. The American Statistician, 52 (1):33–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/sawyer.
html.

Strand:1998:SPM

[SB98] Matthew Strand and Duane Boes. Statistical practice — modeling road racing times of competitive recreational runners using extreme value theory. The American Statistician, 52 (3):205–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/strand.
html.

Stefanski:2002:GCM

[SB02] Leonard A. Stefanski and Dennis D. Boos. General — the calculus of M-estimation. The American Statistician, 56(1): 29–38, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:2009:MCD

[SB09] Marlene A. Smith and Peter G. Bryant. Managing case discussions in introductory business statistics classes: Practical approaches for instructors. The American Statistician, 63 (4):348–355, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Seier:2011:PVL

[SB11] Edith Seier and Douglas G.

Bonett. A polyplot for visualizing location, spread, skewness, and kurtosis. The American Statistician, 65(4):258–261, November 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sellke:2001:CVT

[SBB01] Thomas Sellke, M. J. Bayarri, and James O. Berger. Calibration of p values for testing precise null hypotheses. The American Statistician, 55(1): 62–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Sellke.htm.

Seely:1997:CSS

[SBL97] Justus F. Seely, David Birkes, and Youngjo Lee. Characterizing sums of squares by their distributions. The American Statistician, 51(1): 55–58, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/seely.html.

Samuel-Cahn:1994:CUE

[SC94] Ester Samuel-Cahn. Combining unbiased estimators. The American Statistician, 48(1):

34–36, ???? 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Silver:1997:PSD

[SC97] Edward A. Silver and Daniel Costa. A property of symmetric distributions and a related order statistic result. The American Statistician, 51(1): 32–33, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/silver.html.

Shmueli:1999:ADH

[SC99] Galit Shmueli and Ayala Cohen. Analysis and display of hierarchical life-time data. The American Statistician, 53(2): 140–??, 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shi:2009:IMG

[SC09] Lei Shi and Gemai Chen. Influence measures for general linear models with correlated errors. The American Statistician, 63 (1):40–42, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schmitt:1981:EHS

[Sch81] Robert C. Schmitt. Early Hawaiian statistics. The American Statistician, 35(1):1–3, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwager:1984:BSL

[Sch84] Steven J. Schwager. Bonferroni sometimes loses. The American Statistician, 38(3):192–197, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schey:1985:GDO

[Sch85]  Harry M. Schey.  A geo-

metric description of orthogonal contrasts in one-way analysis of variance. The American Statistician, 39(2):104–106, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheaffer:1990:TMQ

[Sch90]  Richard L. Scheaffer.  To-

ward a more quantitatively literate citizenry. The American Statistician, 44(1):2–??, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schechtman:1991:EAV

[Sch91] Edna Schechtman. On estimating the asymptotic variance of a function of U statistics. The American Statistician, 45 (2):103–106, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schey:1993:RBM

[Sch93a] Harry M. Schey. The relationship between the magnitudes of SSR(x_{2}) and SSR(x_{2}|x_{1}): A geometric description. The American Statistician, 47(1):26–??, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schucany:1993:ER

[Sch93b] William R. Schucany. Editor’s report. The American Statistician, 47(4):239–??, November 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwarz:1993:MMA

[Sch93c] Carl J. Schwarz. The mixedmodel ANOVA: The truth, the computer packages, the books. part I: Balance data. The American Statistician, 47(1): 48–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schervish:1996:VWT

[Sch96a] Mark J. Schervish. P values: What they are and what they are not. The American Statistician, 50(3):203–206, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwertman:1996:CBQ

[Sch96b] Neil C. Schwertman. A connection between quadratic-type confidence limits and fiducial limits. The American Statistician, 50(3):242–243, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheaffer:1997:DBB

[Sch97] Richard L. Scheaffer. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3): 268–??, August 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schou:1999:BRD

[Sch99] Sue B. Schou. Book review:

Data, Statistics, and Decision Models with Excel. The American Statistician, 53(4):389–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schaffner:2000:RSP

[Sch00] Andrew Schaffner. Revealing statistical principles. The American Statistician, 54(3): 225–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2003:Ia

[Sch03a] Fritz Scheuren. Introduction.

The American Statistician, 57 (3):189–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s10/p189.

Scheuren:2003:Ib

[Sch03b] Fritz Scheuren. Introduction.

The American Statistician, 57 (4):253–??, November 2003.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n4/s6/p253.

Scheuren:2003:LBW

[Sch03c] Fritz Scheuren. Looking back so we can look forward. The American Statistician, 57(2): 94–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s7/p94.

Scheuren:2004:Ia

[Sch04a] Fritz Scheuren. Introduction. The American Statistician, 58 (1):12–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s5/p12.

Scheuren:2004:Ib

[Sch04b] Fritz Scheuren. Introduction. The American Statistician, 58 (2):144, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2004:Ic

[Sch04c] Fritz Scheuren. Introduction. The American Statistician, 58 (3):224, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2004:Id

[Sch04d] Fritz Scheuren. Introduction.

The American Statistician, 58 (4):290–291, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schoolfield:2004:ESM

[Sch04e] Clyde H. Schoolfield, Jr. Elements of stochastic modeling. The American Statistician, 58 (4):360–361, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwartzbaum:2004:ACT

[Sch04f] Judith Schwartzbaum. ActivEpi companion textbook: A supplement for use with The ActivEpi CD-ROM. The American Statistician, 58(4):359– 360, November 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Ia

[Sch05a] Fritz Scheuren. Introduction.

The American Statistician, 59 (1):57, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Ib

[Sch05b] Fritz Scheuren. Introduction.

The American Statistician, 59 (2):162–163, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Ic

[Sch05c] Fritz Scheuren. Introduction. The American Statistician, 59 (3):223, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:Id

[Sch05d] Fritz Scheuren. Introduction. The American Statistician, 59 (4):308, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2005:MI

[Sch05e] Fritz Scheuren. Multiple imputation. The American Statistician, 59(4):315–319, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scheuren:2006:I

[Sch06a] Fritz Scheuren. Introduction. The American Statistician, 60 (1):32–33, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2006:UPC

[Sch06b] Mark Schilling. Understanding probability: Chance rules in everyday life. The American Statistician, 60(1):97–98, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schell:2010:IKS

[Sch10a] Michael J. Schell. Identifying key statistical papers from 1985 to 2002 using citation data for applied biostatisticians. The American Statistician, 64(4): 310–317, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwarz:2010:CCD

[Sch10b] Wolf Schwarz. Comparing continuous and discrete birthday coincidences: “same-day” versus “within 24 hours”. The American Statistician, 64(1): 34–36, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scariano:1987:EVI

[SD87] Stephen M. Scariano and James M. Davenport. The effects of violations of independence assumptions in the oneway ANOVA. The American Statistician, 41(2):123–129, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Satten:2001:GKM

[SD01] Glen A. Satten and Somnath Datta. General — the Kaplan–Meier estimator as an inverse-probability-of-censoring weighted average. The American Statistician, 55(3):207–210, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sekar:2004:MEB

[SD04] C. Chandra Sekar and W. Edwards Deming. On a method of estimating birth and death rates and the extent of registration (excerpt). The American Statistician, 58(1):13– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s6/p13.

Searle:1994:AVC

[Sea94] Shayle R. Searle. Analysis of variance computing package output for unbalanced data from fixed-effects models with nested factors. The American Statistician, 48(2):148–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Searle:1997:BRB

[Sea97] Shayle R. Searle. Built-in restrictions on best linear unbiased predictors (BLUP) of random effects in mixed models. The American Statistician, 51 (1):19–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/searle.html.

Senn:1990:RNM

[Sen90] Stephen Senn. Regression: A new mode for an old meaning? The American Statistician, 44 (2):181–??, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Senn:2008:NCS

[Sen08] Stephen Senn. A note concerning a selection “paradox” of Dawid’s. The American Statistician, 62(3):206–210, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Seymann:1991:C

[Sey91] Richard G. Seymann. Comment. The American Statistician, 45(4):287–??, November 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sprott:1993:DBT

[SF93] D. A. Sprott and V. T. Farewell.

The difference between two normal means. The American Statistician, 47(2):126–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sparks:2000:UDF

[SF00] Ross S. Sparks and John B. F.

Field. Using Deming’s funnel experiment to demonstrate effects of violating assumptions underlying Shewhart’s control charts. The American Statistician, 54(4):291–??, November 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Sparks.htm.

Sampson:1991:PTT

[SG91] Paul D. Sampson and Peter Guttorp. Power transformations and tests of environmental impact as interaction effects. The American Statistician, 45(2):83–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1992:BST

[SG92] A. F. M. Smith and A. E. Gelfand. Bayesian statistics without tears: A sampling– resampling perspective. The American Statistician, 46(2): 84–88, May 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spence:1993:RS

[SG93] Ian Spence and Robert F. Garrison. A remarkable scatterplot. The American Statistician, 47 (1):12–??, February 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shettle:1998:LMS

[SG98] Carolyn Shettle and Catherine Gaddy. The labor market for statisticians and other scientists. The American Statistician, 52(4):295–??, November 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/shettle.
html.

Schall:2000:DBP

[SG00] Teddy Schall and Gary Smith General. Do baseball players regress toward the mean? The American Statistician, 54(4): 231–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Schall.htm.

Schenker:2001:SPJ

[SG01] Nathaniel Schenker and Jane F.

Gentleman. Statistical practice — on judging the significance of differences by examining the overlap between confidence intervals. The American Statistician, 55(3):182–186, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwertman:1990:SNP

[SGC90] Neil C. Schwertman, A. J.

Gilks, and J. Cameron. A simple noncalculus proof that the median minimizes the sum of the absolute deviations. The American Statistician, 44(1): 38–39, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stroup:2004:TSP

[SGCS04] Donna F. Stroup, Richard A.

Goodman, Ralph Cordell, and Richard Scheaffer. Teaching statistical principles using epidemiology: Measuring the health of populations. The American Statistician, 58(1): 77–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/

                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s15/p77.

Schabenberger:2000:CSE

[SGK00] Oliver Schabenberger, Timothy G. Gregoire, and Fanzhi Kong. Collections of simple effects and their relationship to main effects and interactions in factorials. The American Statistician, 54(3): 210–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Schabenberger.htm.

Shepherd:2007:ECS

[SGM07] Bryan E. Shepherd, Peter B. Gilbert, and Devan V. Mehrotra. Eliciting a counterfactual sensitivity parameter. The American Statistician, 61(1): 56–63, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shapiro:1982:CCT

[Sha82] Stanley H. Shapiro. Collapsing contingency tables—a geometric approach. The American Statistician, 36(1):43–46, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shaffer:1991:GMT

[Sha91a] Juliet Popper Shaffer. The Gauss–Markov theorem and random regressors. The American Statistician, 45(4):269–273, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shah:1991:RBC

[Sha91b] Arvind K. Shah. Relationship between the coefficients of determination of algebraically related models. The American Statistician, 45(4):300–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sharpe:2001:RBT

[Sha01] Norean Radke Sharpe. Reviews of books and teaching materials — statistics for management and economics, fifth edition. The American Statistician, 55(4):370, 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sheu:1984:PAB

[She84] Shey Shiung Sheu. The Poisson approximation to the binomial distribution. The American Statistician, 38(3):206–207, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sherman:1997:CSM

[She97] Michael Sherman. Comparing the sample mean and the sample median: An exploration in the exponential power family. The American Statistician, 51 (1):52–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.

                                        amstat.org/publications/
                                        tas/abstracts/sherman.html.

Sherman:1998:PAS

[She98] Malcolm J. Sherman. A probabilistic analysis of the Sacco and Vanzetti evidence. The American Statistician, 52(1):83–??, February 1998. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sherman:2007:PLN

[She07] Michael Sherman. Probabilities: The little numbers that rule our lives. The American Statistician, 61(4):372, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shier:1988:MPM

[Shi88a] D. R. Shier. The monotonicity of power means using entropy. The American Statistician, 42 (3):203–204, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shiffler:1988:MSO

[Shi88b] Ronald E. Shiffler. Maximum Z scores and outliers. The American Statistician, 42(1):79–80, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shieh:2001:GIB

[Shi01] Gwowen Shieh. General — the inequality between the coefficient of determination and the sum of squared simple correlation coefficients. The American Statistician, 55(2):121–124, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shoemaker:1995:TDD

[Sho95] Lewis H. Shoemaker. Tests for differences in dispersion based on quantiles. The American Statistician, 49(2):179–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Short:1997:DBB

[Sho97] Thomas H. Short. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):263–??, August 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Short:1998:BRE

[Sho98] Thomas H. Short. Book review:

Elementary Bayesian Statistics. The American Statistician, 52(3):282–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shoemaker:2003:FTE

[Sho03] Lewis H. Shoemaker. Fixing the F test for equal variances. The American Statistician, 57 (2):105–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s9/p105.

Shuster:1982:NOS

[Shu82] Jonathan J. Shuster. Nonparametric optimality of the sample mean and sample variance. The American Statistician, 36 (3):176–178, August 1982. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shuster:1991:SRC

[Shu91] Jonathan J. Shuster. The statistician in a reverse cocaine sting. The American Statistician, 45(2):123–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shuster:2001:GDS

[Shu01] Jonathan Shuster. General — dependence in simulation: A case study from the game of contract bridge. The American Statistician, 55(4):306–309, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shuster:2006:UTP

[Shu06] Jonathan J. Shuster. Using a two-player coin game paradox in the classroom. The American Statistician, 60(1):68–70, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1982:ECT

[SI82] T. M. Smith and B. Iglewicz. An effective classroom technique for comparison of robust estimators. The American Statistician, 36(3):163–168, August 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Searls:1990:NEV

[SI90] Donald T. Searls and Pichai Intarapanich. A note on an estimator for the variance that utilizes the kurtosis. The American Statistician, 44(4):295–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Silvapulle:1997:CEI

[Sil97] Mervyn J. Silvapulle. A curious example involving the likelihood ratio test against one-sided hypotheses. The American Statistician, 51(2): 178–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/silvapulle.
html.

Simon:1994:WSP

[Sim94] Julian L. Simon. What some puzzling problems teach about the theory of simulation and the use of resampling. The American Statistician, 48(4):290–??, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Simonoff:1998:LRC

[Sim98] Jeffrey S. Simonoff. Logistic regression, categorical predictors, and goodness-of-fit: It depends on who you ask. The American Statistician, 52(1): 10–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/simonoff.
                                        html.

Simon:2007:ARM

[Sim07] Gary Simon. Applied regression modeling: A business approach. The American Statistician, 61(4):367–368, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sinha:2008:IBS

[Sin08] Samiran Sinha. Introduction to Bayesian statistics. The American Statistician, 62(3):268, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sisson:2007:G

[Sis07] S. A. Sisson. Genetics. The American Statistician, 61(2): 112–119, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stanley:2010:CIB

[SJD10] T. D. Stanley, Stephen B. Jarrell, and Hristos Doucouliagos. Could it be better to discard 90% of the data? A statistical paradox. The American Statistician, 64(1):70–77, February 2010. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snapinn:2005:IIT

[SJI05] Steven M. Snapinn, Qi Jiang, and Boris Iglewicz. Illustrating the impact of a timevarying covariate with an extended Kaplan–Meier estimator. The American Statistician, 59(4):301–307, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Shah:1993:TMU

[SK93] Arvind K. Shah and K. Krishnamoorthy. Testing means using hypothesis-dependent variance estimates. The American Statistician, 47(2):115–??, May 1993. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwertman:1995:RGP

[SK95] Neil C. Schwertman and Terry L. Kiser. The role of geometry in pairwise and mutual independence. The American Statistician, 49(1):56–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sklar:2007:UBS

[Skl07] Jeffrey C. Sklar. Understanding business statistics. The American Statistician, 61(4):369–370, November 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Skoulakis:2008:RFC

[Sko08] Georgios Skoulakis. A recursive formula for computing central moments of a multivariate lognormal distribution. The American Statistician, 62 (2):147–150, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samuels:1992:SSR

[SL92] Myra L. Samuels and Tai-Fang C. Lu. Sample size requirements for the back-of-the-envelope binomial confidence interval. The American Statistician, 46(3):228–??, August 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Speed:1979:ETM

[SM79] F. M. Speed and Charles J. Monlezun. Exact f tests for the method of unweighted means in a 2^{k} experiment. The American Statistician, 33(??):15–18, 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stroup:1991:NNA

[SM91] W. W. Stroup and D. K. Mulitze. Nearest neighbor adjusted best linear unbiased prediction. The American Statistician, 45(3):194–??, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schau:1997:UMT

[SM97] Candace Schau and Nancy Mattern. Use of map techniques in teaching applied statistics courses. The American Statistician, 51(2):171–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/schau.html.

Steiner:2009:TVR

[SM09] Stefan H. Steiner and R. Jock MacKay. Teaching variation reduction using a virtual manufacturing environment. The American Statistician, 63(4): 361–365, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schenker:2007:CFF

[SMC^{+}07] Nathaniel Schenker, Katherine L. Monti, George W. Cobb, Ronald S. Fecso, and Joan S. Chmiel. Combining features of a frequency table and a stemand-leaf plot to summarize the American Statistical Association’s strategic activities. The American Statistician, 61(3): 245–247, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwertman:1991:PMN

[SMH91] Neil C. Schwertman, Thomas A.

McCready, and Lesley Howard. Probability models for the NCAA regional basketball tournaments. The American Statistician, 45(1):35–??, February 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1995:RFO

[Smi95] Peter J. Smith. A recursive formulation of the old problem of obtaining moments from cumulants and vice versa. The American Statistician, 49(2): 217–218, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1998:BRB

[Smi98] Marlene Smith. Book reviews: Basic Business Statistics: A Casebook and Business Analysis Using Regression: A Casebook. The American Statistician, 52(3):281–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

                                        amstat.org/publications/
                                        tas/ti-83.html;   http://
                                        www.amstat.org/publications/
                                        tas/ti-83.pdf.

Smith:2011:MMR

[Smi11] Marlene A. Smith. Missteps in multiple regression student projects: Beyond associationnot-causation. The American Statistician, 65(3):190–197, August 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stein:1997:RSS

[SMP97] Philip G. Stein, James R. Matey, and Karen Pitts. A review of statistical software for the Apple Macintosh. The American Statistician, 51(1): 67–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/stein.html.

Serroyen:2009:NML

[SMVD09] Jan Serroyen, Geert Molenberghs, Geert Verbeke, and Marie Davidian. Nonlinear models for longitudinal data. The American Statistician, 63 (4):378–388, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snee:1993:WMS

[Sne93] Ronald D. Snee. What’s missing in statistical education? The American Statistician, 47 (2):149–??, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snell:1998:ECS

[Sne98] J. Laurie Snell. An electronic companion to statistics (CD-ROM). The American Statistician, 52(1):88–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snee:2008:WED

[Sne08] Ronald D. Snee. W. Edwards Deming’s “Making AnotherWorld”. The American Statistician, 62(3):251–255, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Snyder:2005:PP

[Sny05] Carl Snyder. The problem of prosperity. The American Statistician, 59(1):62–67, February 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samsa:1994:ISW

[SO94] Gregory Samsa and Eugene Z. Oddone. Integrating scientific writing into a statistics curriculum: A course in statistically based scientific writing. The American Statistician, 48 (2):117–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solow:1993:ERI

[Sol93] Andrew R. Solow. Estimating record inclusion probability. The American Statistician, 47 (3):206–??, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solo:2000:SDS

[Sol00] V. Solo. A simple derivation of the smoothing spline. The American Statistician, 54(1):??, February 2000. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Soler:2010:WTI

[Sol10] Frank P. Soler. Who is teaching introductory statistics? The American Statistician, 64(1): 19–20, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Somes:1986:GMH

[Som86] Grant W. Somes. The generalized Mantel–Haenszel statistic. The American Statistician, 40(??):106–108, 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sowey:2003:GWE

[Sow03]  Eric R. Sowey.  The get-

ting of wisdom: Educating statisticians to enhance their clients’ numeracy. The American Statistician, 57(2): 89–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s5/p89.

Searle:1985:EPS

[SP85] Shayle R. Searle and Friedrich Pukelsheim. Establishing χ^{2} properties of sums of squares using induction. The American Statistician, 39(4):301–303, ???? 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Searle:1986:EIC

[SP86] Shayle R. Searle and Friedrich Pukelsheim. Effects of intraclass correlation on weighted averages. The American Statistician, 40(2):103–105, ???? 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1994:MSM

[SP94] Woollcott Smith and Milton Parnes. Mean streets: The median of a size-biased sample and the population mean. The American Statistician, 48 (2):106–??, May 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Samaranayake:1998:BRM

[SP98] V. A. Samaranayake and Jagdish K. Patel. Book review: Modern Engineering Statistics and Elements of Engineering Probability and Statistics. The American Statistician, 52(2): 199–?? (or 197–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sparks:2008:MD

[Spa08] JonDavid Sparks. Missing data. The American Statistician, 62 (3):269–270, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Speed:2008:SPE

[Spe08] F. Michael Speed. SAS programming for enterprise guide users. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spiegelman:1985:JIG

[Spi85] C. H. Spiegelman. Jensen’s inequality for general location parameter. The American Statistician, 39(1):54, ???? 1985.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spiegelman:1986:TPU

[Spi86] Clifford H. Spiegelman. Two pitfalls of using standard regression diagnostics when both X and Y have measurement error. The American Statistician, 40 (3):245–248, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [Spi87].

Spiegelman:1987:CTP

[Spi87] Clifford H. Spiegelman. Correction: “Two pitfalls of using standard regression diagnostics when both X and Y have measurement error” [amer. statist. 40 (1986), no. 3, 245–248]. The American Statistician, 41 (1):92, ???? 1987. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Spi86].

Spirer:1990:SST

[Spi90] Herbert F. Spirer. Statistical software teaching package: Statistix 3.1. The American Statistician, 44(4):303–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spirling:2007:TPI

[Spi07]  Arthur Spirling.  “turning

points” in the Iraq conflict. The American Statistician, 61(4): 315–320, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spurrier:1998:TCD

[Spu98] John D. Spurrier. Teacher’s corner — discussion — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):241–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Silva:1983:EPU

[SR83] Luis Carlos Silva and Antonio Rubi. On the estimation of proportions using complete cluster sampling. The American Statistician, 37(1):53–54, ???? 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stromberg:1996:EIW

[SR96] Arnold J. Stromberg and Subathra Ramanathan. Easy implementation of writing in introductory statistics courses. The American Statistician, 50 (2):159–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sharpe:1997:RAS

[SR97] N. Radke Sharpe and R. A. Roberts. The relationship among sums of squares, correlation coefficients, and suppression. The American Statistician, 51(1):46–??, February 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts/sharpe.html.

Szekely:2004:SPP

[SR04] Gábor J. Székely and Donald St. P. Richards. The St. Petersburg Paradox and the crash of high-tech stocks in 2000. The American Statistician, 58 (3):225–231, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Szekely:2005:RRS

[SR05] Gabor J. Székely and Donald St. P. Richards. Response: Remain steadfast with the St. Petersburg Paradox to quantify irrational exuberance. The American Statistician, 59(3): 235–239, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Strug:2007:IES

[SRC07] Lisa J. Strug, Charles A. Rohde, and Paul N. Corey. An introduction to evidential sample size calculations. The American Statistician, 61(3):207–212, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Speckman:2008:DPC

[SRMP08] Paul L. Speckman, Jeffrey N.

Rouder, Richard D. Morey, and Michael S. Pratte. Delta plots and coherent distribution ordering. The American Statistician, 62(3):262–266, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sterling:1995:PDR

[SRW95] T. D. Sterling, W. L. Rosenbaum, and J. J. Weinkam. Publication decisions revisited: The effect of the outcome of statistical tests on the decision to publish and vice versa. The American Statistician, 49(1):108–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Suissa:1984:UMP

[SS84] Samy Suissa and Jonathan J. Shuster. Are uniformly most powerful unbiased tests really best. The American Statistician, 38(3):204–206, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwarz:1991:IPS

[SS91] Carl J. Schwarz and M. Samanta. An inductive proof of the sampling distributions for the MLE’s of the parameters in an inverse Gaussian distribution. The American Statistician, 45 (3):223–225, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sellke:1997:CIU

[SS97] Thomas M. Sellke and Sarah H. Sellke. Chebyshev inequalities for unimodal distributions. The American Statistician, 51(1):

34–40, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/sellke.html.

Silva:1998:NTG

[SS98] A. Pedro Duarte Silva and Antonie Stam. Nonparametric two-group classification: Concepts and a SAS-based software package. The American Statistician, 52(2):185–??, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Smith:1999:CNB

[SS99] Tyler Smith and Neil C. Schwertman. Can the NCAA Basketball Tournament seeding be used to predict margin of victory? The American Statistician, 53(2):94–??, 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Singpurwalla:2001:GNR

[SS01] Nozer D. Singpurwalla and Andrew Swift. General — network reliability and Borel’s Paradox. The American Statistician, 55 (3):213–218, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sinharay:2002:SBF

[SS02] Sandip Sinharay and Hal S.

Stern. On the sensitivity of Bayes factors to the prior distributions. The American Statistician, 56(3):196– ??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s6/p196.

Scheaffer:2004:SUE

[SS04] Richard L. Scheaffer and Elizabeth A. Stasny. The state of undergraduate education in statistics. The American Statistician, 58(4):265–271, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Solow:2005:HSN

[SS05] Andrew R. Solow and Woollcott Smith. How surprising is a new record? The American Statistician, 59(2):153–155, May 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Simon:2006:LL

[SS06] Gary A. Simon and Jeffrey S. Simonoff. “last licks”. The American Statistician, 60(1): 13–18, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scott:2008:SHF

[SS08] David W. Scott and Warren R. Scott. Smoothed histograms for frequency data on irregular intervals. The American Statistician, 62(3):256–261, August 2008. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Solow:2009:SPC

[SS09] Andrew R. Solow and Woollcott K. Smith. A statistical problem concerning the Mar Saba letter. The American Statistician, 63(3):254–257, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sackrowitz:1999:VRV

[SSC99] Harold Sackrowitz and Ester Samuel-Cahn. P values as random variables—expected P values. The American Statistician, 53(4):326–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Searle:1981:SCM

[SSH81] S. R. Searle, F. M. Speed, and H. V. Henderson. Some computational and model equivalences in analyses of variance of unequal-subclass-numbers data. The American Statistician, 35(1):16–33, ???? 1981. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schwertman:1996:MPM

[SSH96] Neil C. Schwertman, Kathryn L.

Schenk, and Brett C. Holbrook. More probability models for the NCAA regional basketball tournaments. The American Statistician, 50(1):34–??, ???? 1996.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Searle:1980:PMM

[SSM80] S. R. Searle, F. M. Speed, and G. A. Milliken. Population marginal means in the linear model: an alternative to least squares means. The American Statistician, 34(4):216–221, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

StLaurent:1990:EMG

[St.90] Roy T. St. Laurent. The equivalence of the Milliken-Graybill procedure and the score test. The American Statistician, 44 (1):36–37, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spencer:1995:SSI

[ST95] William J. Spencer and Paul A. Tobias. Statistics in the semiconductor industry: A competitive necessity. The American Statistician, 49(3):245–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stasny:2001:HGJ

[Sta01] Elizabeth A. Stasny. How to get a job in academics. The American Statistician, 55(1): 35–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.

amstat.org/publications/
tas/Stasny.htm.

Stasny:2004:SRT

[Sta04] Elizabeth A. Stasny. Some random thoughts on doing research. The American Statistician, 58(3):200–202, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stamey:2008:MMS

[Sta08] James Stamey. Modern mathematical statistics with applications. The American Statistician, 62(4):358, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stephenson:1990:SSR

[Ste90] Robert W. Stephenson. A study of student reaction to the use of Minitab in an introductory statistics course. The American Statistician, 44(3):231–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stern:1991:PWF

[Ste91] Hal Stern. On the probability of winning a football game. The American Statistician, 45 (3):179–??, August 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stefanski:1992:MLR

[Ste92] Leonard A. Stefanski. Monotone likelihood ratio of a “faulty-inspection” distribution. The American Statistician, 46(2):110–??, May 1992. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stefanski:1996:NAG

[Ste96] L. A. Stefanski. A note on the arithmetic-geometric-harmonic mean inequalities. The American Statistician, 50(3):246–247, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stephenson:2001:IPS

[Ste01] W. Robert Stephenson. Introduction to Probability and Statistics for Engineers and Scientists (2nd ed.). The American Statistician, 55(1):81–??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stern:2004:SCF

[Ste04] Hal S. Stern. Statistics and the college football championship. The American Statistician, 58 (3):179–185, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stefanski:2007:RR

[Ste07] Leonard A. Stefanski. Residual (sur)realism. The American Statistician, 61(2):163–177, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stefanski:2008:NCL

[Ste08] Leonard A. Stefanski. The North Carolina Lottery coincidence. The American Statistician, 62(2):130–134, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:1986:ESC

[Sti86] Stephen M. Stigler. Estimating serial correlation by visual inspection of diagnostic plots. The American Statistician, 40 (2):111–116, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stine:1995:GIV

[Sti95] Robert A. Stine. Graphical interpretation of variance inflation factors. The American Statistician, 49(1):53–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:1999:FSS

[Sti99] Stephen M. Stigler. The foundations of statistics at stanford. The American Statistician, 53 (3):263–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stine:2000:SM

[Sti00] Robert Stine. Statistics with Mathematica. The American Statistician, 54(2):153–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:2008:SHT

[Sti08] Stephen M. Stigler. Statistics at Harvard? A toast to the Harvard Statistics Department. The American Statistician, 62 (3):246–250, August 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stigler:2010:CHR

[Sti10] Stephen M. Stigler. The changing history of robustness. The American Statistician, 64(4): 277–281, November 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stone:1998:BPC

[Sto98]  Richard A. Stone.  The

blind paper cutter: Teaching about variation, bias, stability, and process control. The American Statistician, 52(3): 244–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts_98/stone.html.

Stokes:2001:ER

[Sto01] Lynne Stokes. Editor’s report. The American Statistician, 55 (2):87–88, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stokes:2002:ERa

[Sto02] Lynne Stokes. Editor’s reports. The American Statistician, 56(2):83–??, May 2002.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s1/p83.

Strayhorn:1990:EER

[Str90] Joseph M. Strayhorn. Estimating the errors remaining in a data set: Techniques for quality control. The American Statistician, 44(1):14–??, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Strauss:1992:MFL

[Str92] David Strauss. The many faces of logistic regression. The American Statistician, 46(4): 321–??, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Strickland:2002:RBT

[Str02] Pamela A. Ohman Strickland.

Reviews of books and teaching materials — SPSS for Windows: An introduction to use and interpretation in research. The American Statistician, 56 (1):77–??, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stukel:1990:GME

[Stu90] Therese A. Stukel. A general model for estimating EDsub100p for binary response dose-response data. The American Statistician, 44(1):19–??, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Stufken:2009:IEI

[Stu09] John Stufken. An invitation: Editor’s introduction to “desired and feared — what do we do now and over the next 50 years?”. The American Statistician, 63(3):201, August 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Su:2008:DMM

[Su08] Xiaogang Su. Data mining methods and models. The American Statistician, 62(1): 91, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sukhatme:1992:PTS

[Suk92] Shashikala Sukhatme. Powers of two-sample rank tests under the Lehmann alternatives. The American Statistician, 46 (3):212–214, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Sullivan:1996:UDF

[Sul96] Jerry Sullivan. Understanding the degrees of freedom concept by computer experiments. The American Statistician, 50 (3):234–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Symanzik:2003:TEC

[SV03] Jürgen Symanzik and Natascha Vukasinovic. Teaching experiences with a course on “Web-based statistics”. The American Statistician, 57(1): 46–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s7/p46.

Small:2011:STF

[SVR11] Dylan S. Small, Kevin G.

Volpp, and Paul R. Rosenbaum. Structured testing of 2×2 factorial effects: An analytic plan requiring fewer observations. The American Statistician, 65(1): 11–15, February 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Singer:1990:ITA

[SW90] Judith D. Singer and John B.

Willett. Improving the teaching of applied statistics: Putting the data back into data analysis. The American Statistician, 44(3):223–??, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Scott:1991:TR

[SW91] Alastair Scott and Chris Wild.

Transformations and R^{2}. The American Statistician, 45(2): 127–??, May 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:1994:SSP

[SW94] Mark F. Schilling and Ann E. Watkins. A suggestion for sunflower plots. The American Statistician, 48(4):303–??, November 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spence:2005:WPH

[SW05] Ian Spence and Howard Wainer. William Playfair and his graphical inventions: An excerpt from the introduction to the republication of his atlas and statistical breviary. The American Statistician, 59(3):224–229, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Schilling:2002:HHB

[SWW02] Mark F. Schilling, Ann E. Watkins, and William Watkins. Is human height bimodal? The American Statistician, 56(3): 223–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s10/p223.

Saw:1988:CCI

[SYC88] John G. Saw, Mark C. K. Yang, and Tse Chin. Correction to: “Chebyshev inequality with estimated mean Mo and variance” [Amer. Statist. 38 (1984), no. 2, 130–132, MR 85b:62019]. The American Statistician, 42 (2):166, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [SYM84].

Saw:1984:CIE

[SYM84] John G. Saw, Mark C. K. Yang, and Tse Chin Mo. Chebyshev inequality with estimated mean and variance. The American Statistician, 38(2):130–132, ???? 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [SYC88].

Strasak:2007:USM

[SZM^{+}07] Alexander M. Strasak, Qamruz Zaman, Gerhard Marinell, Karl P. Pfeiffer, and Hanno Ulmer. The use of statistics in medical research. The American Statistician, 61(1):47–55, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Taleb:2007:BSD

[Tal07] Nassim Nicholas Taleb. Black swans and the domains of statistics. The American Statistician, 61(3):198–200, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tam:1984:COS

[Tam84] S. M. Tam. On covariances from overlapping samples. The American Statistician, 38(4): 288–289, ???? 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tanis:1999:BRM

[Tan99] Elliot A. Tanis. Book review: Maple V Student Version: Release 5 . The American Statistician, 53(4):389–??, November 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tang:2002:BRE

[Tan02] Boxin Tang. Book review: Experimental Design With Applications in Management, Engineering, and the Sciences. The American Statistician, 56(4): 333–??, November 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s21/p333.

Tarpey:2000:NPS

[Tar00] Thaddeus Tarpey. A note on the prediction sum of squares statistic for restricted least squares. The American Statistician, 54(2):116–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Tarpey.htm.

Tarpey:2007:LTM

[Tar07] Thaddeus Tarpey. Linear transformations and the k-means clustering algorithm. The American Statistician, 61(1): 34–40, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Theil:1988:ITM

[TC88] Henri Theil and Ching-Fan Chung. Information-theoretic measures of fit for univariate and multivariate linear regressions. The American Statistician, 42(4):249–252, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Terrell:2003:BRA

[Ter03]  George Terrell.  Book re-

view: All the Mathematics You Missed [But Need to Know for Graduate School]. The American Statistician, 57(1): 69–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s17/p69.

Terpstra:2005:SIC

[Ter05] Jeff T. Terpstra. Some illustrative classroom examples regarding sums of discrete random variables with finite support. The American Statistician, 59 (3):258–265, August 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tuyl:2008:CBL

[TGM08] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. A comparison of Bayes–Laplace, Jeffreys, and other priors. The American Statistician, 62(1): 40–44, February 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tarpey:2000:SB

[TH00] Thaddeus Tarpey and John Holcomb. Spline bottles. The American Statistician, 54(2): 129–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/Tarpey2.htm.

Thigpen:1987:SSP

[Thi87] Charles C. Thigpen. A sample size problem in simple linear regression. The American Statistician, 41(3):214–215, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Thompson:1987:SSE

[Tho87] Steven K. Thompson. Sample size for estimating multinomial proportions. The American Statistician, 41(1):42–46, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Thompson:1995:CAS

[Tho95] Paul A. Thompson. Correspondence analysis in statistical package programs. The American Statistician, 49(3):309–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Thompson:2004:IW

[Tho04] Elizabeth Thompson. The importance of “why?”. The American Statistician, 58(3):198, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tibshirani:1997:WFM

[Tib97] Robert Tibshirani. Who is the fastest man in the world? The American Statistician, 51 (2):106–??, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/tibshirani.
html.

Tietjen:1994:RSC

[Tie94]  Gary Tietjen.  Recursive

schemes for calculating cumulative binomial and Poisson probabilities. The American Statistician, 48(2):136–??, May 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tietjen:1998:RSC

[Tie98]  Gary Tietjen.  Recursive

schemes for calculating common cumulative distributions. The American Statistician, 52 (3):263–??, August 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/tietjen.
html.

Timm:2004:EES

[Tim04] Neil H. Timm. Estimating effect sizes in exploratory experimental studies when using a linear model. The American Statistician, 58(3):213–217, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tjur:1998:NRQ

[Tju98] Tue Tjur. Nonlinear regression, quasi likelihood, and overdispersion in generalized linear models. The American Statistician, 52(3):222–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/abstracts_98/tjur.html.

Tjur:2009:CDL

[Tju09] Tue Tjur. Coefficients of determination in logistic regression models—A new proposal: The coefficient of discrimination. The American Statistician, 63(4):366–372, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tarter:1976:IIT

[TK76] M. E. Tarter and R. A. Kronmal. An introduction to the implementation and theory of nonparametric density estimation. The American Statistician, 30(3):105–112, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Troendle:2004:ESC

[TKM04] James F. Troendle, Edward L.

Korn, and Lisa M. McShane. An example of slow convergence of the bootstrap in high dimensions. The American Statistician, 58(1):25– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s8/p25.

Tingley:1993:NOC

[TL93] Maureen Tingley and Chunhui Li. A note on obtaining confidence intervals for discrete parameters. The American Statistician, 47(1):20–23, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Taraldsen:2010:IPI

[TL10] Gunnar Taraldsen and Bo Henry Lindqvist. Improper priors are not improper. The American Statistician, 64(2):154–158, May 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Taylor:1995:CCB

[TM95] Douglas J. Taylor and Keith E.

Muller. Computing confidence bounds for power and sample size of the general linear univariate model. The American Statistician, 49(1):43–47, ???? 1995. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Tweedie:1999:CAR

[TM99] Richard Tweedie and Kerrie Mengersen. Calculating accuracy rates from multiple assessors with limited information. The American Statistician, 53 (3):233–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Thomas:1993:GIP

[TO93] Guy Thomas and John O’Quigley. A geometric interpretation of partial correlation using spherical triangles. The American Statistician, 47(1):30–??, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tobias:2004:AAS

[Tob04] Justin L. Tobias. Assessing assessments of school performance: The case of California. The American Statistician, 58 (1):55–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s12/p55.

Tomkins:1984:APB

[Tom84] R. James Tomkins. Another proof of Borel’s strong law of large numbers. The American Statistician, 38(3):208–209, ???? 1984. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic).

Tortora:1978:NSS

[Tor78] Robert D. Tortora. A note on sample size estimation for multinomial populations. The American Statistician, 32(3): 100–102, ???? 1978. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Torrieri:2007:ACC

[Tor07] Nancy K. Torrieri. America is changing, and so is the census. The American Statistician, 61 (1):16–21, February 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tra:2007:EYC

[Tra07] Yolande Tra. Even you can learn statistics: A guide for everyone who has ever been afraid of statistics. The American Statistician, 61(2):182, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Theoharakis:2003:HDS

[TS03] Vasilis Theoharakis and Mary Skordia. How do statisticians perceive statistics journals? The American Statistician, 57(2):115–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s10/p115.

Tan:2009:SSS

[TTT09] Ming T. Tan, Guo-Liang Tian, and Man-Lai Tang. Sample surveys with sensitive questions: A nonrandomized response approach. The American Statistician, 63(1):9–16, February 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1947:QIO

[Tuk47] John W. Tukey. Question 4: Infinite observations. The American Statistician, 1(3):17– 18, ???? 1947. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1949:AQM

[Tuk49a] John W. Tukey. Answer to question 21: Maximum likelihood estimates, efficiency. The American Statistician, 3(4):12, ???? 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1949:MSFa

[Tuk49b] John W. Tukey. Memorandum on statistics in the Federal Government, Part I. The American Statistician, 3(1):6– 17, ???? 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1949:MSFb

[Tuk49c] John W. Tukey. Memorandum on statistics in the Federal Government, Part II. The American Statistician, 3(2):12– 16, ???? 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1965:TTSb

[Tuk65] John W. Tukey. The technical tools of statistics. The American Statistician, 19(2):23–28, ???? 1965. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1980:WNB

[Tuk80] John W. Tukey. We need both exploratory and confirmatory. The American Statistician, 34 (??):23–25, ???? 1980. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1985:CMI

[Tuk85a] John W. Tukey. Comments on More intelligent statistical software and statistical expert systems: Future directions. The American Statistician, 39(??): 12–14, ???? 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1985:DPP

[Tuk85b] John W. Tukey. Discussion:

Projection pursuit. The American Statistician, 39(2):517–518, June 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/projecteuclid.org/euclid.
aos/1176349533;   http://
www.jstor.org/stable/2241189.

Tukey:1986:SS

[Tuk86] John W. Tukey. Sunset salvo. The American Statistician, 40 (??):72–76, ???? 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Turner:1997:SEI

[Tur97] Danny W. Turner. A simple example illustrating a well-known property of the correlation coefficient. The American Statistician, 51(2):170–??, May 1997. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/turner.html.

Turiel:2007:QRB

[Tur07] Thomas P. Turiel. Quantum random bit generators. The American Statistician, 61(3): 255–259, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tuyl:2009:LE

[Tuy09] Frank Tuyl. Letter to the Editor. The American Statistician, 63(3):294, August 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tepping:1991:MHH

[TW91] Benjamin J. Tepping and Joseph Waksberg. Morris H. Hansen, 1910–1990. The American Statistician, 45(1):2–3, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wu:2001:GCG

[tWGB01] Chi tsung Wu, Marcia L.

Gumpertz, and Dennis D. Boos. General — comparison of GEE, MINQUE, ML, and REML estimating equations for normally distributed data. The American Statistician, 55(2):125–130, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Uyar:1990:RPS

[UE90] Bulent Uyar and Orhan Erdem. Regression procedures in SAS: Problems? The American Statistician, 44(4):296–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Utts:2008:ETB

[UJ08] Jessica Utts and Wesley Johnson. The evolution of teaching Bayesian statistics to nonstatisticians. The American Statistician, 62(3):199–201, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Upton:2000:CIM

[Upt00] Graham J. G. Upton. Conditional independence, Mantel– Haenszel test, and the Yates correction. The American Statistician, 54(2):112–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Upton.htm.

Upton:2001:SCG

[Upt01] Graham J. G. Upton. Statistical computing and graphics — A toroidal scatter diagram for ternary variables. The American Statistician, 55(3): 240–243, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Utts:2002:BRN

[Utt02] Jessica Utts. Book review: News and Numbers: A Guide to Reporting Statistical Claims and Controversies in Health and Related Fields, (2nd Ed.). The American Statistician, 56 (4):330–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n4/s15/p330.

Utts:2003:SSS

[Utt03a] Jessica Utts. Special section on statistical literacy: How you can help. The American Statistician, 57(2):73–??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n2/s1/p73.

Utts:2003:WEC

[Utt03b]  Jessica Utts.  What edu-

cated citizens should know about statistics and probability. The American Statistician, 57(2):74–??, May 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n2/s2/p74.

Verrill:1990:WGP

[VAD90] Steve Verrill, Michael Axelrod, and Mark Durst. We’ve got the positive correlation BLUEs. The American Statistician, 44 (2):171–??, May 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vangel:1996:CIN

[Van96a] Mark G. Vangel. Confidence intervals for a normal coefficient of variation. The American Statistician, 50(1):21–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vanhonacker:1996:MAR

[Van96b] Wilfried R. Vanhonacker. Metaanalysis and response surface extrapolation: A least squares approach. The American Statistician, 50(4):294–??, November 1996. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

VanLeeuwen:1997:NCS

[Van97] Dawn M. VanLeeuwen. A note on the covariance structure in a linear model. The American Statistician, 51(2): 140–144, May 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/vanleeuwen.
                                        html.

Varian:1972:LEB

[Var72] Hal R. Varian. Letter to the Editor: Benford’s Law. The American Statistician, 26(3):65–66, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

                                        links.jstor.org/sici?sici=
                                        0003-1305%28197206%2926%
                                        3A3%3C62%3ALTTE%3E2.0.CO%
                                        3B2-Q.

Vardeman:1992:WAO

[Var92] Stephen B. Vardeman. What about the other intervals? The American Statistician, 46(3): 193–??, August 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:1998:TCD

[Var98] Stephen Vardeman. Teacher’s corner — discussion — A one-semester, laboratory-based, quality-oriented statistics curriculum for engineering students. The American Statistician, 52(3):242–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Viveros:1994:BNB

[VBB94] Roman Viveros, K. Balasubramanian, and N. Balakrishnan. Binomial and negative binomial analogues under correlated Bernoulli trials. The American Statistician, 48(3):243–??, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vonCollani:2010:RDF

[vC10] Elart von Collani. Response to “Desired and Feared — What Do We Do Now and Over the Next 50 Years” by Xiao-Li Meng. The American Statistician, 64(1):23–25, February 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verrill:2005:DFT

[VD05] Steve Verrill and Mark Durst.

The decline and fall of type II error rates. The American Statistician, 59(4):287–291, November 2005. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vanDorp:2002:STS

[vDK02] J. René van Dorp and Samuel Kotz. The standard twosided power distribution and its properties: With applications in financial engineering. The American Statistician, 56 (2):90–??, May 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n2/s3/p90.

Villa:2006:UMG

[VE06] Enrique R. Villa and Luis A. Escobar. Using moment generating functions to derive mixture distributions. The American Statistician, 60(1):75–80, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verrill:1999:WGC

[Ver99] Steve Verrill. When good confidence intervals go bad: Predictor sort experiments and ANOVA. The American Statistician, 53(1):38–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verrall:2001:ISA

[Ver01] Richard Verrall. Introduction statistics with applications in general insurance. The American Statistician, 55(1):84–??, February 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vonHippel:2004:BSM

[vH04] Paul T. von Hippel. Biases in SPSS 12.0 missing value analysis. The American Statistician, 58(2):160–164, May 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vos:2005:ECD

[VH05] Paul W. Vos and Suzanne Hudson. Evaluation criteria for discrete confidence intervals. The American Statistician, 59(2):

137–142, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vickers:2008:DAE

[Vic08] Andrew J. Vickers. Decision analysis for the evaluation of diagnostic tests, prediction models, and molecular markers. The American Statistician, 62(4): 314–320, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vidakovic:1994:ARL

[Vid94] Brani Vidakovic. All roads lead to Rome — even in the honeycomb world. The American Statistician, 48(3):234–236, August 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Venables:1989:CBM

[VM89] W. N. Venables and K. W. Morris. A connection between minimum variance unbiased estimation and maximum likelihood. The American Statistician, 43 (2):118–121, ???? 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Valz:1990:SDV

[VM90] Paul D. Valz and Ian A.

McLeod. A simplified derivation of the variance of Kendall’s rank correlation coefficient. The American Statistician, 44(1): 39–??, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vanBelle:1993:SSF

[vM93] Gerald van Belle and Donald C. Martin. Sample size as a function of coefficient of variation and ratio of means. The American Statistician, 47(3):165–??, August 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Velleman:1996:MTS

[VM96] Paul F. Velleman and David S. Moore. Multimedia for teaching statistics: Promises and pitfalls. The American Statistician, 50(3):217–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:2003:SES

[VM03] Stephen B. Vardeman and Max D. Morris. Statistics and ethics: Some advice for young statisticians. The American Statistician, 57(1): 21–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s3/p21.

Verbeke:2007:WCG

[VM07] Geert Verbeke and Geert Molenberghs. What can go wrong with the score test? The American Statistician, 61(4): 289–290, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Voss:1999:RMM

[Vos99] Daniel T. Voss. Resolving the mixed models controversy. The American Statistician, 53(4): 352–??, November 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

vonEye:1996:SNL

[vS96] Alexander von Eye and Christiane Spiel. Standard and nonstandard log-linear symmetry models for measuring change in categorical variables. The American Statistician, 50(4): 300–??, November 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Verbeke:2001:CLM

[VSL01] Geert Verbeke, Bart Spiessens, and Emmanuel Lesaffre. Conditional linear mixed models. The American Statistician, 55 (1):25–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Verbeke.htm.

Velleman:1993:NOI

[VW93] Paul F. Velleman and Leland Wilkinson. Nominal, ordinal, interval, and ratio typologies are misleading. The American Statistician, 47(1):65–??, February 1993. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vardeman:2010:ESM

[VWB^{+}10] Stephen B. Vardeman, Joanne R. Wendelberger, Tom Burr, Michael S. Hamada, Leslie M. Moore, J. Marcus Jobe, Max D. Morris, and Huaiqing Wu. Elementary statistical methods and measurement error. The American Statistician, 64(1): 46–51, February 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Vaupel:1985:HRS

[VY85] James W. Vaupel and Anatoli I. Yashin. Heterogeneity’s ruses: some surprising effects of selection on population dynamics. The American Statistician, 39 (3):176–185, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wainer:1996:DE

[Wai96] Howard Wainer. Depicting error. The American Statistician, 50(2):101–??, ???? 1996. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Waller:1995:DCF

[Wal95a] Lance Waller. Does the characteristic function numerically distinguish distribution? The American Statistician, 49(2): 150–??, ???? 1995. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Walsh:1995:EPN

[Wal95b] Dennis P. Walsh. Equating Poisson and normal probability functions to derive Stirling’s formula. The American Statistician, 49(3):270–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wallman:1998:DRE

[Wal98] Katherine K. Wallman. Data on race and ethnicity: Revising the Federal standard. The American Statistician, 52(1): 30–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/wallman.
html.

Waldman:2000:EDC

[Wal00a] Donald M. Waldman. Estimation in discrete choice models with choice-based samples. The American Statistician, 54(4): 303–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Waldman.htm.

Wallace:2000:BCS

[Wal00b] Edward C. Wallace. A Basic Course in Statistics, 4th edition. The American Statistician, 54(2):151–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Walker:2002:BRC

[Wal02] John H. Walker. Book review: Contemporary Business Statistics with Microsoft Excel. The American Statistician, 56 (3):250–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s19/p250.

Walker:2005:RAS

[Wal05] Helen M. Walker. The role of the American Statistical Association. The American Statistician, 59(1):68–71, February 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wann:1991:RE

[Wan91] Marie D. Wann. A. Ross Eckler, 1901–1991. The American Statistician, 45(3):173, August 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wand:1997:DBC

[Wan97] M. P. Wand. Data-based choice of histogram bin width. The American Statistician, 51(1): 59–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        tas/abstracts/wand.html.

Wang:2000:FIW

[Wan00] Y. H. Wang. Fiducial intervals: What are they? The American Statistician, 54(2):105–??, May 2000. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/Wang.htm.

Wand:2002:SCG

[Wan02] M. P. Wand. Statistical computing and graphics — vector differential calculus in statistics. The American Statistician, 56(1):55–62, 2002. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2010:CFP

[Wan10] Hsiuying Wang. Closed form prediction intervals applied for disease counts. The American Statistician, 64(3):250–256, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wardrop:1995:SPH

[War95] Robert L. Wardrop. Simpson’s paradox and the hot hand in basketball. The American Statistician, 49(1):24–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wardell:1997:SSI

[War97] Don G. Wardell. Small-sample interval estimation of Bernoulli and Poisson parameters. The American Statistician, 51(4): 321–??, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/wardell.html.

Warner:2002:RBT

[War02] Bradley A. Warner. Reviews of books and teaching materials — basic engineering data collection and analysis. The American Statistician, 56(1):76, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wasserman:2006:ACD

[Was06] Stanley Wasserman. Analyzing categorical data. The American Statistician, 60(3):291–292, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolkewitz:2010:TPS

[WASB10] Martin Wolkewitz, Arthur Allignol, Martin Schumacher, and Jan Beyersmann. Two pitfalls in survival analyses of time-dependent exposure: A case study in a cohort of Oscar nominees. The American Statistician, 64(3):205–211, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Watts:1991:WIS

[Wat91] Donald G. Watts. Why is introductory statistics difficult to learn? and what can we do to make it easier? The American Statistician, 45(4):290–??, November 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wasserstein:1991:BPL

[WB91] Ronald L. Wasserstein and John E. Boyer. Bounds on the power of linear rank tests for scale parameters. The American Statistician, 45(1):10–??, February 1991. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

White:2002:GTP

[WB02] Chris White and Scott Berry.

General — tiered polychotomous regression: Ranking NFL quarterbacks. The American Statistician, 56(1):10–21, 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wainer:2004:TSP

[WB04] Howard Wainer and Lisa M.

Brown. Two statistical paradoxes in the interpretation of group differences. The American Statistician, 58(2):117–123, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wall:2001:GEC

[WBT01] Melanie M. Wall, James Boen, and Richard Tweedie. General — an effective confidence interval for the mean with samples of size one and two. The American Statistician, 55(2): 102–105, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Waters:1987:OAM

[WC87] James R. Waters and Alexander J. Chester. Optimal allocation in multivariate, two-stage sampling designs. The American Statistician, 41(1):46–50, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiseman:1997:MLB

[WC97] Frederick Wiseman and Sangit Chatterjee. Major league baseball player salaries: Bringing realism into introductory statistics courses. The American Statistician, 51(4):350– ??, November 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/wiseman.html.

Wickham:2007:RG

[WC07] Hadley Wickham and Dianne Cook. R graphics. The American Statistician, 61(1):99–100, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Whiteside:2008:BEN

[WE08] Mary M. Whiteside and Mark E. Eakin. A better estimate of the number of valid signatures on a petition. The American Statistician, 62(1):17–21, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Weiss:1982:JW

[Wei82]  Lionel Weiss.  Jacob Wol-

fowitz (1910–1981). The American Statistician, 36(2):126–127, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Weir:2007:ISF

[Wei07] Bruce Weir. Introduction to statistics for forensic scientists. The American Statistician, 61 (3):274, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Weldon:1998:BST

[Wel98] Larry Weldon. Basic statistics: Tales of distributions. The American Statistician, 52 (1):86–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:1990:GPB

[Wes90] Peter H. Westfall. Graphical presentation of a basketball game. The American Statistician, 44(4):305–??, November 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2006:E

[Wes06] Peter Westfall. Editorial. The American Statistician, 60(1):1– 2, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2007:ER

[Wes07] Peter H. Westfall. Editor’s report. The American Statistician, 61(2):103–104, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2008:EN

[Wes08] Peter H. Westfall. Editor’s note. The American Statistician, 62(3):189, August 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilkinson:2009:HCH

[WF09] Leland Wilkinson and Michael Friendly. The history of the cluster heat map. The American Statistician, 63(2):179–184, May 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

West:2011:OCS

[WG11] Brady T. West and Andrzej T. Galecki. An overview of current software procedures for fitting linear mixed models. The American Statistician, 65(4): 274–282, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Woodward:2009:EFB

[WGHE09] Wayne A. Woodward, Henry L. Gray, James R. Haney, and Alan C. Elliott. Examining factors to better understand autoregressive models. The American Statistician, 63(4):

335–342, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:2007:BS

[WH07] Peter H. Westfall and Joseph M.

Hilbe. The black swan. The American Statistician, 61(3): 193–194, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiens:1999:WLN

[Wie99] Brian L. Wiens. When lognormal and gamma models give different results: A case study. The American Statistician, 53 (2):89–??, 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiegand:2006:SAM

[Wie06] Ryan E. Wiegand. Statistical analysis of medical data using SAS. The American Statistician, 60(4):345–346, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiens:2007:HSD

[Wie07] Brian Wiens. Handbook of statistical distributions with applications. The American Statistician, 61(4):370, November 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiegand:2008:DPA

[Wie08] Ryan E. Wiegand. Data preparation for analytics using SAS. The American Statistician, 62 (4):356–357, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilson:1992:SCS

[Wil92] William J. Wilson. Statistical consulting is scholarship. The American Statistician, 46(4): 295–??, November 1992. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wild:1994:EWV

[Wil94] C. J. Wild. Embracing the “wider view” of statistics. The American Statistician, 48(2): 163–??, May 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilkinson:1999:DP

[Wil99] Leland Wilkinson. Dot plots. The American Statistician, 53 (3):276–??, August 1999. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Williams:2002:BRA

[Wil02] Calvin Williams. Book review: Applied Multivariate Data Analysis (2nd Edition). The American Statistician, 56 (3):248–??, August 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v56n3/s17/p248.

Wilkinson:2006:RPC

[Wil06a] Leland Wilkinson. Revising the Pareto chart. The American Statistician, 60(4):332–334, November 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilks:2006:USE

[Wil06b] S. S. Wilks. Undergraduate statistical education. The American Statistician, 60(1):39–45, February 2006. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Willink:2008:UPN

[Wil08] Robin Willink. A unique property of the normal distribution associated with perturbing a general random variable. The American Statistician, 62 (2):144–146, May 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiorkowski:1990:LAB

[Wio90] John J. Wiorkowski. A lightly annotated bibliography of the publications of the American Statistical Association. The American Statistician, 44(2): 106–113, ???? 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Witmer:1997:DBB

[Wit97] Jeffrey A. Witmer. Discussion — Bayes for beginners? some reasons to hesitate. The American Statistician, 51(3):262–??, August 1997. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolter:2003:RUB

[WJM^{+}03] Kirk Wolter, Diana Jergovic, Whitney Moore, Joe Murphy, and Colm O’Muircheartaigh. Reliability of the uncertified ballots in the 2000 Presidential Election in Florida. The American Statistician, 57(1): 1–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s1/p1.

Williford:1995:MCT

[WKB^{+}95] William O. Williford, William F. Krol, Stephen F. Bingham, Joseph F. Collins, and David G. Weiss. The multicenter clinical trials coordination center statistician: “More Than a Consultant”. The American Statistician, 49(2):221–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wilson:1999:USC

[WL99] P. David Wilson and Patricia Langenberg. Usual and shortest confidence intervals on odds ratios from logistic regression. The American Statistician, 53(4):332–??, November 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Warner:1996:UNN

[WM96] Brad Warner and Manavendra Misra. Understanding neural networks as statistical tools. The American Statistician, 50(4):284–??, November 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Williamson:2009:RCO

[WMAY09] Patricia Pepple Williamson, Darcy P. Mays, Ghidewon Abay Asmerom, and Yingying Yang. Revisiting the classical occupancy problem. The American Statistician, 63(4):356–360, November 2009. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Withers:2007:RRM

[WN07] Christopher S. Withers and Saralees Nadarajah. A recurrence relation for moments of the noncentral chi square. The American Statistician, 61(4): 337–338, November 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Womer:1985:ILS

[Wom85] Norman Keith Womer. Identification and least squares in simultaneous equations. The American Statistician, 39(4): 295–297, ???? 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wonnacott:1998:BRU

[Won98] Thomas Wonnacott. Book review: The Ultimate Resource 2 . The American Statistician, 52(2):202–?? (or 200–??), May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wood:2008:CFA

[Woo08] Phil Wood. Confirmatory factor analysis for applied research. The American Statistician, 62(1):91–92, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

West:1998:STW

[WOR98] R. Webster West, R. Todd Ogden, and A. J. Rossini. Statistical tools on the World Wide Web. The American Statistician, 52(3):257–??, August 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

amstat.org/publications/
tas/west.html.

Wright:1990:WZD

[Wri90] Tommy Wright. When zero defectives appear in a sample: Upper bounds on confidence coefficients of upper bounds. The American Statistician, 44 (1):40–??, February 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wright:1992:LIR

[Wri92] Tommy Wright. Lagrange’s identity reveals correlation coefficient and straight-line connection. The American Statistician, 46(2):106–??, May 1992.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolkowicz:1979:ESI

[WS79] Henry Wolkowicz and George P. H. Styan. Extensions of Samuelson’s inequality. The American Statistician, 33(3): 143–144, ???? 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wolkowicz:1980:HSI

[WS80] H. Wolkowicz and G. P. H.

Styan. A history of Samuelson’s inequality. The American Statistician, 34(??):250, ???? 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Whitmore:1987:HDI

[WS87] G. A. Whitmore and V. Seshadri. A heuristic derivation of the inverse Gaussian distribution. The American Statistician, 41(4):280–281, ???? 1987. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wild:1993:CTP

[WS93] C. J. Wild and G. A. F. Seber. Comparing two proportions from the same survey. The American Statistician, 47 (3):178–181, August 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See correction [WS94].

Wild:1994:CCT

[WS94] C. J. Wild and G. A. F. Seber. Correction: “Comparing two proportions from the same survey” [Amer. Statist. 47 (1993), no. 3, 178–181]. The American Statistician, 48 (3):269, ???? 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [WS93].

Wendell:2001:LCI

[WS01] John P. Wendell and Josef Schmee. Likelihood confidence intervals for proportions in finite populations. The American Statistician, 55(1): 55–??, February 2001. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Wendell.htm.

Winkler:2002:SPR

[WSF02] Robert L. Winkler, James E. Smith, and Dennis G. Fryback. Statistical practice — the role of informative priors in zero-numerator problems: Being conservative versus being candid. The American Statistician, 56(1):1–4, 2002. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Webb:2010:PIA

[WSF10] Russell Y. Webb, Peter J. Smith, and Abdulla Firag. On the probability of improved accuracy with increased sample size. The American Statistician, 64(3):257–262, August 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:1993:IDP

[WSS93] Y. H. Wang, Jordan Stoyanov, and Qi Man Shao. On independence and dependence properties of a set of random events. The American Statistician, 47 (2):112–115, May 1993. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Waller:1992:PPC

[WT92] Lance A. Waller and Bruce W.

Turnbull. Probability plotting with censored data. The American Statistician, 46(1):5– ??, February 1992. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Waller:1995:ODF

[WTH95] Lance A. Waller, Bruce W.

Turnbull, and J. Michael Hardin. Obtaining distribution functions by numerical inversion of characteristic functions with applications. The American Statistician, 49(4):346–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wuensch:1998:USI

[Wue98] Karl L. Wuensch. Understanding statistics: An introduction for the social sciences. The American Statistician, 52(1):

87–??, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wulff:2003:BRF

[Wul03] Shaun S. Wulff. Book review: A First Course in Design and Analysis of Experiments. The American Statistician, 57(1): 66–??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n1/s12/p66.

Wulff:2007:SMM

[Wul07] Shaun S. Wulff. SAS for mixed models. The American Statistician, 61(2):184–185, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wuthrich:2008:SPM

[Wüt08] Mario V. Wüthrich. Statistical and probabilistic methods in actuarial science. The American Statistician, 62(4):360, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Westfall:1997:MTD

[WW97] Peter H. Westfall and Russell D. Wolfinger. Multiple tests with discrete distributions. The American Statistician, 51(1): 3–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.

amstat.org/publications/
tas/abstracts/westfall.html.

Wang:2011:UIN

[WW11] S. S. J. Wang and M. P.

Wand. Using Infer.NET for statistical analyses. The American Statistician, 65(2):115–126, May 2011. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wachsmuth:2003:GBP

[WWD03] Amanda Wachsmuth, Leland Wilkinson, and Gerard E. Dallal. Galton’s bend: A previously undiscovered nonlinearity in Galton’s family stature regression data. The American Statistician, 57(3): 190–??, August 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v57n3/s11/p190.

Wu:2004:ECD

[WY04] Samuel S. Wu and Mark C. K.

Yang. Evaluation of the current decision rule in figure skating and possible improvements. The American Statistician, 58 (1):46–??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s11/p46.

Xie:2003:SAD

[XN03] Rongrong Xie and Paul I. Nelson. Separation among distributions related by linear regression. The American Statistician, 57(1):33– ??, February 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v57n1/s5/p33.

Yu:2009:IPI

[YA09] Keming Yu and Abdallah Ally. Improving prediction intervals: Some elementary methods. The American Statistician, 63(1): 17–19, February 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yalta:2007:NRG

[Yal07] A. Talha Yalta. The numerical reliability of GAUSS 8.0. The American Statistician, 61 (3):262–268, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yang:1988:CLT

[Yan88] Shie-Shien Yang. A central limit theorem for the bootstrap mean. The American Statistician, 42(3):202–203, ???? 1988. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yuan:2010:CND

[YB10] Ke-Hai Yuan and Peter M. Bentler. Consistency of normaldistribution-based pseudo maximum likelihood estimates when data are missing at random. The American Statistician, 64 (3):263–267, August 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yu:2006:SES

[YCF06] Philip L. H. Yu, Jennifer S. K.

Chan, and Wing K. Fung. Statistical exploration from SARS. The American Statistician, 60 (1):81–91, February 2006. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yucel:2008:UCI

[YHZ08] Recai M. Yucel, Yulei He, and Alan M. Zaslavsky. Using calibration to improve rounding in imputation. The American Statistician, 62(2):125–129, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Young:2005:NSW

[You05] Allyn A. Young. National statistics in war and peace. The American Statistician, 59 (1):58–61, February 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yang:2001:GES

[YT01] Li Yang and Anastasios A. Tsiatis. General — efficiency study of estimators for a treatment effect in a pretest-posttest trial. The American Statistician, 55 (4):314–321, 2001. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yu:1994:SSP

[Yu94] Philip L. H. Yu. A simple statistical project: Image reconstruction. The American Statistician, 48(1):30–??, February 1994. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Yozgatligil:2009:RMS

[YW09] Ceylan Yozgatligil and William W. S. Wei. Representation of multiplicative seasonal vector autoregressive moving average models. The American Statistician, 63(4):328–334, November 2009. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zamba:2008:SDQ

[Zam08] Gideon (K. D.) Zamba. Statistical development of quality in medicine. The American Statistician, 62(4):359–360, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zelterman:1995:ETS

[ZCM95] Daniel Zelterman, Ivan Siu-Fung Chan, and Paul W. Mielke, Jr. Exact tests of significance in higher dimensional tables. The American Statistician, 49(4):357–??, ???? 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zehna:1991:PSI

[Zeh91] Peter W. Zehna. On proving that simX and S^{2} are independent. The American Statistician, 45(2):121–??, May 1991. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zellner:1982:JN

[Zel82] Arnold Zellner. Is Jeffreys a “necessarist”. The American Statistician, 36(1):28–30, ???? 1982. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zellner:1985:CPD

[Zel85] Arnold Zellner. Correction:

“On prior distributions for binary trials” [Amer. Statist. 38 (1984), no. 4, 244–251, MR 87c:62004a] by S. Geisser. The American Statistician, 39(2): 158, ???? 1985. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). See [Gei84].

Zellner:1988:OIP

[Zel88] Arnold Zellner. Optimal information processing and Bayes’s theorem. with comments and a reply by the author. The American Statistician, 42(4): 278–284, ???? 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zerbe:1978:FTG

[Zer78] Gary O. Zerbe. On Fieller’s theorem and the general linear model. The American Statistician, 32(3):103–105, ???? 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2000:OSC

[ZG00] Xiao-Hua Zhou and Sujuan Gao. One-sided confidence intervals for means of positively skewed distributions. The American Statistician, 54 (2):100–??, May 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Zhou.htm.

Ziegel:1990:ASA

[ZH90] Eric R. Ziegel and J. Stuart Hunter. American Statistical Association’s section on physical and engineering sciences: Past, present, and future. The American Statistician, 44(4): 267–??, November 1990. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2000:PSC

[Zha00] Li-Chun Zhang. Post-stratification and calibration—A synthesis. The American Statistician, 54 (3):178–??, August 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/Li-Chun.htm.

Zhang:2007:SMS

[Zha07a] Lingyun Zhang. Sample mean and sample variance. The American Statistician, 61(2): 159–160, May 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhang:2007:FB

[Zha07b] Ying Zhang. Fundamentals of biostatistics. The American Statistician, 61(2):183, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2001:TCU

[Zho01]  Mai Zhou.  Teacher’s cor-

ner — understanding the Cox regression models with timechange covariates. The American Statistician, 55(2):153–155, 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhu:2008:KE

[Zhu08] Mu Zhu. Kernels and ensembles. The American Statistician, 62(2):97–109, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zimmerman:2000:SCS

[Zim00] Dale L. Zimmerman. Statistical computing software reviews: Viewing the correlation structure of longitudinal data through a PRISM. The American Statistician, 54(4): 310–??, November 2000. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.

                                        amstat.org/publications/
                                        tas/Zimmerman.htm.

Zimmerman:2007:SPD

[Zim07] M. Bridget Zimmerman. SAS programming and data analysis: A theory and programdriven approach. The American Statistician, 61(3):277–278, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2009:RMR

[ZL09] Hua Zhou and Kenneth Lange. Rating movies and rating the raters who rate them. The American Statistician, 63(4): 297–307, November 2009. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zheng:1994:CCR

[ZM94] Qi Zheng and James H. Matis. Correlation coefficient revisited. The American Statistician, 48 (3):240–??, August 1994. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhu:2011:GSM

[ZNY11] Li Zhu, Liyun Ni, and Bin Yao. Group sequential methods and software applications. The American Statistician, 65 (2):127–135, May 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:1999:CSP

[ZPH99] Xiao-Hua Zhou, Anthony J. Perkins, and Siu L. Hui. Comparisons of software packages for generalized linear multilevel models. The American Statistician, 53(3):282–??, August 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zeng:2011:PFD

[ZPM^{+}11] Chan Zeng, Zhaoxing Pan, Samantha MaWhinney, Anna E. Barón, and Gary O. Zerbe. Permutation and F distribution of tests in the multivariate general linear model. The American Statistician, 65(1):31–36, February 2011. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zhou:2010:NBI

[ZR10] Xiang Zhou and Jerome P. Reiter. A note on Bayesian inference after multiple imputation. The American Statistician, 64 (2):159–163, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zubizarreta:2011:MSS

[ZRK^{+}11] José R. Zubizarreta, Caroline E. Reinke, Rachel R. Kelz, Jeffrey H. Silber, and Paul R. Rosenbaum. Matching for several sparse nominal variables in a case-control study of readmission following surgery. The American Statistician, 65(4): 229–238, November 2011. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zuo:2010:CIO

[Zuo10] Yijun Zuo. Is the t confidence interval ^{−}X ± t_{α}(n − 1)(s/\sqrt{\{n\}}) optimal? The American Statistician, 64(2): 170–173, May 2010. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).